### v2

# MILU: A Multi-task Indic Language Understanding Benchmark

Abstract
Evaluating Large Language Models (LLMs) in low-resource and linguistically diverse languages remains a significant challenge in NLP, particularly for languages using non-Latin scripts like those spoken in India. Existing benchmarks predominantly focus on English, leaving substantial gaps in assessing LLM capabilities in these languages. We introduce MILUMulti-task Indic Language Understanding Benchmark—a comprehensive evaluation benchmark designed to address this gap. MILU spans 8 domains and 42 subjects across 11 Indic languages, reflecting both general and culturally specific knowledge. With an India-centric design, MILU incorporates material from regional and state-level examinations, covering topics such as local history, arts, festivals, and laws, alongside standard subjects like science and mathematics. We evaluate over 45 LLMs, and find that current LLMs struggle with MILU, with GPT-4o achieving the highest average accuracy at 72%. Open multilingual models outperform language-specific fine-tuned models, which perform only slightly better than random baselines. Models also perform better in high-resource languages as compared to low-resource ones. Domain-wise analysis indicates that models perform poorly in culturally relevant areas like Arts & Humanities and Law & Governance compared to general fields like STEM. To the best of our knowledge, MILU is the first of its kind benchmark focused on Indic languages, serving as a crucial step towards comprehensive cultural evaluation. All code, benchmarks, and artifacts will be made publicly available to foster open research.
MILU: A Multi-task Indic Language Understanding Benchmark
Sshubam Verma1  Mohammed Safi Ur Rahman Khan1,2 Vishwajeet Kumar3Rudra Murthy3Jaydeep Sen3 1Nilekani Centre at AI4Bharat  2Indian Institute of Technology, Madras 3IBM Research, India Correspondence: {sshubamverma, safikhan}@ai4bharat.org, {vishk024, rmurthyv}@in.ibm.com [Uncaptioned image] https://huggingface.co/datasets/ai4bharat/MILU [Uncaptioned image] https://github.com/AI4Bharat/MILU
1 Introduction
Recent advancements in Large Language Models (LLMs) have reshaped the field of NLP, by enabling these models to perform a variety of tasks across diverse domains Doddapaneni et al. (2023); OpenAI et al. (2023); Team et al. (2024a); Anthropic . While many LLMs now claim to support multiple languages, there is still a huge discrepancy in their performance in English and other languages Liu et al. (2024). Particularly languages using non-Latin scripts, such as those in India, are affected the most by this discrepancy Ahuja et al. (2023). One key reason for this is the absence of high-quality benchmarks for these languages. Well-designed benchmarks are crucial in driving model development by revealing limitations and guiding improvements McIntosh et al. (2024). However, most existing benchmarks focus primarily on English, leaving significant gaps in evaluating LLMs capability in low-resource and linguistically diverse languages.
Figure 1: Average performance of all the evaluated models on MILU. The closed models are shown in Orange, the open models are shown in Green, and the language-specific models are shown in Blue.
India’s diversity presents a unique challenge for these models. With over 1.4 billion people speaking more than 120 languages and around 19,500 dialects across 28 states Javed et al. (2024), many of which are underrepresented in NLP research, the need for not just linguistic but culturally appropriate benchmarks becomes urgent Doddapaneni et al. (2023); Singh et al. (2024b). Standard benchmarks like MMLU Hendrycks et al. (2021) and AGIEval Zhong et al. (2023), while useful for evaluating general world knowledge, falls short in capturing these intricacies of India. Each state in India has its own history, traditions, festivals, and art forms, forming a rich cultural mosaic. Translating existing English benchmarks into Indian languages fails to capture this knowledge, which is required for real-world applications. Given the increasing deployment of LLMs in tasks that directly impact local populations, the need for a benchmark that evaluates both linguistic competence and cultural understanding has become more pressing than ever.
In this work, we introduce MILU-Multi-task Indic Language Understanding Benchmark, a comprehensive evaluation dataset designed to address these gaps. MILU spans 8 domains and 42 subjects across 11 Indic languages, reflecting both general and culturally specific knowledge. We designed MILU with an India-first perspective by collecting questions from various national, state, and regional exams. These questions include culturally relevant subjects such as local history, arts, festivals, and laws, alongside traditional academic subjects like science and mathematics. Following previous efforts Hendrycks et al. (2021); Zhong et al. (2023), we create this benchmark by collecting questions from over 1500 competitive exams from India. We focus on region-specific exams to authentically capture local knowledge in the respective language.
We evaluate 45 different LLMs - a mix of closed proprietary, open-source, and language-specific models- on MILU. Our findings suggest that models struggle with MILU, with GPT-4o achieving the highest average accuracy at 72%. Interestingly, open multilingual models outperform language-specific models, which only achieve slightly better than random scores. Our analysis of in-context learning reveals that adding more examples improves performance in base models, but the effect on instruct models remains inconclusive. We also explore how performance scales with the number of parameters, finding significant improvements as model size increases. Our domain-wise analysis reveals that models perform poorly in culturally relevant areas, such as Arts & Humanities and Social Sciences, compared to more general fields like STEM. All the artifacts will be released publicly.
2 Related Works
Large Language Models (LLMs): Recent LLMs, both proprietary-GPT-4o, GPT-4o-mini, Claude-3, and Gemini Team et al. (2024a)-and open-source-the Llama-3 Dubey et al. (2024) and Gemma Team et al. (2024b) series-have demonstrated significant improvements across various tasks and benchmarks Chiang et al. (2024); Wang et al. (2024). While these models are primarily trained in English, many claim to include a reasonable amount of multilingual data in their pretraining corpus Team et al. (2024b); Dubey et al. (2024); Aryabumi et al. (2024). Additionally, significant progress has been made in developing language-specific models Ustun et al. (2024); Nguyen et al. (2023), including for Indic languages Gala et al. (2024); Balachandran (2023); Kohli et al. (2023). Most of these models are built on top of stronger English base models by (optionally) either continually pretraining on smaller, language-specific datasets Zhao et al. (2024), or by language-specific instruction fine-tuning Ghosh et al. (2024). In this work, we conduct a comprehensive evaluation of the performance of these models on  MILU.
LLM Evaluation Benchmarks: Over the years, various benchmarks have been developed to evaluate the performance of large language models (LLMs). Recent benchmarks such as MMLU Hendrycks et al. (2021), MMLUPro Wang et al. (2024), AGIEval Zhong et al. (2023), BigGenBench Kim et al. (2024), and HellaSwag Zellers et al. (2019) assess these models across a wide range of tasks. However, these benchmarks primarily focus on English, and progress on multilingual benchmarks has been comparatively slower Doddapaneni et al. (2023); Kumar et al. (2022). Some popular multilingual benchmarks include OMGEval Liu et al. (2024), XTREME Hu et al. (2020), and XQuAD Artetxe et al. (2020), though these are largely limited to simple natural language understanding (NLU) tasks. As a result, previous studies often rely on translations of popular English benchmarks to evaluate performance in other languages Dubey et al. (2024); Gala et al. (2024). This approach is suboptimal, as it fails to account for cultural nuances and concepts unique to specific languages.
Indic LLM Evaluation Benchmarks: The first major Indic language benchmarks, IndicGLUE Kakwani et al. (2020) and IndicNLG-Benchmark Kumar et al. (2022), cover 11 languages, focusing on various language understanding and generation tasks, respectively. IndicXTREME Doddapaneni et al. (2023) extends these efforts to include all 22 scheduled Indian languages for NLU evaluation. More recently, IndicGenBench Singh et al. (2024b) offers a more comprehensive evaluation of multilingual language generation by consolidating multiple generation tasks. In addition, the Indic-QA Benchmark Singh et al. (2024a) assesses context-based question-answering (QA) performance in 11 Indian languages, while L3Cube-IndicQuest Rohera et al. (2024) explores regional knowledge through a translation-based approach to data creation. More recently, Pariksha Watts et al. (2024) evaluated more than 25 Indic models and created dynamic leaderboards by collecting more than 90,000 human preferences. Other initiatives, such as Airavata Gala et al. (2024) and the Indic-LLM-Leaderboard Kolavi (2024), focus on translating existing English benchmarks into Indian languages. In contrast, MILU addresses this limitation by centering on India-specific topics and questions, providing a more culturally relevant and context-aware evaluation.
3 MILU: The IndicMMLU Benchmark
In this section, we describe the collection process (§3.1), the cleaning and filtering process (§3.2), and present the analysis (§3.3).
3.1 Questions Curation
MILU is a large, multi-domain test set containing multiple-choice based questions (MCQs) taken from over 42 subjects with an emphasis on India-specific knowledge. This benchmark covers many domains, including Mathematics, Science, Social Sciences, Humanities, Arts, Business Studies, and Law, among others. MILU is designed as a culturally relevant benchmark to assess general problem-solving abilities and India-specific knowledge. These questions were sourced following an approach similar to AGIEval Zhong et al. (2023), collecting the questions from various public exams taken by individuals intending to either pursue higher studies or seek career advancements, such as qualification tests and national and state-level civil services exams, among others.
We gathered exam-specific questions by scraping various online exam portals that offer previously released question papers from various exams in multiple different languages. These portals typically tag questions manually with topic names and language details, and subject experts ensure the accuracy of the answers. Our benchmark includes questions from over 40 different types of exams conducted both at the national and state levels over recent years. Regional state exams are particularly valuable as they cover various state-level topics and emphasize the official language of each state. Further details about the different exams included in our collection are provided in the Appendix A.
In total, we collected more than 150K questions across 11 Indian Languages- Bengali (bn), Gujarati (gu), Hindi (hi), Kannada (kn), Malayalam (ml), Marathi (mr), Odia (or), Punjabi (pn), Tamil (ta), Telugu (te), and English (en)-spanning 42 diverse subjects. English questions are also included as these often address Indian culture-specific content, which is notably missing from existing popular benchmarks. Table 1 presents a brief comparison of the type of questions present in MMLU Hendrycks et al. (2021) and MILU.
MMLU MILU
Which principle was established by the Supreme Court’s decision in Marbury v. Madison? Which of the decisions was taken by Maharashtra government on 24th November 2001 that had a far-reaching impact on the education field?
At breakfast, lunch, and dinner, Joe randomly chooses with equal probabilities either an apple, an orange, or a banana to eat. On a given day, what is the probability that Joe will eat at least two different kinds of fruit? Shyam’s monthly income is Rs. 12,000. He saves Rs. 1200. Find the percent of his savings and his expenditure.
The Space Shuttle orbits 300 km above Earth’s surface; Earth’s radius is 6,400 km. What is the gravitational acceleration experienced by the Space Shuttle? FASTag at the Toll Gates uses which waves?
The primary goal of the Gramm-Rudman Acts of 1985 and 1987 was to Under the Pradhan Mantri Kisan Samman Nidhi Scheme, the financial help provided to the small and marginal farmers is ___________ .
A guitar string creates a sound wave of known frequency. Which of the following describes a correct and practical method of measuring the wavelength of the sound wave with a meterstick? In a Sitar, which type of sound vibrations are produced?
Table 1: Comparison of some question from MMLU Hendrycks et al. (2021) with similar questions from MILU. The Red highlights the global concepts covered in MMLU, while Green highlights the India-centric concepts covered in MILU.
3.2 Data Cleaning and Filtering
Despite our best efforts to maintain the quality of questions collected, some amount of noise or errors may still be present. To address potential noise in the questions, we employ multiple layers of manual and automated cleaning filters. Initially, we manually review a large sample of questions to detect and eliminate potential sources of noise. During the collection process, we exclude any reading-comprehension-style questions, images-based questions, and those with more than four answer options to ensure uniformity and consistency. To remove incorrect language entries, we utilize a combination of IndicLID Madhani et al. (2023) and Unicode-based filtering Khan et al. (2024), ensuring that the questions are in the correct language. To further refine the dataset, we remove any duplicate questions to retain only the unique ones. As a final step, we manually verify a sample of questions from each language to ensure accuracy and correct any remaining errors.
Figure 2: Distribution of the number of questions across different domains, averaged across all languages. Refer to Section (§3.3) for more details.
Upon examination, we found that approximately 45% of questions were accurately labeled with a topic name, while the remaining questions lacked this information. To address this issue, we first translate the untagged questions into English using IndicTrans2 Gala et al. (2023) and then prompt GPT-4o-mini model to assign an appropriate topic name to the question. Finally, in total, we get around 20K tags. However, these tags are highly fine-grained, often having a heavy overlap. To organize them, we embed the tags using the NV-Embed-v2 Lee et al. (2024) model and apply K-means clustering to group tags into 50 clusters. We manually review these clusters and assign appropriate subject labels. Following the manual merging of related clusters, we determine 42 distinct subject names, which fall into eight main domains: Arts and Humanities, Social Sciences, Environmental Sciences, Law and Governance, Health and Medicine, Mathematics and Sciences, Engineering and Technology, and Business Studies. The final distribution of domains within MILU is shown in Figure 2. Detailed analysis of question distribution by topic & language is provided in Appendix B.
Finally, we observed that some topics in certain languages had less than 100 questions. To ensure thorough evaluation across all subjects and languages, we aimed to have at least 100 questions per subject in each language. For subjects with insufficient questions, we sampled questions from the English set from that subject and translated them into the required language using GPT-4o. We chose GPT-4o over specialized translation models for their ability to remain task-aware during translation Ahuja et al. (2024), ensuring the translated content aligns with the intent of the question.
In total, we release around 85K questions across 42 subjects across 8 domains in 11 languages, capping each subject-language pair at 500 questions for feasible evaluations.
lang
Total Qs
Total
Translated Qs
Avg Words
Per Q
bn 7138 1601 15.72
gu 5327 2755 16.69
hi 15450 115 20.63
kn 6734 1522 12.83
ml 4670 1534 12.82
mr 7424 1235 18.8
or 5025 1452 15.63
pa 4363 2341 19.9
ta 7059 1524 13.32
te 7847 1298 16.13
en 14036 - 22.01
total 85073 15377 16.77
Table 2: Overall statistics of MILU. Refer to Section (§3.3) for more details.
3.3 Data Analysis
Table 2 shows the overall statistics of MILU. Of the total 85K questions, only 18% of questions are translated from English, with the remainder preserved in their original source languages. The average question length varies across languages, with languages such as kn, ml, and ta having shorter word counts due to their agglutinative nature.
As shown in Figure 2, MILU  consists of eight domains.
Arts & Humanities: This domain covers topics such as Indian art, literature, dance, festivals, and architecture. Its content varies significantly by language, as it is highly dependent on regional culture.
Science & Math: Includes topics in physics, chemistry, biology, and mathematics, with references to Vedic Math, Ayurveda, and ancient as well as modern scientific achievements.
Health & Medicine: Covers public health policies, modern healthcare, traditional Ayurveda and Unani medicine, and health-related government initiatives.
Business Studies: Focuses on entrepreneurship, trade, and economic policies such as Make in India. It also covers topics like taxation, MSMEs, and global trade.
Law & Governance: Includes topics on the constitution, governance, judiciary, public administration, fundamental rights, and various governmental schemes, including local governance policies.
Environmental Sciences: Covers topics related to biodiversity, environmental policies, and both local and national environmental initiatives.
Social Sciences: Explores topics such as history, geography, and politics. The content in this domain is region-specific and varies by language.
Engineering & Technology: Includes discussions on modern developments in India, such as IT, telecommunications, infrastructure, and space technology, with a focus on government policies.
Figure 2 shows the domain-wise statistics of MILU. We see a fairly large number of questions in the culturally relevant domains of Arts & Humanities and Social Sciences. Detailed statistics per language can be found in the Appendix B.
4 Experimental Setup
We evaluate 42 different models on MILU, including large proprietary models, open-source multilingual models, and popular fine-tuned models specific to Indic languages. Both the base versions and instruction fine-tuned variants of these models, wherever applicable, are evaluated to measure the improvements gained from fine-tuning. All models, except for proprietary models and Llama-3.1-405B, are tested under 0-shot, 1-shot, and 5-shot setups. We maintain a separate validation set of approximately 9,000 questions to serve as examples for few-shot evaluations.
For non-API-based models, we use the LM-Evaluation-Harness Gao et al. (2024); Biderman et al. (2024) to ensure clean and reproducible evaluations. We use the log-likelihood method, where the probability of a given output string is computed by conditioning it on some provided input Brown et al. (2020). Specifically, the log-likelihood of an answer $(a)$ given the question $(x)$, i.e., $logP(a|x)$, is calculated by concatenating the answer $(a)$ with question $(x)$, and then summing up the log probabilities, of each target token. For multiple choice questions, given $k$ possible answer strings, we select the answer string $(a_{i})$ with the highest conditional log probability, i.e., $argmax(logP(a_{1}|x),...,logP(a_{k}|x))$.
The API-based models are evaluated using the generative approach due to the lack of support for prompt log probabilities. We explicitly prompt these models to generate the correct response in a structured JSON format to simplify response parsing. Due to the high costs involved, these models are evaluated only in the zero-shot setup.
5 Results and Discussions
Model bn gu hi kn ml mr or pa ta te en avg
Closed Models
GPT-4o 75.32 72.06 77.16 74.85 70.96 71.79 67.68 71.85 67.16 70.51 80.66 72.73
GPT-4o-mini 59.13 55.94 62.04 55.59 53.27 54.66 46.12 56.7 49.49 52.04 69.4 55.85
Gemini-1.5-Pro 68.11 66.93 69.93 68.83 64.95 65.58 61.68 64.58 65.13 63.75 76.46 66.90
Gemini-1.5-Flash 60.71 60.31 63.86 63.25 58.65 58.71 55.98 61.12 57.06 56.82 68.53 60.45
Krutrim-spectre-v2 42.74 41.61 44.5 44.17 40.66 39.44 35.18 39.81 38.82 38.17 45.83 40.99
Open Multilingual Models
$<$ 4B Models
meta-llama/Llama-3.2-1B 25.28 25.49 26.90 25.24 25.05 26.46 24.93 24.82 25.72 25.75 34.08 26.34
meta-llama/Llama-3.2-1B-Instruct 25.79 25.71 25.31 25.85 25.09 25.56 24.53 25.02 26.51 26.09 25.79 25.57
sarvamai/sarvam-2b-v0.5 29.83 29.06 31.15 28.69 25.87 28.77 26.85 28.72 27.78 28.30 34.47 29.04
sarvamai/sarvam-1 32.99 33.18 32.90 32.98 30.34 31.43 30.48 32.22 30.01 31.69 35.63 32.17
google/gemma-2-2b 29.77 28.89 32.26 29.67 29.05 30.98 25.63 30.48 30.06 28.43 50.18 31.40
google/gemma-2-2b-it 30.58 29.34 34.16 28.70 29.16 32.03 26.80 28.99 30.40 29.29 51.85 31.94
meta-llama/Llama-3.2-3B 32.74 30.94 34.27 31.41 28.97 31.94 27.98 30.30 30.54 29.69 49.59 32.58
meta-llama/Llama-3.2-3B-Instruct 32.97 32.77 30.42 31.61 30.10 28.35 26.64 32.34 30.99 31.17 31.68 30.82
nvidia/Nemotron-4-Mini-Hindi-4B-Base 31.33 25.83 50.53 27.53 27.10 38.60 26.68 26.38 27.96 28.17 50.54 32.79
7B to 27B Models
Telugu-LLM-Labs/Navarasa-2.0 36.01 36.88 39.51 39.42 36.05 36.61 35.00 36.35 36.13 36.25 51.04 38.11
neulab/Pangea-7B 42.14 37.11 41.97 34.18 34.62 36.98 32.89 36.35 32.93 36.17 55.87 38.29
CohereForAI/aya-23-8B 28.04 27.01 35.96 27.39 28.45 30.67 27.20 26.99 28.91 26.30 45.14 30.19
meta-llama/Llama-3.1-8B 37.60 35.44 42.29 36.87 34.53 37.51 32.63 36.16 35.57 34.63 58.83 38.37
meta-llama/Llama-3.1-8B-Instruct 34.19 33.30 35.37 35.69 34.00 32.26 29.91 32.89 33.65 33.08 45.66 34.55
google/gemma-2-9b 52.60 50.61 55.52 53.28 49.50 49.39 40.01 48.81 48.81 47.99 68.87 51.40
google/gemma-2-9b-it 46.07 39.02 44.49 43.31 41.11 38.02 34.68 40.29 38.68 38.91 63.08 42.51
google/gemma-2-27b 62.86 58.96 63.28 62.54 59.35 59.17 45.11 59.15 57.27 57.00 74.00 59.88
google/gemma-2-27b-it 59.47 55.04 60.49 57.55 53.36 54.43 45.65 54.84 54.37 53.57 72.90 56.52
$>$ 27B Models
CohereForAI/aya-23-35B 32.83 30.05 42.96 31.37 32.35 36.71 31.14 27.50 34.62 29.19 52.27 34.64
meta-llama/Llama-3.1-70B 50.53 47.09 47.51 49.01 47.94 46.88 47.7 47.55 47.66 46.33 55.33 48.5
meta-llama/Llama-3.1-70B-Instruct 61.59 58.66 57.92 61.50 57.83 53.05 52.00 59.54 56.45 56.26 53.41 57.11
meta-llama/Llama-3.1-405B 64.31 59.77 50.36 62.14 62.20 49.70 53.97 60.06 57.20 58.40 60.86 58.09
Table 3: Evaluation results of all the Closed and Open Multilingual models supporting all languages on MILU. We report 5-shot accuracies for all open models (except for Llama-3.1-70B, Llama-3.1-70B-Instruct & Llama-3.1-405B for which we report 0-shot accuracy) with the accuracy averaged across all the domains per language. Higher values indicate better model performance for the given language. Refer Section (§5.1) for more details.
In this section, we discuss the results and our findings across all the experiments conducted.
5.1 How do models supporting all languages perform?
Table 3 shows the performance of various models supporting all the languages in the 5-shot setup. Among these, GPT-4o emerges as the strongest model, consistently outperforming all competitors across languages, with an average accuracy of 72.73%. Closed proprietary models, generally achieve higher performance than the open-source alternatives. Among the open-source models, Llama-3.1-70B shows the best performance.
Performance is notably higher in high-resource languages- en, hi, bn -as compared to the low-resource ones. Interestingly, the models also show low accuracies in agglutinative languages, such as ta, te. Notably, models specifically trained for Indian languages such as Krutrim-spectre-v2 and sarvam-2b-v0.5, perform slightly below their English counterparts of comparable scale. Additionally, while several models achieve better results in English, their performance in other languages is often nearly at the random baseline.
5.2 How do language-specific fine-tuned models perform?
lang Model
Business
Studies
Engg.
& Tech
Social
Sciences
Env.
Sciences
Law &
Governance
Math &
Science
Arts &
Humanities
Health &
Medicine
Avg
hi OpenHathi-7B-Hi 29.74 29.46 28.69 27.75 31.52 28.60 27.56 30.58 29.24
Airavata 27.97 29.03 29.89 27.15 27.44 28.10 26.91 33.00 28.69
ProjectIndus 26.19 25.69 26.89 25.04 23.05 25.40 25.44 25.56 25.41
Gajendra-v0.1 29.30 29.15 25.65 29.30 27.21 29.34 28.38 29.47 28.48
open-aditi-v6-llama3 29.29 32.74 31.02 30.15 33.96 32.42 31.96 33.58 31.89
AryaBhatta-GemmaGenZ-Merged 39.39 39.15 38.48 41.14 39.28 40.26 35.87 48.08 40.21
te telugu-llama-7b-instruct-v0.1 32.12 26.67 26.93 24.23 25.34 25.93 25.63 24.52 26.42
Telugu-Llama2-7B-v0-Base 22.17 27.43 25.25 27.77 25.53 26.87 25.32 27.45 25.97
Telugu-Llama2-7B-v0-Instruct 25.30 21.07 24.99 23.54 23.62 26.43 27.55 27.86 25.05
ka Kan-LLaMA-7B-base 28.88 29.19 30.50 27.35 28.30 26.49 29.25 30.50 28.81
Ambari-7B-base-v0.1 31.49 27.08 29.26 28.62 27.82 27.88 27.94 31.00 28.89
Ambari-7B-Instruct-v0.1 28.26 27.91 24.92 26.90 23.54 28.46 26.95 27.00 26.74
mr Misal-7B-base-v0.1 27.09 26.56 25.28 26.70 26.36 25.59 26.79 23.89 26.03
Misal-7B-instruct-v0.1 24.82 20.95 21.51 22.69 20.62 23.78 23.06 22.10 22.44
ta tamil-llama-7b-instruct-v0.2 22.17 23.53 25.03 24.64 28.53 26.29 25.29 22.57 24.76
ml malayalam-llama-7b-instruct-v0.1 30.00 28.44 23.43 27.87 25.00 26.37 23.33 20.00 25.56
Table 4: Evaluation results of all the language-specific fine-tuned models on MILU. We report domain level 5-shot accuracies for all the models on the language supported by the model. Higher values indicate better model performance for the given domain. Refer to Section (§5.2) for more details.
We evaluate around 16 Indic language LLMs on MILU. These models are primarily built by adapting English LLMs, such as Llama-2-7B, by first continually pretraining on small amount Indic language data, followed by optionally instruction fine-tuning them. As seen from Table 4, across languages, the models exhibit average performance comparable to random baselines, with minimal variations among them.
Among the evaluated models, the AryaBhatta-GemmaGenZ-Vikas model stands out, performing better than its counterparts. When analyzed across domains, the models generally perform worse in Arts, Humanities, and Social Sciences than in STEM subjects.
Figure 3: Comparison of Base and Instruct models averaged across all languages for varying number of in-context examples. We plot the average accuracies of the Gemma and Llama series of models, highlighting the performance trend as the number of in-context examples increases. Refer to Section (§5.3) for more details.
(a) Base Models
(b) Instruct Models
Figure 4: Evaluation results of Base models (4(a)) and Instruct models (4(b)) on the different domains supported in MILU. The plot shows the average 5-shot accuracies across all languages for various models. Refer to Section (§5.5)
5.3 How do the models respond to In-Context examples?
We compare the performance of different Base and Instruct models across zero, one, and five-shot setups. As shown in Figure 3, the performance of base models consistently improves with an increasing number of in-context examples, with the 5-shot setup yielding the best results. In contrast, Instruct models exhibit more varied behavior, where models either stagnate or even degrade in performance. This also aligns with expectations, as Instruct models are specifically fine-tuned as conversation assistants and may not respond well to the few-shot in-context examples format. This also correlates with results in Table 3 where most instruction finetuned variants performs worse on English itself. This essentially implies the inferior performance is less about the language shift but due to the change of tasks during instruction finetuning and inference task for MILUevaluation. Understandably, the instruction finetuned variants performs poor on the overall average across languages as well.
5.4 How does the performance vary with scale?
Figure 5: Comparison of performance of Gemma and Llama family of models across different parameter scales. We plot the zero-shot average accuracies of all models across languages. Refer to Section (§5.4) for more details.
We evaluate the Llama and Gemma family of models, ranging from 1B to 405B parameters, to analyze how performance scales with model size. Figure 5 shows that the model performance improves significantly with increasing scale. Notably, instruction-tuned models in the Llama family show more substantial improvements as compared to those in the Gemma family.
5.5 How do the models perform on different domains?
We analyze the performance of various base and instruct models across multiple domains and languages. Similar trends to those in Section (§5.2) are observed where the open models perform poorly in domains specific to Indian culture—such as Arts & Humanities, Social Sciences, and Law & Governance—but demonstrate higher performance in STEM fields. This suggests that the training corpora for these models lack sufficient culturally specific data. Bridging this gap requires a more inclusive data distribution that ensures equitable representation of all cultures and languages.
lang Model CPT IFT 0 1 5
hi OpenHathi Y N -0.36% -0.28% 0.59%
Airavata Y Y 1.54% 2.14% 2.93%
Gajendra-v0.1 Y Y 0.65% 0.85% 3.14%
kn Ambari-base Y N 0.75% 0.86% 2.07%
Ambari-Instruct Y Y 0.70% 0.46% 1.84%
Kan-7B-base Y N 0.76% 1.85% 2.11%
te telugu-llama-instruct Y Y 0.83% 0.56% 1.31%
Telugu-7B-v0-Base Y N 0.43% -0.79% 0.81%
Telugu-7B-v0-Instruct Y Y -1.02% -0.16% 0.32%
mr Misal-7B-base-v0.1 Y N 0.07% -0.44% -1.11%
Misal-7B-instruct-v0.1 Y Y -2.45% -2.04% -2.43%
ta tamil-llama-instruct Y Y -0.06% 0.22% 0.08%
ml malayalam-llama-instruct Y Y 0.62% 1.05% 0.45%
Table 5: Performance differences of Indian language models compared to the Llama-2-7B baseline at different in-context example settings. Each value shows the relative improvement or decline (difference) over the baseline, with a +ve value indicating improvement and a -ve value indicating a drop in the performance. Refer to Section (§5.6) for more details.
5.6 Does language adaptation on models help?
As most Indic LLMs are built on English base models like Llama-2-7b, we assess the impact of language adaptation on their performance. Table 5 compares language-specific models with the original Llama-2-7b, and instruction-tuned models with Llama-2-7b-chat. Our findings show minimal gains, with some models even underperforming post-adaptation.
Given the varied training data, this is not a direct "apples-to-apples" comparison but highlights key challenges. We conjecture that limited performance gains may result from small language-specific datasets and reliance on parameter-efficient methods like LoRA Hu et al. (2022). Another contributing factor could be the lack of diversity in instruction fine-tuning datasets. Models like Airavata, which utilize more diverse data Gala et al. (2024), exhibit noticeably better performance. Further investigation is required to fully understand the limitations and opportunities in this area.
6 Conclusion
In this paper, we introduced MILUMultilingual Indic Language Understanding Benchmark-a comprehensive benchmark specifically designed to evaluate LLMs across 11 Indic languages, spanning diverse domains and culturally relevant subjects. We evaluate 45 different LLMs and find that the majority of LLMs struggle on MILU, with GPT-4o achieving the highest average accuracy. The analysis also shows that models perform significantly better in high-resource languages than low-resource ones, highlighting the need for more robust multilingual strategies. Additionally, the domain-specific analysis indicates that models perform better in general fields such as STEM while facing challenges in culturally relevant subjects like Arts, Humanities, and Law, highlighting the lack of this knowledge in the current models and datasets. To foster open research, we release MILU alongside all code, and other artifacts. As LLMs continue to become pivotal in modern applications, we hope that MILU offers a foundational benchmark for developing more inclusive, culturally aware models that perform well across both general and culturally relevant domains.
Limitations
This work has a few limitations. First, we restricted our study to the top 11 languages due to the lack of readily available questions in low-resource languages, which we aim to address in future work. Second, limited computational resources prevented a thorough evaluation of larger models, such as Llama-3.1-70B and Llama-3.1-405B. Third, the scarcity of questions necessitated translating a portion of the dataset. Finally, our evaluation primarily relies on the log-likelihood approach, which may yield different results compared to other established methods, such as generation-based evaluation and chain-of-thought (CoT) prompting.
Ethics
All data described in this work was scraped from publicly available resources. The datasets used in this paper will be made available under permissible licenses. Additionally, the code used for our evaluations will be made publicly available under the MIT License. We only used ChatGPT for assistance purely with the language of the paper, e.g., paraphrasing, spell-checking, or polishing the author’s original content, without suggesting new content.
Appendix
Appendix A Details about different Exams
We collected our questions from over 40 exam types ranging from various National and state level civil service examinations to examinations conducted by various government and private organizations. Tables 6, 7 and 8
Organization Years Covered
Railway Recruitment Board 2018-2019
Railway Recruitment Cell 2018
Dedicated Freight Corridor Corporation of India Limited 2016-2021
Intelligence Bureau 2012-2021
Union Public Service Commission 2010-2023
Table 6: Overview of various national-level exams and the corresponding years of coverage considered in MILU.
Organization Years Covered
Punjab Police 2016-2022
Punjab State Power Corporation Limited 2018
Chandigarh Police 2018
Telangana Police 2015-2022
Andhra Pradesh Police 2016-2018
Tamil Nadu Public Service Commission 2010-2021
Tamil Nadu Uniformed Services Recruitment Board 2010-2022
Odisha Police 2022
Karnataka State Police 2019-2023
Madhya Pradesh Police 2016-2023
Delhi Police 2014-2022
Haryana Police 2021
Lucknow Metro Rail Corporation 2018
Delhi Metro Rail Corporation 2017-2020
Punjab Subordinate Service Selection Board 2022
Punjab State Transmission Corporation Limited 2016
Gujarat Metro Rail Corporation 2022
Kolkata Police 2023
Maharashtra Police 2021
Jharkhand Police 2015-2021
Table 7: Overview of various government and private organization exams and the corresponding years of coverage considered in MILU.
Organization Years Covered
Staff Selection Commission 2017-2023
West Bengal Civil Service 2015-2021
Bihar Public Service Commission 2015-2020
Maharashtra Public Service Commission 2013-2023
Rajasthan Subordinate and Ministerial Services Selection Board 2016-2022
Odisha Public Service Commission 2016-2022
Uttar Pradesh Public Service Commission 2012-2023
Haryana Public Service Commission 2014-2022
Andhra Pradesh Public Service Commission 2017-2019
Chhattisgarh Public Service Commission 2014-2022
Jammu & Kashmir Public Service Commission 2021
Himachal Pradesh Public Service Commission 2015-2022
Jharkhand Public Service Commission 2021
Delhi Subordinate Services Selection Board 2019-2022
Table 8: Overview of various State-level civil services exams and the corresponding years of coverage considered in MILU.
Appendix B Details about subject and language distribution.
Detailed analysis of topic and language distribution across languages can be found in Table 9 and Figure 6
Topic bn en gu hi kn ml mr or pa ta te Total
Agriculture 100 497 99 497 100 99 100 100 100 100 100 1892
Anthropology 100 100 100 99 100 100 100 100 100 100 100 1099
Architecture and Design 100 100 100 100 100 100 100 100 100 100 100 1100
Arts and Culture 165 500 100 499 163 100 183 100 100 165 199 2274
Astronomy and Astrophysics 100 100 100 128 100 100 100 100 100 100 100 1128
Biology 284 499 100 499 244 99 335 100 100 314 318 2892
Business and Management 100 323 100 427 100 100 100 100 100 100 106 1656
Chemistry 364 499 116 499 306 100 379 104 100 337 401 3205
Computer Science 192 500 100 500 169 100 169 100 100 172 223 2325
Defense and Security 100 100 100 211 100 100 100 100 100 100 100 1211
Earth Sciences 100 440 100 499 100 100 100 99 100 100 100 1838
Economics 139 498 100 496 100 100 142 100 100 126 167 2068
Education 100 314 100 338 100 100 157 100 100 100 100 1609
Energy and Power 100 148 100 295 100 100 100 100 100 100 100 1343
Engineering 500 500 500 500 500 234 500 336 100 500 500 4670
Environmental Science 100 499 100 499 128 100 190 99 100 100 179 2094
Ethics and Human Rights 100 100 100 100 100 100 100 100 100 100 100 1100
Finance and Investment 202 500 100 500 162 100 177 100 100 179 280 2400
Food Science 100 100 100 99 100 100 100 100 100 100 100 1099
Geography 179 500 100 498 219 99 169 100 100 175 192 2331
Health and Medicine 100 499 100 499 100 100 112 100 100 116 107 1933
History 375 498 100 496 273 100 269 100 100 275 391 2977
Information Technology 100 100 100 179 100 100 100 100 100 100 100 1179
International Relations 100 120 100 257 100 100 100 100 100 100 100 1277
Language Studies 100 500 100 499 129 100 149 100 100 267 174 2218
Law and Ethics 148 500 100 500 157 100 232 100 100 127 160 2224
Literature and Linguistics 111 499 100 499 120 100 113 100 100 247 142 2131
Logical Reasoning 407 500 154 499 241 100 359 100 100 251 404 3115
Materials Science 100 263 100 458 100 100 100 100 100 100 100 1621
Mathematics 500 500 500 499 500 349 500 500 264 500 500 5112
Media and Communication 100 100 100 147 100 100 100 100 100 100 100 1147
Music and Performing Arts 100 100 100 151 100 100 100 100 100 100 100 1151
Physics 500 500 359 500 500 190 500 288 100 500 500 4437
Politics and Governance 255 499 100 498 247 100 356 100 100 241 424 2920
Psychology 100 100 100 129 100 100 100 100 100 100 100 1129
Public Administration 100 99 99 99 99 100 99 100 99 100 100 1094
Religion and Spirituality 100 100 100 248 100 100 100 100 100 100 100 1248
Social Welfare and Development 100 112 100 193 100 100 100 100 100 100 100 1205
Sociology 115 500 100 500 100 100 159 100 100 111 170 2055
Sports and Recreation 202 500 100 500 178 100 177 100 100 156 210 2323
Technology and Innovation 100 500 100 500 100 100 100 100 100 100 100 1900
Transportation and Logistics 100 130 100 317 99 100 98 99 100 100 100 1343
TOTAL 7138 14036 5327 15450 6734 4670 7424 5025 4363 7059 7847 85073
Table 9: Detailed subject level statistics of MILU across different languages.
Figure 6: Distribution of tags across languages in MILU.
Appendix C Details about the various models evaluated
Model details about the different models evaluated in this work is present in Table 10.
Model Name Type #Params Link
GPT-4o Instruct - Link
GPT-4o-mini Instruct - Link
Gemini-1.5-Pro Instruct - Link
Gemini-1.5-Flash Instruct - Link
Krutrim-spectre-v2 Instruct - Link
meta-llama/Llama-3.2-1B Base 1B Link
meta-llama/Llama-3.2-1B-Instruct Instruct 1B Link
sarvamai/sarvam-2b-v0.5 Base 2B Link
google/gemma-2-2b Base 2B Link
google/gemma-2-2b-it Instruct 2B Link
meta-llama/Llama-3.2-3B Base 3B Link
meta-llama/Llama-3.2-3B-Instruct Instruct 3B Link
Telugu-LLM-Labs/Indic-gemma-7b-finetuned-sft-Navarasa-2.0 Instruct 7B Link
CohereForAI/aya-23-8B Base 8B Link
meta-llama/Llama-3.1-8B Base 8B Link
meta-llama/Llama-3.1-8B-Instruct Instruct 8B Link
google/gemma-2-9b Base 9B Link
google/gemma-2-9b-it Instruct 9B Link
google/gemma-2-27b Base 27B Link
google/gemma-2-27b-it Instruct 27B Link
CohereForAI/aya-23-35B Base 35B Link
meta-llama/Llama-3.1-70B Base 70B Link
meta-llama/Llama-3.1-70B-Instruct Instruct 70B Link
meta-llama/Llama-3.1-405B Base 405B Link
sarvamai/OpenHathi-7B-Hi-v0.1-Base Base 7B Link
ai4bharat/Airavata Instruct 7B Link
BhabhaAI/Gajendra-v0.1 Instruct 7B Link
manishiitg/open-aditi-v6-llama3 Instruct 8B Link
GenVRadmin/AryaBhatta-GemmaGenZ-Vikas-Merged Instruct 7B Link
nickmalhotra/ProjectIndus Base 1B Link
abhinand/telugu-llama-7b-instruct-v0.1 Instruct 7B Link
Telugu-LLM-Labs/Telugu-Llama2-7B-v0-Base Base 7B Link
Telugu-LLM-Labs/Telugu-Llama2-7B-v0-Instruct Instruct 7B Link
Tensoic/Kan-LLaMA-7B-base Base 7B Link
Cognitive-Lab/Ambari-7B-base-v0.1 Base 7B Link
Cognitive-Lab/Ambari-7B-Instruct-v0.1 Instruct 7B Link
smallstepai/Misal-7B-base-v0.1 Base 7B Link
smallstepai/Misal-7B-instruct-v0.1 Instruct 7B Link
abhinand/tamil-llama-7b-instruct-v0.2 Instruct 7B Link
abhinand/malayalam-llama-7b-instruct-v0.1 Instruct 7B Link
Table 10: Details about the different models evaluated on MILU.
Subject ta
Agriculture 29.00/ 28.00/ 26.00
Anthropology 21.00/ 23.00/ 26.00
Architecture and Design 19.00/ 24.00/ 20.00
Arts and Culture 31.52/ 32.73/ 29.70
Astronomy and Astrophysics 22.00/ 32.00/ 32.00
Biology 24.20/ 19.43/ 24.52
Business and Management 36.00/ 28.00/ 24.00
Chemistry 23.74/ 24.04/ 23.44
Computer Science 28.49/ 25.00/ 26.74
Defense and Security 29.00/ 24.00/ 32.00
Earth Sciences 28.00/ 25.00/ 28.00
Economics 24.60/ 20.63/ 19.05
Education 28.00/ 24.00/ 26.00
Energy and Power 29.00/ 20.00/ 17.00
Engineering 26.60/ 26.00/ 25.20
Environmental Science 23.00/ 19.00/ 20.00
Ethics and Human Rights 31.00/ 27.00/ 29.00
Finance and Investment 27.93/ 27.93/ 23.46
Food Science 27.00/ 21.00/ 21.00
Geography 22.86/ 26.29/ 24.57
Health and Medicine 22.41/ 24.14/ 24.14
History 26.18/ 24.73/ 24.00
Information Technology 28.00/ 23.00/ 25.00
International Relations 26.00/ 28.00/ 28.00
Language Studies 24.72/ 24.34/ 25.47
Law and Ethics 21.26/ 26.77/ 22.83
Literature and Linguistics 26.32/ 25.51/ 23.48
Logical Reasoning 20.72/ 21.12/ 25.10
Materials Science 25.00/ 19.00/ 18.00
Mathematics 25.40/ 25.60/ 28.00
Media and Communication 18.00/ 19.00/ 19.00
Music and Performing Arts 25.00/ 27.00/ 27.00
Physics 22.60/ 24.40/ 24.20
Politics and Governance 29.05/ 31.12/ 30.29
Psychology 32.00/ 28.00/ 27.00
Public Administration 26.00/ 21.00/ 19.00
Religion and Spirituality 29.00/ 31.00/ 33.00
Social Welfare and Development 21.00/ 23.00/ 21.00
Sociology 19.82/ 22.52/ 23.42
Sports and Recreation 32.05/ 30.13/ 30.77
Technology and Innovation 21.00/ 24.00/ 29.00
Transportation and Logistics 23.00/ 29.00/ 27.00
Table 11: Detailed subject-wise evaluation for abhinand/tamil-llama-7b-instruct-v0.2 on MILU across different languages. The results reported are X / Y / Z where X denote the 0-shot, Y denotes 1-shot and Z denotes 5-shot performance respectively.
Subject hi
Agriculture 26.56/ 26.96/ 29.98
Anthropology 30.30/ 26.26/ 30.30
Architecture and Design 28.00/ 31.00/ 26.00
Arts and Culture 25.45/ 24.65/ 27.45
Astronomy and Astrophysics 24.22/ 29.69/ 29.69
Biology 25.05/ 32.06/ 32.06
Business and Management 27.87/ 29.51/ 26.70
Chemistry 29.66/ 31.46/ 31.26
Computer Science 28.60/ 30.60/ 30.60
Defense and Security 28.91/ 22.27/ 24.64
Earth Sciences 26.05/ 27.45/ 26.25
Economics 26.81/ 25.20/ 27.02
Education 27.22/ 25.15/ 26.04
Energy and Power 31.53/ 29.83/ 31.53
Engineering 24.00/ 25.60/ 26.40
Environmental Science 23.65/ 25.85/ 25.65
Ethics and Human Rights 33.00/ 25.00/ 29.00
Finance and Investment 26.60/ 28.20/ 30.20
Food Science 24.24/ 31.31/ 33.33
Geography 26.31/ 27.11/ 26.71
Health and Medicine 29.06/ 32.87/ 32.67
History 27.02/ 28.63/ 27.62
Information Technology 32.96/ 34.64/ 32.96
International Relations 31.91/ 31.91/ 31.52
Language Studies 25.05/ 28.86/ 25.25
Law and Ethics 29.20/ 29.40/ 29.20
Literature and Linguistics 32.87/ 31.86/ 30.06
Logical Reasoning 26.45/ 27.45/ 27.25
Materials Science 24.67/ 28.60/ 25.55
Mathematics 24.25/ 23.05/ 22.44
Media and Communication 23.81/ 25.85/ 24.49
Music and Performing Arts 25.17/ 24.50/ 25.83
Physics 26.20/ 25.20/ 23.40
Politics and Governance 28.71/ 27.71/ 26.91
Psychology 19.38/ 23.26/ 23.26
Public Administration 39.39/ 41.41/ 40.40
Religion and Spirituality 27.82/ 29.03/ 29.44
Social Welfare and Development 22.28/ 23.32/ 27.98
Sociology 23.40/ 23.20/ 24.20
Sports and Recreation 29.40/ 30.20/ 31.60
Technology and Innovation 26.20/ 28.60/ 30.60
Transportation and Logistics 25.87/ 26.50/ 27.13
Table 12: Detailed subject-wise evaluation for AI4Bharat/Airavata on MILU across different languages. The results reported are X / Y / Z where X denote the 0-shot, Y denotes 1-shot and Z denotes 5-shot performance respectively.
Subject hi
Agriculture 27.97/ 27.36/ 28.97
Anthropology 27.27/ 26.26/ 20.20
Architecture and Design 23.00/ 20.00/ 25.00
Arts and Culture 28.06/ 28.26/ 24.25
Astronomy and Astrophysics 27.34/ 30.47/ 35.94
Biology 31.06/ 30.06/ 30.46
Business and Management 26.23/ 24.12/ 29.27
Chemistry 28.26/ 25.65/ 30.06
Computer Science 25.80/ 28.80/ 29.80
Defense and Security 23.22/ 30.81/ 28.91
Earth Sciences 23.65/ 26.45/ 29.46
Economics 27.02/ 27.82/ 29.44
Education 26.33/ 28.99/ 31.07
Energy and Power 27.80/ 26.44/ 29.83
Engineering 20.40/ 20.20/ 24.00
Environmental Science 25.45/ 26.25/ 30.06
Ethics and Human Rights 27.00/ 24.00/ 24.00
Finance and Investment 23.60/ 25.00/ 29.20
Food Science 32.32/ 37.37/ 28.28
Geography 26.71/ 25.90/ 28.71
Health and Medicine 30.46/ 30.46/ 30.66
History 26.01/ 28.02/ 29.03
Information Technology 27.93/ 31.28/ 36.87
International Relations 25.68/ 25.29/ 24.12
Language Studies 23.05/ 24.85/ 26.65
Law and Ethics 26.60/ 26.60/ 28.80
Literature and Linguistics 27.45/ 27.25/ 29.26
Logical Reasoning 22.24/ 27.05/ 26.05
Materials Science 26.86/ 29.26/ 29.48
Mathematics 25.65/ 26.85/ 25.25
Media and Communication 23.81/ 26.53/ 32.65
Music and Performing Arts 25.17/ 22.52/ 28.48
Physics 24.20/ 25.40/ 27.80
Politics and Governance 23.09/ 24.70/ 27.11
Psychology 26.36/ 25.58/ 25.58
Public Administration 36.36/ 34.34/ 32.32
Religion and Spirituality 25.81/ 22.98/ 29.03
Social Welfare and Development 25.91/ 24.87/ 25.91
Sociology 26.20/ 25.60/ 24.80
Sports and Recreation 26.00/ 26.40/ 26.60
Technology and Innovation 28.80/ 28.20/ 30.40
Transportation and Logistics 25.24/ 23.97/ 24.29
Table 13: Detailed subject-wise evaluation for BhabhaAI/Gajendra-v0.1 on MILU across different languages. The results reported are X / Y / Z where X denote the 0-shot, Y denotes 1-shot and Z denotes 5-shot performance respectively.
Subject kn
Agriculture 25.00/ 28.00/ 23.00
Anthropology 25.00/ 26.00/ 26.00
Architecture and Design 25.00/ 31.00/ 28.00
Arts and Culture 26.99/ 29.45/ 24.54
Astronomy and Astrophysics 28.00/ 30.00/ 30.00
Biology 29.10/ 26.23/ 28.28
Business and Management 36.00/ 25.00/ 26.00
Chemistry 31.37/ 25.82/ 28.43
Computer Science 28.40/ 33.14/ 33.14
Defense and Security 27.00/ 31.00/ 33.00
Earth Sciences 29.00/ 31.00/ 33.00
Economics 29.00/ 32.00/ 37.00
Education 28.00/ 30.00/ 29.00
Energy and Power 26.00/ 22.00/ 19.00
Engineering 24.60/ 27.20/ 28.20
Environmental Science 21.88/ 23.44/ 29.69
Ethics and Human Rights 29.00/ 19.00/ 26.00
Finance and Investment 29.63/ 28.40/ 31.48
Food Science 23.00/ 30.00/ 38.00
Geography 27.85/ 27.40/ 28.77
Health and Medicine 23.00/ 29.00/ 24.00
History 24.54/ 28.57/ 29.67
Information Technology 26.00/ 35.00/ 31.00
International Relations 32.00/ 36.00/ 26.00
Language Studies 28.68/ 19.38/ 27.91
Law and Ethics 25.48/ 22.29/ 26.75
Literature and Linguistics 18.33/ 23.33/ 23.33
Logical Reasoning 21.99/ 21.58/ 25.73
Materials Science 28.00/ 24.00/ 23.00
Mathematics 23.00/ 24.00/ 23.60
Media and Communication 32.00/ 22.00/ 28.00
Music and Performing Arts 32.00/ 29.00/ 30.00
Physics 26.40/ 26.60/ 26.00
Politics and Governance 27.53/ 23.89/ 25.51
Psychology 23.00/ 34.00/ 31.00
Public Administration 36.36/ 30.30/ 32.32
Religion and Spirituality 27.00/ 27.00/ 31.00
Social Welfare and Development 28.00/ 26.00/ 32.00
Sociology 24.00/ 33.00/ 30.00
Sports and Recreation 22.47/ 29.78/ 27.53
Technology and Innovation 25.00/ 29.00/ 32.00
Transportation and Logistics 25.25/ 26.26/ 29.29
Table 14: Detailed subject-wise evaluation for Cognitive-Lab/Ambari-7B-base-v0.1 on MILU across different languages. The results reported are X / Y / Z where X denote the 0-shot, Y denotes 1-shot and Z denotes 5-shot performance respectively.
Subject kn
Agriculture 31.00/ 26.00/ 23.00
Anthropology 29.00/ 23.00/ 22.00
Architecture and Design 29.00/ 21.00/ 23.00
Arts and Culture 25.15/ 28.83/ 30.06
Astronomy and Astrophysics 29.00/ 24.00/ 33.00
Biology 27.87/ 25.82/ 29.10
Business and Management 38.00/ 31.00/ 28.00
Chemistry 31.37/ 30.07/ 31.05
Computer Science 24.26/ 25.44/ 30.77
Defense and Security 27.00/ 22.00/ 19.00
Earth Sciences 26.00/ 28.00/ 27.00
Economics 31.00/ 32.00/ 29.00
Education 32.00/ 30.00/ 32.00
Energy and Power 26.00/ 21.00/ 26.00
Engineering 23.20/ 26.20/ 28.20
Environmental Science 25.78/ 28.12/ 26.56
Ethics and Human Rights 29.00/ 25.00/ 28.00
Finance and Investment 23.46/ 25.31/ 27.78
Food Science 22.00/ 21.00/ 33.00
Geography 31.51/ 24.20/ 31.05
Health and Medicine 17.00/ 22.00/ 21.00
History 27.11/ 28.57/ 30.04
Information Technology 33.00/ 32.00/ 35.00
International Relations 22.00/ 27.00/ 26.00
Language Studies 25.58/ 30.23/ 23.26
Law and Ethics 28.03/ 23.57/ 21.66
Literature and Linguistics 23.33/ 22.50/ 24.17
Logical Reasoning 21.16/ 21.16/ 24.07
Materials Science 22.00/ 29.00/ 26.00
Mathematics 23.00/ 26.80/ 27.00
Media and Communication 30.00/ 25.00/ 26.00
Music and Performing Arts 29.00/ 27.00/ 25.00
Physics 26.40/ 23.60/ 24.20
Politics and Governance 22.67/ 29.15/ 25.51
Psychology 17.00/ 18.00/ 25.00
Public Administration 24.24/ 23.23/ 28.28
Religion and Spirituality 27.00/ 26.00/ 29.00
Social Welfare and Development 29.00/ 25.00/ 27.00
Sociology 18.00/ 19.00/ 22.00
Sports and Recreation 19.10/ 24.16/ 24.16
Technology and Innovation 32.00/ 23.00/ 25.00
Transportation and Logistics 32.32/ 31.31/ 27.27
Table 15: Detailed subject-wise evaluation for Cognitive-Lab/Ambari-7B-Instruct-v0.1 on MILU across different languages. The results reported are X / Y / Z where X denote the 0-shot, Y denotes 1-shot and Z denotes 5-shot performance respectively.
Subject hi
Agriculture 29.38/ 30.18/ 36.62
Anthropology 37.37/ 32.32/ 40.40
Architecture and Design 26.00/ 28.00/ 35.00
Arts and Culture 26.25/ 29.46/ 40.48
Astronomy and Astrophysics 37.50/ 42.19/ 46.88
Biology 35.27/ 35.07/ 50.30
Business and Management 31.15/ 32.32/ 38.64
Chemistry 28.86/ 38.68/ 48.30
Computer Science 32.60/ 34.00/ 36.60
Defense and Security 28.91/ 36.02/ 35.55
Earth Sciences 32.06/ 28.66/ 40.48
Economics 27.42/ 32.06/ 41.53
Education 23.96/ 27.51/ 30.77
Energy and Power 26.10/ 31.19/ 43.39
Engineering 25.80/ 26.60/ 33.20
Environmental Science 29.66/ 30.26/ 44.89
Ethics and Human Rights 30.00/ 31.00/ 41.00
Finance and Investment 27.40/ 34.20/ 38.00
Food Science 24.24/ 33.33/ 47.47
Geography 31.33/ 30.72/ 42.57
Health and Medicine 34.27/ 34.67/ 48.70
History 27.02/ 32.86/ 36.90
Information Technology 34.64/ 37.99/ 50.84
International Relations 31.52/ 31.13/ 42.41
Language Studies 22.65/ 27.05/ 29.26
Law and Ethics 28.00/ 29.80/ 38.60
Literature and Linguistics 24.85/ 28.86/ 33.47
Logical Reasoning 25.45/ 27.25/ 29.66
Materials Science 31.88/ 30.35/ 35.59
Mathematics 26.85/ 24.25/ 30.06
Media and Communication 25.85/ 26.53/ 40.82
Music and Performing Arts 21.19/ 30.46/ 33.77
Physics 32.60/ 32.80/ 40.00
Politics and Governance 25.50/ 27.71/ 41.97
Psychology 23.26/ 22.48/ 42.64
Public Administration 30.30/ 27.27/ 36.36
Religion and Spirituality 28.23/ 33.87/ 42.34
Social Welfare and Development 26.94/ 27.98/ 35.75
Sociology 23.80/ 26.60/ 28.00
Sports and Recreation 29.00/ 34.80/ 43.80
Technology and Innovation 31.80/ 30.80/ 41.60
Transportation and Logistics 23.97/ 23.97/ 30.28
Table 16: Detailed subject-wise evaluation for GenVRadmin/AryaBhatta-GemmaGenZ-Vikas-Merged on MILU across different languages. The results reported are X / Y / Z where X denote the 0-shot, Y denotes 1-shot and Z denotes 5-shot performance respectively.
Subject hi
Agriculture 29.38/ 28.97/ 27.16
Anthropology 36.36/ 36.36/ 28.28
Architecture and Design 34.00/ 31.00/ 33.00
Arts and Culture 34.07/ 33.67/ 34.87
Astronomy and Astrophysics 42.97/ 41.41/ 37.50
Biology 36.67/ 38.08/ 36.07
Business and Management 32.55/ 31.15/ 29.04
Chemistry 37.47/ 34.67/ 31.66
Computer Science 37.40/ 34.00/ 35.80
Defense and Security 30.33/ 28.44/ 27.49
Earth Sciences 31.86/ 31.46/ 33.07
Economics 34.68/ 33.06/ 30.44
Education 31.36/ 29.88/ 30.77
Energy and Power 34.58/ 32.20/ 32.54
Engineering 28.60/ 27.60/ 27.80
Environmental Science 31.86/ 31.26/ 30.26
Ethics and Human Rights 36.00/ 35.00/ 36.00
Finance and Investment 29.80/ 30.20/ 28.40
Food Science 36.36/ 36.36/ 28.28
Geography 33.33/ 32.73/ 30.12
Health and Medicine 44.09/ 40.28/ 38.88
History 39.92/ 38.10/ 36.49
Information Technology 47.49/ 47.49/ 43.02
International Relations 40.08/ 40.47/ 34.63
Language Studies 29.86/ 29.06/ 27.45
Law and Ethics 41.60/ 39.20/ 40.00
Literature and Linguistics 29.66/ 30.26/ 29.86
Logical Reasoning 30.46/ 28.46/ 29.06
Materials Science 29.26/ 31.00/ 30.35
Mathematics 28.06/ 28.26/ 27.86
Media and Communication 36.73/ 34.01/ 31.29
Music and Performing Arts 36.42/ 32.45/ 34.44
Physics 33.60/ 29.80/ 29.00
Politics and Governance 36.95/ 36.75/ 32.33
Psychology 41.86/ 37.21/ 33.33
Public Administration 32.32/ 33.33/ 34.34
Religion and Spirituality 29.84/ 31.45/ 29.44
Social Welfare and Development 30.05/ 30.57/ 30.57
Sociology 30.00/ 30.00/ 27.40
Sports and Recreation 32.00/ 30.80/ 28.60
Technology and Innovation 36.60/ 36.80/ 34.00
Transportation and Logistics 30.60/ 29.65/ 28.71
Table 17: Detailed subject-wise evaluation for manishiitg/open-aditi-v6-llama3 on MILU across different languages. The results reported are X / Y / Z where X denote the 0-shot, Y denotes 1-shot and Z denotes 5-shot performance respectively.
Subject hi
Agriculture 22.74/ 22.74/ 21.93
Anthropology 25.25/ 24.24/ 24.24
Architecture and Design 32.00/ 28.00/ 29.00
Arts and Culture 24.85/ 26.05/ 26.25
Astronomy and Astrophysics 35.94/ 35.94/ 32.03
Biology 24.85/ 24.45/ 24.05
Business and Management 25.29/ 25.06/ 25.76
Chemistry 25.85/ 25.25/ 24.45
Computer Science 26.40/ 27.00/ 25.20
Defense and Security 27.01/ 24.64/ 25.12
Earth Sciences 25.45/ 24.05/ 23.85
Economics 28.02/ 27.42/ 27.42
Education 26.04/ 24.85/ 24.85
Energy and Power 23.73/ 24.07/ 25.08
Engineering 23.60/ 23.40/ 24.00
Environmental Science 28.86/ 28.26/ 27.86
Ethics and Human Rights 30.00/ 31.00/ 22.00
Finance and Investment 24.00/ 25.00/ 25.40
Food Science 27.27/ 29.29/ 27.27
Geography 27.11/ 25.90/ 26.51
Health and Medicine 25.45/ 26.05/ 23.85
History 24.80/ 25.81/ 26.21
Information Technology 26.26/ 27.37/ 25.14
International Relations 26.85/ 28.40/ 29.96
Language Studies 21.44/ 21.44/ 22.24
Law and Ethics 24.80/ 24.80/ 22.80
Literature and Linguistics 23.65/ 23.85/ 22.24
Logical Reasoning 27.05/ 27.86/ 27.45
Materials Science 28.17/ 29.04/ 28.60
Mathematics 23.25/ 22.44/ 23.05
Media and Communication 28.57/ 31.29/ 26.53
Music and Performing Arts 21.19/ 22.52/ 23.84
Physics 24.40/ 24.00/ 21.60
Politics and Governance 26.71/ 25.30/ 22.29
Psychology 27.91/ 27.91/ 26.36
Public Administration 26.26/ 27.27/ 28.28
Religion and Spirituality 26.21/ 29.44/ 27.82
Social Welfare and Development 26.94/ 27.98/ 26.42
Sociology 27.00/ 28.00/ 26.40
Sports and Recreation 27.80/ 27.40/ 26.60
Technology and Innovation 23.80/ 25.20/ 26.40
Transportation and Logistics 23.97/ 26.81/ 24.92
Table 18: Detailed subject-wise evaluation for nickmalhotra/projectindus on MILU across different languages. The results reported are X / Y / Z where X denote the 0-shot, Y denotes 1-shot and Z denotes 5-shot performance respectively.
Subject hi
Agriculture 24.35/ 30.18/ 29.98
Anthropology 20.20/ 25.25/ 26.26
Architecture and Design 23.00/ 32.00/ 27.00
Arts and Culture 27.66/ 27.05/ 28.06
Astronomy and Astrophysics 24.22/ 31.25/ 33.59
Biology 27.45/ 26.65/ 28.86
Business and Management 26.70/ 25.53/ 29.74
Chemistry 26.65/ 26.25/ 28.86
Computer Science 26.80/ 29.20/ 25.80
Defense and Security 27.96/ 24.64/ 27.96
Earth Sciences 22.04/ 24.65/ 26.85
Economics 26.21/ 30.85/ 33.27
Education 24.26/ 26.92/ 25.15
Energy and Power 30.17/ 31.19/ 32.54
Engineering 22.20/ 21.40/ 25.40
Environmental Science 25.05/ 25.65/ 27.05
Ethics and Human Rights 31.00/ 33.00/ 36.00
Finance and Investment 27.00/ 27.80/ 26.20
Food Science 20.20/ 24.24/ 30.30
Geography 24.30/ 27.91/ 27.11
Health and Medicine 33.07/ 28.66/ 30.86
History 29.44/ 30.24/ 32.66
Information Technology 25.70/ 30.17/ 35.20
International Relations 26.46/ 28.40/ 32.30
Language Studies 25.45/ 24.85/ 28.46
Law and Ethics 28.60/ 28.80/ 34.20
Literature and Linguistics 26.25/ 25.85/ 27.45
Logical Reasoning 28.86/ 23.85/ 28.26
Materials Science 22.93/ 24.45/ 25.76
Mathematics 24.25/ 27.05/ 25.25
Media and Communication 29.93/ 27.89/ 24.49
Music and Performing Arts 27.15/ 24.50/ 24.50
Physics 26.00/ 27.20/ 29.60
Politics and Governance 23.49/ 24.90/ 27.91
Psychology 24.81/ 23.26/ 26.36
Public Administration 29.29/ 33.33/ 32.32
Religion and Spirituality 28.63/ 26.21/ 30.24
Social Welfare and Development 27.46/ 25.91/ 26.42
Sociology 28.80/ 28.40/ 25.80
Sports and Recreation 28.80/ 29.00/ 31.40
Technology and Innovation 29.00/ 29.60/ 30.40
Transportation and Logistics 29.02/ 24.29/ 27.44
Table 19: Detailed subject-wise evaluation for sarvamai/OpenHathi-7B-Hi-v0.1-Base on MILU across different languages. The results reported are X / Y / Z where X denote the 0-shot, Y denotes 1-shot and Z denotes 5-shot performance respectively.
Subject kn
Agriculture 23.00/ 29.00/ 28.00
Anthropology 30.00/ 41.00/ 38.00
Architecture and Design 26.00/ 27.00/ 25.00
Arts and Culture 26.38/ 26.99/ 29.45
Astronomy and Astrophysics 26.00/ 30.00/ 21.00
Biology 23.77/ 29.51/ 28.69
Business and Management 30.00/ 28.00/ 28.00
Chemistry 28.10/ 27.12/ 25.82
Computer Science 29.59/ 36.09/ 36.69
Defense and Security 25.00/ 26.00/ 28.00
Earth Sciences 24.00/ 24.00/ 24.00
Economics 30.00/ 32.00/ 29.00
Education 28.00/ 38.00/ 36.00
Energy and Power 28.00/ 22.00/ 25.00
Engineering 23.00/ 23.60/ 24.80
Environmental Science 23.44/ 32.81/ 30.47
Ethics and Human Rights 24.00/ 28.00/ 30.00
Finance and Investment 26.54/ 31.48/ 29.63
Food Science 27.00/ 29.00/ 31.00
Geography 32.88/ 31.96/ 26.94
Health and Medicine 17.00/ 27.00/ 30.00
History 27.11/ 26.01/ 31.14
Information Technology 33.00/ 37.00/ 35.00
International Relations 25.00/ 25.00/ 29.00
Language Studies 26.36/ 19.38/ 24.03
Law and Ethics 26.11/ 26.75/ 24.84
Literature and Linguistics 22.50/ 25.00/ 21.67
Logical Reasoning 22.41/ 23.24/ 22.82
Materials Science 25.00/ 26.00/ 24.00
Mathematics 24.60/ 22.60/ 23.80
Media and Communication 28.00/ 32.00/ 32.00
Music and Performing Arts 30.00/ 28.00/ 29.00
Physics 26.20/ 28.20/ 26.60
Politics and Governance 26.72/ 29.55/ 30.36
Psychology 29.00/ 29.00/ 25.00
Public Administration 31.31/ 29.29/ 31.31
Religion and Spirituality 26.00/ 27.00/ 35.00
Social Welfare and Development 32.00/ 34.00/ 31.00
Sociology 29.00/ 28.00/ 30.00
Sports and Recreation 26.40/ 30.90/ 29.21
Technology and Innovation 29.00/ 30.00/ 34.00
Transportation and Logistics 36.36/ 31.31/ 32.32
Table 20: Detailed subject-wise evaluation for Tensoic/Kan-LLaMA-7B-base on MILU across different languages. The results reported are X / Y / Z where X denote the 0-shot, Y denotes 1-shot and Z denotes 5-shot performance respectively.
Subject ml
Agriculture 22.22/ 18.18/ 25.25
Anthropology 24.00/ 20.00/ 21.00
Architecture and Design 22.00/ 21.00/ 23.00
Arts and Culture 25.00/ 20.00/ 23.00
Astronomy and Astrophysics 27.00/ 28.00/ 26.00
Biology 24.24/ 27.27/ 26.26
Business and Management 26.00/ 24.00/ 28.00
Chemistry 25.00/ 26.00/ 26.00
Computer Science 24.00/ 28.00/ 27.00
Defense and Security 18.00/ 18.00/ 18.00
Earth Sciences 30.00/ 31.00/ 34.00
Economics 23.00/ 26.00/ 27.00
Education 23.00/ 21.00/ 22.00
Energy and Power 32.00/ 28.00/ 30.00
Engineering 27.35/ 27.78/ 25.64
Environmental Science 30.00/ 31.00/ 28.00
Ethics and Human Rights 29.00/ 19.00/ 24.00
Finance and Investment 35.00/ 31.00/ 35.00
Food Science 27.00/ 22.00/ 20.00
Geography 30.30/ 27.27/ 24.24
Health and Medicine 23.00/ 23.00/ 20.00
History 30.00/ 22.00/ 25.00
Information Technology 35.00/ 37.00/ 37.00
International Relations 23.00/ 25.00/ 24.00
Language Studies 21.00/ 17.00/ 20.00
Law and Ethics 27.00/ 30.00/ 27.00
Literature and Linguistics 28.00/ 24.00/ 25.00
Logical Reasoning 29.00/ 26.00/ 27.00
Materials Science 28.00/ 25.00/ 20.00
Mathematics 20.92/ 22.92/ 21.78
Media and Communication 24.00/ 24.00/ 22.00
Music and Performing Arts 25.00/ 21.00/ 24.00
Physics 25.79/ 27.89/ 30.53
Politics and Governance 30.00/ 28.00/ 31.00
Psychology 26.00/ 25.00/ 24.00
Public Administration 23.00/ 25.00/ 20.00
Religion and Spirituality 27.00/ 24.00/ 26.00
Social Welfare and Development 34.00/ 33.00/ 30.00
Sociology 20.00/ 21.00/ 20.00
Sports and Recreation 31.00/ 27.00/ 25.00
Technology and Innovation 25.00/ 26.00/ 26.00
Transportation and Logistics 24.00/ 25.00/ 32.00
Table 21: Detailed subject-wise evaluation for abhinand/malayalam-llama-7b-instruct-v0.1 on MILU across different languages. The results reported are X / Y / Z where X denote the 0-shot, Y denotes 1-shot and Z denotes 5-shot performance respectively.
Subject te
Agriculture 25.00/ 23.00/ 21.00
Anthropology 33.00/ 30.00/ 33.00
Architecture and Design 24.00/ 20.00/ 20.00
Arts and Culture 28.64/ 29.65/ 28.14
Astronomy and Astrophysics 25.00/ 25.00/ 25.00
Biology 24.53/ 23.27/ 22.01
Business and Management 36.79/ 36.79/ 39.62
Chemistry 25.69/ 27.43/ 27.93
Computer Science 25.56/ 30.04/ 28.70
Defense and Security 26.00/ 25.00/ 25.00
Earth Sciences 27.00/ 29.00/ 29.00
Economics 24.55/ 28.74/ 29.94
Education 22.00/ 22.00/ 28.00
Energy and Power 26.00/ 26.00/ 24.00
Engineering 30.40/ 29.20/ 28.00
Environmental Science 21.23/ 23.46/ 23.46
Ethics and Human Rights 21.00/ 22.00/ 20.00
Finance and Investment 26.79/ 24.64/ 26.79
Food Science 28.00/ 25.00/ 21.00
Geography 28.65/ 28.12/ 23.44
Health and Medicine 28.04/ 30.84/ 28.04
History 31.71/ 32.48/ 29.67
Information Technology 30.00/ 38.00/ 33.00
International Relations 27.00/ 29.00/ 31.00
Language Studies 27.01/ 28.16/ 27.59
Law and Ethics 27.50/ 24.37/ 26.87
Literature and Linguistics 25.35/ 21.83/ 18.31
Logical Reasoning 24.75/ 24.50/ 25.25
Materials Science 26.00/ 24.00/ 27.00
Mathematics 20.40/ 23.00/ 24.40
Media and Communication 23.00/ 23.00/ 25.00
Music and Performing Arts 28.00/ 23.00/ 30.00
Physics 22.80/ 25.20/ 28.20
Politics and Governance 27.59/ 27.59/ 29.48
Psychology 28.00/ 22.00/ 22.00
Public Administration 35.00/ 31.00/ 28.00
Religion and Spirituality 22.00/ 24.00/ 24.00
Social Welfare and Development 30.00/ 27.00/ 29.00
Sociology 21.18/ 20.59/ 18.82
Sports and Recreation 24.76/ 25.71/ 26.67
Technology and Innovation 29.00/ 24.00/ 22.00
Transportation and Logistics 26.00/ 22.00/ 26.00
Table 22: Detailed subject-wise evaluation for abhinand/telugu-llama-7b-instruct-v0.1 on MILU across different languages. The results reported are X / Y / Z where X denote the 0-shot, Y denotes 1-shot and Z denotes 5-shot performance respectively.
Subject mr
Agriculture 20.00/ 20.00/ 17.00
Anthropology 15.00/ 20.00/ 20.00
Architecture and Design 18.00/ 19.00/ 19.00
Arts and Culture 20.22/ 20.77/ 22.95
Astronomy and Astrophysics 20.00/ 19.00/ 19.00
Biology 24.18/ 23.58/ 23.28
Business and Management 20.00/ 21.00/ 22.00
Chemistry 23.75/ 25.33/ 25.59
Computer Science 26.63/ 25.44/ 27.22
Defense and Security 24.00/ 22.00/ 19.00
Earth Sciences 23.00/ 29.00/ 28.00
Economics 28.17/ 26.76/ 28.17
Education 24.20/ 23.57/ 22.29
Energy and Power 27.00/ 25.00/ 22.00
Engineering 21.60/ 21.60/ 22.20
Environmental Science 23.16/ 23.16/ 26.84
Ethics and Human Rights 20.00/ 18.00/ 17.00
Finance and Investment 23.16/ 24.29/ 24.29
Food Science 25.00/ 27.00/ 21.00
Geography 21.30/ 20.71/ 18.93
Health and Medicine 20.54/ 27.68/ 23.21
History 24.91/ 24.54/ 25.65
Information Technology 25.00/ 24.00/ 23.00
International Relations 24.00/ 27.00/ 23.00
Language Studies 21.48/ 23.49/ 23.49
Law and Ethics 21.12/ 21.55/ 23.71
Literature and Linguistics 20.35/ 23.01/ 22.12
Logical Reasoning 26.18/ 24.79/ 26.74
Materials Science 16.00/ 19.00/ 16.00
Mathematics 22.00/ 23.00/ 21.20
Media and Communication 20.00/ 17.00/ 17.00
Music and Performing Arts 27.00/ 27.00/ 31.00
Physics 25.00/ 24.40/ 23.40
Politics and Governance 21.07/ 20.22/ 22.75
Psychology 24.00/ 25.00/ 24.00
Public Administration 24.24/ 25.25/ 18.18
Religion and Spirituality 19.00/ 22.00/ 24.00
Social Welfare and Development 18.00/ 19.00/ 22.00
Sociology 18.24/ 19.50/ 21.38
Sports and Recreation 25.99/ 22.03/ 22.03
Technology and Innovation 18.00/ 20.00/ 18.00
Transportation and Logistics 21.43/ 26.53/ 24.49
Table 23: Detailed subject-wise evaluation for smallstepai/Misal-7B-instruct-v0.1 on MILU across different languages. The results reported are X / Y / Z where X denote the 0-shot, Y denotes 1-shot and Z denotes 5-shot performance respectively.
Subject mr
Agriculture 34.00/ 28.00/ 26.00
Anthropology 27.00/ 35.00/ 37.00
Architecture and Design 21.00/ 18.00/ 19.00
Arts and Culture 25.68/ 25.14/ 26.23
Astronomy and Astrophysics 25.00/ 23.00/ 24.00
Biology 23.88/ 25.97/ 25.07
Business and Management 19.00/ 27.00/ 26.00
Chemistry 29.82/ 27.70/ 29.02
Computer Science 20.12/ 24.26/ 23.08
Defense and Security 23.00/ 26.00/ 22.00
Earth Sciences 30.00/ 28.00/ 25.00
Economics 27.46/ 29.58/ 30.99
Education 38.22/ 28.66/ 28.66
Energy and Power 25.00/ 24.00/ 28.00
Engineering 21.60/ 28.00/ 26.80
Environmental Science 26.32/ 30.53/ 32.11
Ethics and Human Rights 20.00/ 22.00/ 25.00
Finance and Investment 25.99/ 22.03/ 24.29
Food Science 19.00/ 17.00/ 21.00
Geography 24.85/ 21.30/ 23.67
Health and Medicine 29.46/ 33.93/ 26.79
History 26.77/ 27.14/ 28.25
Information Technology 19.00/ 17.00/ 21.00
International Relations 30.00/ 30.00/ 25.00
Language Studies 22.15/ 22.15/ 20.81
Law and Ethics 25.00/ 29.31/ 32.33
Literature and Linguistics 22.12/ 27.43/ 29.20
Logical Reasoning 30.64/ 30.64/ 30.36
Materials Science 37.00/ 27.00/ 24.00
Mathematics 22.80/ 26.40/ 24.60
Media and Communication 24.00/ 31.00/ 31.00
Music and Performing Arts 23.00/ 23.00/ 30.00
Physics 26.00/ 24.00/ 23.00
Politics and Governance 30.62/ 25.84/ 26.12
Psychology 30.00/ 22.00/ 21.00
Public Administration 29.29/ 25.25/ 25.25
Religion and Spirituality 18.00/ 26.00/ 28.00
Social Welfare and Development 25.00/ 18.00/ 23.00
Sociology 25.79/ 19.50/ 16.35
Sports and Recreation 26.55/ 24.29/ 29.38
Technology and Innovation 28.00/ 30.00/ 31.00
Transportation and Logistics 27.55/ 25.51/ 28.57
Table 24: Detailed subject-wise evaluation for smallstepai/Misal-7B-base-v0.1 on MILU across different languages. The results reported are X / Y / Z where X denote the 0-shot, Y denotes 1-shot and Z denotes 5-shot performance respectively.
Subject te
Agriculture 25.00/ 21.00/ 29.00
Anthropology 22.00/ 19.00/ 21.00
Architecture and Design 19.00/ 14.00/ 17.00
Arts and Culture 24.62/ 21.11/ 22.11
Astronomy and Astrophysics 31.00/ 27.00/ 28.00
Biology 19.81/ 19.50/ 22.33
Business and Management 17.92/ 16.98/ 18.87
Chemistry 27.93/ 27.43/ 29.43
Computer Science 27.80/ 27.35/ 30.49
Defense and Security 35.00/ 29.00/ 32.00
Earth Sciences 31.00/ 32.00/ 31.00
Economics 26.95/ 23.35/ 23.35
Education 30.00/ 24.00/ 23.00
Energy and Power 24.00/ 20.00/ 26.00
Engineering 26.80/ 27.00/ 26.60
Environmental Science 26.82/ 23.46/ 24.02
Ethics and Human Rights 17.00/ 18.00/ 19.00
Finance and Investment 25.00/ 22.50/ 24.29
Food Science 22.00/ 24.00/ 25.00
Geography 26.56/ 26.56/ 27.08
Health and Medicine 25.23/ 28.04/ 29.91
History 29.16/ 26.85/ 26.09
Information Technology 32.00/ 30.00/ 33.00
International Relations 16.00/ 21.00/ 21.00
Language Studies 28.74/ 27.59/ 29.31
Law and Ethics 27.50/ 20.62/ 23.75
Literature and Linguistics 28.17/ 23.94/ 25.35
Logical Reasoning 26.49/ 26.73/ 26.24
Materials Science 24.00/ 27.00/ 31.00
Mathematics 25.00/ 23.80/ 24.20
Media and Communication 33.00/ 36.00/ 34.00
Music and Performing Arts 23.00/ 17.00/ 25.00
Physics 26.80/ 25.40/ 27.40
Politics and Governance 29.25/ 28.77/ 27.36
Psychology 26.00/ 26.00/ 23.00
Public Administration 27.00/ 32.00/ 34.00
Religion and Spirituality 20.00/ 27.00/ 26.00
Social Welfare and Development 32.00/ 27.00/ 30.00
Sociology 25.29/ 24.71/ 25.88
Sports and Recreation 21.43/ 22.38/ 21.90
Technology and Innovation 22.00/ 21.00/ 21.00
Transportation and Logistics 22.00/ 24.00/ 27.00
Table 25: Detailed subject-wise evaluation for Telugu-LLM-Labs/Telugu-Llama2-7B-v0-Base on MILU across different languages. The results reported are X / Y / Z where X denote the 0-shot, Y denotes 1-shot and Z denotes 5-shot performance respectively.
Topic bn en gu hi kn ml mr or pa ta te
Agriculture 70 79.5 79.8 75.7 74 63.6 67 63 74 65 72
Anthropology 79 80 69 76.8 73 73 83 72 72 76 80
Architecture and Design 77 80 80 77 70 75 72 74 69 66 81
Arts and Culture 86.1 85.8 83 82 86.5 80 83.6 78 76 74.5 82.9
Astronomy and Astrophysics 88 95 92 88.3 88 85 84 87 85 88 88
Biology 87.3 92.4 83 90 82 82.8 79.1 82 89 76.8 81.4
Business and Management 67 73.4 62 73.8 64 61 67 65 67 63 59.4
Chemistry 78.6 91.2 76.7 89.2 75.5 79 78.1 78.8 82 73.6 78.6
Computer Science 56.8 65.6 63 61.2 65.7 62 55.6 61 71 57 56.1
Defense and Security 88 87 82 89.1 92 80 90 82 81 76 86
Earth Sciences 81 90 78 83.8 85 79 78 63.6 74 84 80
Economics 79.1 81.3 75 81 77 63 70.4 71 80 77.8 70.7
Education 70 72.6 58 66.6 65 58 65 50 57 59 62
Energy and Power 81 86.5 81 80.7 78 78 81 77 83 65 76
Engineering 73 72.6 75.4 66 76.4 73.9 72.6 64.3 56 60.2 74.6
Environmental Science 81 86.6 80 82 84.4 74 80 71.7 74 78 77.7
Ethics and Human Rights 69 82 71 71 68 66 64 67 62 69 72
Finance and Investment 60.9 66 62 65.2 66 53 59.3 62 56 52.5 51.1
Food Science 81 88 80 86.9 80 76 80 74 82 82 86
Geography 84.4 85.4 80 82.9 86.8 74.7 74 81 77 71.4 77.1
Health and Medicine 82 89.8 82 90 86 83 71.4 79 87 81 81.3
History 90.7 80.7 84 85.1 86.4 72 76.6 73 80 74.9 75.4
Information Technology 90 92 86 87.2 90 91 92 89 86 81 85
International Relations 85 91.7 80 84 82 82 80 77 76 84 80
Language Studies 57 92 57 61.5 58.9 56 59.7 58 60 52.1 56.3
Law and Ethics 91.9 86.6 76 86.6 80.3 74 75 73 84 81.1 78.1
Literature and Linguistics 84.7 83.2 82 79.2 83.3 80 82.3 69 78 51 70.4
Logical Reasoning 66.1 62.7 62.3 58.3 66.4 55 61.3 58 64 61.4 57.9
Materials Science 84 84.8 73 78.6 71 75 71 65 67 58 73
Mathematics 45.2 51.2 47.2 45.5 49 51 49 46 40.9 47 43.6
Media and Communication 78 84 79 82.3 84 77 75 78 78 81 81
Music and Performing Arts 76 86 80 85.4 79 78 80 72 81 71 84
Physics 72.8 82.2 65.2 75.8 70 68.4 71.4 62.2 64 62.4 65
Politics and Governance 90.2 79.4 78 84.7 87 78 70.5 75 71 80.1 74.3
Psychology 77 91 80 81.4 72 76 74 77 75 66 77
Public Administration 60 73.7 65.7 76.8 61.6 62 51.5 59 66.7 66 66
Religion and Spirituality 87 93 85 89.5 88 82 85 78 80 80 81
Social Welfare and Development 73 77.7 75 80.8 69 68 78 69 78 64 72
Sociology 65.2 68.6 55 63.6 61 56 72.3 49 53 59.5 56.5
Sports and Recreation 86.6 90.4 83 86.4 86.5 85 84.2 81 83 80.8 85.7
Technology and Innovation 84 86.8 83 82.8 86 84 84 75 82 73 74
Transportation and Logistics 69 72.3 66 69.4 66.7 58 61.2 60.6 67 64 56
Table 26: Detailed subject-wise evaluation for GPT-4o on MILU across different languages. The results reported are for 0-shot experiments.
Topic bn en gu hi kn ml mr or pa ta te
Agriculture 60 64 60.6 61 56 43.4 45 43 51 51 53
Anthropology 55 67 49 63.6 48 49 60 39 58 52 60
Architecture and Design 55 60 52 53 48 50 53 39 51 51 56
Arts and Culture 65.5 76 54 65.9 57.7 57 57.9 45 62 53.9 55.8
Astronomy and Astrophysics 76 92 78 76.6 73 68 68 62 67 67 71
Biology 75 86.8 68 79.6 57.8 59.6 57.9 49 68 53.8 58.8
Business and Management 62 63.8 56 60.2 58 46 54 44 57 49 44.3
Chemistry 62.1 84.6 64.7 75.2 58.2 62 58.6 53.8 70 46.9 54.4
Computer Science 47.9 56.8 51 50.6 54.4 51 45.6 44 58 43.6 45.7
Defense and Security 68 81 56 66.8 58 55 61 52 54 52 55
Earth Sciences 61 77.5 61 62.9 56 60 65 46.5 56 60 57
Economics 59.7 70.9 54 62.3 65 51 47.2 48 62 57.1 54.5
Education 49 58.3 47 51.8 55 47 55.4 36 41 48 54
Energy and Power 62 77 68 65.8 59 60 59 49 56 47 59
Engineering 57.2 57.6 59 52.6 54.2 53.8 57.4 47.3 47 42.8 54.4
Environmental Science 66 73.5 64 67.5 68.8 53 64.2 49.5 58 55 58.7
Ethics and Human Rights 64 71 61 62 61 62 60 55 57 50 63
Finance and Investment 54 56.8 45 53 51.9 50 48.6 40 51 44.7 36.8
Food Science 65 77 63 66.7 63 64 71 60 64 60 67
Geography 60.9 69.6 59 67.5 61.2 49.5 53.3 49 55 50.3 59.4
Health and Medicine 78 83.2 68 81.8 73 73 58.9 62 80 61.2 62.6
History 69.9 65.7 63 64.1 61.9 56 52 49 71 55.3 47.1
Information Technology 88 88 81 87.7 80 88 80 80 76 70 84
International Relations 66 84.2 66 69.6 65 60 60 53 65 65 62
Language Studies 37 87.2 48 50.9 44.2 41 47 46 53 41.2 38.5
Law and Ethics 71.6 71 53 68 56.1 50 54.7 35 46 48 48.1
Literature and Linguistics 63.1 69.7 55 61.1 50 55 56.6 38 56 40.1 51.4
Logical Reasoning 47.9 47.7 40.3 41.9 40.2 38 45.5 39 47 39 43.8
Materials Science 60 73 46 59.4 51 55 53 38 59 38 55
Mathematics 38.4 39.2 37.6 38.5 39 37.8 39.4 35.4 38.3 40.2 34.4
Media and Communication 66 82 71 73.5 69 60 65 58 66 64 64
Music and Performing Arts 57 69 56 61.6 60 50 56 34 49 45 61
Physics 54.6 76 52.4 60.4 49.4 52.1 55.6 39.2 52 46 46.8
Politics and Governance 69.8 63.5 49 66.3 64.8 57 50.6 44 50 58.5 53.3
Psychology 59 87 65 66.7 58 57 57 52 67 57 68
Public Administration 49 66.7 53.5 62.6 43.4 44 41.4 53 58.6 47 50
Religion and Spirituality 65 86 68 74.2 67 68 61 52 57 57 71
Social Welfare and Development 60 64.3 59 64.2 50 41 61 39 57 52 48
Sociology 47.8 56.8 42 49.2 53 40 50.6 41 51 44.1 47.1
Sports and Recreation 64.9 75 58 66.2 62.4 53 65 45 55 55.8 58.6
Technology and Innovation 60 76 69 67.6 67 64 65 60 63 58 58
Transportation and Logistics 45 60 56 49.8 42.4 45 44.9 46.5 52 55 38
Table 27: Detailed subject-wise evaluation for GPT-4o-mini on MILU across different languages. The results reported are for 0-shot experiments.
Topic bn en gu hi kn ml mr or pa ta te
Agriculture 60 76.7 66.7 72.2 66 51.5 53 52 59 50 60
Anthropology 67 73 57 68.7 60 49 69 50 58 60 69
Architecture and Design 59 69 46 65 48 53 67 51 52 57 57
Arts and Culture 58.8 77.8 58 69.9 61.3 51 62.3 51 56 56.4 58.3
Astronomy and Astrophysics 81 94 89 82.8 82 73 81 81 74 78 74
Biology 80.3 90.2 73 85.8 83.2 77.8 71.9 70 81 74.8 70.1
Business and Management 61 71.8 56 67.9 61 62 61 44 62 60 51.9
Chemistry 80.5 90.8 75 88.2 81 76 76.5 69.2 83 76 78.1
Computer Science 55.2 58.8 61 53.4 59.8 56 49.1 55 61 52.9 49.3
Defense and Security 66 83 59 73 63 55 71 63 64 56 65
Earth Sciences 67 87.5 67 75.6 73 71 70 50.5 59 68 66
Economics 73.4 75.3 64 74.2 75 60 60.6 57 61 77.8 59.3
Education 54 67.8 54 60.4 55 56 55.4 44 48 59 56
Energy and Power 68 82.4 71 70.5 69 76 67 59 62 60 64
Engineering 65.8 69.2 68.8 60.4 70.4 64.5 69.6 53.9 54 59.6 67.8
Environmental Science 73 80 67 78.8 73.4 66 63.7 56.6 69 67 70.9
Ethics and Human Rights 58 72 60 61 56 60 56 58 59 61 61
Finance and Investment 70.3 73 73 67.4 75.3 58 67.8 59 52 76 61.4
Food Science 73 85 73 77.8 69 76 78 67 79 77 78
Geography 64.8 78 67 71.5 74.9 65.7 63.9 62 68 66.3 63
Health and Medicine 76 87.4 66 85.8 84 79 68.8 72 83 75 74.8
History 72 75.1 69 74.2 75.5 54 60.6 61 68 67.6 54.5
Information Technology 79 88 78 77.7 79 78 80 70 71 74 70
International Relations 63 81.7 68 70 70 71 72 55 68 70 70
Language Studies 52 89 49 57.1 55.8 43 58.4 47 62 49.1 42
Law and Ethics 70.9 76.8 60 71.4 72.6 56 62.5 45 71 69.3 55.6
Literature and Linguistics 63.1 74.3 52 65.1 58.3 59 59.3 50 57 36.4 55.6
Logical Reasoning 60.9 57.8 63 55.9 60.6 51 54.9 55 50 52.2 55.2
Materials Science 80 77.9 71 69.2 65 68 68 56 69 61 77
Mathematics 58.8 64.2 60.6 56.5 58.2 58.2 65 56.4 52.7 60.2 53
Media and Communication 70 81 68 74.1 70 59 65 51 68 69 65
Music and Performing Arts 51 69 60 65.6 58 59 65 46 58 56 60
Physics 78 84.2 75.5 74 79.8 74.2 80.4 64.6 70 73.8 72.4
Politics and Governance 71 69.7 51 71.1 66.4 60 52.5 53 60 70.1 54.7
Psychology 72 92 71 69 66 63 69 66 73 62 73
Public Administration 51 62.6 51.5 62.6 52.5 44 38.4 44 50.5 53 50
Religion and Spirituality 68 85 65 75.4 73 68 68 61 61 72 68
Social Welfare and Development 59 69.6 65 73.1 56 48 65 43 68 59 57
Sociology 60 66.4 49 57.2 50 41 55.3 47 51 46.8 53.5
Sports and Recreation 66.3 80 58 72.4 67.4 71 72.3 52 73 60.3 61.4
Technology and Innovation 72 79.2 73 73.4 69 72 75 63 64 66 64
Transportation and Logistics 43 66.9 66 59 48.5 47 45.9 47.5 57 51 43
Table 28: Detailed subject-wise evaluation for Gemini-1.5-Pro on MILU across different languages. The results reported are for 0-shot experiments.
Topic bn en gu hi kn ml mr or pa ta te
Agriculture 64 65.6 58.6 62.6 60 55.6 57 59 57 52 63
Anthropology 58 66 49 57.6 58 54 62 59 65 54 56
Architecture and Design 58 60 57 60 51 52 64 46 58 52 51
Arts and Culture 58.2 71 60 63.9 62.6 54 59.6 46 54 60 54.8
Astronomy and Astrophysics 79 91 84 83.6 78 71 83 80 76 80 71
Biology 75.4 87 72 84.6 78.6 72.7 69.3 66 82 69.7 73
Business and Management 50 64.7 46 59.3 56 53 51 56 62 52 46.2
Chemistry 71.4 87.6 74.1 84.4 70.6 73 66.5 73.1 77 66.2 70.8
Computer Science 49.5 56.6 57 51 55 59 48.5 53 53 54.1 46.2
Defense and Security 53 76 52 56.9 56 49 53 53 48 51 50
Earth Sciences 66 79.3 65 70.9 81 69 66 54.5 65 66 68
Economics 61.2 71.1 67 65.9 66 51 54.2 63 70 62.7 55.7
Education 50 57 48 52.4 51 49 54.1 37 48 48 47
Energy and Power 62 75 74 66.1 70 71 71 62 67 63 69
Engineering 59.6 59.6 66.2 59.6 62.4 65.4 59.2 61.3 57 55 62.8
Environmental Science 63 74.5 68 70.9 75 60 67.9 66.7 73 61 62
Ethics and Human Rights 61 65 66 60 63 63 59 58 65 61 59
Finance and Investment 49 57.4 46 51.2 55.6 47 49.2 44 57 50.3 44.3
Food Science 65 84 71 81.8 77 73 75 71 72 71 79
Geography 62.6 69.2 59 71.5 68.5 57.6 60.9 63 64 60 61.5
Health and Medicine 74 82.4 72 81 81 80 63.4 71 80 70.7 76.6
History 64.5 62.9 58 66.5 71.1 64 55.8 52 70 56.7 51.7
Information Technology 83 83 78 81.6 76 82 75 79 74 67 75
International Relations 58 72.5 67 68.1 65 62 63 62 59 70 58
Language Studies 44 85.2 54 51.9 51.9 44 51.7 48 61 44.6 45.4
Law and Ethics 70.9 69 57 66.2 59.9 57 59.5 46 62 54.3 53.1
Literature and Linguistics 52.3 65.1 51 59.1 55.8 50 50.4 44 62 43.7 52.8
Logical Reasoning 54.8 53 56.5 53.9 58.9 55 52.4 51 51 51.8 52.7
Materials Science 72 73.4 58 65.3 66 59 68 54 63 55 58
Mathematics 44.6 44.4 48.6 43.3 51.2 47.3 48.6 47.8 43.9 45.4 39.6
Media and Communication 68 76 67 72.8 71 61 59 60 62 65 63
Music and Performing Arts 40 59 53 57.6 53 48 53 46 56 43 44
Physics 68 75.8 66 69.4 68.2 70 65.8 58.7 62 66 61.8
Politics and Governance 70.6 63.1 52 63.9 64 52 51.1 44 46 61 51.7
Psychology 70 88 63 65.1 67 64 65 70 67 60 68
Public Administration 48 55.6 51.5 51.5 44.4 48 40.4 55 54.5 47 45
Religion and Spirituality 68 79 70 69 69 66 66 54 61 65 65
Social Welfare and Development 57 65.2 59 58 50 41 51 48 63 54 58
Sociology 55.7 54 45 55.6 56 47 50.3 42 52 46.8 48.2
Sports and Recreation 58.9 66.6 53 58.4 62.9 55 61.6 53 56 54.5 58.6
Technology and Innovation 66 73.2 68 67.4 61 73 64 64 67 55 57
Transportation and Logistics 44 56.9 54 55.5 52.5 48 38.8 44.4 53 48 51
Table 29: Detailed subject-wise evaluation for Gemini-1.5-Flash on MILU across different languages. The results reported are for 0-shot experiments.
Topic bn en gu hi kn ml mr or pa ta te
Agriculture 51 45.3 52 42.7 49 36.7 38 38 37.9 42 37
Anthropology 47 42 43 43.4 41 40.2 44 38 40 36 35
Architecture and Design 38 50 41.9 43 45 42 49.5 32.3 34 34 43.4
Arts and Culture 55.8 57.4 54 54.8 52.8 52 46.1 37 44 48.8 47.7
Astronomy and Astrophysics 53 53 57 55.9 53 50 57.6 43 37 51 45
Biology 57.4 63.7 61 57.6 60.7 45.5 47.6 40 48 48.1 46.1
Business and Management 40 41.2 46 43.7 43 38 42.9 35 41 36 29.2
Chemistry 32.4 49.7 40 49.2 34.3 40 34.9 30.8 44 34.5 34.8
Computer Science 26 35.1 24 31.1 36.1 34 28 34 29 24.4 23.8
Defense and Security 52 57 45 43.9 39 36 49 41 42 36 43
Earth Sciences 48 50.5 45 45.7 53 49 42 33.3 30 49 46
Economics 46 47.5 49 45.1 54 38 45.5 33 39 43.7 38.9
Education 29 36.3 32 32 32 33 34.7 27 30 31.2 37
Energy and Power 57 54.7 47.9 47.8 54 54.5 47.2 48 52 41 48
Engineering 41.8 36.6 39.4 33.4 41.8 40.3 39 34.3 32 38.2 38
Environmental Science 49 52.9 48.3 50.3 51.2 43 40.6 40.8 48 42.1 47.5
Ethics and Human Rights 43 48.5 44.7 53 40 39.8 38 41.5 34 39 37.4
Finance and Investment 37.6 33 32 31.8 40.7 38 35 25 35 32.6 26.4
Food Science 46 62 49 54.5 52 52 52 41 48 52 56
Geography 54.2 46 60 52.8 58.9 45.5 48.8 41 43.8 46.5 44.5
Health and Medicine 63 62 56 62.3 67 57 48.4 46 53 50.9 46.7
History 61.3 43.8 59.6 56.1 62.5 46 45.5 43 50 50.6 41
Information Technology 63 64 54.9 61.2 60 61 56 55 46 47 62
International Relations 48 51.7 46 46.3 46 45 43 36.8 41 53 47
Language Studies 19 56.4 34 38.7 27.9 32 26.8 27 37 28.1 28.2
Law and Ethics 46.6 39.2 38.4 46.2 42.7 41 37.7 33 35 41.7 32.9
Literature and Linguistics 51.4 45.9 42 50.3 50 39 44.2 39 38 41.1 47.2
Logical Reasoning 27 24.7 26.1 27.4 29 21 27.1 22 24 24 28.1
Materials Science 39 45.6 38 41 49 44 39 33 45 31 32
Mathematics 19.6 25.4 23.7 23.4 20.8 25.3 19.8 20.6 28 22 24.4
Media and Communication 48 52 48 45.6 51 50.5 48 35 49 45 55
Music and Performing Arts 39 45 39 43.7 48 40 52 43 42 36 46
Physics 34.2 46.2 32.5 42.1 37.4 32.1 34.9 30.3 35 32.6 32.9
Politics and Governance 59.2 42.9 48 48.2 55.9 44.4 38.9 35 45 51.9 39.8
Psychology 57 61 54 53.5 44 47.9 44.4 42 47 52 53
Public Administration 40 46.5 41.4 45.7 34.3 32.3 30.2 41.4 37.8 34 40
Religion and Spirituality 62 62 64 61.5 66 59 61 51 57 55 59
Social Welfare and Development 38 47.7 48 38.1 41 34 46.5 33 41 41 40.8
Sociology 35.7 38 35 35.8 34 32 34.2 34 34 37.3 34.3
Sports and Recreation 54.5 50.6 42 51.8 51.1 40 49.4 35 47 45.8 45.7
Technology and Innovation 48 50.8 38 42.2 56 49 48 45 38 44 38
Transportation and Logistics 28 36.2 41 35.1 31.3 33 35.8 33.3 33 40 28
Table 30: Detailed subject-wise evaluation for Krutrim-spectre-v2 on MILU across different languages. The results reported are for 0-shot experiments.
Topic bn en gu hi kn ml mr or pa ta te
Agriculture 32 35.4 38.4 31 30 33.3 27 31 30 28 28
Anthropology 41 40 27 38.4 31 27 31 29 34 26 35
Architecture and Design 31 35 32 30 30 20 19 30 31 30 34
Arts and Culture 28.5 37.2 29 31.3 32.5 24 31.1 27 23 29.7 30.1
Astronomy and Astrophysics 46 45 42 42.2 39 24 41 31 37 40 37
Biology 40.5 46.9 36 38.7 38.9 30.3 31.3 30 39 31.2 34.9
Business and Management 31 37.8 30 34 27 27 29 25 28 34 25.5
Chemistry 34.1 41.3 33.6 40.9 33.3 38 33.8 26 30 27.9 34.9
Computer Science 29.2 31.6 42 31.2 36.1 34 30.2 28 33 30.2 34.5
Defense and Security 29 24 35 31.8 32 29 26 31 37 23 31
Earth Sciences 41 44.1 31 35.7 37 37 42 32.3 36 33 36
Economics 37.4 36.1 39 31.1 43 30 37.3 34 30 36.5 32.9
Education 31 23.9 19 25.1 26 29 35 32 31 28 38
Energy and Power 30 28.4 40 35.9 42 27 36 28 30 36 27
Engineering 30.4 32.6 32.2 29.8 29.4 30.8 30.4 30.4 24 28.4 30.2
Environmental Science 43 39.9 26 38.3 35.2 29 36.3 36.4 45 26 36.3
Ethics and Human Rights 34 30 41 34 29 32 29 27 29 34 24
Finance and Investment 34.2 32.6 38 31.4 31.5 32 28.2 27 29 30.2 28.2
Food Science 39 46 42 45.5 36 31 43 37 40 39 51
Geography 40.8 35.2 34 34.3 37 31.3 31.9 28 33 34.9 30.2
Health and Medicine 33 49.3 43 43.7 45 34 39.3 40 53 39.7 35.5
History 28.5 36.5 42 34.9 36.3 31 37.2 29 41 34.5 32.7
Information Technology 50 42 46 43.6 45 45 46 45 36 38 42
International Relations 36 42.5 34 38.5 34 36 43 40 40 26 40
Language Studies 35 33.8 31 24.6 26.4 25 27.5 32 39 27.3 22.4
Law and Ethics 30.4 39 38 36.6 33.8 37 28.4 30 37 32.3 29.4
Literature and Linguistics 38.7 31.1 33 28.5 31.7 35 25.7 30 30 24.7 34.5
Logical Reasoning 29.2 22.2 19.5 25.1 23.6 28 25.1 28 27 25.1 26.7
Materials Science 32 42.2 26 34.3 36 34 29 24 26 25 19
Mathematics 28.8 27.8 29 30.3 29.2 24.6 28.6 26.2 23.5 27.6 29.2
Media and Communication 39 35 37 27.9 42 36 40 37 35 30 44
Music and Performing Arts 34 39 32 20.5 32 28 24 27 24 25 24
Physics 27.4 40.4 32.3 34.6 27.2 25.8 32.4 31.2 37 28.8 31.4
Politics and Governance 36.5 31.9 32 31.3 34.8 26 29.5 30 19 33.6 34
Psychology 34 36 41 29.5 39 37 31 39 36 31 37
Public Administration 25 30.3 28.3 30.3 35.4 27 21.2 28 36.4 31 32
Religion and Spirituality 34 41 40 33.9 36 47 29 39 29 39 37
Social Welfare and Development 25 31.2 32 26.9 35 31 33 23 34 20 24
Sociology 32.2 28.2 32 30 31 30 24.5 31 23 27 34.1
Sports and Recreation 36.6 33.8 29 30.2 35.4 21 35 31 28 33.3 29
Technology and Innovation 38 37 30 35.6 35 26 43 28 38 35 27
Transportation and Logistics 25 26.9 30 26.5 27.3 32 17.3 28.3 27 22 22
Table 31: Detailed subject-wise evaluation for sarvamai/sarvam-1 on MILU across different languages. The results reported are for 5-shot experiments.
Topic bn en gu hi kn ml mr or pa ta te
Agriculture 28.0/ 34.0/ 29.0 24.3/ 29.0/ 29.8 29.3/ 30.3/ 23.2 22.1/ 26.8/ 26.0 21.0/ 27.0/ 26.0 38.4/ 31.3/ 31.3 27.0/ 26.0/ 27.0 29.0/ 26.0/ 30.0 19.0/ 26.0/ 25.0 26.0/ 29.0/ 26.0 26.0/ 25.0/ 26.0
Anthropology 21.0/ 15.0/ 25.0 24.0/ 37.0/ 39.0 21.0/ 23.0/ 28.0 21.2/ 24.2/ 28.3 31.0/ 27.0/ 28.0 25.0/ 25.0/ 29.0 23.0/ 25.0/ 21.0 28.0/ 29.0/ 28.0 19.0/ 20.0/ 23.0 27.0/ 23.0/ 23.0 19.0/ 24.0/ 19.0
Architecture and Design 25.0/ 19.0/ 25.0 27.0/ 23.0/ 39.0 30.0/ 25.0/ 20.0 22.0/ 29.0/ 38.0 28.0/ 24.0/ 22.0 28.0/ 27.0/ 25.0 23.0/ 24.0/ 25.0 17.0/ 23.0/ 22.0 25.0/ 23.0/ 27.0 29.0/ 32.0/ 29.0 28.0/ 26.0/ 24.0
Arts and Culture 24.2/ 24.9/ 29.7 31.0/ 32.8/ 36.8 30.0/ 24.0/ 22.0 23.2/ 23.6/ 27.1 23.3/ 20.9/ 24.5 31.0/ 27.0/ 27.0 21.9/ 24.6/ 24.0 26.0/ 24.0/ 21.0 29.0/ 25.0/ 27.0 29.7/ 26.1/ 23.6 27.1/ 28.1/ 23.1
Astronomy and Astrophysics 22.0/ 24.0/ 20.0 27.0/ 37.0/ 35.0 30.0/ 34.0/ 33.0 24.2/ 28.9/ 28.9 24.0/ 22.0/ 23.0 24.0/ 24.0/ 23.0 24.0/ 27.0/ 35.0 25.0/ 30.0/ 29.0 24.0/ 23.0/ 24.0 21.0/ 20.0/ 16.0 20.0/ 31.0/ 19.0
Biology 23.6/ 24.3/ 21.5 33.7/ 40.7/ 44.7 21.0/ 17.0/ 21.0 26.1/ 28.3/ 30.7 23.8/ 28.3/ 25.8 24.2/ 22.2/ 24.2 34.3/ 29.5/ 31.9 26.0/ 20.0/ 27.0 21.0/ 24.0/ 18.0 25.2/ 25.5/ 25.2 20.8/ 24.2/ 23.6
Business and Management 28.0/ 22.0/ 21.0 27.2/ 36.2/ 35.0 20.0/ 17.0/ 20.0 26.2/ 27.9/ 29.3 25.0/ 24.0/ 24.0 30.0/ 22.0/ 26.0 20.0/ 22.0/ 16.0 16.0/ 21.0/ 18.0 27.0/ 26.0/ 24.0 24.0/ 21.0/ 23.0 17.0/ 21.7/ 20.8
Chemistry 23.6/ 27.2/ 27.2 32.9/ 34.1/ 35.9 26.7/ 24.1/ 24.1 26.9/ 30.3/ 26.9 26.1/ 31.4/ 29.1 27.0/ 22.0/ 26.0 31.1/ 28.2/ 29.5 26.9/ 26.0/ 23.1 19.0/ 22.0/ 19.0 27.6/ 28.8/ 28.2 25.9/ 28.9/ 28.2
Computer Science 24.5/ 27.1/ 24.0 27.8/ 30.2/ 29.2 28.0/ 21.0/ 27.0 27.2/ 26.8/ 24.8 28.4/ 29.0/ 27.8 23.0/ 32.0/ 33.0 28.4/ 33.1/ 31.9 28.0/ 29.0/ 26.0 25.0/ 20.0/ 21.0 23.3/ 26.7/ 26.2 29.1/ 31.4/ 30.9
Defense and Security 25.0/ 25.0/ 19.0 27.0/ 36.0/ 37.0 23.0/ 21.0/ 22.0 24.6/ 24.2/ 22.3 23.0/ 22.0/ 20.0 22.0/ 24.0/ 24.0 21.0/ 22.0/ 20.0 28.0/ 23.0/ 18.0 33.0/ 27.0/ 26.0 29.0/ 24.0/ 29.0 26.0/ 31.0/ 24.0
Earth Sciences 28.0/ 20.0/ 26.0 29.3/ 36.1/ 39.8 30.0/ 27.0/ 33.0 26.9/ 29.3/ 28.3 29.0/ 27.0/ 26.0 18.0/ 17.0/ 21.0 24.0/ 24.0/ 27.0 29.3/ 26.3/ 24.2 26.0/ 27.0/ 23.0 28.0/ 27.0/ 27.0 26.0/ 28.0/ 32.0
Economics 28.8/ 24.5/ 30.2 29.1/ 33.5/ 32.1 21.0/ 18.0/ 18.0 29.8/ 29.8/ 28.8 22.0/ 15.0/ 25.0 29.0/ 28.0/ 33.0 29.6/ 31.0/ 32.4 24.0/ 27.0/ 27.0 20.0/ 25.0/ 24.0 28.6/ 24.6/ 24.6 26.4/ 28.7/ 23.9
Education 24.0/ 30.0/ 28.0 25.8/ 29.6/ 25.5 22.0/ 24.0/ 27.0 29.6/ 24.3/ 24.6 24.0/ 20.0/ 23.0 20.0/ 24.0/ 25.0 32.5/ 28.0/ 28.0 22.0/ 23.0/ 28.0 24.0/ 27.0/ 24.0 24.0/ 26.0/ 29.0 24.0/ 27.0/ 32.0
Energy and Power 28.0/ 22.0/ 21.0 23.6/ 31.1/ 32.4 33.0/ 26.0/ 28.0 29.1/ 27.8/ 28.1 26.0/ 22.0/ 33.0 27.0/ 26.0/ 30.0 26.0/ 31.0/ 28.0 21.0/ 27.0/ 25.0 27.0/ 29.0/ 31.0 33.0/ 23.0/ 25.0 36.0/ 37.0/ 35.0
Engineering 23.2/ 24.8/ 25.2 30.0/ 29.2/ 30.8 20.6/ 24.0/ 22.2 26.8/ 26.6/ 26.4 22.8/ 21.4/ 23.4 24.8/ 23.9/ 25.6 25.0/ 24.2/ 21.4 25.3/ 26.8/ 26.5 23.0/ 27.0/ 29.0 23.0/ 27.0/ 24.8 27.2/ 23.4/ 23.4
Environmental Science 26.0/ 30.0/ 31.0 32.9/ 36.3/ 39.7 31.0/ 24.0/ 23.0 25.2/ 27.3/ 29.3 21.1/ 25.0/ 25.8 23.0/ 29.0/ 24.0 22.1/ 24.2/ 23.7 22.2/ 24.2/ 25.2 31.0/ 32.0/ 27.0 25.0/ 27.0/ 27.0 27.9/ 28.5/ 26.8
Ethics and Human Rights 24.0/ 21.0/ 31.0 30.0/ 33.0/ 34.0 23.0/ 24.0/ 25.0 25.0/ 30.0/ 22.0 26.0/ 30.0/ 31.0 19.0/ 18.0/ 18.0 27.0/ 29.0/ 29.0 25.0/ 23.0/ 18.0 23.0/ 20.0/ 20.0 21.0/ 21.0/ 19.0 26.0/ 31.0/ 29.0
Finance and Investment 26.7/ 28.2/ 30.2 29.8/ 30.6/ 32.0 22.0/ 24.0/ 23.0 27.2/ 28.0/ 26.0 25.9/ 25.9/ 28.4 25.0/ 28.0/ 21.0 25.4/ 25.4/ 27.1 22.0/ 30.0/ 23.0 18.0/ 28.0/ 21.0 24.0/ 25.1/ 22.9 25.0/ 28.2/ 27.5
Food Science 21.0/ 25.0/ 28.0 34.0/ 39.0/ 45.0 26.0/ 26.0/ 27.0 21.2/ 23.2/ 25.2 19.0/ 23.0/ 30.0 22.0/ 23.0/ 21.0 22.0/ 21.0/ 25.0 20.0/ 23.0/ 26.0 22.0/ 20.0/ 25.0 20.0/ 18.0/ 26.0 20.0/ 26.0/ 26.0
Geography 25.1/ 23.5/ 27.9 26.6/ 32.2/ 33.8 24.0/ 31.0/ 35.0 28.7/ 29.7/ 27.5 27.9/ 21.0/ 22.4 31.3/ 24.2/ 22.2 31.9/ 31.9/ 31.9 32.0/ 31.0/ 27.0 21.0/ 23.0/ 30.0 26.3/ 26.9/ 29.1 27.6/ 22.9/ 25.5
Health and Medicine 20.0/ 20.0/ 23.0 34.7/ 40.3/ 43.5 27.0/ 26.0/ 23.0 25.2/ 29.9/ 30.1 15.0/ 22.0/ 26.0 34.0/ 31.0/ 30.0 21.4/ 24.1/ 29.5 26.0/ 22.0/ 19.0 24.0/ 24.0/ 22.0 27.6/ 23.3/ 25.9 29.9/ 27.1/ 27.1
History 24.3/ 22.4/ 21.9 29.7/ 31.1/ 33.5 26.0/ 26.0/ 27.0 27.2/ 28.0/ 26.2 26.0/ 21.2/ 20.5 24.0/ 24.0/ 25.0 23.4/ 24.9/ 24.2 32.0/ 25.0/ 28.0 29.0/ 29.0/ 27.0 23.6/ 24.7/ 26.6 27.4/ 25.1/ 29.7
Information Technology 26.0/ 28.0/ 30.0 26.0/ 32.0/ 36.0 32.0/ 22.0/ 32.0 33.5/ 25.1/ 29.0 27.0/ 27.0/ 29.0 28.0/ 23.0/ 30.0 29.0/ 34.0/ 36.0 30.0/ 27.0/ 20.0 31.0/ 31.0/ 26.0 22.0/ 22.0/ 27.0 32.0/ 30.0/ 30.0
International Relations 25.0/ 24.0/ 27.0 33.3/ 39.2/ 36.7 28.0/ 20.0/ 32.0 29.2/ 31.1/ 31.1 25.0/ 33.0/ 27.0 31.0/ 28.0/ 30.0 25.0/ 29.0/ 31.0 25.0/ 25.0/ 28.0 28.0/ 25.0/ 26.0 37.0/ 33.0/ 32.0 16.0/ 19.0/ 16.0
Language Studies 30.0/ 21.0/ 29.0 30.2/ 31.6/ 31.2 31.0/ 25.0/ 32.0 24.9/ 25.9/ 27.3 27.9/ 24.8/ 21.7 23.0/ 27.0/ 30.0 28.9/ 30.9/ 30.9 23.0/ 24.0/ 21.0 26.0/ 28.0/ 29.0 26.6/ 23.2/ 27.7 26.4/ 27.6/ 22.4
Law and Ethics 31.8/ 31.1/ 24.3 34.2/ 35.2/ 35.0 28.0/ 26.0/ 26.0 29.6/ 26.0/ 25.4 19.1/ 17.2/ 22.3 28.0/ 22.0/ 18.0 26.3/ 24.1/ 23.3 25.0/ 26.0/ 22.0 27.0/ 30.0/ 34.0 22.1/ 23.6/ 19.7 26.9/ 26.2/ 26.9
Literature and Linguistics 27.9/ 31.5/ 31.5 28.3/ 33.1/ 31.7 22.0/ 21.0/ 24.0 23.6/ 25.2/ 25.4 27.5/ 25.0/ 27.5 27.0/ 29.0/ 24.0 23.9/ 25.7/ 28.3 19.0/ 26.0/ 26.0 33.0/ 27.0/ 27.0 30.4/ 30.0/ 27.5 26.8/ 28.9/ 26.1
Logical Reasoning 25.3/ 27.0/ 25.1 27.6/ 26.8/ 26.6 26.0/ 26.6/ 25.3 24.9/ 25.7/ 24.6 29.9/ 26.6/ 29.5 29.0/ 27.0/ 27.0 23.7/ 25.9/ 24.5 24.0/ 26.0/ 33.0 26.0/ 32.0/ 20.0 27.1/ 25.9/ 26.3 26.0/ 23.0/ 23.5
Materials Science 26.0/ 28.0/ 33.0 28.1/ 33.1/ 33.1 21.0/ 24.0/ 24.0 23.6/ 24.7/ 26.9 24.0/ 23.0/ 25.0 25.0/ 31.0/ 29.0 27.0/ 29.0/ 31.0 26.0/ 27.0/ 29.0 22.0/ 26.0/ 26.0 31.0/ 25.0/ 26.0 20.0/ 19.0/ 26.0
Mathematics 28.0/ 22.6/ 24.2 25.6/ 25.8/ 27.0 28.4/ 27.8/ 25.8 23.2/ 25.7/ 23.2 26.2/ 25.4/ 22.0 30.1/ 27.5/ 22.4 24.2/ 24.6/ 26.0 28.2/ 26.0/ 24.2 25.8/ 26.9/ 29.5 23.2/ 24.8/ 28.6 26.4/ 26.6/ 25.8
Media and Communication 32.0/ 28.0/ 25.0 31.0/ 41.0/ 39.0 26.0/ 25.0/ 23.0 25.9/ 23.8/ 23.8 34.0/ 30.0/ 24.0 24.0/ 23.0/ 23.0 27.0/ 30.0/ 21.0 35.0/ 31.0/ 25.0 18.0/ 18.0/ 25.0 25.0/ 26.0/ 31.0 29.0/ 25.0/ 27.0
Music and Performing Arts 30.0/ 34.0/ 26.0 33.0/ 41.0/ 41.0 27.0/ 20.0/ 20.0 28.5/ 29.1/ 31.1 26.0/ 29.0/ 25.0 35.0/ 29.0/ 26.0 26.0/ 26.0/ 28.0 25.0/ 23.0/ 29.0 25.0/ 28.0/ 28.0 30.0/ 30.0/ 27.0 32.0/ 30.0/ 27.0
Physics 25.4/ 25.6/ 23.6 27.2/ 29.8/ 32.8 23.4/ 30.6/ 26.5 23.2/ 25.6/ 24.2 26.4/ 25.0/ 27.0 28.4/ 29.5/ 23.2 24.4/ 26.2/ 24.8 23.6/ 24.0/ 28.8 28.0/ 27.0/ 23.0 25.4/ 21.4/ 21.4 23.4/ 24.6/ 23.4
Politics and Governance 18.8/ 24.7/ 25.1 30.5/ 34.1/ 31.9 37.0/ 35.0/ 28.0 27.5/ 27.1/ 28.1 22.3/ 26.3/ 25.5 22.0/ 25.0/ 22.0 23.3/ 23.6/ 24.4 30.0/ 21.0/ 25.0 22.0/ 29.0/ 28.0 27.0/ 27.4/ 27.8 22.2/ 28.8/ 28.5
Psychology 20.0/ 19.0/ 19.0 36.0/ 32.0/ 39.0 20.0/ 21.0/ 23.0 24.0/ 30.2/ 27.9 26.0/ 14.0/ 23.0 27.0/ 27.0/ 31.0 26.0/ 27.0/ 30.0 21.0/ 23.0/ 20.0 23.0/ 25.0/ 23.0 22.0/ 20.0/ 22.0 29.0/ 16.0/ 21.0
Public Administration 26.0/ 20.0/ 21.0 33.3/ 44.4/ 44.4 25.2/ 21.2/ 28.3 31.3/ 30.3/ 28.3 26.3/ 29.3/ 36.4 23.0/ 22.0/ 29.0 27.3/ 26.3/ 28.3 31.0/ 34.0/ 28.0 29.3/ 26.3/ 26.3 26.0/ 26.0/ 32.0 19.0/ 28.0/ 31.0
Religion and Spirituality 24.0/ 23.0/ 18.0 27.0/ 35.0/ 37.0 32.0/ 28.0/ 30.0 25.4/ 27.0/ 21.8 25.0/ 27.0/ 24.0 26.0/ 27.0/ 19.0 31.0/ 31.0/ 31.0 24.0/ 20.0/ 19.0 26.0/ 24.0/ 24.0 24.0/ 25.0/ 27.0 24.0/ 27.0/ 25.0
Social Welfare and Development 21.0/ 21.0/ 19.0 24.1/ 29.5/ 34.8 23.0/ 23.0/ 26.0 25.9/ 26.9/ 27.5 20.0/ 26.0/ 22.0 28.0/ 26.0/ 22.0 26.0/ 36.0/ 30.0 27.0/ 24.0/ 29.0 27.0/ 24.0/ 25.0 24.0/ 23.0/ 21.0 27.0/ 22.0/ 25.0
Sociology 24.3/ 28.7/ 27.0 31.8/ 33.2/ 32.4 32.0/ 23.0/ 33.0 24.6/ 27.4/ 27.6 22.0/ 20.0/ 21.0 28.0/ 30.0/ 22.0 25.8/ 26.4/ 23.9 29.0/ 29.0/ 27.0 23.0/ 25.0/ 15.0 25.2/ 22.5/ 24.3 21.8/ 22.9/ 25.9
Sports and Recreation 24.8/ 25.2/ 29.2 29.6/ 35.0/ 35.4 25.0/ 17.0/ 21.0 23.4/ 26.2/ 28.2 21.9/ 23.0/ 23.0 29.0/ 26.0/ 27.0 26.0/ 23.7/ 24.3 20.0/ 18.0/ 17.0 27.0/ 26.0/ 25.0 26.3/ 30.1/ 24.4 29.5/ 26.2/ 23.8
Technology and Innovation 21.0/ 20.0/ 21.0 30.2/ 33.8/ 34.0 26.0/ 26.0/ 25.0 27.2/ 27.0/ 26.8 26.0/ 27.0/ 25.0 30.0/ 23.0/ 19.0 24.0/ 25.0/ 24.0 26.0/ 21.0/ 27.0 23.0/ 23.0/ 20.0 23.0/ 30.0/ 29.0 27.0/ 24.0/ 26.0
Transportation and Logistics 24.0/ 19.0/ 28.0 20.8/ 23.1/ 30.0 20.0/ 24.0/ 24.0 25.2/ 23.3/ 23.3 23.2/ 27.3/ 26.3 27.0/ 25.0/ 22.0 24.5/ 21.4/ 21.4 23.2/ 23.2/ 22.2 19.0/ 20.0/ 18.0 24.0/ 22.0/ 20.0 14.0/ 20.0/ 25.0
Table 32: Detailed subject-wise evaluation for meta-llama/Llama-3.2-1B on MILU across different languages. The results reported are X / Y / Z where X denote the 0-shot, Y denotes 1-shot and Z denotes 5-shot performance respectively.
Topic bn en gu hi kn ml mr or pa ta te
Agriculture 31.0/ 25.0/ 24.0 26.8/ 26.8/ 25.6 31.3/ 29.3/ 24.2 27.4/ 24.8/ 23.3 31.0/ 24.0/ 25.0 30.3/ 24.2/ 23.2 29.0/ 23.0/ 28.0 27.0/ 30.0/ 24.0 27.0/ 22.0/ 23.0 22.0/ 23.0/ 26.0 18.0/ 22.0/ 26.0
Anthropology 26.0/ 33.0/ 30.0 26.0/ 27.0/ 23.0 29.0/ 30.0/ 25.0 33.3/ 37.4/ 25.2 29.0/ 26.0/ 31.0 24.0/ 25.0/ 21.0 29.0/ 27.0/ 24.0 31.0/ 29.0/ 20.0 21.0/ 21.0/ 18.0 21.0/ 25.0/ 30.0 27.0/ 27.0/ 26.0
Architecture and Design 19.0/ 18.0/ 19.0 23.0/ 23.0/ 19.0 23.0/ 28.0/ 20.0 36.0/ 25.0/ 22.0 26.0/ 23.0/ 21.0 28.0/ 30.0/ 28.0 32.0/ 28.0/ 21.0 26.0/ 27.0/ 25.0 32.0/ 27.0/ 27.0 31.0/ 25.0/ 27.0 21.0/ 26.0/ 22.0
Arts and Culture 32.7/ 26.7/ 31.5 32.4/ 25.8/ 27.4 22.0/ 25.0/ 25.0 25.4/ 28.1/ 28.1 23.9/ 22.7/ 21.5 29.0/ 26.0/ 26.0 22.4/ 23.5/ 25.1 23.0/ 27.0/ 23.0 27.0/ 28.0/ 29.0 29.7/ 27.9/ 27.9 22.6/ 25.6/ 29.1
Astronomy and Astrophysics 26.0/ 30.0/ 32.0 34.0/ 28.0/ 29.0 27.0/ 28.0/ 30.0 21.1/ 27.3/ 25.8 26.0/ 29.0/ 30.0 26.0/ 22.0/ 23.0 25.0/ 34.0/ 35.0 25.0/ 18.0/ 24.0 19.0/ 25.0/ 26.0 24.0/ 18.0/ 28.0 27.0/ 22.0/ 26.0
Biology 29.2/ 23.6/ 21.5 30.1/ 26.1/ 23.4 26.0/ 24.0/ 28.0 28.7/ 28.7/ 28.5 27.9/ 26.6/ 26.2 24.2/ 23.2/ 26.3 25.1/ 27.8/ 21.8 28.0/ 26.0/ 26.0 23.0/ 23.0/ 24.0 25.5/ 27.7/ 27.1 23.9/ 23.3/ 20.1
Business and Management 23.0/ 20.0/ 22.0 29.4/ 29.4/ 27.9 26.0/ 31.0/ 20.0 28.8/ 26.5/ 27.9 27.0/ 25.0/ 24.0 22.0/ 29.0/ 31.0 19.0/ 26.0/ 24.0 28.0/ 29.0/ 23.0 28.0/ 24.0/ 26.0 22.0/ 24.0/ 24.0 22.6/ 21.7/ 27.4
Chemistry 28.3/ 29.4/ 27.2 30.9/ 29.5/ 26.1 22.4/ 26.7/ 22.4 33.1/ 27.9/ 27.9 25.5/ 25.5/ 25.2 30.0/ 21.0/ 21.0 28.2/ 26.7/ 25.9 26.9/ 24.0/ 23.1 29.0/ 33.0/ 31.0 27.0/ 24.9/ 27.6 27.9/ 25.4/ 27.2
Computer Science 24.5/ 24.5/ 27.6 29.8/ 27.4/ 26.0 16.0/ 22.0/ 28.0 25.6/ 27.0/ 22.6 24.9/ 22.5/ 27.2 34.0/ 26.0/ 21.0 30.8/ 32.5/ 27.8 23.0/ 32.0/ 25.0 32.0/ 28.0/ 30.0 23.8/ 29.6/ 27.9 27.4/ 26.0/ 26.0
Defense and Security 27.0/ 23.0/ 27.0 29.0/ 25.0/ 23.0 30.0/ 25.0/ 31.0 27.5/ 23.7/ 27.0 30.0/ 35.0/ 35.0 21.0/ 21.0/ 27.0 34.0/ 31.0/ 22.0 25.0/ 20.0/ 24.0 30.0/ 24.0/ 31.0 33.0/ 31.0/ 24.0 34.0/ 37.0/ 36.0
Earth Sciences 39.0/ 28.0/ 23.0 29.1/ 26.4/ 28.0 28.0/ 23.0/ 24.0 26.7/ 25.2/ 25.2 22.0/ 28.0/ 24.0 25.0/ 21.0/ 29.0 27.0/ 26.0/ 27.0 23.2/ 22.2/ 24.2 20.0/ 22.0/ 25.0 29.0/ 21.0/ 24.0 28.0/ 22.0/ 28.0
Economics 27.3/ 25.2/ 27.3 25.1/ 22.7/ 22.7 25.0/ 24.0/ 25.0 26.0/ 28.0/ 25.4 20.0/ 21.0/ 26.0 28.0/ 23.0/ 25.0 27.5/ 24.6/ 21.1 33.0/ 24.0/ 27.0 19.0/ 23.0/ 20.0 24.6/ 21.4/ 25.4 22.2/ 25.8/ 25.1
Education 33.0/ 29.0/ 35.0 28.3/ 23.6/ 23.9 30.0/ 30.0/ 27.0 28.1/ 28.1/ 27.2 33.0/ 33.0/ 33.0 23.0/ 29.0/ 22.0 27.4/ 19.8/ 24.2 23.0/ 26.0/ 26.0 20.0/ 20.0/ 21.0 34.0/ 34.0/ 29.0 31.0/ 27.0/ 36.0
Energy and Power 24.0/ 23.0/ 23.0 31.1/ 29.0/ 24.3 32.0/ 30.0/ 28.0 26.8/ 25.4/ 25.1 26.0/ 25.0/ 17.0 20.0/ 25.0/ 24.0 22.0/ 28.0/ 25.0 24.0/ 25.0/ 25.0 31.0/ 29.0/ 26.0 26.0/ 30.0/ 30.0 24.0/ 20.0/ 23.0
Engineering 25.4/ 22.0/ 23.6 21.8/ 25.6/ 26.6 24.0/ 22.6/ 22.6 26.6/ 25.6/ 24.0 21.8/ 22.8/ 25.2 22.7/ 20.5/ 25.2 25.0/ 25.0/ 23.6 23.2/ 21.7/ 25.6 27.0/ 20.0/ 21.0 28.2/ 25.2/ 27.6 25.4/ 26.6/ 24.8
Environmental Science 28.0/ 31.0/ 29.0 29.5/ 24.2/ 25.7 28.0/ 29.0/ 33.0 24.6/ 27.5/ 29.3 27.3/ 27.3/ 28.1 23.0/ 22.0/ 24.0 28.4/ 23.2/ 27.9 16.2/ 21.2/ 21.2 25.0/ 25.0/ 22.0 28.0/ 30.0/ 29.0 30.7/ 24.6/ 26.8
Ethics and Human Rights 27.0/ 27.0/ 27.0 31.0/ 32.0/ 26.0 29.0/ 27.0/ 26.0 21.0/ 23.0/ 23.0 29.0/ 26.0/ 24.0 25.0/ 16.0/ 21.0 25.0/ 27.0/ 29.0 29.0/ 25.0/ 23.0 30.0/ 25.0/ 21.0 16.0/ 16.0/ 25.0 22.0/ 25.0/ 24.0
Finance and Investment 22.3/ 21.8/ 21.3 28.2/ 27.4/ 26.4 26.0/ 25.0/ 25.0 25.8/ 26.0/ 22.4 22.8/ 24.1/ 27.8 26.0/ 29.0/ 25.0 21.5/ 20.3/ 20.9 15.0/ 13.0/ 22.0 19.0/ 19.0/ 21.0 30.2/ 26.8/ 26.8 26.8/ 25.7/ 23.2
Food Science 26.0/ 21.0/ 19.0 23.0/ 25.0/ 27.0 34.0/ 28.0/ 27.0 27.3/ 26.3/ 22.2 27.0/ 23.0/ 24.0 35.0/ 25.0/ 30.0 29.0/ 30.0/ 26.0 21.0/ 25.0/ 26.0 24.0/ 32.0/ 20.0 21.0/ 17.0/ 19.0 20.0/ 32.0/ 28.0
Geography 27.4/ 21.8/ 21.8 30.0/ 27.4/ 26.4 33.0/ 26.0/ 26.0 28.3/ 26.3/ 26.1 25.1/ 21.0/ 22.8 24.2/ 22.2/ 27.3 27.2/ 23.7/ 23.7 30.0/ 29.0/ 26.0 28.0/ 28.0/ 28.0 33.1/ 33.7/ 32.0 29.2/ 27.6/ 27.6
Health and Medicine 33.0/ 29.0/ 24.0 41.1/ 33.7/ 30.7 24.0/ 26.0/ 30.0 31.7/ 27.9/ 25.9 37.0/ 26.0/ 22.0 26.0/ 22.0/ 18.0 25.9/ 31.2/ 29.5 22.0/ 26.0/ 27.0 29.0/ 27.0/ 29.0 25.9/ 30.2/ 28.4 26.2/ 30.8/ 31.8
History 26.7/ 25.6/ 28.5 31.7/ 28.3/ 27.9 31.0/ 23.0/ 24.0 26.6/ 27.0/ 24.2 28.6/ 24.2/ 27.1 34.0/ 27.0/ 33.0 23.1/ 23.1/ 24.5 34.0/ 29.0/ 29.0 35.0/ 34.0/ 25.0 26.2/ 24.0/ 24.7 28.4/ 25.1/ 26.6
Information Technology 38.0/ 36.0/ 30.0 26.0/ 26.0/ 26.0 30.0/ 24.0/ 24.0 36.3/ 30.7/ 33.0 25.0/ 24.0/ 29.0 27.0/ 24.0/ 23.0 33.0/ 29.0/ 28.0 24.0/ 24.0/ 26.0 33.0/ 25.0/ 27.0 34.0/ 26.0/ 30.0 28.0/ 33.0/ 39.0
International Relations 28.0/ 33.0/ 25.0 24.2/ 30.0/ 23.3 32.0/ 27.0/ 28.0 31.5/ 26.1/ 22.2 29.0/ 35.0/ 25.0 30.0/ 32.0/ 23.0 33.0/ 29.0/ 27.0 30.0/ 26.0/ 27.0 32.0/ 33.0/ 28.0 33.0/ 29.0/ 22.0 22.0/ 17.0/ 19.0
Language Studies 26.0/ 32.0/ 27.0 30.8/ 27.8/ 26.8 33.0/ 33.0/ 30.0 27.7/ 26.2/ 27.9 27.1/ 25.6/ 33.3 21.0/ 23.0/ 22.0 25.5/ 31.5/ 28.9 25.0/ 24.0/ 22.0 32.0/ 30.0/ 27.0 28.5/ 22.9/ 24.3 29.3/ 29.3/ 26.4
Law and Ethics 24.3/ 18.9/ 22.3 30.4/ 27.8/ 28.4 15.0/ 15.0/ 16.0 27.8/ 24.8/ 25.2 25.5/ 23.6/ 22.3 23.0/ 16.0/ 19.0 32.3/ 31.5/ 30.6 24.0/ 23.0/ 23.0 24.0/ 18.0/ 14.0 25.2/ 31.5/ 33.9 20.6/ 22.5/ 22.5
Literature and Linguistics 27.9/ 23.4/ 25.2 25.9/ 27.9/ 24.2 23.0/ 26.0/ 21.0 31.1/ 30.7/ 28.9 26.7/ 29.2/ 29.2 23.0/ 25.0/ 32.0 30.1/ 31.9/ 29.2 22.0/ 23.0/ 16.0 21.0/ 22.0/ 24.0 28.3/ 24.7/ 25.5 28.9/ 28.9/ 27.5
Logical Reasoning 26.8/ 25.8/ 27.5 24.2/ 24.6/ 24.4 27.9/ 29.9/ 27.9 26.5/ 25.2/ 23.8 25.7/ 28.2/ 28.2 26.0/ 24.0/ 24.0 21.4/ 22.6/ 25.4 23.0/ 15.0/ 19.0 36.0/ 28.0/ 30.0 30.3/ 29.5/ 29.9 25.0/ 25.7/ 25.7
Materials Science 33.0/ 27.0/ 33.0 32.7/ 25.9/ 23.2 23.0/ 26.0/ 28.0 26.2/ 22.7/ 25.3 29.0/ 25.0/ 21.0 26.0/ 27.0/ 27.0 38.0/ 29.0/ 24.0 25.0/ 27.0/ 25.0 26.0/ 33.0/ 21.0 26.0/ 22.0/ 23.0 22.0/ 28.0/ 25.0
Mathematics 25.2/ 24.2/ 26.4 27.8/ 25.2/ 24.6 25.6/ 27.4/ 25.0 26.5/ 24.4/ 27.1 23.4/ 19.8/ 25.2 21.8/ 23.8/ 26.1 22.6/ 23.6/ 22.0 27.4/ 26.0/ 25.8 23.9/ 25.4/ 29.5 22.6/ 22.6/ 23.0 24.6/ 25.0/ 25.0
Media and Communication 28.0/ 26.0/ 32.0 32.0/ 21.0/ 26.0 30.0/ 28.0/ 26.0 30.6/ 23.8/ 20.4 26.0/ 27.0/ 28.0 29.0/ 31.0/ 27.0 34.0/ 30.0/ 32.0 23.0/ 30.0/ 27.0 26.0/ 29.0/ 25.0 23.0/ 23.0/ 21.0 33.0/ 37.0/ 33.0
Music and Performing Arts 27.0/ 20.0/ 20.0 33.0/ 31.0/ 31.0 26.0/ 25.0/ 26.0 24.5/ 22.5/ 24.5 25.0/ 28.0/ 26.0 30.0/ 24.0/ 22.0 29.0/ 33.0/ 26.0 34.0/ 35.0/ 38.0 25.0/ 21.0/ 23.0 29.0/ 19.0/ 24.0 30.0/ 21.0/ 25.0
Physics 26.4/ 25.8/ 25.8 32.2/ 28.2/ 27.8 23.4/ 22.8/ 21.7 26.2/ 25.0/ 24.2 25.4/ 27.6/ 25.2 29.5/ 27.9/ 26.8 29.4/ 28.2/ 25.4 22.9/ 22.9/ 24.0 24.0/ 26.0/ 32.0 22.8/ 23.0/ 29.0 26.8/ 24.2/ 26.8
Politics and Governance 27.8/ 27.1/ 25.9 30.9/ 28.9/ 25.9 24.0/ 24.0/ 23.0 25.9/ 25.9/ 24.9 27.9/ 26.3/ 23.5 22.0/ 19.0/ 27.0 27.5/ 24.2/ 25.8 33.0/ 26.0/ 30.0 28.0/ 30.0/ 30.0 26.6/ 25.3/ 23.2 27.1/ 24.5/ 26.9
Psychology 22.0/ 22.0/ 23.0 28.0/ 22.0/ 23.0 30.0/ 27.0/ 31.0 31.8/ 35.7/ 20.2 26.0/ 26.0/ 30.0 25.0/ 27.0/ 21.0 28.0/ 34.0/ 31.0 25.0/ 22.0/ 25.0 31.0/ 27.0/ 32.0 26.0/ 24.0/ 30.0 28.0/ 23.0/ 33.0
Public Administration 21.0/ 23.0/ 16.0 24.2/ 20.2/ 24.2 26.3/ 24.2/ 27.3 33.3/ 35.4/ 30.3 27.3/ 21.2/ 27.3 24.0/ 23.0/ 22.0 29.3/ 27.3/ 19.2 27.0/ 23.0/ 20.0 29.3/ 24.2/ 22.2 22.0/ 13.0/ 24.0 33.0/ 29.0/ 32.0
Religion and Spirituality 26.0/ 24.0/ 25.0 27.0/ 30.0/ 26.0 34.0/ 23.0/ 27.0 25.4/ 18.9/ 22.6 26.0/ 18.0/ 18.0 30.0/ 25.0/ 26.0 35.0/ 30.0/ 30.0 24.0/ 28.0/ 24.0 28.0/ 31.0/ 32.0 25.0/ 26.0/ 24.0 30.0/ 22.0/ 25.0
Social Welfare and Development 22.0/ 21.0/ 20.0 24.1/ 29.5/ 23.2 32.0/ 30.0/ 29.0 23.8/ 26.9/ 25.9 27.0/ 24.0/ 27.0 23.0/ 19.0/ 22.0 32.0/ 26.0/ 25.0 28.0/ 22.0/ 24.0 24.0/ 23.0/ 14.0 28.0/ 25.0/ 28.0 24.0/ 24.0/ 27.0
Sociology 20.9/ 23.5/ 21.7 22.4/ 23.4/ 24.4 27.0/ 28.0/ 29.0 27.0/ 23.4/ 22.8 20.0/ 30.0/ 26.0 18.0/ 19.0/ 24.0 25.8/ 27.7/ 32.1 24.0/ 21.0/ 16.0 25.0/ 21.0/ 25.0 27.0/ 26.1/ 27.9 21.8/ 22.9/ 24.7
Sports and Recreation 24.3/ 21.8/ 16.8 27.8/ 26.0/ 24.4 18.0/ 17.0/ 18.0 21.8/ 21.2/ 18.8 25.8/ 28.6/ 27.0 24.0/ 28.0/ 27.0 26.6/ 26.6/ 23.7 21.0/ 24.0/ 27.0 25.0/ 22.0/ 23.0 26.3/ 25.6/ 25.6 24.8/ 23.3/ 20.5
Technology and Innovation 30.0/ 26.0/ 24.0 31.2/ 26.2/ 24.4 26.0/ 26.0/ 27.0 27.8/ 24.0/ 26.0 25.0/ 33.0/ 31.0 28.0/ 27.0/ 30.0 33.0/ 27.0/ 29.0 26.0/ 23.0/ 18.0 31.0/ 31.0/ 26.0 27.0/ 29.0/ 25.0 20.0/ 14.0/ 20.0
Transportation and Logistics 20.0/ 28.0/ 25.0 27.7/ 31.5/ 24.6 20.0/ 20.0/ 20.0 24.9/ 25.9/ 24.6 24.2/ 21.2/ 25.2 24.0/ 22.0/ 28.0 22.4/ 26.5/ 27.6 23.2/ 26.3/ 28.3 18.0/ 22.0/ 15.0 26.0/ 20.0/ 26.0 23.0/ 23.0/ 24.0
Table 33: Detailed subject-wise evaluation for meta-llama/Llama-3.2-1B-Instruct on MILU across different languages. The results reported are X / Y / Z where X denote the 0-shot, Y denotes 1-shot and Z denotes 5-shot performance respectively.
Topic bn en gu hi kn ml mr or pa ta te
Agriculture 25.0/ 21.0/ 20.0 42.9/ 48.1/ 49.3 27.3/ 26.3/ 27.3 26.2/ 28.0/ 27.4 16.0/ 20.0/ 25.0 24.2/ 28.3/ 25.2 20.0/ 27.0/ 24.0 25.0/ 32.0/ 27.0 23.0/ 29.0/ 34.0 22.0/ 30.0/ 26.0 28.0/ 24.0/ 28.0
Anthropology 33.0/ 33.0/ 36.0 48.0/ 49.0/ 58.0 31.0/ 34.0/ 36.0 29.3/ 32.3/ 38.4 27.0/ 34.0/ 34.0 30.0/ 38.0/ 39.0 32.0/ 41.0/ 38.0 23.0/ 23.0/ 20.0 23.0/ 28.0/ 27.0 33.0/ 29.0/ 27.0 28.0/ 31.0/ 28.0
Architecture and Design 27.0/ 18.0/ 21.0 37.0/ 47.0/ 46.0 28.0/ 23.0/ 19.0 31.0/ 30.0/ 27.0 33.0/ 31.0/ 24.0 29.0/ 24.0/ 24.0 30.0/ 20.0/ 20.0 26.0/ 22.0/ 22.0 36.0/ 35.0/ 36.0 26.0/ 21.0/ 24.0 24.0/ 25.0/ 22.0
Arts and Culture 20.6/ 25.4/ 21.8 48.4/ 55.8/ 57.8 29.0/ 23.0/ 27.0 28.3/ 28.5/ 29.1 26.4/ 27.0/ 28.2 22.0/ 31.0/ 28.0 25.7/ 26.8/ 23.5 19.0/ 20.0/ 21.0 25.0/ 27.0/ 33.0 32.1/ 32.7/ 33.3 22.1/ 26.6/ 24.6
Astronomy and Astrophysics 30.0/ 23.0/ 27.0 54.0/ 64.0/ 68.0 24.0/ 35.0/ 30.0 35.2/ 39.8/ 37.5 31.0/ 24.0/ 34.0 33.0/ 29.0/ 30.0 35.0/ 41.0/ 37.0 29.0/ 19.0/ 27.0 28.0/ 21.0/ 27.0 27.0/ 32.0/ 29.0 28.0/ 26.0/ 31.0
Biology 28.9/ 28.9/ 31.7 62.3/ 67.3/ 69.9 20.0/ 22.0/ 24.0 30.9/ 37.7/ 36.9 29.1/ 34.8/ 31.6 24.2/ 28.3/ 31.3 28.4/ 30.4/ 32.5 25.0/ 24.0/ 25.0 27.0/ 25.0/ 29.0 26.8/ 29.0/ 29.3 25.5/ 31.8/ 30.2
Business and Management 27.0/ 32.0/ 25.0 41.5/ 48.9/ 47.7 30.0/ 28.0/ 26.0 34.4/ 34.2/ 36.3 25.0/ 27.0/ 24.0 20.0/ 30.0/ 28.0 27.0/ 30.0/ 33.0 22.0/ 27.0/ 28.0 24.0/ 28.0/ 34.0 36.0/ 31.0/ 34.0 32.1/ 30.2/ 23.6
Chemistry 29.9/ 36.0/ 38.2 52.3/ 59.1/ 58.3 25.0/ 26.7/ 26.7 33.3/ 37.7/ 36.3 29.7/ 34.6/ 33.3 32.0/ 33.0/ 32.0 31.7/ 32.7/ 34.0 30.8/ 27.9/ 33.7 26.0/ 36.0/ 36.0 31.4/ 33.8/ 36.2 29.2/ 30.7/ 33.2
Computer Science 28.1/ 30.7/ 33.3 33.0/ 35.8/ 37.6 30.0/ 30.0/ 34.0 29.2/ 31.8/ 32.8 29.0/ 32.5/ 33.1 27.0/ 22.0/ 30.0 28.4/ 38.5/ 34.9 23.0/ 24.0/ 20.0 22.0/ 30.0/ 37.0 29.6/ 33.1/ 37.2 31.8/ 33.2/ 32.7
Defense and Security 33.0/ 26.0/ 27.0 39.0/ 58.0/ 60.0 23.0/ 27.0/ 24.0 24.2/ 25.6/ 24.2 30.0/ 29.0/ 24.0 30.0/ 24.0/ 23.0 29.0/ 35.0/ 35.0 25.0/ 27.0/ 28.0 36.0/ 32.0/ 31.0 37.0/ 31.0/ 30.0 24.0/ 28.0/ 23.0
Earth Sciences 33.0/ 39.0/ 43.0 47.9/ 53.2/ 57.0 31.0/ 37.0/ 32.0 29.3/ 32.1/ 33.5 27.0/ 25.0/ 30.0 23.0/ 29.0/ 33.0 32.0/ 35.0/ 35.0 22.2/ 22.2/ 21.2 32.0/ 29.0/ 28.0 26.0/ 36.0/ 29.0 33.0/ 30.0/ 28.0
Economics 30.9/ 25.9/ 36.0 45.4/ 51.0/ 51.2 24.0/ 29.0/ 31.0 31.6/ 32.9/ 35.3 37.0/ 37.0/ 36.0 36.0/ 30.0/ 28.0 24.6/ 29.6/ 29.6 27.0/ 27.0/ 24.0 30.0/ 30.0/ 30.0 28.6/ 28.6/ 30.2 32.3/ 26.4/ 28.7
Education 23.0/ 26.0/ 25.0 36.9/ 40.5/ 43.0 23.0/ 23.0/ 23.0 29.3/ 28.7/ 29.6 32.0/ 33.0/ 32.0 25.0/ 28.0/ 28.0 25.5/ 31.2/ 31.2 21.0/ 26.0/ 29.0 32.0/ 35.0/ 35.0 31.0/ 29.0/ 28.0 30.0/ 32.0/ 23.0
Energy and Power 34.0/ 30.0/ 28.0 40.5/ 51.3/ 53.4 29.0/ 33.0/ 28.0 33.2/ 39.0/ 40.0 25.0/ 28.0/ 31.0 25.0/ 30.0/ 30.0 27.0/ 38.0/ 40.0 24.0/ 23.0/ 26.0 29.0/ 29.0/ 33.0 28.0/ 32.0/ 38.0 27.0/ 28.0/ 26.0
Engineering 25.2/ 28.4/ 27.0 34.4/ 39.4/ 39.8 26.0/ 25.6/ 27.0 27.0/ 31.6/ 32.0 26.6/ 27.2/ 29.0 24.8/ 24.8/ 26.9 32.6/ 28.0/ 29.4 20.5/ 22.6/ 23.2 32.0/ 30.0/ 29.0 29.8/ 29.0/ 30.8 24.4/ 27.2/ 27.6
Environmental Science 22.0/ 23.0/ 25.0 50.3/ 56.3/ 55.7 30.0/ 28.0/ 29.0 29.9/ 35.7/ 37.5 21.9/ 24.2/ 26.6 30.0/ 38.0/ 35.0 27.4/ 33.7/ 32.6 29.3/ 25.2/ 31.3 27.0/ 30.0/ 42.0 30.0/ 31.0/ 34.0 29.0/ 25.7/ 28.5
Ethics and Human Rights 25.0/ 20.0/ 33.0 32.0/ 51.0/ 52.0 34.0/ 29.0/ 35.0 28.0/ 32.0/ 34.0 32.0/ 36.0/ 36.0 28.0/ 28.0/ 32.0 28.0/ 28.0/ 26.0 15.0/ 14.0/ 22.0 31.0/ 33.0/ 27.0 24.0/ 29.0/ 32.0 30.0/ 28.0/ 34.0
Finance and Investment 26.2/ 28.7/ 35.1 41.4/ 43.6/ 46.4 32.0/ 32.0/ 34.0 30.8/ 30.4/ 30.2 28.4/ 32.7/ 31.5 26.0/ 26.0/ 32.0 24.3/ 27.1/ 33.3 20.0/ 23.0/ 22.0 31.0/ 25.0/ 28.0 33.5/ 34.6/ 33.5 22.5/ 24.6/ 25.4
Food Science 19.0/ 25.0/ 27.0 46.0/ 54.0/ 56.0 27.0/ 30.0/ 28.0 31.3/ 28.3/ 36.4 29.0/ 23.0/ 30.0 20.0/ 26.0/ 27.0 21.0/ 29.0/ 23.0 30.0/ 28.0/ 32.0 20.0/ 19.0/ 22.0 25.0/ 20.0/ 26.0 23.0/ 27.0/ 29.0
Geography 22.9/ 29.6/ 26.3 42.0/ 48.4/ 51.2 30.0/ 35.0/ 36.0 31.1/ 28.7/ 29.9 26.5/ 27.9/ 28.8 22.2/ 29.3/ 25.2 31.4/ 30.8/ 30.8 24.0/ 23.0/ 28.0 24.0/ 22.0/ 23.0 29.1/ 27.4/ 28.6 29.2/ 25.5/ 24.0
Health and Medicine 28.0/ 28.0/ 31.0 59.1/ 66.7/ 68.3 23.0/ 23.0/ 21.0 31.9/ 34.1/ 38.3 21.0/ 24.0/ 29.0 22.0/ 22.0/ 28.0 21.4/ 41.1/ 40.2 22.0/ 25.0/ 29.0 30.0/ 38.0/ 31.0 25.9/ 27.6/ 29.3 28.0/ 31.8/ 35.5
History 24.8/ 25.6/ 28.0 40.6/ 44.2/ 45.2 22.0/ 29.0/ 28.0 29.4/ 30.4/ 32.7 28.2/ 27.8/ 29.3 24.0/ 34.0/ 31.0 26.8/ 27.1/ 29.0 26.0/ 28.0/ 29.0 31.0/ 35.0/ 30.0 26.9/ 29.8/ 28.0 29.2/ 30.2/ 29.7
Information Technology 41.0/ 44.0/ 42.0 53.0/ 62.0/ 63.0 46.0/ 35.0/ 36.0 35.8/ 39.1/ 41.3 38.0/ 36.0/ 39.0 24.0/ 35.0/ 36.0 31.0/ 41.0/ 35.0 28.0/ 30.0/ 26.0 33.0/ 36.0/ 35.0 35.0/ 43.0/ 41.0 40.0/ 39.0/ 40.0
International Relations 28.0/ 23.0/ 25.0 58.3/ 60.0/ 59.2 34.0/ 43.0/ 33.0 31.9/ 33.9/ 37.4 21.0/ 27.0/ 31.0 35.0/ 35.0/ 32.0 33.0/ 34.0/ 34.0 30.0/ 24.0/ 16.0 27.0/ 31.0/ 36.0 26.0/ 33.0/ 34.0 27.0/ 26.0/ 26.0
Language Studies 24.0/ 29.0/ 33.0 45.0/ 52.8/ 54.4 35.0/ 25.0/ 24.0 23.2/ 29.5/ 29.3 22.5/ 24.0/ 27.1 22.0/ 26.0/ 30.0 24.8/ 29.5/ 32.2 22.0/ 22.0/ 18.0 33.0/ 31.0/ 29.0 31.5/ 24.7/ 25.1 24.1/ 23.6/ 25.3
Law and Ethics 31.8/ 35.8/ 40.5 41.4/ 52.8/ 53.0 33.0/ 31.0/ 28.0 31.0/ 33.0/ 35.4 21.0/ 22.3/ 26.8 31.0/ 28.0/ 34.0 25.0/ 33.6/ 33.6 25.0/ 29.0/ 27.0 24.0/ 28.0/ 21.0 33.1/ 29.9/ 33.1 30.0/ 24.4/ 25.0
Literature and Linguistics 26.1/ 27.0/ 25.2 39.9/ 40.3/ 42.9 22.0/ 28.0/ 26.0 24.1/ 26.7/ 27.7 30.8/ 26.7/ 25.8 25.0/ 21.0/ 22.0 26.6/ 23.9/ 22.1 21.0/ 23.0/ 23.0 24.0/ 33.0/ 30.0 30.4/ 32.8/ 30.4 26.1/ 26.1/ 30.3
Logical Reasoning 29.5/ 28.0/ 27.0 28.6/ 32.4/ 30.8 27.3/ 26.6/ 26.6 26.7/ 28.7/ 27.1 29.9/ 27.0/ 28.2 23.0/ 24.0/ 25.0 26.7/ 29.0/ 27.3 30.0/ 21.0/ 17.0 17.0/ 24.0/ 25.0 25.9/ 26.7/ 26.7 27.5/ 29.9/ 31.4
Materials Science 27.0/ 31.0/ 27.0 50.2/ 53.2/ 56.3 26.0/ 25.0/ 25.0 27.3/ 31.2/ 30.8 22.0/ 23.0/ 28.0 24.0/ 31.0/ 33.0 33.0/ 33.0/ 33.0 29.0/ 24.0/ 23.0 21.0/ 30.0/ 33.0 32.0/ 26.0/ 33.0 24.0/ 25.0/ 30.0
Mathematics 29.0/ 28.0/ 30.4 30.8/ 32.8/ 31.8 29.0/ 27.4/ 29.6 29.7/ 29.9/ 31.9 27.8/ 24.6/ 28.4 28.6/ 30.7/ 33.2 25.8/ 28.0/ 30.8 31.6/ 29.0/ 28.8 27.7/ 23.1/ 28.0 25.4/ 29.8/ 29.6 27.8/ 27.6/ 26.4
Media and Communication 30.0/ 26.0/ 30.0 51.0/ 61.0/ 60.0 32.0/ 38.0/ 42.0 29.9/ 36.0/ 29.2 30.0/ 32.0/ 30.0 41.0/ 31.0/ 33.0 30.0/ 41.0/ 36.0 32.0/ 33.0/ 33.0 28.0/ 26.0/ 33.0 26.0/ 29.0/ 27.0 27.0/ 31.0/ 33.0
Music and Performing Arts 22.0/ 25.0/ 24.0 43.0/ 44.0/ 45.0 31.0/ 25.0/ 27.0 27.8/ 27.8/ 23.2 28.0/ 25.0/ 27.0 26.0/ 27.0/ 22.0 30.0/ 27.0/ 31.0 27.0/ 25.0/ 24.0 23.0/ 25.0/ 27.0 24.0/ 29.0/ 26.0 31.0/ 28.0/ 31.0
Physics 30.8/ 33.0/ 32.6 50.8/ 51.4/ 54.2 29.8/ 33.4/ 30.6 29.8/ 31.2/ 32.2 26.0/ 32.0/ 31.0 25.3/ 23.7/ 28.4 30.8/ 30.6/ 32.6 23.6/ 24.3/ 25.7 34.0/ 38.0/ 35.0 28.2/ 26.0/ 29.8 28.0/ 29.0/ 30.4
Politics and Governance 27.1/ 26.3/ 29.4 42.7/ 48.1/ 46.9 27.0/ 28.0/ 29.0 27.7/ 32.5/ 31.9 23.1/ 22.3/ 22.7 23.0/ 25.0/ 24.0 24.4/ 27.5/ 29.5 22.0/ 35.0/ 30.0 20.0/ 26.0/ 25.0 25.3/ 29.5/ 29.9 27.8/ 26.2/ 25.7
Psychology 24.0/ 27.0/ 23.0 53.0/ 60.0/ 58.0 33.0/ 24.0/ 34.0 29.5/ 37.2/ 38.0 25.0/ 27.0/ 32.0 25.0/ 29.0/ 28.0 27.0/ 26.0/ 27.0 34.0/ 40.0/ 34.0 26.0/ 35.0/ 37.0 21.0/ 28.0/ 26.0 29.0/ 35.0/ 41.0
Public Administration 25.0/ 26.0/ 26.0 36.4/ 39.4/ 40.4 34.3/ 33.3/ 34.3 23.2/ 29.3/ 28.3 24.2/ 26.3/ 28.3 30.0/ 27.0/ 30.0 22.2/ 32.3/ 26.3 24.0/ 24.0/ 25.0 33.3/ 28.3/ 32.3 29.0/ 27.0/ 32.0 36.0/ 26.0/ 24.0
Religion and Spirituality 29.0/ 27.0/ 26.0 47.0/ 55.0/ 56.0 28.0/ 30.0/ 25.0 24.6/ 25.4/ 25.8 26.0/ 29.0/ 29.0 30.0/ 24.0/ 28.0 31.0/ 30.0/ 34.0 27.0/ 24.0/ 24.0 32.0/ 33.0/ 24.0 27.0/ 20.0/ 18.0 28.0/ 28.0/ 28.0
Social Welfare and Development 30.0/ 30.0/ 38.0 42.9/ 46.4/ 45.5 32.0/ 33.0/ 36.0 31.1/ 36.8/ 39.4 30.0/ 28.0/ 33.0 22.0/ 26.0/ 28.0 24.0/ 33.0/ 30.0 24.0/ 25.0/ 26.0 32.0/ 26.0/ 37.0 22.0/ 28.0/ 28.0 29.0/ 26.0/ 25.0
Sociology 26.1/ 24.3/ 20.9 36.2/ 38.4/ 36.6 17.0/ 22.0/ 29.0 24.4/ 28.6/ 28.2 26.0/ 30.0/ 29.0 19.0/ 22.0/ 23.0 26.4/ 22.6/ 27.0 27.0/ 25.0/ 28.0 22.0/ 25.0/ 26.0 21.6/ 30.6/ 23.4 20.0/ 25.9/ 24.1
Sports and Recreation 22.3/ 27.7/ 25.2 53.0/ 52.8/ 53.6 32.0/ 29.0/ 25.0 30.4/ 28.4/ 29.4 25.8/ 25.3/ 29.2 27.0/ 25.0/ 25.0 22.6/ 28.8/ 29.9 27.0/ 28.0/ 29.0 19.0/ 32.0/ 30.0 27.6/ 28.2/ 27.6 24.8/ 25.2/ 24.8
Technology and Innovation 28.0/ 31.0/ 34.0 47.4/ 52.6/ 54.2 28.0/ 27.0/ 27.0 30.4/ 32.4/ 34.2 33.0/ 34.0/ 41.0 22.0/ 29.0/ 26.0 29.0/ 35.0/ 37.0 20.0/ 19.0/ 19.0 28.0/ 30.0/ 32.0 29.0/ 32.0/ 35.0 29.0/ 26.0/ 30.0
Transportation and Logistics 31.0/ 29.0/ 29.0 36.9/ 40.8/ 41.5 23.0/ 26.0/ 22.0 23.3/ 26.8/ 27.1 24.2/ 23.2/ 27.3 30.0/ 23.0/ 26.0 25.5/ 21.4/ 25.5 22.2/ 24.2/ 24.2 27.0/ 30.0/ 27.0 32.0/ 33.0/ 30.0 33.0/ 27.0/ 27.0
Table 34: Detailed subject-wise evaluation for google/gemma-2-2b on MILU across different languages. The results reported are X / Y / Z where X denote the 0-shot, Y denotes 1-shot and Z denotes 5-shot performance respectively.
Topic bn en gu hi kn ml mr or pa ta te
Agriculture 34.0/ 31.0/ 32.0 41.6/ 47.9/ 51.1 29.3/ 35.4/ 34.3 27.4/ 32.2/ 33.6 17.0/ 19.0/ 16.0 19.2/ 19.2/ 22.2 25.0/ 24.0/ 26.0 26.0/ 27.0/ 30.0 25.0/ 28.0/ 26.0 27.0/ 28.0/ 29.0 28.0/ 31.0/ 28.0
Anthropology 28.0/ 28.0/ 29.0 47.0/ 55.0/ 53.0 31.0/ 30.0/ 32.0 44.4/ 39.4/ 42.4 27.0/ 26.0/ 28.0 33.0/ 29.0/ 40.0 41.0/ 39.0/ 38.0 27.0/ 26.0/ 29.0 24.0/ 29.0/ 27.0 33.0/ 27.0/ 28.0 30.0/ 29.0/ 29.0
Architecture and Design 23.0/ 27.0/ 27.0 40.0/ 51.0/ 48.0 32.0/ 31.0/ 28.0 37.0/ 30.0/ 34.0 19.0/ 24.0/ 27.0 31.0/ 24.0/ 21.0 40.0/ 35.0/ 36.0 20.0/ 25.0/ 20.0 26.0/ 28.0/ 31.0 24.0/ 27.0/ 23.0 23.0/ 27.0/ 25.0
Arts and Culture 26.1/ 27.3/ 29.1 44.6/ 54.4/ 57.0 17.0/ 21.0/ 20.0 29.9/ 32.9/ 32.5 21.5/ 22.7/ 19.6 30.0/ 33.0/ 29.0 30.0/ 29.5/ 26.8 23.0/ 20.0/ 21.0 31.0/ 28.0/ 32.0 24.9/ 24.2/ 30.9 21.6/ 20.6/ 19.6
Astronomy and Astrophysics 31.0/ 28.0/ 32.0 66.0/ 71.0/ 76.0 23.0/ 25.0/ 26.0 35.2/ 38.3/ 42.2 29.0/ 27.0/ 28.0 30.0/ 33.0/ 33.0 37.0/ 37.0/ 38.0 25.0/ 20.0/ 22.0 24.0/ 23.0/ 28.0 33.0/ 32.0/ 31.0 32.0/ 29.0/ 31.0
Biology 31.0/ 29.9/ 26.8 60.1/ 71.3/ 71.5 25.0/ 22.0/ 21.0 34.9/ 40.5/ 40.5 25.8/ 26.2/ 27.9 29.3/ 34.3/ 34.3 31.0/ 31.3/ 32.5 26.0/ 27.0/ 31.0 36.0/ 30.0/ 33.0 32.8/ 31.5/ 31.2 30.5/ 33.0/ 30.5
Business and Management 32.0/ 30.0/ 32.0 41.5/ 52.0/ 48.6 15.0/ 21.0/ 27.0 35.6/ 38.9/ 38.4 27.0/ 33.0/ 29.0 33.0/ 32.0/ 34.0 24.0/ 31.0/ 29.0 36.0/ 33.0/ 30.0 38.0/ 29.0/ 35.0 30.0/ 29.0/ 39.0 27.4/ 28.3/ 30.2
Chemistry 35.2/ 33.2/ 36.3 52.5/ 58.5/ 60.3 28.4/ 31.9/ 33.6 38.5/ 39.3/ 39.3 30.1/ 30.7/ 29.1 32.0/ 30.0/ 32.0 33.5/ 35.6/ 36.7 33.7/ 32.7/ 29.8 30.0/ 30.0/ 29.0 36.5/ 33.5/ 37.7 29.4/ 31.4/ 35.7
Computer Science 33.3/ 36.5/ 33.3 41.0/ 40.4/ 39.0 29.0/ 29.0/ 31.0 29.2/ 36.4/ 35.2 31.9/ 36.7/ 34.9 35.0/ 41.0/ 37.0 35.5/ 32.5/ 34.9 31.0/ 30.0/ 29.0 35.0/ 36.0/ 37.0 34.9/ 33.1/ 37.2 34.1/ 39.0/ 37.2
Defense and Security 29.0/ 30.0/ 27.0 47.0/ 60.0/ 58.0 18.0/ 26.0/ 22.0 26.1/ 24.6/ 26.5 21.0/ 24.0/ 23.0 23.0/ 22.0/ 19.0 24.0/ 34.0/ 32.0 26.0/ 29.0/ 30.0 31.0/ 27.0/ 32.0 27.0/ 31.0/ 34.0 30.0/ 37.0/ 33.0
Earth Sciences 32.0/ 28.0/ 32.0 48.9/ 57.7/ 58.0 39.0/ 41.0/ 39.0 31.5/ 34.1/ 36.3 27.0/ 35.0/ 33.0 34.0/ 37.0/ 35.0 30.0/ 33.0/ 32.0 24.2/ 17.2/ 22.2 32.0/ 28.0/ 31.0 30.0/ 31.0/ 27.0 31.0/ 30.0/ 31.0
Economics 30.2/ 30.2/ 30.9 45.8/ 51.6/ 53.4 24.0/ 25.0/ 27.0 33.5/ 39.1/ 35.5 32.0/ 34.0/ 38.0 29.0/ 33.0/ 31.0 31.0/ 34.5/ 33.1 30.0/ 25.0/ 28.0 31.0/ 29.0/ 27.0 29.4/ 30.9/ 33.3 26.4/ 29.9/ 26.4
Education 31.0/ 26.0/ 27.0 36.3/ 39.8/ 42.4 18.0/ 17.0/ 25.0 31.7/ 30.8/ 33.7 27.0/ 29.0/ 33.0 28.0/ 37.0/ 30.0 32.5/ 39.5/ 36.3 29.0/ 31.0/ 29.0 26.0/ 23.0/ 23.0 23.0/ 28.0/ 25.0 28.0/ 27.0/ 29.0
Energy and Power 31.0/ 28.0/ 25.0 47.3/ 52.0/ 51.3 26.0/ 28.0/ 28.0 31.2/ 36.6/ 35.2 24.0/ 26.0/ 31.0 26.0/ 25.0/ 23.0 27.0/ 31.0/ 32.0 19.0/ 18.0/ 29.0 24.0/ 31.0/ 24.0 26.0/ 26.0/ 25.0 30.0/ 29.0/ 29.0
Engineering 25.2/ 25.4/ 28.8 36.4/ 37.4/ 43.4 28.8/ 27.8/ 29.4 28.8/ 29.6/ 29.0 27.2/ 26.2/ 27.6 26.5/ 27.8/ 26.5 27.8/ 29.0/ 33.0 22.3/ 25.3/ 27.7 29.0/ 22.0/ 25.0 28.8/ 29.0/ 30.6 25.4/ 28.4/ 28.4
Environmental Science 31.0/ 34.0/ 35.0 47.7/ 57.1/ 57.9 32.0/ 33.0/ 32.0 35.1/ 38.5/ 39.5 26.6/ 28.1/ 27.3 30.0/ 29.0/ 29.0 37.9/ 33.2/ 32.1 32.3/ 30.3/ 29.3 30.0/ 32.0/ 34.0 31.0/ 26.0/ 23.0 29.6/ 29.0/ 29.6
Ethics and Human Rights 31.0/ 29.0/ 29.0 43.0/ 47.0/ 49.0 35.0/ 31.0/ 32.0 34.0/ 33.0/ 31.0 29.0/ 33.0/ 27.0 32.0/ 34.0/ 35.0 24.0/ 25.0/ 27.0 21.0/ 21.0/ 23.0 26.0/ 26.0/ 26.0 28.0/ 34.0/ 33.0 38.0/ 36.0/ 32.0
Finance and Investment 27.7/ 30.7/ 33.2 42.4/ 47.6/ 47.6 29.0/ 31.0/ 26.0 31.0/ 31.2/ 35.0 25.3/ 27.8/ 30.2 17.0/ 24.0/ 19.0 24.3/ 29.9/ 30.5 12.0/ 19.0/ 18.0 27.0/ 27.0/ 22.0 25.1/ 30.7/ 33.0 21.1/ 24.3/ 27.9
Food Science 21.0/ 19.0/ 24.0 54.0/ 63.0/ 62.0 31.0/ 27.0/ 26.0 34.3/ 34.3/ 38.4 26.0/ 27.0/ 24.0 22.0/ 29.0/ 31.0 39.0/ 33.0/ 40.0 24.0/ 29.0/ 30.0 23.0/ 25.0/ 24.0 31.0/ 27.0/ 26.0 31.0/ 32.0/ 29.0
Geography 27.9/ 24.6/ 24.6 44.0/ 49.6/ 50.4 31.0/ 33.0/ 32.0 34.7/ 32.1/ 29.7 24.7/ 25.6/ 26.0 22.2/ 25.2/ 28.3 37.3/ 36.1/ 35.5 28.0/ 30.0/ 30.0 25.0/ 20.0/ 21.0 33.1/ 30.9/ 31.4 22.9/ 26.0/ 29.7
Health and Medicine 32.0/ 33.0/ 36.0 57.9/ 67.5/ 68.9 25.0/ 22.0/ 22.0 35.3/ 41.1/ 39.3 36.0/ 33.0/ 30.0 32.0/ 26.0/ 25.0 28.6/ 29.5/ 28.6 23.0/ 27.0/ 25.0 24.0/ 24.0/ 28.0 30.2/ 26.7/ 31.9 34.6/ 36.4/ 31.8
History 25.1/ 25.3/ 25.1 38.1/ 43.8/ 44.0 28.0/ 33.0/ 33.0 29.2/ 33.9/ 33.9 24.9/ 25.6/ 27.8 21.0/ 33.0/ 32.0 27.5/ 24.2/ 26.4 27.0/ 23.0/ 26.0 25.0/ 28.0/ 29.0 28.4/ 30.2/ 29.1 28.4/ 28.4/ 27.6
Information Technology 35.0/ 32.0/ 39.0 56.0/ 65.0/ 64.0 40.0/ 43.0/ 43.0 36.9/ 44.1/ 48.0 31.0/ 36.0/ 41.0 32.0/ 37.0/ 30.0 36.0/ 43.0/ 40.0 31.0/ 31.0/ 24.0 33.0/ 36.0/ 38.0 39.0/ 38.0/ 42.0 43.0/ 46.0/ 43.0
International Relations 28.0/ 22.0/ 28.0 51.7/ 62.5/ 65.0 28.0/ 34.0/ 34.0 32.7/ 36.6/ 39.3 33.0/ 26.0/ 28.0 25.0/ 34.0/ 38.0 41.0/ 41.0/ 41.0 35.0/ 32.0/ 26.0 20.0/ 33.0/ 26.0 27.0/ 29.0/ 29.0 27.0/ 25.0/ 25.0
Language Studies 29.0/ 31.0/ 36.0 51.4/ 59.0/ 62.0 37.0/ 38.0/ 33.0 27.5/ 31.5/ 30.3 22.5/ 28.7/ 27.1 27.0/ 31.0/ 25.0 33.6/ 36.9/ 31.5 27.0/ 27.0/ 25.0 37.0/ 34.0/ 32.0 34.1/ 28.8/ 25.5 27.0/ 31.0/ 27.6
Law and Ethics 34.5/ 37.2/ 35.1 43.8/ 48.0/ 51.6 42.0/ 29.0/ 33.0 33.4/ 34.6/ 35.2 29.9/ 32.5/ 23.6 23.0/ 32.0/ 31.0 23.7/ 24.6/ 25.0 18.0/ 16.0/ 22.0 24.0/ 24.0/ 29.0 27.6/ 25.2/ 25.2 22.5/ 24.4/ 21.9
Literature and Linguistics 27.0/ 27.0/ 26.1 41.5/ 50.3/ 48.1 29.0/ 29.0/ 30.0 25.1/ 29.9/ 29.5 28.3/ 20.0/ 25.8 19.0/ 16.0/ 19.0 30.1/ 25.7/ 29.2 33.0/ 25.0/ 25.0 28.0/ 32.0/ 34.0 23.5/ 25.1/ 28.7 26.1/ 26.1/ 26.1
Logical Reasoning 29.7/ 30.2/ 31.2 29.4/ 31.6/ 33.0 21.4/ 24.7/ 25.3 28.9/ 33.1/ 31.1 26.1/ 25.3/ 23.6 28.0/ 31.0/ 32.0 27.6/ 29.2/ 28.4 23.0/ 26.0/ 25.0 24.0/ 33.0/ 29.0 26.3/ 30.7/ 28.7 24.3/ 28.2/ 28.0
Materials Science 29.0/ 32.0/ 33.0 41.4/ 50.2/ 53.2 30.0/ 30.0/ 36.0 28.0/ 29.9/ 32.5 34.0/ 26.0/ 32.0 32.0/ 41.0/ 34.0 18.0/ 32.0/ 33.0 21.0/ 23.0/ 23.0 23.0/ 27.0/ 27.0 28.0/ 29.0/ 34.0 28.0/ 29.0/ 30.0
Mathematics 30.0/ 29.6/ 30.4 35.2/ 35.4/ 37.2 26.4/ 29.0/ 30.2 32.3/ 34.1/ 32.1 28.8/ 31.2/ 31.0 29.5/ 32.1/ 33.5 30.0/ 31.4/ 30.0 27.8/ 29.2/ 31.2 24.6/ 31.4/ 28.8 30.8/ 29.4/ 33.2 30.6/ 31.2/ 30.0
Media and Communication 36.0/ 32.0/ 35.0 52.0/ 65.0/ 62.0 30.0/ 31.0/ 32.0 25.2/ 32.0/ 34.7 35.0/ 40.0/ 39.0 37.0/ 36.0/ 38.0 23.0/ 29.0/ 33.0 28.0/ 33.0/ 35.0 28.0/ 28.0/ 23.0 26.0/ 30.0/ 29.0 31.0/ 39.0/ 39.0
Music and Performing Arts 25.0/ 23.0/ 31.0 41.0/ 35.0/ 43.0 18.0/ 30.0/ 25.0 25.8/ 22.5/ 23.8 22.0/ 26.0/ 27.0 22.0/ 24.0/ 23.0 28.0/ 27.0/ 24.0 26.0/ 26.0/ 26.0 27.0/ 25.0/ 23.0 29.0/ 31.0/ 32.0 31.0/ 22.0/ 24.0
Physics 29.2/ 32.6/ 36.6 49.6/ 53.0/ 57.4 28.1/ 31.2/ 33.1 35.0/ 37.6/ 36.4 26.8/ 30.8/ 30.6 24.7/ 26.8/ 26.8 30.8/ 33.0/ 35.8 22.9/ 25.7/ 24.6 30.0/ 32.0/ 34.0 23.6/ 29.4/ 31.8 28.2/ 31.0/ 33.0
Politics and Governance 29.4/ 27.1/ 27.1 40.5/ 48.9/ 47.3 32.0/ 24.0/ 25.0 28.5/ 32.5/ 32.1 31.2/ 26.7/ 29.1 30.0/ 28.0/ 34.0 26.4/ 28.6/ 28.6 26.0/ 18.0/ 23.0 27.0/ 26.0/ 25.0 26.1/ 22.8/ 26.1 27.6/ 27.8/ 26.7
Psychology 31.0/ 30.0/ 25.0 48.0/ 63.0/ 68.0 29.0/ 26.0/ 25.0 34.1/ 41.1/ 41.1 26.0/ 28.0/ 31.0 33.0/ 40.0/ 36.0 34.0/ 35.0/ 32.0 33.0/ 37.0/ 42.0 31.0/ 30.0/ 30.0 32.0/ 27.0/ 31.0 32.0/ 38.0/ 38.0
Public Administration 33.0/ 29.0/ 32.0 42.4/ 48.5/ 45.5 33.3/ 30.3/ 33.3 31.3/ 37.4/ 31.3 23.2/ 24.2/ 26.3 32.0/ 26.0/ 20.0 23.2/ 25.2/ 27.3 15.0/ 20.0/ 15.0 29.3/ 33.3/ 34.3 35.0/ 26.0/ 28.0 29.0/ 29.0/ 32.0
Religion and Spirituality 26.0/ 27.0/ 29.0 52.0/ 56.0/ 59.0 24.0/ 33.0/ 29.0 29.8/ 29.0/ 28.6 21.0/ 25.0/ 24.0 20.0/ 32.0/ 27.0 26.0/ 33.0/ 34.0 26.0/ 27.0/ 26.0 35.0/ 32.0/ 28.0 28.0/ 28.0/ 25.0 30.0/ 32.0/ 29.0
Social Welfare and Development 31.0/ 26.0/ 31.0 34.8/ 48.2/ 46.4 30.0/ 26.0/ 24.0 32.1/ 34.7/ 37.3 30.0/ 30.0/ 31.0 16.0/ 23.0/ 21.0 35.0/ 44.0/ 44.0 30.0/ 26.0/ 27.0 22.0/ 28.0/ 34.0 16.0/ 19.0/ 21.0 31.0/ 34.0/ 32.0
Sociology 20.0/ 21.7/ 24.3 36.0/ 36.0/ 38.6 23.0/ 27.0/ 27.0 24.8/ 28.0/ 27.6 24.0/ 29.0/ 27.0 22.0/ 27.0/ 24.0 30.2/ 29.6/ 28.9 20.0/ 23.0/ 24.0 31.0/ 33.0/ 31.0 25.2/ 25.2/ 24.3 20.6/ 20.6/ 18.8
Sports and Recreation 29.2/ 28.2/ 29.2 48.6/ 55.4/ 55.0 29.0/ 27.0/ 30.0 28.8/ 28.4/ 30.2 33.1/ 31.5/ 31.5 32.0/ 26.0/ 24.0 35.0/ 31.6/ 28.8 26.0/ 29.0/ 23.0 28.0/ 28.0/ 30.0 28.2/ 28.2/ 29.5 24.8/ 26.2/ 26.2
Technology and Innovation 31.0/ 34.0/ 34.0 47.2/ 53.8/ 54.8 26.0/ 26.0/ 22.0 33.0/ 38.8/ 38.8 34.0/ 37.0/ 39.0 38.0/ 33.0/ 33.0 31.0/ 38.0/ 41.0 24.0/ 19.0/ 30.0 30.0/ 27.0/ 33.0 31.0/ 30.0/ 34.0 28.0/ 28.0/ 30.0
Transportation and Logistics 28.0/ 30.0/ 32.0 45.4/ 47.7/ 46.2 25.0/ 26.0/ 24.0 28.1/ 28.7/ 27.8 21.2/ 25.2/ 22.2 24.0/ 29.0/ 24.0 28.6/ 25.5/ 28.6 29.3/ 29.3/ 24.2 26.0/ 27.0/ 25.0 27.0/ 25.0/ 28.0 25.0/ 24.0/ 23.0
Table 35: Detailed subject-wise evaluation for google/gemma-2-2b-it on MILU across different languages. The results reported are X / Y / Z where X denote the 0-shot, Y denotes 1-shot and Z denotes 5-shot performance respectively.
Topic bn en gu hi kn ml mr or pa ta te
Agriculture 23.0/ 23.0/ 28.0 27.2/ 33.6/ 33.2 31.3/ 27.3/ 24.2 28.4/ 28.2/ 32.8 19.0/ 26.0/ 29.0 25.2/ 28.3/ 32.3 27.0/ 22.0/ 27.0 31.0/ 26.0/ 25.0 22.0/ 27.0/ 26.0 31.0/ 23.0/ 29.0 32.0/ 24.0/ 27.0
Anthropology 25.0/ 19.0/ 27.0 32.0/ 29.0/ 34.0 27.0/ 31.0/ 35.0 25.2/ 30.3/ 34.3 28.0/ 20.0/ 29.0 24.0/ 13.0/ 24.0 22.0/ 22.0/ 24.0 29.0/ 33.0/ 30.0 21.0/ 30.0/ 27.0 22.0/ 19.0/ 25.0 25.0/ 30.0/ 28.0
Architecture and Design 24.0/ 20.0/ 26.0 31.0/ 34.0/ 30.0 26.0/ 30.0/ 26.0 31.0/ 27.0/ 30.0 30.0/ 29.0/ 26.0 24.0/ 18.0/ 19.0 27.0/ 28.0/ 25.0 29.0/ 27.0/ 22.0 29.0/ 33.0/ 30.0 21.0/ 25.0/ 31.0 22.0/ 28.0/ 31.0
Arts and Culture 26.7/ 24.9/ 29.1 29.2/ 36.8/ 37.0 30.0/ 26.0/ 30.0 22.7/ 24.6/ 27.5 21.5/ 28.8/ 28.2 22.0/ 23.0/ 24.0 25.1/ 25.7/ 25.1 21.0/ 39.0/ 32.0 28.0/ 31.0/ 30.0 29.1/ 25.4/ 26.1 19.6/ 22.6/ 24.1
Astronomy and Astrophysics 26.0/ 33.0/ 34.0 27.0/ 39.0/ 48.0 33.0/ 32.0/ 39.0 35.2/ 31.2/ 43.8 37.0/ 32.0/ 35.0 32.0/ 21.0/ 25.0 30.0/ 27.0/ 32.0 25.0/ 32.0/ 29.0 26.0/ 35.0/ 31.0 29.0/ 24.0/ 29.0 27.0/ 40.0/ 34.0
Biology 22.9/ 26.8/ 30.3 33.1/ 40.1/ 47.3 27.0/ 24.0/ 25.0 28.7/ 27.1/ 34.3 24.2/ 26.2/ 32.0 31.3/ 28.3/ 31.3 23.6/ 26.0/ 30.4 25.0/ 25.0/ 23.0 32.0/ 35.0/ 32.0 29.9/ 26.1/ 28.7 28.9/ 26.7/ 32.1
Business and Management 32.0/ 23.0/ 22.0 24.5/ 26.0/ 29.7 22.0/ 18.0/ 17.0 25.8/ 29.0/ 31.6 27.0/ 26.0/ 28.0 28.0/ 25.0/ 21.0 21.0/ 24.0/ 24.0 26.0/ 22.0/ 22.0 27.0/ 26.0/ 26.0 25.0/ 30.0/ 31.0 19.8/ 20.8/ 19.8
Chemistry 29.7/ 32.7/ 34.9 30.3/ 32.9/ 36.3 20.7/ 37.1/ 38.8 29.3/ 29.1/ 37.7 27.1/ 32.7/ 31.4 29.0/ 39.0/ 38.0 30.6/ 31.4/ 34.3 32.7/ 24.0/ 27.9 34.0/ 32.0/ 34.0 26.4/ 28.8/ 31.4 29.9/ 32.4/ 34.4
Computer Science 28.1/ 25.0/ 26.6 28.0/ 31.8/ 35.6 32.0/ 29.0/ 23.0 27.2/ 31.2/ 31.8 33.1/ 29.0/ 30.8 26.0/ 36.0/ 33.0 32.5/ 32.5/ 27.8 30.0/ 27.0/ 29.0 23.0/ 31.0/ 33.0 33.1/ 31.4/ 27.3 28.7/ 30.5/ 28.2
Defense and Security 23.0/ 21.0/ 22.0 35.0/ 31.0/ 31.0 24.0/ 23.0/ 23.0 21.8/ 22.3/ 28.0 27.0/ 28.0/ 30.0 30.0/ 31.0/ 31.0 29.0/ 23.0/ 26.0 25.0/ 29.0/ 26.0 23.0/ 24.0/ 25.0 23.0/ 24.0/ 22.0 20.0/ 23.0/ 25.0
Earth Sciences 34.0/ 26.0/ 41.0 26.8/ 37.5/ 42.5 22.0/ 27.0/ 32.0 23.4/ 30.5/ 37.1 30.0/ 21.0/ 31.0 21.0/ 26.0/ 29.0 24.0/ 32.0/ 39.0 27.3/ 29.3/ 30.3 30.0/ 31.0/ 30.0 28.0/ 28.0/ 27.0 29.0/ 31.0/ 37.0
Economics 36.7/ 38.1/ 41.0 31.3/ 35.5/ 38.4 33.0/ 36.0/ 33.0 27.8/ 30.2/ 30.6 25.0/ 28.0/ 33.0 23.0/ 29.0/ 30.0 26.8/ 26.1/ 30.3 31.0/ 32.0/ 29.0 32.0/ 32.0/ 34.0 19.1/ 28.6/ 30.9 28.1/ 28.1/ 27.5
Education 25.0/ 26.0/ 28.0 22.9/ 26.1/ 28.0 25.0/ 25.0/ 29.0 25.4/ 24.0/ 27.5 30.0/ 26.0/ 24.0 16.0/ 26.0/ 19.0 27.4/ 22.3/ 31.2 29.0/ 27.0/ 30.0 18.0/ 30.0/ 35.0 23.0/ 26.0/ 34.0 25.0/ 18.0/ 24.0
Energy and Power 22.0/ 24.0/ 30.0 31.8/ 40.5/ 40.5 26.0/ 27.0/ 35.0 30.5/ 32.9/ 34.6 36.0/ 30.0/ 36.0 24.0/ 22.0/ 25.0 21.0/ 28.0/ 27.0 22.0/ 21.0/ 24.0 24.0/ 27.0/ 30.0 27.0/ 27.0/ 32.0 29.0/ 25.0/ 29.0
Engineering 27.2/ 24.2/ 27.8 26.0/ 27.8/ 30.4 24.6/ 28.2/ 28.0 26.0/ 25.8/ 26.4 24.6/ 23.2/ 27.6 24.4/ 26.1/ 25.6 25.4/ 26.8/ 28.8 25.6/ 26.2/ 26.8 20.0/ 24.0/ 20.0 28.2/ 24.8/ 25.2 26.8/ 23.6/ 29.6
Environmental Science 26.0/ 26.0/ 33.0 29.5/ 31.9/ 36.7 21.0/ 27.0/ 39.0 24.1/ 29.7/ 35.1 23.4/ 21.9/ 19.5 24.0/ 27.0/ 25.0 27.4/ 24.2/ 24.7 20.2/ 22.2/ 35.4 24.0/ 35.0/ 41.0 34.0/ 32.0/ 28.0 27.9/ 26.3/ 26.8
Ethics and Human Rights 21.0/ 25.0/ 28.0 33.0/ 33.0/ 33.0 21.0/ 29.0/ 37.0 28.0/ 26.0/ 31.0 24.0/ 31.0/ 32.0 23.0/ 29.0/ 35.0 23.0/ 27.0/ 35.0 25.0/ 29.0/ 26.0 33.0/ 31.0/ 33.0 27.0/ 28.0/ 28.0 23.0/ 24.0/ 23.0
Finance and Investment 27.7/ 30.2/ 32.7 30.6/ 30.4/ 29.6 27.0/ 36.0/ 36.0 27.8/ 29.8/ 31.6 29.0/ 30.9/ 32.7 35.0/ 27.0/ 32.0 28.8/ 29.4/ 27.7 24.0/ 21.0/ 30.0 27.0/ 32.0/ 36.0 26.3/ 25.7/ 25.7 25.4/ 27.5/ 29.3
Food Science 26.0/ 30.0/ 33.0 26.0/ 38.0/ 42.0 39.0/ 28.0/ 27.0 27.3/ 26.3/ 33.3 29.0/ 26.0/ 27.0 26.0/ 24.0/ 26.0 27.0/ 24.0/ 28.0 22.0/ 24.0/ 25.0 18.0/ 30.0/ 25.0 27.0/ 25.0/ 27.0 28.0/ 26.0/ 30.0
Geography 24.6/ 24.0/ 27.4 24.2/ 32.0/ 34.2 28.0/ 33.0/ 31.0 21.7/ 26.7/ 28.7 24.2/ 29.2/ 32.9 21.2/ 23.2/ 27.3 23.7/ 23.7/ 26.0 24.0/ 24.0/ 30.0 24.0/ 33.0/ 30.0 22.9/ 20.6/ 27.4 23.4/ 29.2/ 25.5
Health and Medicine 23.0/ 28.0/ 38.0 26.1/ 40.7/ 42.1 27.0/ 28.0/ 28.0 30.7/ 29.5/ 38.1 24.0/ 37.0/ 38.0 27.0/ 32.0/ 31.0 28.6/ 24.1/ 25.9 27.0/ 21.0/ 30.0 32.0/ 42.0/ 44.0 25.9/ 28.4/ 28.4 34.6/ 32.7/ 34.6
History 24.0/ 27.7/ 30.7 28.3/ 33.7/ 35.7 31.0/ 28.0/ 33.0 28.0/ 27.0/ 30.2 25.3/ 26.7/ 30.8 21.0/ 24.0/ 25.0 25.3/ 30.9/ 29.0 34.0/ 26.0/ 29.0 35.0/ 29.0/ 32.0 27.6/ 24.7/ 29.1 27.6/ 28.4/ 28.1
Information Technology 32.0/ 34.0/ 45.0 26.0/ 30.0/ 45.0 32.0/ 39.0/ 40.0 29.6/ 30.2/ 33.5 28.0/ 33.0/ 38.0 29.0/ 36.0/ 34.0 33.0/ 36.0/ 42.0 34.0/ 29.0/ 31.0 31.0/ 33.0/ 34.0 35.0/ 38.0/ 37.0 31.0/ 26.0/ 34.0
International Relations 29.0/ 32.0/ 36.0 25.0/ 33.3/ 36.7 21.0/ 32.0/ 28.0 26.5/ 30.0/ 33.1 31.0/ 29.0/ 30.0 34.0/ 23.0/ 22.0 24.0/ 32.0/ 40.0 27.0/ 30.0/ 22.0 33.0/ 28.0/ 29.0 28.0/ 24.0/ 33.0 28.0/ 34.0/ 30.0
Language Studies 39.0/ 29.0/ 34.0 30.8/ 31.2/ 31.4 28.0/ 30.0/ 37.0 24.2/ 26.1/ 26.1 26.4/ 24.8/ 26.4 22.0/ 24.0/ 26.0 28.2/ 30.9/ 26.9 37.0/ 34.0/ 35.0 29.0/ 36.0/ 36.0 27.7/ 26.2/ 26.6 31.0/ 32.2/ 25.9
Law and Ethics 27.7/ 33.1/ 33.8 27.0/ 30.8/ 33.2 30.0/ 26.0/ 27.0 28.0/ 30.6/ 34.4 27.4/ 21.0/ 18.5 22.0/ 24.0/ 23.0 26.3/ 23.7/ 28.0 29.0/ 30.0/ 31.0 23.0/ 23.0/ 26.0 31.5/ 29.9/ 31.5 25.6/ 25.0/ 23.8
Literature and Linguistics 26.1/ 27.0/ 25.2 27.1/ 27.1/ 29.3 27.0/ 27.0/ 27.0 25.1/ 23.6/ 26.9 25.0/ 30.8/ 32.5 20.0/ 19.0/ 21.0 21.2/ 22.1/ 24.8 19.0/ 20.0/ 23.0 15.0/ 28.0/ 29.0 30.0/ 29.1/ 32.0 31.7/ 25.4/ 29.6
Logical Reasoning 31.4/ 32.2/ 29.5 25.6/ 24.4/ 29.0 29.2/ 25.3/ 25.3 26.7/ 25.9/ 28.9 29.5/ 22.4/ 21.2 29.0/ 28.0/ 24.0 31.5/ 26.2/ 27.6 27.0/ 24.0/ 19.0 32.0/ 29.0/ 22.0 29.1/ 28.3/ 25.5 26.2/ 22.3/ 24.0
Materials Science 22.0/ 25.0/ 31.0 26.6/ 31.2/ 36.1 16.0/ 24.0/ 22.0 24.9/ 27.1/ 31.4 25.0/ 30.0/ 30.0 24.0/ 29.0/ 30.0 23.0/ 23.0/ 22.0 15.0/ 25.0/ 24.0 29.0/ 22.0/ 26.0 19.0/ 15.0/ 24.0 29.0/ 29.0/ 30.0
Mathematics 28.8/ 28.6/ 29.8 23.2/ 26.8/ 27.4 21.6/ 24.2/ 24.8 24.6/ 24.9/ 27.5 24.8/ 24.0/ 24.4 22.6/ 24.9/ 24.9 22.8/ 25.6/ 26.2 24.0/ 26.0/ 24.6 26.9/ 20.1/ 23.5 25.0/ 29.2/ 28.0 21.8/ 22.8/ 23.2
Media and Communication 39.0/ 26.0/ 33.0 28.0/ 35.0/ 31.0 22.0/ 34.0/ 32.0 21.8/ 23.1/ 28.6 37.0/ 30.0/ 36.0 25.0/ 15.0/ 19.0 23.0/ 33.0/ 34.0 32.0/ 31.0/ 31.0 22.0/ 20.0/ 22.0 34.0/ 28.0/ 32.0 32.0/ 32.0/ 37.0
Music and Performing Arts 31.0/ 23.0/ 23.0 30.0/ 27.0/ 31.0 25.0/ 29.0/ 33.0 21.9/ 28.5/ 25.8 22.0/ 23.0/ 24.0 29.0/ 29.0/ 34.0 22.0/ 30.0/ 32.0 20.0/ 19.0/ 22.0 24.0/ 21.0/ 18.0 21.0/ 21.0/ 17.0 27.0/ 26.0/ 22.0
Physics 21.4/ 23.4/ 28.4 28.6/ 33.4/ 35.6 22.8/ 25.6/ 28.4 26.0/ 28.0/ 32.6 23.2/ 27.8/ 27.6 24.7/ 24.2/ 19.5 27.8/ 29.0/ 33.2 23.3/ 26.4/ 24.3 25.0/ 26.0/ 23.0 24.6/ 24.6/ 26.4 25.4/ 28.8/ 27.0
Politics and Governance 23.1/ 25.9/ 31.8 26.9/ 34.5/ 36.1 27.0/ 28.0/ 26.0 28.1/ 29.5/ 29.3 27.1/ 25.1/ 25.9 20.0/ 19.0/ 15.0 26.7/ 26.7/ 25.0 27.0/ 26.0/ 22.0 23.0/ 21.0/ 17.0 22.0/ 25.3/ 27.0 25.5/ 27.4/ 28.8
Psychology 17.0/ 16.0/ 29.0 26.0/ 25.0/ 27.0 32.0/ 22.0/ 27.0 22.5/ 25.6/ 26.4 32.0/ 33.0/ 35.0 23.0/ 18.0/ 21.0 22.0/ 17.0/ 27.0 28.0/ 29.0/ 28.0 24.0/ 18.0/ 21.0 18.0/ 15.0/ 19.0 26.0/ 33.0/ 34.0
Public Administration 19.0/ 22.0/ 25.0 19.2/ 33.3/ 32.3 22.2/ 31.3/ 24.2 28.3/ 27.3/ 40.4 29.3/ 29.3/ 28.3 18.0/ 21.0/ 19.0 22.2/ 18.2/ 31.3 21.0/ 16.0/ 29.0 24.2/ 30.3/ 33.3 24.0/ 22.0/ 32.0 20.0/ 21.0/ 29.0
Religion and Spirituality 24.0/ 32.0/ 31.0 20.0/ 28.0/ 34.0 29.0/ 32.0/ 33.0 24.6/ 26.2/ 27.4 14.0/ 22.0/ 23.0 22.0/ 24.0/ 25.0 22.0/ 24.0/ 25.0 17.0/ 24.0/ 36.0 22.0/ 27.0/ 21.0 23.0/ 28.0/ 23.0 21.0/ 25.0/ 27.0
Social Welfare and Development 21.0/ 20.0/ 25.0 24.1/ 30.4/ 33.0 20.0/ 30.0/ 30.0 21.8/ 25.4/ 35.2 24.0/ 28.0/ 30.0 25.0/ 26.0/ 26.0 30.0/ 27.0/ 33.0 20.0/ 24.0/ 24.0 22.0/ 34.0/ 37.0 23.0/ 26.0/ 32.0 37.0/ 36.0/ 33.0
Sociology 21.7/ 20.0/ 21.7 28.0/ 27.8/ 30.8 20.0/ 24.0/ 34.0 22.2/ 24.6/ 26.6 19.0/ 27.0/ 26.0 16.0/ 23.0/ 26.0 24.5/ 23.3/ 20.8 22.0/ 23.0/ 24.0 23.0/ 22.0/ 19.0 15.3/ 18.9/ 23.4 27.7/ 33.5/ 27.7
Sports and Recreation 22.8/ 24.8/ 21.3 23.8/ 31.2/ 34.2 23.0/ 23.0/ 30.0 22.6/ 23.6/ 29.4 25.3/ 26.4/ 27.5 29.0/ 19.0/ 25.0 23.2/ 23.2/ 27.1 23.0/ 22.0/ 28.0 25.0/ 33.0/ 32.0 25.6/ 25.0/ 26.3 18.1/ 26.2/ 30.0
Technology and Innovation 27.0/ 20.0/ 31.0 26.0/ 31.0/ 34.2 22.0/ 25.0/ 23.0 26.0/ 27.8/ 34.2 30.0/ 31.0/ 36.0 23.0/ 26.0/ 26.0 33.0/ 37.0/ 35.0 26.0/ 26.0/ 26.0 20.0/ 31.0/ 38.0 25.0/ 25.0/ 30.0 28.0/ 27.0/ 32.0
Transportation and Logistics 27.0/ 28.0/ 20.0 26.2/ 23.1/ 23.1 22.0/ 28.0/ 23.0 25.2/ 27.4/ 25.2 27.3/ 21.2/ 28.3 23.0/ 24.0/ 26.0 22.4/ 23.5/ 21.4 30.3/ 27.3/ 26.3 28.0/ 21.0/ 24.0 24.0/ 28.0/ 23.0 23.0/ 29.0/ 28.0
Table 36: Detailed subject-wise evaluation for sarvamai/sarvam-2b-v0.5 on MILU across different languages. The results reported are X / Y / Z where X denote the 0-shot, Y denotes 1-shot and Z denotes 5-shot performance respectively.
Topic bn en gu hi kn ml mr or pa ta te
Agriculture 31.0/ 32.0/ 35.0 34.0/ 46.1/ 47.5 20.2/ 27.3/ 29.3 25.4/ 29.4/ 35.2 27.0/ 33.0/ 29.0 28.3/ 27.3/ 27.3 31.0/ 34.0/ 34.0 33.0/ 33.0/ 29.0 29.0/ 25.0/ 22.0 28.0/ 28.0/ 28.0 18.0/ 23.0/ 26.0
Anthropology 34.0/ 34.0/ 38.0 37.0/ 57.0/ 58.0 36.0/ 41.0/ 33.0 33.3/ 31.3/ 25.2 25.0/ 24.0/ 27.0 23.0/ 27.0/ 24.0 27.0/ 32.0/ 34.0 26.0/ 38.0/ 32.0 30.0/ 27.0/ 27.0 30.0/ 27.0/ 34.0 26.0/ 25.0/ 19.0
Architecture and Design 27.0/ 32.0/ 34.0 38.0/ 60.0/ 62.0 27.0/ 21.0/ 24.0 27.0/ 35.0/ 34.0 32.0/ 38.0/ 34.0 24.0/ 25.0/ 28.0 30.0/ 38.0/ 36.0 24.0/ 24.0/ 27.0 25.0/ 27.0/ 31.0 34.0/ 30.0/ 26.0 23.0/ 30.0/ 28.0
Arts and Culture 27.9/ 29.1/ 32.7 40.0/ 59.6/ 59.6 30.0/ 31.0/ 30.0 25.9/ 31.1/ 33.7 27.6/ 28.8/ 32.5 26.0/ 34.0/ 31.0 22.9/ 26.2/ 31.1 15.0/ 15.0/ 21.0 24.0/ 26.0/ 29.0 27.3/ 27.9/ 26.7 26.1/ 23.6/ 28.1
Astronomy and Astrophysics 29.0/ 28.0/ 34.0 43.0/ 56.0/ 69.0 29.0/ 39.0/ 35.0 31.2/ 33.6/ 34.4 37.0/ 39.0/ 35.0 31.0/ 34.0/ 37.0 33.0/ 34.0/ 38.0 24.0/ 28.0/ 22.0 30.0/ 34.0/ 31.0 27.0/ 35.0/ 37.0 34.0/ 38.0/ 34.0
Biology 25.7/ 28.5/ 33.1 45.5/ 62.5/ 66.7 25.0/ 28.0/ 27.0 32.7/ 34.9/ 39.5 33.2/ 34.4/ 33.6 28.3/ 28.3/ 25.2 30.8/ 32.8/ 35.8 28.0/ 35.0/ 33.0 29.0/ 34.0/ 35.0 28.3/ 33.4/ 29.9 27.4/ 31.4/ 32.4
Business and Management 29.0/ 31.0/ 31.0 36.2/ 48.9/ 50.8 31.0/ 30.0/ 38.0 30.9/ 34.4/ 37.7 33.0/ 32.0/ 29.0 30.0/ 27.0/ 27.0 31.0/ 35.0/ 38.0 24.0/ 18.0/ 23.0 32.0/ 30.0/ 36.0 34.0/ 33.0/ 30.0 30.2/ 31.1/ 27.4
Chemistry 28.3/ 29.4/ 33.5 39.1/ 54.9/ 56.7 29.3/ 33.6/ 30.2 33.3/ 35.3/ 38.1 29.4/ 25.5/ 30.1 33.0/ 25.0/ 31.0 30.1/ 31.4/ 33.5 25.0/ 30.8/ 29.8 27.0/ 24.0/ 34.0 26.4/ 27.9/ 29.4 32.7/ 30.4/ 33.2
Computer Science 29.2/ 30.2/ 34.9 34.4/ 38.8/ 39.4 28.0/ 33.0/ 37.0 32.6/ 30.6/ 32.8 30.2/ 32.5/ 32.5 26.0/ 29.0/ 30.0 27.2/ 33.1/ 34.9 27.0/ 25.0/ 24.0 30.0/ 31.0/ 22.0 28.5/ 30.2/ 36.6 30.9/ 31.8/ 36.3
Defense and Security 23.0/ 24.0/ 30.0 30.0/ 52.0/ 50.0 24.0/ 23.0/ 24.0 25.6/ 24.6/ 25.6 25.0/ 24.0/ 25.0 21.0/ 20.0/ 24.0 16.0/ 20.0/ 25.0 27.0/ 26.0/ 29.0 33.0/ 32.0/ 34.0 28.0/ 33.0/ 32.0 26.0/ 26.0/ 24.0
Earth Sciences 32.0/ 36.0/ 32.0 36.6/ 55.9/ 59.6 30.0/ 36.0/ 34.0 30.5/ 27.9/ 34.7 30.0/ 30.0/ 25.0 28.0/ 28.0/ 31.0 21.0/ 31.0/ 28.0 25.2/ 32.3/ 28.3 30.0/ 39.0/ 30.0 24.0/ 26.0/ 28.0 25.0/ 21.0/ 25.0
Economics 30.2/ 34.5/ 35.2 36.4/ 48.6/ 51.4 35.0/ 29.0/ 32.0 28.6/ 34.1/ 38.1 28.0/ 29.0/ 32.0 29.0/ 30.0/ 29.0 31.0/ 30.3/ 28.2 25.0/ 31.0/ 29.0 29.0/ 25.0/ 27.0 39.7/ 36.5/ 33.3 32.3/ 38.9/ 35.3
Education 32.0/ 26.0/ 31.0 31.5/ 39.5/ 38.2 25.0/ 29.0/ 28.0 28.7/ 33.1/ 33.7 33.0/ 36.0/ 37.0 23.0/ 26.0/ 32.0 36.3/ 30.6/ 38.2 25.0/ 22.0/ 27.0 31.0/ 29.0/ 25.0 27.0/ 30.0/ 32.0 24.0/ 29.0/ 27.0
Energy and Power 28.0/ 29.0/ 25.0 37.8/ 54.7/ 54.7 29.0/ 37.0/ 41.0 26.8/ 30.9/ 40.7 31.0/ 32.0/ 27.0 32.0/ 27.0/ 18.0 19.0/ 20.0/ 30.0 28.0/ 28.0/ 30.0 27.0/ 30.0/ 29.0 33.0/ 29.0/ 41.0 33.0/ 33.0/ 32.0
Engineering 30.0/ 25.6/ 32.4 31.2/ 36.8/ 40.8 25.8/ 26.2/ 28.2 25.6/ 27.4/ 28.2 25.2/ 29.0/ 29.8 29.1/ 28.2/ 30.3 27.6/ 25.0/ 28.0 29.5/ 29.5/ 28.6 28.0/ 34.0/ 32.0 27.8/ 29.4/ 30.6 24.0/ 27.0/ 26.2
Environmental Science 37.0/ 38.0/ 31.0 40.7/ 54.9/ 55.5 27.0/ 34.0/ 31.0 28.7/ 33.1/ 36.5 27.3/ 27.3/ 32.8 35.0/ 29.0/ 32.0 25.3/ 32.1/ 34.2 30.3/ 29.3/ 27.3 43.0/ 38.0/ 34.0 34.0/ 28.0/ 31.0 31.8/ 29.6/ 25.1
Ethics and Human Rights 35.0/ 30.0/ 41.0 40.0/ 52.0/ 49.0 32.0/ 30.0/ 31.0 35.0/ 44.0/ 39.0 38.0/ 33.0/ 32.0 24.0/ 29.0/ 28.0 33.0/ 36.0/ 39.0 29.0/ 28.0/ 32.0 24.0/ 26.0/ 25.0 29.0/ 37.0/ 35.0 36.0/ 35.0/ 36.0
Finance and Investment 27.7/ 29.2/ 30.7 36.8/ 42.2/ 43.2 35.0/ 37.0/ 35.0 33.8/ 33.0/ 34.0 30.9/ 37.0/ 38.9 18.0/ 24.0/ 29.0 27.1/ 33.3/ 33.9 20.0/ 26.0/ 23.0 36.0/ 35.0/ 31.0 32.4/ 32.4/ 34.1 31.4/ 31.4/ 32.9
Food Science 21.0/ 29.0/ 31.0 50.0/ 60.0/ 65.0 30.0/ 32.0/ 32.0 27.3/ 33.3/ 39.4 31.0/ 33.0/ 40.0 28.0/ 36.0/ 32.0 21.0/ 27.0/ 32.0 32.0/ 27.0/ 24.0 26.0/ 28.0/ 26.0 31.0/ 32.0/ 28.0 33.0/ 28.0/ 29.0
Geography 31.3/ 35.2/ 35.2 36.6/ 46.6/ 50.0 37.0/ 31.0/ 29.0 26.5/ 28.3/ 31.3 28.3/ 34.7/ 36.5 31.3/ 41.4/ 45.5 28.4/ 30.8/ 34.3 22.0/ 29.0/ 30.0 30.0/ 30.0/ 28.0 30.9/ 32.6/ 30.9 25.0/ 27.1/ 28.1
Health and Medicine 31.0/ 34.0/ 31.0 43.3/ 65.7/ 68.7 30.0/ 30.0/ 35.0 33.5/ 37.3/ 42.9 25.0/ 28.0/ 28.0 31.0/ 33.0/ 31.0 35.7/ 33.0/ 32.1 23.0/ 25.0/ 23.0 35.0/ 37.0/ 42.0 32.8/ 32.8/ 29.3 29.9/ 29.9/ 32.7
History 30.9/ 28.8/ 31.2 38.8/ 44.8/ 44.6 35.0/ 35.0/ 37.0 32.7/ 32.9/ 32.9 31.9/ 30.0/ 33.3 34.0/ 27.0/ 30.0 32.7/ 30.5/ 29.4 36.0/ 32.0/ 30.0 29.0/ 30.0/ 33.0 31.6/ 33.8/ 32.7 32.2/ 31.7/ 29.9
Information Technology 40.0/ 40.0/ 42.0 43.0/ 58.0/ 57.0 37.0/ 35.0/ 37.0 36.9/ 41.3/ 45.2 38.0/ 37.0/ 41.0 31.0/ 33.0/ 35.0 43.0/ 46.0/ 44.0 29.0/ 29.0/ 31.0 43.0/ 48.0/ 40.0 30.0/ 29.0/ 31.0 43.0/ 51.0/ 46.0
International Relations 23.0/ 25.0/ 27.0 43.3/ 53.3/ 57.5 31.0/ 31.0/ 32.0 31.9/ 34.2/ 38.5 21.0/ 30.0/ 30.0 30.0/ 31.0/ 30.0 30.0/ 33.0/ 43.0 27.0/ 35.0/ 32.0 35.0/ 35.0/ 35.0 39.0/ 37.0/ 39.0 24.0/ 22.0/ 28.0
Language Studies 28.0/ 29.0/ 33.0 38.2/ 46.8/ 49.6 25.0/ 26.0/ 31.0 25.1/ 28.1/ 28.5 30.2/ 24.0/ 24.8 25.0/ 21.0/ 22.0 29.5/ 33.6/ 34.9 32.0/ 31.0/ 30.0 38.0/ 38.0/ 37.0 26.6/ 30.0/ 28.8 24.7/ 29.3/ 29.3
Law and Ethics 40.5/ 39.2/ 43.2 44.6/ 52.8/ 51.2 29.0/ 32.0/ 30.0 35.2/ 37.2/ 38.8 29.3/ 25.5/ 26.8 32.0/ 32.0/ 31.0 29.3/ 26.7/ 29.7 32.0/ 32.0/ 26.0 28.0/ 37.0/ 34.0 26.0/ 28.3/ 28.3 29.4/ 25.6/ 26.9
Literature and Linguistics 28.8/ 30.6/ 36.0 36.9/ 46.7/ 44.9 29.0/ 21.0/ 26.0 24.1/ 25.7/ 31.7 27.5/ 30.8/ 30.0 24.0/ 27.0/ 28.0 22.1/ 29.2/ 26.6 21.0/ 19.0/ 25.0 28.0/ 33.0/ 32.0 30.4/ 32.4/ 29.1 22.5/ 23.9/ 28.9
Logical Reasoning 28.5/ 34.9/ 35.9 31.0/ 34.2/ 31.2 26.6/ 29.2/ 29.2 27.1/ 26.2/ 29.3 31.1/ 25.7/ 31.9 31.0/ 26.0/ 25.0 22.8/ 26.2/ 27.0 19.0/ 23.0/ 26.0 34.0/ 37.0/ 27.0 28.7/ 29.9/ 29.5 25.7/ 28.0/ 27.5
Materials Science 25.0/ 28.0/ 30.0 33.8/ 48.7/ 51.7 26.0/ 23.0/ 22.0 26.4/ 31.2/ 32.3 27.0/ 26.0/ 28.0 32.0/ 32.0/ 34.0 29.0/ 30.0/ 26.0 27.0/ 19.0/ 21.0 26.0/ 28.0/ 27.0 27.0/ 26.0/ 28.0 30.0/ 26.0/ 24.0
Mathematics 27.6/ 32.0/ 32.6 30.6/ 30.0/ 31.6 28.6/ 29.2/ 27.2 28.3/ 27.5/ 26.9 26.8/ 25.0/ 27.2 27.5/ 28.1/ 27.8 26.4/ 27.4/ 25.4 26.2/ 27.4/ 27.2 31.1/ 29.2/ 28.4 24.2/ 25.6/ 27.8 27.6/ 29.6/ 28.2
Media and Communication 33.0/ 35.0/ 33.0 41.0/ 57.0/ 48.0 35.0/ 36.0/ 32.0 34.0/ 32.0/ 35.4 38.0/ 41.0/ 32.0 36.0/ 31.0/ 34.0 26.0/ 29.0/ 28.0 37.0/ 32.0/ 34.0 26.0/ 26.0/ 28.0 30.0/ 25.0/ 30.0 35.0/ 35.0/ 34.0
Music and Performing Arts 32.0/ 31.0/ 28.0 38.0/ 47.0/ 46.0 27.0/ 34.0/ 32.0 27.8/ 33.1/ 30.5 27.0/ 27.0/ 29.0 23.0/ 23.0/ 20.0 28.0/ 32.0/ 35.0 21.0/ 32.0/ 30.0 23.0/ 32.0/ 34.0 27.0/ 30.0/ 29.0 31.0/ 32.0/ 28.0
Physics 28.8/ 31.8/ 31.8 37.8/ 49.4/ 52.6 32.0/ 29.5/ 28.1 27.6/ 30.6/ 32.8 29.4/ 32.0/ 33.0 26.3/ 25.3/ 27.9 27.4/ 30.8/ 35.4 28.8/ 28.1/ 28.5 27.0/ 27.0/ 27.0 29.0/ 28.8/ 29.2 28.8/ 33.2/ 32.0
Politics and Governance 27.8/ 31.8/ 30.6 37.9/ 47.1/ 45.3 25.0/ 31.0/ 30.0 27.5/ 31.5/ 34.9 23.9/ 25.1/ 26.7 26.0/ 33.0/ 33.0 27.5/ 27.8/ 29.5 23.0/ 23.0/ 21.0 28.0/ 28.0/ 31.0 36.1/ 31.9/ 33.2 30.2/ 25.2/ 29.7
Psychology 28.0/ 30.0/ 34.0 37.0/ 57.0/ 59.0 28.0/ 35.0/ 31.0 28.7/ 34.9/ 34.9 34.0/ 40.0/ 43.0 25.0/ 27.0/ 32.0 29.0/ 28.0/ 31.0 37.0/ 37.0/ 44.0 29.0/ 35.0/ 27.0 37.0/ 39.0/ 34.0 29.0/ 34.0/ 29.0
Public Administration 25.0/ 22.0/ 27.0 35.4/ 39.4/ 43.4 29.3/ 30.3/ 34.3 32.3/ 35.4/ 37.4 27.3/ 25.2/ 26.3 23.0/ 24.0/ 23.0 26.3/ 27.3/ 28.3 21.0/ 27.0/ 29.0 27.3/ 30.3/ 29.3 30.0/ 29.0/ 30.0 27.0/ 24.0/ 30.0
Religion and Spirituality 17.0/ 27.0/ 28.0 33.0/ 55.0/ 59.0 37.0/ 41.0/ 43.0 22.6/ 31.4/ 31.4 35.0/ 38.0/ 34.0 34.0/ 35.0/ 39.0 28.0/ 36.0/ 30.0 27.0/ 30.0/ 28.0 27.0/ 27.0/ 31.0 18.0/ 25.0/ 25.0 29.0/ 39.0/ 37.0
Social Welfare and Development 29.0/ 34.0/ 37.0 31.2/ 50.0/ 48.2 38.0/ 35.0/ 37.0 30.0/ 29.0/ 31.1 28.0/ 34.0/ 36.0 18.0/ 20.0/ 18.0 32.0/ 35.0/ 44.0 34.0/ 29.0/ 29.0 26.0/ 29.0/ 29.0 28.0/ 37.0/ 28.0 31.0/ 30.0/ 24.0
Sociology 21.7/ 21.7/ 24.3 36.4/ 38.2/ 38.6 36.0/ 34.0/ 33.0 28.2/ 30.2/ 30.2 31.0/ 27.0/ 29.0 18.0/ 23.0/ 26.0 27.7/ 25.2/ 26.4 20.0/ 29.0/ 26.0 28.0/ 28.0/ 30.0 21.6/ 20.7/ 25.2 24.7/ 20.0/ 22.4
Sports and Recreation 27.7/ 26.7/ 29.7 40.0/ 53.2/ 52.2 31.0/ 25.0/ 36.0 30.4/ 32.2/ 37.8 28.1/ 28.6/ 32.0 30.0/ 27.0/ 26.0 27.7/ 32.2/ 38.4 26.0/ 32.0/ 28.0 28.0/ 34.0/ 33.0 32.0/ 35.3/ 30.8 26.7/ 29.0/ 29.0
Technology and Innovation 42.0/ 42.0/ 34.0 38.6/ 49.8/ 50.6 24.0/ 30.0/ 33.0 31.6/ 33.4/ 39.4 31.0/ 35.0/ 36.0 27.0/ 32.0/ 28.0 29.0/ 33.0/ 31.0 28.0/ 36.0/ 32.0 25.0/ 31.0/ 29.0 29.0/ 36.0/ 39.0 30.0/ 27.0/ 29.0
Transportation and Logistics 27.0/ 28.0/ 32.0 33.9/ 39.2/ 43.1 29.0/ 33.0/ 29.0 28.4/ 27.4/ 27.8 31.3/ 34.3/ 33.3 19.0/ 29.0/ 27.0 25.5/ 29.6/ 29.6 25.2/ 22.2/ 26.3 21.0/ 23.0/ 24.0 31.0/ 29.0/ 27.0 23.0/ 24.0/ 34.0
Table 37: Detailed subject-wise evaluation for meta-llama/Llama-3.2-3B on MILU across different languages. The results reported are X / Y / Z where X denote the 0-shot, Y denotes 1-shot and Z denotes 5-shot performance respectively.
Topic bn en gu hi kn ml mr or pa ta te
Agriculture 27.0/ 22.0/ 32.0 36.4/ 33.8/ 32.6 32.3/ 29.3/ 27.3 31.4/ 27.4/ 25.6 32.0/ 21.0/ 32.0 25.2/ 29.3/ 33.3 35.0/ 30.0/ 30.0 27.0/ 25.0/ 33.0 31.0/ 25.0/ 27.0 31.0/ 29.0/ 33.0 32.0/ 33.0/ 35.0
Anthropology 31.0/ 30.0/ 38.0 32.0/ 43.0/ 30.0 33.0/ 33.0/ 33.0 35.4/ 37.4/ 36.4 29.0/ 37.0/ 40.0 24.0/ 27.0/ 30.0 31.0/ 24.0/ 31.0 31.0/ 32.0/ 26.0 34.0/ 33.0/ 37.0 34.0/ 31.0/ 40.0 26.0/ 28.0/ 39.0
Architecture and Design 31.0/ 30.0/ 40.0 40.0/ 33.0/ 32.0 30.0/ 33.0/ 27.0 33.0/ 29.0/ 35.0 34.0/ 30.0/ 42.0 36.0/ 34.0/ 31.0 39.0/ 33.0/ 31.0 31.0/ 25.0/ 26.0 32.0/ 30.0/ 33.0 29.0/ 39.0/ 35.0 33.0/ 33.0/ 31.0
Arts and Culture 27.9/ 29.1/ 29.1 39.2/ 36.6/ 28.8 27.0/ 22.0/ 27.0 29.7/ 30.3/ 27.5 27.0/ 23.3/ 22.7 32.0/ 32.0/ 36.0 25.1/ 32.2/ 24.0 24.0/ 20.0/ 23.0 27.0/ 33.0/ 28.0 30.3/ 25.4/ 27.9 25.6/ 29.1/ 27.6
Astronomy and Astrophysics 36.0/ 32.0/ 36.0 57.0/ 51.0/ 36.0 39.0/ 40.0/ 41.0 39.1/ 39.8/ 39.1 39.0/ 35.0/ 38.0 35.0/ 37.0/ 36.0 39.0/ 35.0/ 36.0 30.0/ 33.0/ 32.0 38.0/ 29.0/ 41.0 31.0/ 34.0/ 44.0 32.0/ 37.0/ 31.0
Biology 32.4/ 27.5/ 29.2 43.5/ 40.1/ 33.7 31.0/ 25.0/ 27.0 36.7/ 37.1/ 33.9 30.7/ 32.8/ 30.7 30.3/ 34.3/ 22.2 29.5/ 29.0/ 25.7 24.0/ 25.0/ 23.0 31.0/ 32.0/ 38.0 27.1/ 22.9/ 27.7 30.5/ 34.0/ 32.4
Business and Management 25.0/ 30.0/ 40.0 37.8/ 37.5/ 31.3 34.0/ 35.0/ 34.0 23.9/ 24.6/ 29.0 27.0/ 26.0/ 26.0 32.0/ 25.0/ 32.0 28.0/ 24.0/ 31.0 28.0/ 25.0/ 23.0 27.0/ 23.0/ 27.0 29.0/ 33.0/ 36.0 26.4/ 26.4/ 31.1
Chemistry 33.0/ 27.5/ 38.2 48.1/ 40.3/ 34.9 27.6/ 29.3/ 39.7 32.9/ 33.7/ 34.9 26.5/ 28.1/ 27.5 23.0/ 21.0/ 30.0 26.9/ 30.9/ 33.0 26.0/ 28.8/ 26.9 26.0/ 31.0/ 35.0 27.0/ 30.0/ 30.6 27.4/ 31.2/ 32.9
Computer Science 27.1/ 27.1/ 27.6 32.6/ 32.6/ 32.6 31.0/ 26.0/ 27.0 26.2/ 28.4/ 29.6 29.0/ 22.5/ 36.1 36.0/ 27.0/ 38.0 24.9/ 24.3/ 27.8 31.0/ 25.0/ 21.0 22.0/ 21.0/ 21.0 28.5/ 27.9/ 33.7 32.3/ 30.0/ 35.4
Defense and Security 29.0/ 28.0/ 30.0 45.0/ 40.0/ 33.0 36.0/ 29.0/ 26.0 28.4/ 24.6/ 27.5 29.0/ 36.0/ 40.0 25.0/ 35.0/ 30.0 33.0/ 27.0/ 23.0 26.0/ 27.0/ 22.0 28.0/ 35.0/ 30.0 28.0/ 29.0/ 36.0 33.0/ 41.0/ 32.0
Earth Sciences 35.0/ 38.0/ 46.0 43.2/ 40.7/ 35.9 34.0/ 33.0/ 35.0 32.1/ 29.3/ 28.9 28.0/ 38.0/ 46.0 29.0/ 28.0/ 30.0 33.0/ 33.0/ 24.0 23.2/ 22.2/ 23.2 29.0/ 34.0/ 30.0 29.0/ 31.0/ 32.0 33.0/ 33.0/ 35.0
Economics 33.8/ 36.7/ 33.1 38.1/ 37.5/ 29.9 22.0/ 30.0/ 28.0 29.2/ 32.1/ 28.4 33.0/ 27.0/ 41.0 32.0/ 26.0/ 35.0 33.1/ 27.5/ 28.2 28.0/ 31.0/ 26.0 31.0/ 29.0/ 34.0 32.5/ 30.9/ 28.6 25.1/ 32.3/ 28.1
Education 33.0/ 34.0/ 38.0 33.4/ 28.7/ 27.4 28.0/ 33.0/ 31.0 33.7/ 36.1/ 30.8 39.0/ 32.0/ 34.0 22.0/ 24.0/ 30.0 37.6/ 33.8/ 36.3 29.0/ 31.0/ 33.0 23.0/ 28.0/ 32.0 31.0/ 37.0/ 23.0 41.0/ 39.0/ 39.0
Energy and Power 27.0/ 25.0/ 35.0 43.9/ 37.2/ 29.7 27.0/ 32.0/ 36.0 31.9/ 32.9/ 29.1 33.0/ 30.0/ 30.0 24.0/ 25.0/ 29.0 31.0/ 37.0/ 31.0 27.0/ 26.0/ 31.0 35.0/ 39.0/ 40.0 30.0/ 33.0/ 30.0 27.0/ 32.0/ 25.0
Engineering 25.8/ 28.4/ 25.8 32.8/ 26.4/ 26.0 23.4/ 29.2/ 30.4 27.4/ 27.4/ 28.6 30.0/ 26.4/ 30.0 31.2/ 29.9/ 29.5 26.0/ 23.2/ 25.4 26.5/ 27.1/ 28.3 29.0/ 25.0/ 30.0 26.2/ 26.4/ 29.2 29.2/ 28.4/ 31.8
Environmental Science 32.0/ 25.0/ 36.0 43.9/ 40.5/ 31.5 30.0/ 31.0/ 38.0 33.1/ 32.1/ 31.7 30.5/ 25.8/ 37.5 29.0/ 30.0/ 34.0 28.4/ 29.5/ 32.6 20.2/ 20.2/ 24.2 35.0/ 34.0/ 39.0 33.0/ 32.0/ 33.0 34.1/ 30.7/ 34.6
Ethics and Human Rights 32.0/ 29.0/ 28.0 38.0/ 47.0/ 33.0 30.0/ 33.0/ 28.0 27.0/ 27.0/ 27.0 27.0/ 25.0/ 25.0 28.0/ 27.0/ 31.0 35.0/ 35.0/ 35.0 19.0/ 20.0/ 18.0 24.0/ 28.0/ 23.0 27.0/ 35.0/ 38.0 31.0/ 29.0/ 27.0
Finance and Investment 32.2/ 31.2/ 33.7 30.8/ 30.4/ 29.2 30.0/ 29.0/ 33.0 26.4/ 26.6/ 26.8 34.0/ 34.0/ 30.2 24.0/ 41.0/ 26.0 28.2/ 29.4/ 30.5 27.0/ 33.0/ 25.0 26.0/ 27.0/ 35.0 25.7/ 32.4/ 31.8 28.2/ 28.6/ 31.4
Food Science 28.0/ 23.0/ 35.0 38.0/ 37.0/ 24.0 31.0/ 28.0/ 37.0 27.3/ 33.3/ 26.3 31.0/ 21.0/ 29.0 31.0/ 30.0/ 26.0 35.0/ 36.0/ 37.0 26.0/ 27.0/ 31.0 25.0/ 30.0/ 28.0 21.0/ 28.0/ 28.0 30.0/ 35.0/ 33.0
Geography 28.5/ 29.0/ 30.2 40.6/ 36.2/ 32.0 31.0/ 38.0/ 35.0 30.3/ 31.1/ 31.1 31.5/ 26.0/ 33.3 32.3/ 29.3/ 31.3 32.5/ 28.4/ 28.4 27.0/ 24.0/ 27.0 26.0/ 28.0/ 23.0 30.3/ 24.6/ 31.4 28.1/ 28.1/ 32.8
Health and Medicine 25.0/ 21.0/ 33.0 54.1/ 49.3/ 38.7 27.0/ 35.0/ 29.0 37.3/ 37.1/ 34.7 37.0/ 27.0/ 43.0 27.0/ 20.0/ 24.0 28.6/ 21.4/ 25.0 26.0/ 30.0/ 25.0 29.0/ 36.0/ 36.0 26.7/ 29.3/ 39.7 31.8/ 29.9/ 33.6
History 33.1/ 32.0/ 36.3 37.5/ 35.3/ 29.9 42.0/ 36.0/ 39.0 31.9/ 34.7/ 33.1 34.8/ 28.9/ 36.3 28.0/ 35.0/ 29.0 29.7/ 29.7/ 32.0 27.0/ 28.0/ 26.0 29.0/ 32.0/ 35.0 34.2/ 34.5/ 34.5 28.1/ 29.9/ 34.0
Information Technology 38.0/ 30.0/ 41.0 34.0/ 45.0/ 35.0 45.0/ 42.0/ 44.0 41.3/ 43.0/ 38.6 36.0/ 30.0/ 39.0 39.0/ 34.0/ 37.0 42.0/ 38.0/ 39.0 30.0/ 31.0/ 31.0 36.0/ 42.0/ 41.0 35.0/ 26.0/ 30.0 41.0/ 44.0/ 43.0
International Relations 29.0/ 24.0/ 35.0 46.7/ 45.8/ 35.0 36.0/ 38.0/ 40.0 33.5/ 32.7/ 28.8 33.0/ 32.0/ 30.0 43.0/ 35.0/ 40.0 38.0/ 28.0/ 30.0 28.0/ 23.0/ 26.0 36.0/ 41.0/ 42.0 34.0/ 37.0/ 40.0 26.0/ 28.0/ 25.0
Language Studies 30.0/ 31.0/ 31.0 42.6/ 38.6/ 35.8 29.0/ 26.0/ 29.0 29.9/ 27.5/ 30.7 32.6/ 26.4/ 30.2 23.0/ 23.0/ 25.0 30.2/ 28.2/ 26.9 28.0/ 20.0/ 21.0 32.0/ 30.0/ 33.0 25.5/ 22.1/ 24.7 27.0/ 21.3/ 24.7
Law and Ethics 45.3/ 45.3/ 47.3 41.0/ 41.6/ 39.0 26.0/ 29.0/ 33.0 36.4/ 37.2/ 36.4 29.9/ 28.0/ 27.4 24.0/ 28.0/ 31.0 27.2/ 28.9/ 28.0 25.0/ 30.0/ 29.0 31.0/ 29.0/ 30.0 29.1/ 26.8/ 35.4 23.8/ 23.8/ 28.7
Literature and Linguistics 27.0/ 25.2/ 25.2 35.1/ 35.7/ 27.5 32.0/ 31.0/ 30.0 32.3/ 30.7/ 30.3 31.7/ 27.5/ 29.2 34.0/ 28.0/ 27.0 25.7/ 27.4/ 24.8 20.0/ 23.0/ 26.0 25.0/ 29.0/ 30.0 27.9/ 27.1/ 28.7 26.8/ 23.9/ 31.0
Logical Reasoning 25.8/ 29.5/ 32.9 24.0/ 26.6/ 27.8 23.4/ 24.0/ 29.2 29.3/ 32.1/ 28.9 23.2/ 27.8/ 27.4 26.0/ 27.0/ 32.0 24.2/ 24.2/ 27.0 17.0/ 17.0/ 21.0 30.0/ 30.0/ 32.0 27.1/ 23.9/ 25.9 26.7/ 24.8/ 27.5
Materials Science 31.0/ 25.0/ 32.0 39.5/ 36.1/ 33.8 27.0/ 33.0/ 34.0 30.1/ 29.0/ 30.3 27.0/ 34.0/ 31.0 32.0/ 31.0/ 40.0 33.0/ 36.0/ 36.0 26.0/ 29.0/ 24.0 35.0/ 29.0/ 30.0 29.0/ 32.0/ 33.0 25.0/ 23.0/ 25.0
Mathematics 30.0/ 27.8/ 34.0 26.0/ 29.8/ 30.4 25.6/ 28.0/ 32.4 26.7/ 26.7/ 27.7 26.6/ 26.2/ 29.0 28.4/ 26.9/ 26.9 26.2/ 23.4/ 25.8 28.0/ 27.4/ 28.2 25.0/ 29.2/ 32.2 23.6/ 24.0/ 30.2 23.2/ 24.0/ 31.4
Media and Communication 36.0/ 35.0/ 39.0 46.0/ 39.0/ 36.0 32.0/ 34.0/ 33.0 32.6/ 28.6/ 32.6 41.0/ 33.0/ 39.0 32.0/ 27.0/ 30.0 30.0/ 35.0/ 34.0 34.0/ 31.0/ 38.0 27.0/ 30.0/ 32.0 32.0/ 30.0/ 36.0 38.0/ 39.0/ 39.0
Music and Performing Arts 28.0/ 30.0/ 36.0 37.0/ 34.0/ 30.0 37.0/ 28.0/ 32.0 27.2/ 26.5/ 24.5 30.0/ 26.0/ 31.0 31.0/ 30.0/ 28.0 25.0/ 28.0/ 28.0 21.0/ 24.0/ 25.0 32.0/ 31.0/ 40.0 23.0/ 23.0/ 27.0 35.0/ 33.0/ 34.0
Physics 25.8/ 24.8/ 31.6 39.8/ 36.4/ 35.8 27.9/ 28.1/ 31.8 29.4/ 28.0/ 33.0 28.2/ 25.4/ 29.6 25.8/ 26.8/ 30.5 27.6/ 27.8/ 27.6 25.7/ 26.0/ 28.1 26.0/ 33.0/ 35.0 32.4/ 28.0/ 30.6 27.8/ 30.6/ 31.4
Politics and Governance 29.4/ 29.8/ 35.7 40.1/ 36.3/ 32.5 27.0/ 38.0/ 38.0 33.7/ 34.3/ 31.5 33.2/ 26.3/ 34.0 28.0/ 26.0/ 28.0 24.4/ 24.2/ 24.2 26.0/ 21.0/ 30.0 33.0/ 32.0/ 29.0 24.9/ 29.0/ 29.9 29.2/ 30.4/ 30.2
Psychology 40.0/ 27.0/ 44.0 43.0/ 47.0/ 30.0 33.0/ 30.0/ 30.0 38.8/ 34.9/ 34.1 34.0/ 30.0/ 42.0 32.0/ 28.0/ 27.0 29.0/ 27.0/ 25.0 25.0/ 28.0/ 29.0 27.0/ 21.0/ 27.0 36.0/ 36.0/ 40.0 32.0/ 30.0/ 32.0
Public Administration 24.0/ 28.0/ 20.0 33.3/ 30.3/ 29.3 22.2/ 28.3/ 31.3 28.3/ 28.3/ 27.3 28.3/ 25.2/ 27.3 27.0/ 30.0/ 30.0 28.3/ 26.3/ 25.2 21.0/ 22.0/ 25.0 31.3/ 30.3/ 32.3 23.0/ 28.0/ 32.0 27.0/ 24.0/ 29.0
Religion and Spirituality 26.0/ 25.0/ 25.0 46.0/ 43.0/ 31.0 37.0/ 42.0/ 41.0 30.2/ 31.1/ 25.8 31.0/ 23.0/ 25.0 32.0/ 33.0/ 26.0 39.0/ 33.0/ 25.0 24.0/ 27.0/ 26.0 31.0/ 34.0/ 37.0 32.0/ 31.0/ 28.0 30.0/ 27.0/ 29.0
Social Welfare and Development 22.0/ 28.0/ 32.0 36.6/ 35.7/ 33.9 38.0/ 39.0/ 40.0 28.5/ 32.1/ 32.6 23.0/ 33.0/ 33.0 29.0/ 30.0/ 32.0 35.0/ 31.0/ 34.0 25.0/ 24.0/ 20.0 28.0/ 36.0/ 31.0 37.0/ 35.0/ 32.0 32.0/ 34.0/ 37.0
Sociology 26.1/ 22.6/ 24.3 30.0/ 29.0/ 26.0 35.0/ 35.0/ 40.0 25.0/ 24.0/ 26.4 31.0/ 26.0/ 27.0 18.0/ 20.0/ 22.0 25.8/ 30.2/ 28.3 14.0/ 15.0/ 22.0 29.0/ 28.0/ 32.0 34.2/ 35.1/ 25.2 30.0/ 25.3/ 25.9
Sports and Recreation 24.8/ 23.8/ 25.7 44.0/ 37.2/ 29.8 27.0/ 25.0/ 23.0 33.8/ 31.0/ 29.2 26.4/ 27.5/ 24.2 21.0/ 31.0/ 27.0 32.2/ 29.4/ 23.7 21.0/ 26.0/ 25.0 25.0/ 28.0/ 27.0 33.3/ 29.5/ 25.6 26.7/ 27.6/ 25.2
Technology and Innovation 30.0/ 34.0/ 40.0 41.6/ 37.0/ 32.0 30.0/ 36.0/ 43.0 32.6/ 34.0/ 32.2 35.0/ 28.0/ 32.0 24.0/ 25.0/ 31.0 32.0/ 29.0/ 29.0 32.0/ 30.0/ 33.0 37.0/ 35.0/ 36.0 28.0/ 28.0/ 35.0 29.0/ 29.0/ 27.0
Transportation and Logistics 26.0/ 25.0/ 25.0 37.7/ 38.5/ 26.2 25.0/ 32.0/ 28.0 30.9/ 29.6/ 29.6 26.3/ 24.2/ 28.3 24.0/ 20.0/ 29.0 23.5/ 17.3/ 17.3 20.2/ 28.3/ 25.2 20.0/ 22.0/ 28.0 27.0/ 24.0/ 27.0 22.0/ 26.0/ 26.0
Table 38: Detailed subject-wise evaluation for meta-llama/Llama-3.2-3B-Instruct on MILU across different languages. The results reported are X / Y / Z where X denote the 0-shot, Y denotes 1-shot and Z denotes 5-shot performance respectively.
Topic bn en gu hi kn ml mr or pa ta te
Agriculture 25 46.7 22.2 52.5 36 31.3 29 20 27 32 28
Anthropology 27 48 22 51.5 26 23 37 24 25 24 19
Architecture and Design 34 54 28 50 26 25 50 22 30 27 26
Arts and Culture 26.1 58.6 28 60.1 28.8 26 48.6 24 37 26.1 24.1
Astronomy and Astrophysics 31 72 26 58.6 24 25 56 23 27 30 29
Biology 33.1 66.3 22 65.9 30.3 28.3 36.4 32 28 27.7 31.4
Business and Management 33 43.3 25 47.8 29 28 40 31 21 37 28.3
Chemistry 33.8 57.7 23.3 57.3 24.8 28 40.1 31.7 30 26.7 26.9
Computer Science 36.5 39.8 21 41.4 30.8 28 36.1 26 27 37.2 40.4
Defense and Security 37 54 23 51.2 29 22 43 36 30 33 24
Earth Sciences 39 59.6 29 55.5 25 33 43 24.2 23 25 37
Economics 33.8 54.4 24 51.8 28 23 41.5 32 24 25.4 31.1
Education 34 46.2 20 38.8 31 29 40.8 31 24 28 28
Energy and Power 28 56.8 29 50.2 28 28 46 30 29 31 31
Engineering 25 35.6 26.6 36 25.2 25.2 34.6 25.9 21 28 25.8
Environmental Science 32 51.3 23 58.1 28.1 24 45.3 16.2 26 31 25.7
Ethics and Human Rights 34 43 23 54 25 26 46 19 25 30 28
Finance and Investment 26.7 39.6 27 45 30.9 37 34.5 24 32 31.8 29.3
Food Science 32 61 24 52.5 29 27 53 27 25 29 20
Geography 24 52 31 53.4 26.9 32.3 39.1 30 26 23.4 26.6
Health and Medicine 39 63.1 25 69.3 21 22 40.2 34 29 33.6 29
History 32.3 51 25 60.7 27.5 29 45.7 33 22 27.6 23.3
Information Technology 44 60 32 56.4 37 33 53 32 31 45 39
International Relations 30 57.5 30 54.5 22 31 54 25 31 25 33
Language Studies 32 59.8 28 40.3 23.3 29 29.5 21 29 29.2 27
Law and Ethics 43.2 56.8 27 60.8 29.9 23 36.2 27 22 23.6 26.2
Literature and Linguistics 29.7 48.3 15 49.5 34.2 22 39.8 23 23 27.5 23.9
Logical Reasoning 27 34.6 26 30.7 27 21 29.8 22 30 28.7 28
Materials Science 33 48.3 28 44.8 26 29 38 23 20 29 27
Mathematics 30.4 28.2 25.6 29.3 29.2 28.4 28 24.8 25 30 27.4
Media and Communication 42 59 21 53.7 43 30 47 37 25 37 30
Music and Performing Arts 28 47 25 51 26 24 43 33 28 20 29
Physics 32.2 55.2 29.5 45.8 26.2 24.7 36 24.6 28 25 29
Politics and Governance 25.9 50.9 28 54.4 20.6 28 37.9 26 24 28.2 25.2
Psychology 30 69 21 56.6 28 23 33 33 36 31 38
Public Administration 32 48.5 23.2 50.5 29.3 33 23.2 27 25.2 24 24
Religion and Spirituality 43 54 31 58.5 27 25 51 26 27 26 33
Social Welfare and Development 36 45.5 27 53.9 40 24 45 24 24 25 32
Sociology 30.4 42.2 26 40.8 22 28 29.6 35 29 17.1 30
Sports and Recreation 26.7 57 26 58.6 27 23 40.1 18 22 17.9 32.9
Technology and Innovation 35 51.2 29 51.8 21 32 53 27 27 22 27
Transportation and Logistics 32 40.8 28 40.4 22.2 29 29.6 29.3 16 26 26
Table 39: Detailed subject-wise evaluation for nvidia/Nemotron-4-Mini-Hindi-4B-Base on MILU across different languages. The results reported are for 5-shot experiments.
Topic bn en gu hi kn ml mr or pa ta te
Agriculture 29.0/ 22.0/ 30.0 35.8/ 46.1/ 46.5 21.2/ 23.2/ 22.2 26.2/ 27.4/ 29.0 19.0/ 26.0/ 23.0 28.3/ 27.3/ 28.3 28.0/ 26.0/ 27.0 27.0/ 23.0/ 24.0 26.0/ 30.0/ 32.0 29.0/ 21.0/ 22.0 26.0/ 25.0/ 25.0
Anthropology 33.0/ 21.0/ 25.0 46.0/ 50.0/ 48.0 28.0/ 30.0/ 26.0 25.2/ 24.2/ 27.3 23.0/ 22.0/ 23.0 20.0/ 20.0/ 21.0 28.0/ 32.0/ 25.0 24.0/ 26.0/ 23.0 28.0/ 24.0/ 23.0 25.0/ 25.0/ 27.0 24.0/ 27.0/ 24.0
Architecture and Design 34.0/ 30.0/ 26.0 37.0/ 53.0/ 44.0 27.0/ 23.0/ 25.0 27.0/ 34.0/ 35.0 19.0/ 14.0/ 17.0 27.0/ 25.0/ 27.0 26.0/ 24.0/ 24.0 27.0/ 26.0/ 19.0 21.0/ 24.0/ 24.0 19.0/ 23.0/ 26.0 19.0/ 16.0/ 16.0
Arts and Culture 26.7/ 24.9/ 21.8 48.4/ 58.8/ 57.4 24.0/ 27.0/ 28.0 25.7/ 27.7/ 28.1 23.9/ 27.0/ 27.0 24.0/ 28.0/ 29.0 22.9/ 22.4/ 22.4 15.0/ 14.0/ 23.0 19.0/ 26.0/ 23.0 26.1/ 25.4/ 26.1 23.1/ 24.1/ 20.6
Astronomy and Astrophysics 16.0/ 28.0/ 21.0 53.0/ 69.0/ 66.0 24.0/ 26.0/ 23.0 25.0/ 28.1/ 30.5 30.0/ 23.0/ 27.0 22.0/ 25.0/ 19.0 32.0/ 28.0/ 29.0 17.0/ 17.0/ 17.0 32.0/ 24.0/ 26.0 27.0/ 26.0/ 28.0 21.0/ 28.0/ 25.0
Biology 23.6/ 23.9/ 22.9 49.5/ 62.7/ 62.5 24.0/ 25.0/ 22.0 27.5/ 26.7/ 24.4 27.9/ 23.8/ 23.8 29.3/ 29.3/ 31.3 26.3/ 28.7/ 29.5 24.0/ 24.0/ 32.0 31.0/ 22.0/ 27.0 27.4/ 28.0/ 25.5 18.9/ 20.1/ 22.6
Business and Management 25.0/ 23.0/ 29.0 37.1/ 48.6/ 45.5 28.0/ 25.0/ 22.0 28.3/ 26.7/ 30.4 27.0/ 19.0/ 18.0 21.0/ 21.0/ 24.0 20.0/ 24.0/ 27.0 24.0/ 27.0/ 25.0 26.0/ 31.0/ 29.0 30.0/ 28.0/ 29.0 21.7/ 20.8/ 17.0
Chemistry 28.8/ 31.0/ 31.3 44.9/ 51.1/ 51.1 26.7/ 22.4/ 23.3 28.3/ 27.3/ 29.1 30.1/ 30.7/ 30.4 24.0/ 26.0/ 25.0 29.8/ 30.9/ 29.5 22.1/ 24.0/ 26.9 24.0/ 21.0/ 20.0 25.5/ 26.4/ 29.1 28.9/ 28.7/ 29.7
Computer Science 30.2/ 35.9/ 33.9 37.0/ 41.2/ 40.4 36.0/ 37.0/ 33.0 28.2/ 32.0/ 28.6 31.4/ 40.8/ 40.2 31.0/ 28.0/ 29.0 33.1/ 32.5/ 31.4 31.0/ 29.0/ 25.0 30.0/ 21.0/ 24.0 36.0/ 36.0/ 36.6 33.2/ 32.3/ 31.8
Defense and Security 25.0/ 28.0/ 25.0 49.0/ 57.0/ 58.0 25.0/ 29.0/ 21.0 27.5/ 24.6/ 23.7 26.0/ 27.0/ 22.0 30.0/ 27.0/ 30.0 26.0/ 27.0/ 27.0 27.0/ 35.0/ 29.0 27.0/ 30.0/ 27.0 34.0/ 22.0/ 23.0 29.0/ 30.0/ 24.0
Earth Sciences 35.0/ 27.0/ 43.0 43.6/ 54.1/ 54.3 22.0/ 19.0/ 25.0 25.7/ 27.1/ 27.1 24.0/ 21.0/ 19.0 19.0/ 21.0/ 23.0 24.0/ 26.0/ 24.0 18.2/ 24.2/ 28.3 25.0/ 21.0/ 26.0 23.0/ 26.0/ 29.0 32.0/ 32.0/ 24.0
Economics 25.2/ 28.1/ 26.6 41.4/ 51.0/ 50.2 23.0/ 28.0/ 29.0 29.4/ 31.9/ 33.7 30.0/ 21.0/ 22.0 27.0/ 29.0/ 27.0 29.6/ 29.6/ 30.3 26.0/ 26.0/ 28.0 31.0/ 22.0/ 26.0 24.6/ 28.6/ 28.6 18.6/ 19.8/ 24.6
Education 22.0/ 20.0/ 20.0 35.7/ 44.3/ 41.7 25.0/ 25.0/ 24.0 26.9/ 26.3/ 26.3 35.0/ 24.0/ 27.0 27.0/ 25.0/ 24.0 33.8/ 28.7/ 37.6 29.0/ 30.0/ 28.0 22.0/ 25.0/ 24.0 29.0/ 30.0/ 24.0 34.0/ 20.0/ 27.0
Energy and Power 28.0/ 28.0/ 27.0 33.1/ 51.3/ 53.4 29.0/ 35.0/ 36.0 23.7/ 26.1/ 30.5 30.0/ 24.0/ 28.0 37.0/ 29.0/ 30.0 27.0/ 24.0/ 35.0 25.0/ 23.0/ 25.0 32.0/ 31.0/ 26.0 26.0/ 14.0/ 23.0 23.0/ 21.0/ 25.0
Engineering 26.8/ 28.4/ 27.8 35.2/ 35.8/ 35.8 24.2/ 25.0/ 22.2 25.2/ 24.8/ 28.4 26.0/ 29.4/ 29.2 26.5/ 26.5/ 28.6 25.0/ 27.0/ 28.8 26.2/ 25.0/ 25.0 25.0/ 25.0/ 27.0 26.8/ 29.4/ 27.6 25.8/ 26.0/ 25.8
Environmental Science 24.0/ 27.0/ 34.0 41.3/ 53.5/ 51.3 29.0/ 28.0/ 22.0 27.7/ 26.9/ 30.9 21.9/ 23.4/ 19.5 26.0/ 26.0/ 29.0 18.9/ 25.3/ 26.3 32.3/ 28.3/ 32.3 22.0/ 21.0/ 22.0 22.0/ 23.0/ 31.0 25.1/ 22.9/ 28.5
Ethics and Human Rights 18.0/ 23.0/ 27.0 35.0/ 47.0/ 45.0 23.0/ 27.0/ 26.0 29.0/ 26.0/ 23.0 29.0/ 27.0/ 28.0 21.0/ 21.0/ 21.0 29.0/ 28.0/ 35.0 21.0/ 21.0/ 16.0 23.0/ 27.0/ 27.0 30.0/ 32.0/ 30.0 20.0/ 22.0/ 22.0
Finance and Investment 20.8/ 25.2/ 28.2 36.0/ 42.0/ 38.0 26.0/ 24.0/ 22.0 28.6/ 27.0/ 28.8 24.7/ 25.9/ 21.0 30.0/ 24.0/ 31.0 24.3/ 24.9/ 25.4 22.0/ 23.0/ 21.0 32.0/ 27.0/ 23.0 24.0/ 27.4/ 29.0 25.4/ 23.2/ 21.1
Food Science 37.0/ 25.0/ 29.0 53.0/ 51.0/ 57.0 28.0/ 32.0/ 30.0 21.2/ 28.3/ 27.3 27.0/ 34.0/ 26.0 26.0/ 28.0/ 26.0 26.0/ 30.0/ 24.0 31.0/ 25.0/ 25.0 28.0/ 22.0/ 15.0 21.0/ 26.0/ 27.0 22.0/ 26.0/ 26.0
Geography 25.7/ 30.7/ 27.9 42.0/ 52.2/ 49.4 29.0/ 19.0/ 21.0 27.5/ 28.3/ 26.9 23.7/ 28.8/ 31.5 22.2/ 23.2/ 18.2 24.9/ 30.2/ 29.6 28.0/ 27.0/ 19.0 31.0/ 29.0/ 30.0 23.4/ 30.3/ 28.6 26.6/ 30.2/ 27.1
Health and Medicine 34.0/ 28.0/ 33.0 51.3/ 64.1/ 60.7 25.0/ 29.0/ 29.0 29.5/ 26.7/ 28.7 26.0/ 26.0/ 29.0 21.0/ 25.0/ 15.0 26.8/ 31.2/ 30.4 30.0/ 27.0/ 18.0 26.0/ 28.0/ 24.0 21.6/ 20.7/ 17.2 19.6/ 25.2/ 29.0
History 24.3/ 24.8/ 26.7 37.4/ 45.0/ 42.6 26.0/ 20.0/ 23.0 29.0/ 32.5/ 33.1 27.5/ 24.2/ 22.7 27.0/ 26.0/ 31.0 27.5/ 23.1/ 26.4 31.0/ 31.0/ 29.0 20.0/ 26.0/ 21.0 24.0/ 27.3/ 29.4 28.1/ 24.0/ 26.1
Information Technology 27.0/ 36.0/ 35.0 44.0/ 57.0/ 55.0 29.0/ 30.0/ 27.0 27.9/ 25.7/ 27.9 26.0/ 34.0/ 33.0 26.0/ 29.0/ 30.0 31.0/ 36.0/ 30.0 30.0/ 32.0/ 34.0 33.0/ 37.0/ 36.0 19.0/ 26.0/ 26.0 27.0/ 34.0/ 31.0
International Relations 30.0/ 33.0/ 34.0 58.3/ 61.7/ 61.7 29.0/ 27.0/ 20.0 26.9/ 29.6/ 28.8 28.0/ 22.0/ 23.0 24.0/ 28.0/ 24.0 27.0/ 26.0/ 32.0 27.0/ 26.0/ 28.0 30.0/ 27.0/ 25.0 28.0/ 25.0/ 27.0 18.0/ 16.0/ 14.0
Language Studies 28.0/ 29.0/ 26.0 42.4/ 52.0/ 52.8 16.0/ 28.0/ 28.0 21.6/ 25.2/ 25.2 31.0/ 23.3/ 20.9 26.0/ 21.0/ 23.0 22.8/ 25.5/ 30.9 21.0/ 26.0/ 18.0 39.0/ 33.0/ 38.0 24.3/ 24.3/ 25.1 27.6/ 29.3/ 28.7
Law and Ethics 29.7/ 27.0/ 25.0 45.2/ 55.4/ 53.6 21.0/ 23.0/ 22.0 33.8/ 31.4/ 34.0 22.3/ 19.8/ 21.7 24.0/ 30.0/ 30.0 26.3/ 24.1/ 30.2 22.0/ 25.0/ 25.0 25.0/ 25.0/ 21.0 23.6/ 27.6/ 26.8 26.2/ 18.1/ 20.0
Literature and Linguistics 27.0/ 27.9/ 25.2 40.9/ 49.5/ 45.7 16.0/ 25.0/ 20.0 27.1/ 25.7/ 26.5 26.7/ 25.8/ 25.0 28.0/ 27.0/ 25.0 22.1/ 22.1/ 23.9 33.0/ 38.0/ 35.0 26.0/ 24.0/ 21.0 24.7/ 27.5/ 27.5 23.9/ 27.5/ 21.8
Logical Reasoning 22.6/ 24.6/ 22.9 33.0/ 32.6/ 32.8 19.5/ 27.9/ 26.6 24.1/ 25.7/ 25.9 24.5/ 25.3/ 29.9 32.0/ 21.0/ 25.0 23.7/ 27.6/ 24.2 25.0/ 27.0/ 23.0 15.0/ 28.0/ 22.0 24.7/ 27.5/ 25.9 30.7/ 29.5/ 25.5
Materials Science 24.0/ 17.0/ 22.0 39.9/ 47.9/ 43.7 26.0/ 22.0/ 18.0 28.0/ 24.4/ 26.9 37.0/ 33.0/ 31.0 41.0/ 27.0/ 31.0 31.0/ 29.0/ 29.0 25.0/ 21.0/ 19.0 23.0/ 20.0/ 18.0 29.0/ 22.0/ 27.0 29.0/ 29.0/ 26.0
Mathematics 25.8/ 26.4/ 29.4 25.0/ 27.4/ 27.0 26.2/ 25.2/ 23.0 24.1/ 26.2/ 23.8 22.0/ 27.0/ 26.4 31.2/ 28.4/ 27.8 24.0/ 24.6/ 28.0 24.4/ 24.8/ 27.2 25.0/ 26.9/ 28.0 23.8/ 18.8/ 23.2 25.2/ 27.4/ 24.8
Media and Communication 24.0/ 23.0/ 24.0 37.0/ 56.0/ 55.0 26.0/ 23.0/ 27.0 26.5/ 29.9/ 33.3 27.0/ 24.0/ 23.0 23.0/ 25.0/ 23.0 24.0/ 28.0/ 24.0 29.0/ 28.0/ 23.0 24.0/ 20.0/ 22.0 26.0/ 29.0/ 29.0 30.0/ 31.0/ 34.0
Music and Performing Arts 28.0/ 13.0/ 22.0 43.0/ 50.0/ 51.0 28.0/ 20.0/ 25.0 27.2/ 29.1/ 28.5 21.0/ 23.0/ 24.0 22.0/ 22.0/ 18.0 20.0/ 25.0/ 21.0 31.0/ 28.0/ 33.0 32.0/ 27.0/ 22.0 23.0/ 27.0/ 23.0 20.0/ 24.0/ 26.0
Physics 21.2/ 25.6/ 24.6 36.2/ 47.0/ 44.4 22.0/ 28.1/ 27.3 25.6/ 25.2/ 26.8 22.8/ 27.0/ 24.4 22.6/ 24.7/ 23.2 23.6/ 26.0/ 24.0 20.1/ 25.0/ 27.4 22.0/ 33.0/ 33.0 20.6/ 23.4/ 22.2 25.6/ 27.2/ 24.2
Politics and Governance 25.9/ 27.5/ 27.1 39.1/ 48.9/ 44.3 26.0/ 26.0/ 21.0 25.5/ 27.5/ 26.1 26.7/ 19.0/ 23.1 19.0/ 23.0/ 23.0 21.6/ 20.8/ 23.9 24.0/ 19.0/ 20.0 27.0/ 31.0/ 30.0 31.1/ 22.8/ 26.1 25.2/ 25.5/ 23.1
Psychology 27.0/ 26.0/ 24.0 47.0/ 52.0/ 56.0 27.0/ 25.0/ 26.0 28.7/ 23.3/ 21.7 26.0/ 30.0/ 28.0 26.0/ 23.0/ 28.0 35.0/ 23.0/ 26.0 33.0/ 28.0/ 24.0 29.0/ 30.0/ 34.0 25.0/ 21.0/ 21.0 28.0/ 26.0/ 25.0
Public Administration 31.0/ 25.0/ 21.0 38.4/ 48.5/ 48.5 25.2/ 26.3/ 23.2 33.3/ 33.3/ 28.3 37.4/ 30.3/ 26.3 29.0/ 27.0/ 29.0 28.3/ 19.2/ 24.2 29.0/ 25.0/ 31.0 23.2/ 26.3/ 29.3 28.0/ 25.0/ 29.0 32.0/ 28.0/ 31.0
Religion and Spirituality 15.0/ 24.0/ 29.0 44.0/ 63.0/ 57.0 34.0/ 30.0/ 32.0 23.8/ 28.6/ 25.0 17.0/ 33.0/ 23.0 28.0/ 27.0/ 26.0 16.0/ 22.0/ 26.0 28.0/ 27.0/ 29.0 28.0/ 30.0/ 25.0 22.0/ 24.0/ 26.0 18.0/ 24.0/ 31.0
Social Welfare and Development 33.0/ 29.0/ 28.0 38.4/ 47.3/ 42.9 26.0/ 20.0/ 26.0 25.9/ 23.3/ 29.0 22.0/ 32.0/ 28.0 23.0/ 25.0/ 26.0 34.0/ 30.0/ 33.0 25.0/ 26.0/ 24.0 27.0/ 34.0/ 29.0 25.0/ 33.0/ 28.0 32.0/ 30.0/ 32.0
Sociology 21.7/ 21.7/ 20.0 33.6/ 39.6/ 37.2 29.0/ 20.0/ 23.0 25.2/ 28.8/ 29.0 16.0/ 22.0/ 22.0 26.0/ 29.0/ 26.0 30.2/ 25.8/ 25.8 19.0/ 23.0/ 28.0 19.0/ 24.0/ 30.0 22.5/ 28.8/ 26.1 24.7/ 27.7/ 30.0
Sports and Recreation 23.8/ 29.2/ 27.2 43.0/ 58.8/ 56.8 20.0/ 20.0/ 17.0 26.0/ 28.4/ 26.6 25.3/ 20.8/ 24.7 23.0/ 29.0/ 29.0 23.7/ 29.9/ 26.0 18.0/ 25.0/ 21.0 21.0/ 26.0/ 25.0 25.6/ 26.9/ 28.2 23.8/ 22.9/ 22.4
Technology and Innovation 32.0/ 26.0/ 21.0 43.2/ 52.2/ 52.0 27.0/ 26.0/ 26.0 29.4/ 25.4/ 29.6 30.0/ 27.0/ 31.0 25.0/ 29.0/ 32.0 30.0/ 27.0/ 28.0 24.0/ 23.0/ 34.0 25.0/ 24.0/ 27.0 32.0/ 28.0/ 26.0 19.0/ 19.0/ 17.0
Transportation and Logistics 22.0/ 30.0/ 20.0 32.3/ 38.5/ 34.6 24.0/ 17.0/ 25.0 25.9/ 26.5/ 25.2 24.2/ 21.2/ 24.2 30.0/ 29.0/ 27.0 23.5/ 21.4/ 20.4 25.2/ 20.2/ 29.3 23.0/ 29.0/ 27.0 23.0/ 27.0/ 31.0 24.0/ 25.0/ 29.0
Table 40: Detailed subject-wise evaluation for meta-llama/Llama-2-7b-hf on MILU across different languages. The results reported are X / Y / Z where X denote the 0-shot, Y denotes 1-shot and Z denotes 5-shot performance respectively.
Topic bn en gu hi kn ml mr or pa ta te
Agriculture 21.0/ 32.0/ 23.0 28.0/ 28.0/ 27.4 29.3/ 31.3/ 26.3 28.2/ 27.6/ 26.8 24.0/ 27.0/ 21.0 28.3/ 32.3/ 36.4 31.0/ 35.0/ 27.0 24.0/ 26.0/ 29.0 20.0/ 29.0/ 25.0 28.0/ 29.0/ 25.0 24.0/ 25.0/ 23.0
Anthropology 29.0/ 29.0/ 29.0 28.0/ 30.0/ 27.0 25.0/ 30.0/ 17.0 34.3/ 24.2/ 25.2 31.0/ 27.0/ 26.0 26.0/ 22.0/ 24.0 23.0/ 24.0/ 28.0 20.0/ 27.0/ 31.0 23.0/ 21.0/ 29.0 31.0/ 26.0/ 33.0 30.0/ 18.0/ 20.0
Architecture and Design 33.0/ 31.0/ 26.0 38.0/ 24.0/ 25.0 27.0/ 28.0/ 18.0 31.0/ 25.0/ 24.0 22.0/ 24.0/ 22.0 31.0/ 33.0/ 28.0 29.0/ 23.0/ 19.0 24.0/ 27.0/ 17.0 24.0/ 18.0/ 26.0 20.0/ 19.0/ 22.0 22.0/ 20.0/ 19.0
Arts and Culture 29.1/ 24.9/ 24.9 31.8/ 30.6/ 30.0 20.0/ 24.0/ 24.0 22.9/ 21.6/ 23.1 25.8/ 25.1/ 19.0 13.0/ 21.0/ 22.0 25.7/ 23.5/ 20.2 26.0/ 22.0/ 22.0 24.0/ 24.0/ 24.0 26.7/ 24.9/ 29.1 21.1/ 23.1/ 29.1
Astronomy and Astrophysics 29.0/ 24.0/ 25.0 37.0/ 30.0/ 35.0 25.0/ 24.0/ 30.0 21.1/ 28.1/ 26.6 22.0/ 25.0/ 27.0 27.0/ 24.0/ 32.0 26.0/ 34.0/ 29.0 20.0/ 21.0/ 24.0 21.0/ 30.0/ 27.0 28.0/ 29.0/ 34.0 28.0/ 31.0/ 25.0
Biology 21.1/ 24.3/ 20.1 33.5/ 26.7/ 29.5 26.0/ 26.0/ 24.0 23.1/ 25.4/ 22.9 27.9/ 27.9/ 28.3 28.3/ 24.2/ 26.3 25.4/ 23.3/ 26.3 26.0/ 30.0/ 21.0 25.0/ 19.0/ 26.0 29.3/ 28.3/ 29.6 20.4/ 23.3/ 24.8
Business and Management 22.0/ 28.0/ 19.0 28.5/ 27.6/ 27.9 26.0/ 24.0/ 23.0 23.6/ 25.8/ 23.6 25.0/ 23.0/ 25.0 24.0/ 21.0/ 28.0 24.0/ 27.0/ 27.0 33.0/ 30.0/ 27.0 29.0/ 26.0/ 22.0 34.0/ 30.0/ 26.0 25.5/ 15.1/ 23.6
Chemistry 26.9/ 28.8/ 26.7 36.9/ 30.5/ 29.1 23.3/ 22.4/ 24.1 25.7/ 26.1/ 23.2 23.2/ 23.2/ 26.5 21.0/ 19.0/ 16.0 23.5/ 24.3/ 23.0 20.2/ 25.0/ 24.0 20.0/ 20.0/ 17.0 22.9/ 21.1/ 21.4 26.9/ 27.7/ 25.9
Computer Science 26.0/ 25.0/ 29.2 27.2/ 28.2/ 24.8 28.0/ 26.0/ 31.0 27.6/ 25.4/ 28.2 26.0/ 29.6/ 29.6 29.0/ 23.0/ 23.0 29.0/ 25.4/ 34.3 24.0/ 32.0/ 26.0 24.0/ 25.0/ 25.0 27.3/ 29.1/ 31.4 26.0/ 28.2/ 30.0
Defense and Security 24.0/ 24.0/ 19.0 34.0/ 30.0/ 25.0 33.0/ 34.0/ 25.0 27.5/ 28.9/ 29.9 29.0/ 25.0/ 29.0 28.0/ 26.0/ 22.0 26.0/ 27.0/ 32.0 17.0/ 20.0/ 19.0 37.0/ 26.0/ 34.0 24.0/ 28.0/ 20.0 28.0/ 29.0/ 23.0
Earth Sciences 37.0/ 31.0/ 33.0 29.8/ 29.1/ 25.9 25.0/ 26.0/ 23.0 29.5/ 28.1/ 25.9 33.0/ 24.0/ 22.0 25.0/ 22.0/ 22.0 32.0/ 34.0/ 30.0 30.3/ 28.3/ 25.2 27.0/ 26.0/ 31.0 23.0/ 25.0/ 29.0 24.0/ 24.0/ 16.0
Economics 18.0/ 23.7/ 28.1 32.5/ 29.5/ 30.5 24.0/ 21.0/ 19.0 24.2/ 24.6/ 24.0 30.0/ 26.0/ 27.0 27.0/ 23.0/ 22.0 29.6/ 23.9/ 24.6 25.0/ 18.0/ 33.0 19.0/ 20.0/ 18.0 20.6/ 19.1/ 22.2 25.1/ 27.5/ 26.4
Education 27.0/ 20.0/ 24.0 26.8/ 29.3/ 29.9 25.0/ 21.0/ 28.0 28.7/ 28.4/ 25.7 39.0/ 33.0/ 27.0 25.0/ 20.0/ 26.0 22.9/ 26.1/ 23.6 27.0/ 26.0/ 23.0 24.0/ 24.0/ 30.0 25.0/ 22.0/ 26.0 29.0/ 33.0/ 26.0
Energy and Power 33.0/ 26.0/ 29.0 20.9/ 30.4/ 20.3 27.0/ 28.0/ 33.0 24.1/ 28.8/ 24.8 24.0/ 20.0/ 18.0 29.0/ 24.0/ 23.0 28.0/ 19.0/ 21.0 22.0/ 23.0/ 24.0 24.0/ 23.0/ 25.0 27.0/ 24.0/ 16.0 26.0/ 28.0/ 25.0
Engineering 26.4/ 26.0/ 25.8 28.2/ 27.6/ 28.2 22.4/ 25.0/ 21.4 23.2/ 25.2/ 24.6 27.2/ 26.8/ 26.6 29.1/ 26.5/ 27.8 27.8/ 26.4/ 29.0 27.1/ 22.0/ 22.3 26.0/ 27.0/ 23.0 26.2/ 23.8/ 28.8 26.4/ 26.4/ 28.2
Environmental Science 18.0/ 22.0/ 22.0 30.1/ 29.5/ 28.5 29.0/ 27.0/ 24.0 25.4/ 23.6/ 24.9 25.0/ 24.2/ 28.1 24.0/ 23.0/ 21.0 19.5/ 25.8/ 24.2 21.2/ 18.2/ 25.2 31.0/ 27.0/ 35.0 28.0/ 28.0/ 33.0 24.0/ 25.1/ 26.8
Ethics and Human Rights 26.0/ 28.0/ 32.0 34.0/ 31.0/ 31.0 15.0/ 24.0/ 15.0 21.0/ 27.0/ 29.0 23.0/ 27.0/ 19.0 20.0/ 18.0/ 25.0 33.0/ 28.0/ 29.0 27.0/ 27.0/ 24.0 18.0/ 26.0/ 24.0 20.0/ 18.0/ 27.0 19.0/ 16.0/ 9.0
Finance and Investment 24.8/ 23.8/ 26.7 31.2/ 30.0/ 30.2 28.0/ 23.0/ 30.0 23.6/ 26.0/ 24.4 29.0/ 26.5/ 23.5 27.0/ 32.0/ 26.0 26.6/ 24.3/ 22.0 23.0/ 28.0/ 33.0 34.0/ 27.0/ 27.0 27.9/ 32.4/ 28.5 25.0/ 26.8/ 26.1
Food Science 32.0/ 29.0/ 31.0 31.0/ 32.0/ 34.0 23.0/ 24.0/ 23.0 27.3/ 29.3/ 26.3 22.0/ 19.0/ 17.0 28.0/ 22.0/ 21.0 22.0/ 29.0/ 28.0 24.0/ 20.0/ 21.0 22.0/ 22.0/ 24.0 24.0/ 18.0/ 19.0 20.0/ 30.0/ 21.0
Geography 29.0/ 25.7/ 24.6 29.0/ 29.6/ 26.0 35.0/ 33.0/ 37.0 24.5/ 28.3/ 23.3 18.7/ 22.4/ 24.2 16.2/ 16.2/ 27.3 22.5/ 24.9/ 23.1 24.0/ 21.0/ 23.0 30.0/ 21.0/ 18.0 28.0/ 24.6/ 27.4 24.5/ 24.5/ 25.0
Health and Medicine 24.0/ 28.0/ 22.0 35.1/ 29.9/ 30.5 22.0/ 19.0/ 21.0 23.6/ 24.6/ 24.1 21.0/ 24.0/ 28.0 31.0/ 30.0/ 24.0 22.3/ 24.1/ 17.9 26.0/ 19.0/ 20.0 23.0/ 27.0/ 21.0 20.7/ 19.8/ 23.3 27.1/ 29.9/ 30.8
History 30.7/ 30.1/ 24.5 31.5/ 30.9/ 33.1 33.0/ 27.0/ 18.0 27.0/ 24.8/ 26.8 24.5/ 20.9/ 27.5 20.0/ 18.0/ 22.0 26.0/ 20.8/ 23.4 31.0/ 32.0/ 32.0 25.0/ 22.0/ 24.0 25.8/ 28.0/ 21.1 26.1/ 25.6/ 23.8
Information Technology 30.0/ 25.0/ 33.0 29.0/ 32.0/ 34.0 31.0/ 27.0/ 26.0 26.3/ 28.5/ 31.8 30.0/ 28.0/ 29.0 20.0/ 23.0/ 20.0 17.0/ 19.0/ 23.0 26.0/ 28.0/ 24.0 27.0/ 24.0/ 30.0 18.0/ 23.0/ 19.0 29.0/ 28.0/ 25.0
International Relations 22.0/ 25.0/ 25.0 38.3/ 30.8/ 38.3 32.0/ 27.0/ 30.0 26.1/ 23.4/ 23.0 18.0/ 24.0/ 20.0 22.0/ 23.0/ 22.0 20.0/ 23.0/ 25.0 22.0/ 30.0/ 28.0 33.0/ 30.0/ 31.0 24.0/ 22.0/ 23.0 20.0/ 22.0/ 18.0
Language Studies 28.0/ 33.0/ 33.0 33.0/ 32.8/ 33.4 21.0/ 24.0/ 21.0 23.4/ 26.9/ 24.1 27.1/ 31.8/ 32.6 25.0/ 21.0/ 32.0 26.9/ 24.8/ 24.8 26.0/ 26.0/ 26.0 30.0/ 36.0/ 32.0 23.6/ 26.6/ 25.5 28.7/ 29.9/ 31.0
Law and Ethics 24.3/ 29.7/ 20.9 37.0/ 36.2/ 33.4 24.0/ 25.0/ 28.0 29.6/ 27.6/ 27.0 24.8/ 23.6/ 22.3 25.0/ 29.0/ 26.0 28.4/ 23.7/ 29.3 17.0/ 19.0/ 21.0 21.0/ 20.0/ 20.0 30.7/ 29.1/ 31.5 26.9/ 23.8/ 28.7
Literature and Linguistics 26.1/ 27.0/ 24.3 28.7/ 25.7/ 26.5 26.0/ 23.0/ 29.0 27.1/ 29.3/ 28.1 22.5/ 27.5/ 22.5 26.0/ 27.0/ 28.0 20.3/ 23.0/ 27.4 23.0/ 30.0/ 29.0 28.0/ 24.0/ 21.0 28.3/ 23.5/ 22.3 26.1/ 23.9/ 31.0
Logical Reasoning 25.3/ 24.1/ 23.3 26.6/ 24.8/ 26.6 26.0/ 29.9/ 31.8 28.3/ 27.9/ 26.7 29.5/ 27.4/ 29.5 30.0/ 32.0/ 31.0 25.1/ 25.9/ 26.7 30.0/ 31.0/ 28.0 25.0/ 23.0/ 23.0 27.5/ 27.5/ 25.5 29.5/ 29.0/ 29.5
Materials Science 29.0/ 18.0/ 16.0 29.3/ 28.9/ 28.5 26.0/ 24.0/ 27.0 24.2/ 26.9/ 25.3 33.0/ 25.0/ 29.0 31.0/ 26.0/ 38.0 29.0/ 22.0/ 22.0 25.0/ 31.0/ 35.0 28.0/ 25.0/ 24.0 22.0/ 19.0/ 18.0 27.0/ 27.0/ 28.0
Mathematics 27.8/ 27.6/ 26.8 25.4/ 26.0/ 28.8 21.0/ 24.2/ 26.8 25.9/ 26.2/ 22.2 21.0/ 26.6/ 25.4 22.1/ 22.4/ 19.8 23.0/ 23.0/ 24.2 24.0/ 26.6/ 27.2 27.7/ 27.7/ 26.5 25.0/ 22.4/ 24.8 22.6/ 25.2/ 20.6
Media and Communication 27.0/ 28.0/ 22.0 28.0/ 28.0/ 31.0 31.0/ 23.0/ 28.0 31.3/ 29.9/ 28.6 24.0/ 24.0/ 20.0 31.0/ 29.0/ 24.0 27.0/ 25.0/ 27.0 29.0/ 30.0/ 37.0 27.0/ 25.0/ 19.0 17.0/ 18.0/ 14.0 29.0/ 38.0/ 26.0
Music and Performing Arts 27.0/ 19.0/ 22.0 36.0/ 36.0/ 36.0 29.0/ 21.0/ 25.0 21.2/ 21.9/ 27.2 20.0/ 27.0/ 22.0 25.0/ 24.0/ 30.0 23.0/ 24.0/ 28.0 33.0/ 30.0/ 27.0 25.0/ 31.0/ 29.0 27.0/ 28.0/ 27.0 25.0/ 19.0/ 28.0
Physics 23.6/ 25.8/ 23.4 30.4/ 29.4/ 28.0 24.2/ 23.1/ 21.2 23.0/ 23.8/ 25.6 26.6/ 26.4/ 24.0 26.8/ 24.2/ 25.8 25.2/ 27.4/ 24.8 29.2/ 26.0/ 29.2 27.0/ 33.0/ 23.0 23.4/ 24.0/ 20.2 26.4/ 25.6/ 24.0
Politics and Governance 27.5/ 28.6/ 27.8 34.1/ 35.7/ 31.7 24.0/ 24.0/ 21.0 24.1/ 25.3/ 25.7 24.7/ 23.1/ 25.1 20.0/ 22.0/ 28.0 23.0/ 26.1/ 24.4 30.0/ 31.0/ 33.0 23.0/ 28.0/ 28.0 28.2/ 28.6/ 29.0 25.9/ 25.5/ 22.4
Psychology 28.0/ 34.0/ 28.0 35.0/ 27.0/ 30.0 23.0/ 31.0/ 25.0 26.4/ 20.2/ 18.6 29.0/ 22.0/ 27.0 30.0/ 23.0/ 26.0 25.0/ 21.0/ 23.0 27.0/ 29.0/ 25.0 25.0/ 22.0/ 29.0 32.0/ 23.0/ 30.0 26.0/ 21.0/ 22.0
Public Administration 25.0/ 24.0/ 24.0 31.3/ 29.3/ 31.3 17.2/ 21.2/ 23.2 31.3/ 29.3/ 30.3 29.3/ 32.3/ 28.3 24.0/ 25.0/ 30.0 25.2/ 20.2/ 19.2 27.0/ 28.0/ 30.0 29.3/ 28.3/ 30.3 24.0/ 20.0/ 23.0 26.0/ 26.0/ 31.0
Religion and Spirituality 16.0/ 13.0/ 20.0 35.0/ 32.0/ 28.0 31.0/ 24.0/ 24.0 21.0/ 23.8/ 25.8 20.0/ 22.0/ 20.0 24.0/ 18.0/ 16.0 19.0/ 26.0/ 27.0 33.0/ 27.0/ 29.0 29.0/ 29.0/ 24.0 20.0/ 24.0/ 25.0 21.0/ 20.0/ 27.0
Social Welfare and Development 16.0/ 22.0/ 16.0 33.0/ 32.1/ 33.9 27.0/ 29.0/ 23.0 25.4/ 22.3/ 24.9 25.0/ 26.0/ 25.0 24.0/ 26.0/ 26.0 23.0/ 28.0/ 27.0 22.0/ 22.0/ 23.0 23.0/ 26.0/ 27.0 26.0/ 23.0/ 28.0 33.0/ 30.0/ 27.0
Sociology 20.0/ 27.0/ 23.5 31.0/ 26.0/ 28.8 26.0/ 27.0/ 28.0 25.4/ 20.8/ 22.4 23.0/ 28.0/ 25.0 22.0/ 23.0/ 22.0 22.0/ 18.9/ 20.8 24.0/ 22.0/ 19.0 28.0/ 21.0/ 25.0 23.4/ 24.3/ 32.4 28.8/ 32.9/ 23.5
Sports and Recreation 23.8/ 28.7/ 28.2 33.2/ 27.2/ 27.0 28.0/ 27.0/ 25.0 25.2/ 23.4/ 25.4 23.0/ 20.2/ 22.5 38.0/ 25.0/ 26.0 26.0/ 27.7/ 26.6 17.0/ 16.0/ 17.0 21.0/ 28.0/ 26.0 24.4/ 22.4/ 19.2 18.6/ 22.4/ 24.8
Technology and Innovation 34.0/ 31.0/ 25.0 30.0/ 30.2/ 31.6 18.0/ 19.0/ 23.0 26.2/ 27.6/ 24.4 22.0/ 21.0/ 21.0 23.0/ 29.0/ 23.0 24.0/ 30.0/ 27.0 24.0/ 24.0/ 26.0 26.0/ 25.0/ 26.0 24.0/ 36.0/ 24.0 27.0/ 25.0/ 30.0
Transportation and Logistics 20.0/ 24.0/ 26.0 30.0/ 26.9/ 26.9 28.0/ 23.0/ 24.0 23.0/ 27.8/ 24.6 25.2/ 23.2/ 27.3 23.0/ 19.0/ 18.0 30.6/ 26.5/ 26.5 25.2/ 24.2/ 23.2 29.0/ 28.0/ 24.0 29.0/ 25.0/ 22.0 23.0/ 20.0/ 22.0
Table 41: Detailed subject-wise evaluation for meta-llama/Llama-2-7b-chat-hf on MILU across different languages. The results reported are X / Y / Z where X denote the 0-shot, Y denotes 1-shot and Z denotes 5-shot performance respectively.
Topic bn en gu hi kn ml mr or pa ta te
Agriculture 32 48.5 37.4 37.6 35 33.3 31 35 39 32 32
Anthropology 43 50 35 36.4 26 32 31 27 39 30 27
Architecture and Design 48 52 36 34 30 39 32 25 31 27 35
Arts and Culture 36.4 55.2 33 37.1 33.7 25 34.4 22 36 35.8 30.6
Astronomy and Astrophysics 56 82 46 46.1 36 38 44 39 38 42 42
Biology 46.5 71.3 44 55.1 36.5 36.4 35.5 26 39 28 42.1
Business and Management 33 52.6 40 43.1 37 38 35 31 37 39 20.8
Chemistry 46.4 69.3 45.7 53.7 33 34 40.4 35.6 39 30.9 40.9
Computer Science 39.1 45.6 44 37.6 42 33 42.6 35 38 29.6 34.5
Defense and Security 46 65 32 37.4 36 31 29 29 39 25 37
Earth Sciences 45 67.7 33 41.1 30 38 39 25.2 30 32 41
Economics 49.6 58.8 33 40.9 36 32 39.4 28 32 29.4 39.5
Education 43 45.5 36 39.1 33 36 43.3 40 36 29 36
Energy and Power 48 56.1 42 43.4 28 34 41 34 35 26 40
Engineering 41.2 49.6 35.8 37.4 34.4 30.8 38.4 32.4 33 32.8 40.6
Environmental Science 43 61.7 40 43.5 35.9 34 35.8 38.4 42 34 37.4
Ethics and Human Rights 48 59 38 49 37 38 42 31 39 43 45
Finance and Investment 44.5 50.2 39 41.6 40.7 30 37.9 32 39 41.9 36.8
Food Science 48 76 46 53.5 36 41 49 35 48 39 53
Geography 36.9 50.4 40 41.2 32.4 23.2 34.3 26 30 30.3 28.6
Health and Medicine 49 71.7 39 50.5 41 41 42.9 35 47 35.3 44.9
History 38.7 46.2 40 41.9 37 35 32.3 35 34 26.9 32.7
Information Technology 71 64 53 64.2 45 60 58 47 44 45 54
International Relations 45 60.8 42 47.9 36 47 44 47 43 48 31
Language Studies 33 71.8 38 36.1 33.3 34 30.9 38 37 30.7 35.1
Law and Ethics 46.6 53.4 34 43.2 24.8 24 31 37 38 33.1 30.6
Literature and Linguistics 34.2 51.7 27 37.7 30.8 32 30.1 37 36 30 19
Logical Reasoning 38.3 40.8 22.1 35.1 26.1 27 27 25 34 30.7 31.9
Materials Science 40 58.6 26 43 34 36 45 28 31 36 46
Mathematics 43.2 41.4 40 37.3 38.6 44.1 40.6 39.4 34.8 38 34.2
Media and Communication 55 73 44 49 36 41 35 42 41 43 50
Music and Performing Arts 31 47 28 32.5 32 29 36 29 32 20 29
Physics 43.6 63.2 36.2 45.4 35.2 37.4 40 33.7 39 34.4 42.6
Politics and Governance 39.6 52.3 30 37 35.6 22 32 28 35 28.6 31.8
Psychology 45 71 42 46.5 35 40 42 32 34 38 51
Public Administration 31 54.5 32.3 36.4 27.3 31 30.3 26 40.4 40 27
Religion and Spirituality 39 53 45 41.1 33 29 43 25 33 25 36
Social Welfare and Development 42 53.6 34 40.9 35 33 46 31 32 29 28
Sociology 38.3 41.2 28 36.4 23 20 23.3 24 32 41.4 28.2
Sports and Recreation 30.7 57.4 25 41.2 28.1 25 35.6 31 30 26.9 28.1
Technology and Innovation 50 60 43 49.6 40 43 48 38 39 35 45
Transportation and Logistics 28 49.2 37 37.2 24.2 26 29.6 20.2 24 28 34
Table 42: Detailed subject-wise evaluation for neulab/Pangea-7B on MILU across different languages. The results reported are for 5-shot experiments.
Topic bn en gu hi kn ml mr or pa ta te
Agriculture 26.0/ 24.0/ 28.0 41.0/ 43.5/ 41.4 19.2/ 22.2/ 23.2 37.0/ 36.2/ 36.4 31.0/ 32.0/ 34.0 34.3/ 30.3/ 32.3 23.0/ 27.0/ 27.0 29.0/ 30.0/ 32.0 25.0/ 29.0/ 32.0 29.0/ 29.0/ 33.0 28.0/ 27.0/ 27.0
Anthropology 37.0/ 32.0/ 34.0 49.0/ 45.0/ 43.0 27.0/ 32.0/ 37.0 36.4/ 41.4/ 40.4 39.0/ 30.0/ 34.0 32.0/ 31.0/ 33.0 25.0/ 31.0/ 30.0 31.0/ 31.0/ 33.0 27.0/ 38.0/ 32.0 29.0/ 27.0/ 29.0 30.0/ 29.0/ 24.0
Architecture and Design 31.0/ 30.0/ 35.0 46.0/ 52.0/ 50.0 26.0/ 23.0/ 21.0 33.0/ 29.0/ 31.0 27.0/ 23.0/ 32.0 25.0/ 25.0/ 27.0 33.0/ 31.0/ 30.0 19.0/ 28.0/ 29.0 29.0/ 24.0/ 26.0 26.0/ 28.0/ 31.0 20.0/ 21.0/ 19.0
Arts and Culture 25.4/ 24.2/ 27.3 46.2/ 48.4/ 46.2 28.0/ 28.0/ 27.0 39.3/ 39.5/ 37.9 28.8/ 28.8/ 26.4 26.0/ 19.0/ 20.0 27.9/ 29.0/ 27.9 27.0/ 30.0/ 34.0 31.0/ 36.0/ 36.0 30.3/ 31.5/ 30.3 24.1/ 27.1/ 29.1
Astronomy and Astrophysics 29.0/ 25.0/ 27.0 63.0/ 62.0/ 62.0 25.0/ 32.0/ 25.0 43.0/ 41.4/ 43.8 27.0/ 26.0/ 28.0 27.0/ 25.0/ 27.0 33.0/ 43.0/ 41.0 33.0/ 28.0/ 30.0 28.0/ 29.0/ 31.0 32.0/ 39.0/ 37.0 28.0/ 26.0/ 25.0
Biology 25.0/ 25.0/ 25.7 52.3/ 53.9/ 54.1 30.0/ 28.0/ 28.0 38.1/ 41.3/ 43.9 27.1/ 24.6/ 25.4 26.3/ 30.3/ 23.2 34.9/ 34.6/ 33.4 24.0/ 18.0/ 20.0 28.0/ 36.0/ 33.0 26.4/ 25.5/ 23.2 24.2/ 24.2/ 24.2
Business and Management 27.0/ 31.0/ 27.0 41.8/ 43.6/ 44.0 25.0/ 34.0/ 29.0 34.0/ 38.2/ 38.2 26.0/ 26.0/ 30.0 37.0/ 30.0/ 30.0 24.0/ 30.0/ 28.0 27.0/ 31.0/ 25.0 27.0/ 23.0/ 18.0 36.0/ 30.0/ 30.0 18.9/ 16.0/ 15.1
Chemistry 25.6/ 25.6/ 23.1 50.3/ 48.9/ 51.5 28.4/ 24.1/ 22.4 36.3/ 39.3/ 38.7 33.7/ 29.7/ 29.4 28.0/ 23.0/ 27.0 28.5/ 29.5/ 31.7 31.7/ 30.8/ 30.8 23.0/ 28.0/ 25.0 25.5/ 27.0/ 24.3 27.4/ 27.2/ 28.7
Computer Science 28.6/ 31.8/ 31.2 40.2/ 36.8/ 37.6 25.0/ 27.0/ 26.0 30.6/ 33.0/ 29.8 26.0/ 28.4/ 24.3 23.0/ 27.0/ 30.0 26.6/ 31.9/ 29.0 34.0/ 31.0/ 32.0 18.0/ 19.0/ 22.0 30.2/ 28.5/ 27.3 25.6/ 28.7/ 24.7
Defense and Security 27.0/ 26.0/ 27.0 47.0/ 40.0/ 39.0 30.0/ 30.0/ 27.0 31.8/ 30.3/ 31.3 27.0/ 31.0/ 31.0 25.0/ 27.0/ 32.0 29.0/ 31.0/ 30.0 26.0/ 34.0/ 31.0 33.0/ 33.0/ 29.0 24.0/ 30.0/ 29.0 25.0/ 24.0/ 16.0
Earth Sciences 35.0/ 32.0/ 32.0 49.5/ 51.1/ 50.0 30.0/ 21.0/ 25.0 34.9/ 38.3/ 38.9 24.0/ 23.0/ 23.0 29.0/ 30.0/ 29.0 27.0/ 33.0/ 33.0 22.2/ 22.2/ 21.2 23.0/ 18.0/ 20.0 23.0/ 26.0/ 25.0 33.0/ 30.0/ 27.0
Economics 30.9/ 25.9/ 23.0 44.4/ 43.0/ 43.4 35.0/ 39.0/ 34.0 36.1/ 35.3/ 38.7 33.0/ 25.0/ 28.0 27.0/ 24.0/ 24.0 29.6/ 38.0/ 31.0 24.0/ 26.0/ 22.0 27.0/ 25.0/ 26.0 30.9/ 31.8/ 29.4 27.0/ 28.7/ 31.7
Education 28.0/ 29.0/ 28.0 39.2/ 36.6/ 35.4 23.0/ 18.0/ 22.0 31.1/ 31.4/ 32.8 27.0/ 25.0/ 24.0 31.0/ 25.0/ 23.0 30.6/ 32.5/ 32.5 17.0/ 22.0/ 29.0 23.0/ 27.0/ 19.0 31.0/ 27.0/ 32.0 32.0/ 29.0/ 33.0
Energy and Power 33.0/ 31.0/ 35.0 49.3/ 48.6/ 50.0 29.0/ 34.0/ 37.0 34.6/ 36.3/ 35.9 22.0/ 26.0/ 27.0 21.0/ 23.0/ 24.0 28.0/ 33.0/ 36.0 27.0/ 32.0/ 31.0 29.0/ 27.0/ 24.0 24.0/ 25.0/ 29.0 26.0/ 26.0/ 26.0
Engineering 27.2/ 27.4/ 28.4 39.4/ 40.4/ 40.6 28.8/ 29.6/ 28.0 34.8/ 34.8/ 33.0 27.6/ 30.6/ 30.8 24.8/ 23.9/ 26.5 26.4/ 28.6/ 29.6 27.7/ 24.4/ 27.7 29.0/ 26.0/ 24.0 27.4/ 26.0/ 26.8 27.4/ 27.8/ 29.4
Environmental Science 32.0/ 32.0/ 29.0 48.5/ 49.7/ 50.3 25.0/ 22.0/ 21.0 36.3/ 36.7/ 39.5 28.1/ 30.5/ 27.3 31.0/ 26.0/ 29.0 28.4/ 31.1/ 32.1 29.3/ 30.3/ 29.3 26.0/ 22.0/ 26.0 28.0/ 26.0/ 29.0 26.8/ 27.4/ 27.9
Ethics and Human Rights 31.0/ 31.0/ 25.0 46.0/ 51.0/ 52.0 24.0/ 21.0/ 20.0 40.0/ 38.0/ 39.0 20.0/ 29.0/ 22.0 24.0/ 33.0/ 33.0 35.0/ 32.0/ 34.0 22.0/ 22.0/ 27.0 25.0/ 23.0/ 25.0 32.0/ 26.0/ 26.0 30.0/ 26.0/ 26.0
Finance and Investment 27.2/ 30.2/ 31.7 37.0/ 40.2/ 39.4 27.0/ 34.0/ 39.0 28.8/ 32.6/ 32.0 38.3/ 34.6/ 40.7 24.0/ 25.0/ 25.0 29.9/ 26.6/ 29.9 24.0/ 17.0/ 22.0 34.0/ 30.0/ 32.0 32.4/ 31.8/ 34.6 26.4/ 27.1/ 27.9
Food Science 33.0/ 35.0/ 39.0 50.0/ 57.0/ 60.0 24.0/ 31.0/ 31.0 43.4/ 45.5/ 47.5 29.0/ 36.0/ 34.0 26.0/ 33.0/ 34.0 37.0/ 30.0/ 30.0 30.0/ 29.0/ 36.0 34.0/ 45.0/ 44.0 32.0/ 27.0/ 28.0 22.0/ 22.0/ 22.0
Geography 21.2/ 23.5/ 26.3 40.8/ 44.8/ 43.8 24.0/ 30.0/ 29.0 32.1/ 30.9/ 32.7 27.4/ 25.1/ 26.0 33.3/ 31.3/ 33.3 30.8/ 26.6/ 26.6 25.0/ 26.0/ 30.0 18.0/ 27.0/ 28.0 29.1/ 32.6/ 25.1 30.2/ 27.6/ 25.5
Health and Medicine 32.0/ 35.0/ 39.0 55.5/ 58.9/ 57.1 22.0/ 27.0/ 29.0 44.9/ 45.5/ 46.9 28.0/ 27.0/ 29.0 31.0/ 32.0/ 32.0 25.9/ 28.6/ 25.9 19.0/ 22.0/ 27.0 30.0/ 31.0/ 32.0 30.2/ 29.3/ 27.6 24.3/ 27.1/ 25.2
History 25.3/ 23.2/ 25.6 36.5/ 40.8/ 39.0 24.0/ 23.0/ 23.0 36.3/ 34.3/ 34.5 24.2/ 23.8/ 24.9 29.0/ 31.0/ 31.0 30.5/ 26.4/ 29.7 27.0/ 25.0/ 26.0 31.0/ 28.0/ 30.0 29.8/ 34.2/ 34.5 24.6/ 27.4/ 25.8
Information Technology 37.0/ 33.0/ 41.0 47.0/ 50.0/ 49.0 30.0/ 29.0/ 25.0 44.7/ 51.4/ 45.8 25.0/ 29.0/ 32.0 35.0/ 36.0/ 31.0 41.0/ 44.0/ 46.0 29.0/ 27.0/ 27.0 24.0/ 27.0/ 27.0 31.0/ 34.0/ 32.0 39.0/ 35.0/ 34.0
International Relations 30.0/ 34.0/ 27.0 56.7/ 58.3/ 55.8 21.0/ 22.0/ 21.0 44.0/ 39.7/ 37.7 24.0/ 22.0/ 27.0 27.0/ 27.0/ 26.0 35.0/ 28.0/ 32.0 27.0/ 32.0/ 25.0 27.0/ 29.0/ 26.0 29.0/ 29.0/ 26.0 25.0/ 30.0/ 26.0
Language Studies 25.0/ 30.0/ 26.0 51.8/ 53.8/ 55.2 24.0/ 25.0/ 24.0 36.7/ 31.9/ 31.7 28.7/ 26.4/ 27.9 29.0/ 27.0/ 28.0 29.5/ 26.2/ 26.9 31.0/ 33.0/ 32.0 25.0/ 26.0/ 29.0 25.8/ 29.2/ 26.2 27.6/ 28.2/ 26.4
Law and Ethics 16.9/ 21.6/ 20.9 43.6/ 46.0/ 47.2 23.0/ 26.0/ 27.0 39.2/ 37.4/ 38.4 22.9/ 26.8/ 26.8 23.0/ 21.0/ 21.0 32.8/ 34.1/ 29.3 29.0/ 28.0/ 25.0 27.0/ 30.0/ 30.0 28.3/ 26.0/ 29.9 30.0/ 26.9/ 28.7
Literature and Linguistics 27.0/ 31.5/ 31.5 42.1/ 41.7/ 41.1 30.0/ 30.0/ 33.0 35.1/ 33.7/ 34.7 32.5/ 25.0/ 25.8 26.0/ 25.0/ 27.0 36.3/ 37.2/ 36.3 30.0/ 33.0/ 29.0 14.0/ 21.0/ 23.0 31.6/ 32.4/ 34.0 25.4/ 23.9/ 21.8
Logical Reasoning 29.0/ 31.0/ 29.0 34.0/ 34.2/ 34.0 26.0/ 25.3/ 28.6 28.1/ 25.7/ 26.2 27.8/ 24.5/ 25.3 31.0/ 31.0/ 32.0 25.6/ 27.9/ 27.9 21.0/ 24.0/ 20.0 25.0/ 30.0/ 23.0 31.5/ 31.1/ 30.7 28.5/ 25.7/ 27.5
Materials Science 29.0/ 30.0/ 26.0 46.0/ 44.9/ 46.4 19.0/ 25.0/ 23.0 34.1/ 34.7/ 35.4 22.0/ 26.0/ 24.0 33.0/ 41.0/ 35.0 32.0/ 34.0/ 33.0 28.0/ 30.0/ 32.0 36.0/ 27.0/ 29.0 28.0/ 32.0/ 33.0 29.0/ 29.0/ 27.0
Mathematics 29.8/ 30.0/ 30.6 29.8/ 27.6/ 28.6 25.6/ 28.4/ 27.8 27.1/ 28.3/ 27.7 23.6/ 27.4/ 25.8 24.6/ 29.2/ 29.2 25.4/ 25.6/ 26.6 25.6/ 26.0/ 24.4 26.9/ 29.5/ 27.3 28.0/ 30.4/ 28.0 27.0/ 27.6/ 26.4
Media and Communication 27.0/ 31.0/ 30.0 58.0/ 57.0/ 63.0 28.0/ 31.0/ 29.0 40.1/ 39.5/ 40.8 30.0/ 26.0/ 23.0 42.0/ 33.0/ 34.0 36.0/ 29.0/ 26.0 23.0/ 24.0/ 21.0 23.0/ 24.0/ 24.0 26.0/ 26.0/ 24.0 30.0/ 29.0/ 28.0
Music and Performing Arts 27.0/ 19.0/ 24.0 45.0/ 43.0/ 41.0 20.0/ 23.0/ 19.0 28.5/ 31.1/ 27.2 30.0/ 26.0/ 26.0 21.0/ 26.0/ 27.0 34.0/ 30.0/ 28.0 31.0/ 32.0/ 31.0 20.0/ 23.0/ 30.0 32.0/ 27.0/ 26.0 29.0/ 30.0/ 30.0
Physics 28.0/ 25.8/ 26.6 43.8/ 48.4/ 50.2 26.2/ 26.7/ 25.9 34.2/ 35.4/ 35.2 23.4/ 25.4/ 25.2 26.3/ 27.9/ 26.3 28.8/ 28.8/ 30.0 30.6/ 26.0/ 27.1 15.0/ 30.0/ 23.0 26.6/ 28.4/ 29.6 25.6/ 26.2/ 26.6
Politics and Governance 29.0/ 28.6/ 25.9 44.3/ 44.5/ 44.9 28.0/ 31.0/ 28.0 36.5/ 34.5/ 34.7 30.4/ 27.5/ 28.3 36.0/ 41.0/ 33.0 29.8/ 29.8/ 30.1 26.0/ 26.0/ 22.0 22.0/ 28.0/ 24.0 30.3/ 33.2/ 30.3 27.6/ 26.4/ 27.1
Psychology 27.0/ 34.0/ 36.0 60.0/ 58.0/ 62.0 30.0/ 27.0/ 25.0 44.2/ 37.2/ 43.4 30.0/ 26.0/ 19.0 30.0/ 33.0/ 30.0 25.0/ 28.0/ 36.0 33.0/ 29.0/ 30.0 25.0/ 23.0/ 25.0 32.0/ 33.0/ 32.0 27.0/ 29.0/ 28.0
Public Administration 23.0/ 26.0/ 25.0 45.5/ 44.4/ 41.4 28.3/ 26.3/ 22.2 34.3/ 39.4/ 39.4 21.2/ 33.3/ 26.3 32.0/ 28.0/ 29.0 27.3/ 26.3/ 25.2 31.0/ 28.0/ 30.0 18.2/ 19.2/ 22.2 36.0/ 31.0/ 33.0 34.0/ 31.0/ 27.0
Religion and Spirituality 24.0/ 24.0/ 27.0 42.0/ 45.0/ 38.0 17.0/ 20.0/ 24.0 31.9/ 33.5/ 32.7 25.0/ 24.0/ 27.0 35.0/ 34.0/ 33.0 39.0/ 30.0/ 32.0 30.0/ 27.0/ 22.0 35.0/ 24.0/ 30.0 28.0/ 25.0/ 30.0 37.0/ 25.0/ 31.0
Social Welfare and Development 25.0/ 22.0/ 28.0 42.0/ 43.8/ 44.6 28.0/ 28.0/ 28.0 32.6/ 38.3/ 38.9 26.0/ 35.0/ 32.0 28.0/ 31.0/ 28.0 35.0/ 37.0/ 36.0 25.0/ 25.0/ 23.0 25.0/ 18.0/ 20.0 30.0/ 35.0/ 32.0 26.0/ 25.0/ 16.0
Sociology 23.5/ 21.7/ 20.9 34.6/ 36.2/ 36.2 32.0/ 35.0/ 36.0 32.8/ 33.2/ 32.4 27.0/ 33.0/ 27.0 25.0/ 30.0/ 30.0 27.7/ 32.1/ 29.6 23.0/ 20.0/ 24.0 24.0/ 27.0/ 28.0 28.8/ 27.0/ 27.0 24.1/ 19.4/ 23.5
Sports and Recreation 23.8/ 25.7/ 25.7 46.2/ 47.6/ 46.2 22.0/ 23.0/ 20.0 37.2/ 36.8/ 38.2 27.5/ 23.6/ 23.0 28.0/ 25.0/ 27.0 30.5/ 33.9/ 36.7 25.0/ 28.0/ 26.0 21.0/ 32.0/ 32.0 27.6/ 23.7/ 24.4 21.4/ 20.9/ 21.9
Technology and Innovation 30.0/ 32.0/ 27.0 48.6/ 52.8/ 51.8 30.0/ 29.0/ 33.0 37.4/ 39.0/ 39.4 25.0/ 28.0/ 25.0 30.0/ 19.0/ 21.0 34.0/ 37.0/ 39.0 33.0/ 33.0/ 33.0 23.0/ 20.0/ 22.0 29.0/ 30.0/ 30.0 23.0/ 27.0/ 19.0
Transportation and Logistics 27.0/ 25.0/ 25.0 43.9/ 43.9/ 44.6 32.0/ 27.0/ 27.0 29.6/ 34.1/ 33.8 35.4/ 30.3/ 29.3 31.0/ 25.0/ 25.0 31.6/ 29.6/ 31.6 29.3/ 26.3/ 24.2 26.0/ 26.0/ 25.0 31.0/ 34.0/ 31.0 30.0/ 25.0/ 27.0
Table 43: Detailed subject-wise evaluation for CohereForAI/aya-23-8B on MILU across different languages. The results reported are X / Y / Z where X denote the 0-shot, Y denotes 1-shot and Z denotes 5-shot performance respectively.
Topic bn en gu hi kn ml mr or pa ta te
Agriculture 37.0/ 35.0/ 32.0 38.8/ 56.5/ 53.5 32.3/ 34.3/ 30.3 27.6/ 35.8/ 38.0 33.0/ 38.0/ 41.0 31.3/ 31.3/ 35.4 29.0/ 34.0/ 38.0 35.0/ 37.0/ 38.0 28.0/ 29.0/ 30.0 25.0/ 35.0/ 36.0 36.0/ 32.0/ 34.0
Anthropology 37.0/ 37.0/ 38.0 46.0/ 66.0/ 62.0 35.0/ 38.0/ 40.0 34.3/ 42.4/ 39.4 32.0/ 34.0/ 34.0 28.0/ 35.0/ 31.0 37.0/ 42.0/ 39.0 36.0/ 39.0/ 38.0 33.0/ 35.0/ 35.0 29.0/ 36.0/ 32.0 30.0/ 34.0/ 32.0
Architecture and Design 28.0/ 30.0/ 35.0 52.0/ 57.0/ 60.0 33.0/ 35.0/ 38.0 31.0/ 41.0/ 42.0 39.0/ 40.0/ 35.0 31.0/ 35.0/ 33.0 31.0/ 39.0/ 41.0 29.0/ 29.0/ 27.0 28.0/ 38.0/ 37.0 29.0/ 34.0/ 35.0 38.0/ 30.0/ 32.0
Arts and Culture 29.1/ 32.1/ 31.5 48.4/ 64.6/ 63.8 26.0/ 45.0/ 42.0 27.7/ 39.1/ 41.3 30.1/ 33.7/ 36.8 38.0/ 44.0/ 39.0 22.9/ 36.1/ 40.4 17.0/ 18.0/ 20.0 46.0/ 40.0/ 36.0 30.3/ 43.6/ 37.0 28.6/ 35.7/ 35.2
Astronomy and Astrophysics 28.0/ 39.0/ 48.0 60.0/ 80.0/ 79.0 35.0/ 49.0/ 48.0 40.6/ 44.5/ 47.7 35.0/ 43.0/ 48.0 40.0/ 46.0/ 45.0 33.0/ 56.0/ 50.0 36.0/ 38.0/ 38.0 40.0/ 46.0/ 53.0 41.0/ 42.0/ 39.0 33.0/ 39.0/ 42.0
Biology 29.2/ 37.0/ 33.8 53.3/ 73.4/ 74.2 29.0/ 33.0/ 32.0 34.9/ 46.7/ 50.1 35.2/ 41.4/ 40.6 36.4/ 32.3/ 32.3 31.6/ 36.1/ 38.8 26.0/ 32.0/ 31.0 42.0/ 47.0/ 51.0 34.1/ 41.1/ 36.0 30.5/ 36.5/ 34.3
Business and Management 32.0/ 35.0/ 29.0 39.3/ 50.5/ 51.4 26.0/ 39.0/ 38.0 37.9/ 38.6/ 42.4 34.0/ 35.0/ 37.0 43.0/ 33.0/ 34.0 30.0/ 36.0/ 36.0 31.0/ 27.0/ 27.0 25.0/ 34.0/ 33.0 29.0/ 33.0/ 34.0 28.3/ 26.4/ 22.6
Chemistry 34.1/ 40.1/ 40.9 54.1/ 67.5/ 69.7 35.3/ 38.8/ 43.1 38.5/ 48.3/ 49.5 33.0/ 35.3/ 37.9 32.0/ 43.0/ 42.0 36.1/ 42.2/ 43.0 25.0/ 33.7/ 28.8 32.0/ 35.0/ 38.0 32.3/ 34.1/ 35.0 37.9/ 40.6/ 38.9
Computer Science 27.6/ 26.0/ 30.2 39.4/ 44.8/ 43.6 37.0/ 37.0/ 41.0 33.2/ 41.2/ 39.4 36.1/ 43.8/ 43.8 19.0/ 33.0/ 31.0 30.2/ 33.7/ 34.3 26.0/ 33.0/ 35.0 33.0/ 32.0/ 39.0 32.6/ 38.4/ 38.4 29.1/ 37.7/ 39.5
Defense and Security 31.0/ 34.0/ 38.0 45.0/ 66.0/ 64.0 32.0/ 35.0/ 29.0 33.7/ 36.0/ 40.3 33.0/ 38.0/ 37.0 34.0/ 33.0/ 37.0 30.0/ 41.0/ 37.0 31.0/ 33.0/ 29.0 32.0/ 40.0/ 42.0 32.0/ 32.0/ 30.0 30.0/ 33.0/ 37.0
Earth Sciences 42.0/ 41.0/ 40.0 49.1/ 62.1/ 65.9 34.0/ 33.0/ 30.0 31.1/ 41.1/ 42.7 38.0/ 37.0/ 38.0 31.0/ 38.0/ 38.0 32.0/ 39.0/ 40.0 34.3/ 34.3/ 33.3 33.0/ 36.0/ 36.0 38.0/ 35.0/ 38.0 32.0/ 35.0/ 34.0
Economics 36.0/ 36.7/ 41.7 48.4/ 58.4/ 60.6 31.0/ 36.0/ 36.0 30.0/ 40.7/ 48.0 36.0/ 42.0/ 40.0 31.0/ 27.0/ 23.0 29.6/ 33.8/ 34.5 29.0/ 28.0/ 27.0 25.0/ 32.0/ 36.0 34.1/ 38.9/ 36.5 28.7/ 34.7/ 37.7
Education 38.0/ 38.0/ 40.0 38.5/ 49.7/ 51.6 28.0/ 36.0/ 38.0 33.1/ 33.7/ 39.1 28.0/ 33.0/ 35.0 23.0/ 27.0/ 30.0 31.9/ 38.2/ 38.2 28.0/ 36.0/ 33.0 25.0/ 30.0/ 24.0 37.0/ 43.0/ 39.0 25.0/ 31.0/ 29.0
Energy and Power 35.0/ 44.0/ 47.0 46.0/ 63.5/ 66.2 39.0/ 41.0/ 39.0 34.6/ 44.1/ 52.2 28.0/ 34.0/ 35.0 34.0/ 42.0/ 39.0 25.0/ 44.0/ 42.0 25.0/ 34.0/ 34.0 33.0/ 36.0/ 38.0 39.0/ 38.0/ 36.0 37.0/ 38.0/ 38.0
Engineering 26.4/ 32.8/ 33.0 38.2/ 48.2/ 49.4 27.6/ 29.0/ 30.6 31.0/ 36.4/ 34.2 29.4/ 28.8/ 29.4 27.8/ 30.3/ 31.6 25.8/ 29.2/ 31.8 29.2/ 25.0/ 29.8 29.0/ 33.0/ 33.0 29.2/ 36.0/ 34.0 28.2/ 31.0/ 33.8
Environmental Science 26.0/ 34.0/ 36.0 45.7/ 65.3/ 61.3 33.0/ 35.0/ 35.0 31.1/ 43.9/ 45.9 36.7/ 32.8/ 34.4 30.0/ 37.0/ 39.0 25.8/ 37.4/ 41.6 39.4/ 38.4/ 43.4 40.0/ 43.0/ 37.0 32.0/ 42.0/ 42.0 25.7/ 31.8/ 32.4
Ethics and Human Rights 33.0/ 35.0/ 45.0 44.0/ 62.0/ 64.0 36.0/ 41.0/ 38.0 41.0/ 42.0/ 49.0 32.0/ 41.0/ 37.0 33.0/ 38.0/ 41.0 36.0/ 43.0/ 40.0 36.0/ 33.0/ 38.0 30.0/ 35.0/ 28.0 30.0/ 38.0/ 38.0 30.0/ 34.0/ 39.0
Finance and Investment 32.2/ 39.1/ 37.1 42.6/ 49.6/ 50.0 39.0/ 40.0/ 36.0 33.4/ 43.2/ 42.4 37.0/ 42.6/ 37.0 36.0/ 32.0/ 31.0 29.9/ 41.8/ 36.7 26.0/ 27.0/ 26.0 35.0/ 35.0/ 35.0 39.7/ 42.5/ 37.4 33.9/ 35.0/ 35.0
Food Science 29.0/ 32.0/ 37.0 53.0/ 74.0/ 74.0 32.0/ 34.0/ 33.0 28.3/ 48.5/ 53.5 29.0/ 34.0/ 40.0 36.0/ 38.0/ 37.0 22.0/ 35.0/ 41.0 29.0/ 36.0/ 38.0 32.0/ 36.0/ 38.0 30.0/ 31.0/ 36.0 33.0/ 42.0/ 48.0
Geography 36.3/ 38.6/ 41.3 43.0/ 58.0/ 59.2 44.0/ 42.0/ 49.0 29.1/ 38.8/ 41.0 32.4/ 40.6/ 36.1 38.4/ 43.4/ 44.4 30.2/ 34.9/ 36.7 41.0/ 43.0/ 41.0 30.0/ 42.0/ 37.0 29.1/ 30.9/ 26.3 30.7/ 33.9/ 36.5
Health and Medicine 31.0/ 38.0/ 46.0 50.7/ 75.0/ 74.2 31.0/ 47.0/ 41.0 37.1/ 50.3/ 55.9 41.0/ 47.0/ 45.0 37.0/ 41.0/ 42.0 34.8/ 33.9/ 35.7 33.0/ 44.0/ 42.0 49.0/ 46.0/ 47.0 35.3/ 37.9/ 36.2 32.7/ 40.2/ 42.1
History 34.9/ 41.6/ 42.1 42.2/ 59.0/ 59.6 43.0/ 51.0/ 47.0 33.5/ 43.8/ 46.0 38.8/ 42.5/ 45.1 31.0/ 37.0/ 33.0 34.9/ 39.4/ 37.5 35.0/ 38.0/ 36.0 32.0/ 36.0/ 32.0 35.6/ 41.1/ 41.1 37.3/ 35.5/ 38.6
Information Technology 35.0/ 41.0/ 48.0 53.0/ 68.0/ 71.0 43.0/ 41.0/ 41.0 43.0/ 57.5/ 54.8 36.0/ 43.0/ 47.0 41.0/ 45.0/ 44.0 43.0/ 49.0/ 53.0 29.0/ 38.0/ 44.0 39.0/ 47.0/ 44.0 39.0/ 50.0/ 45.0 42.0/ 51.0/ 49.0
International Relations 30.0/ 37.0/ 42.0 55.0/ 70.0/ 69.2 34.0/ 40.0/ 39.0 33.9/ 40.9/ 45.1 29.0/ 33.0/ 39.0 37.0/ 34.0/ 35.0 30.0/ 40.0/ 40.0 36.0/ 40.0/ 47.0 49.0/ 47.0/ 48.0 38.0/ 48.0/ 50.0 30.0/ 31.0/ 32.0
Language Studies 29.0/ 26.0/ 25.0 53.0/ 69.6/ 70.0 27.0/ 27.0/ 25.0 30.1/ 35.1/ 33.9 30.2/ 33.3/ 33.3 26.0/ 29.0/ 24.0 26.2/ 30.2/ 30.9 28.0/ 33.0/ 34.0 38.0/ 42.0/ 37.0 32.2/ 30.0/ 30.0 20.7/ 29.9/ 29.3
Law and Ethics 38.5/ 48.6/ 49.3 51.0/ 65.4/ 64.8 22.0/ 37.0/ 35.0 41.8/ 51.8/ 52.4 32.5/ 38.2/ 40.8 25.0/ 29.0/ 31.0 36.6/ 40.1/ 39.2 32.0/ 29.0/ 34.0 42.0/ 43.0/ 42.0 34.6/ 37.0/ 35.4 26.9/ 30.0/ 33.1
Literature and Linguistics 29.7/ 37.8/ 41.4 45.5/ 62.9/ 60.1 29.0/ 32.0/ 31.0 26.7/ 32.5/ 35.9 32.5/ 37.5/ 37.5 33.0/ 32.0/ 33.0 30.1/ 31.9/ 31.9 28.0/ 30.0/ 31.0 27.0/ 33.0/ 31.0 31.6/ 31.6/ 36.0 30.3/ 33.8/ 33.1
Logical Reasoning 31.2/ 34.6/ 34.6 27.6/ 32.4/ 31.8 22.7/ 29.2/ 29.9 28.3/ 31.5/ 32.5 30.3/ 31.9/ 29.5 32.0/ 22.0/ 29.0 25.6/ 25.1/ 26.2 27.0/ 27.0/ 24.0 27.0/ 23.0/ 26.0 32.7/ 31.1/ 31.9 29.5/ 29.5/ 27.0
Materials Science 31.0/ 32.0/ 40.0 43.4/ 56.3/ 59.3 25.0/ 19.0/ 18.0 31.4/ 34.3/ 34.5 29.0/ 28.0/ 30.0 41.0/ 43.0/ 35.0 33.0/ 39.0/ 44.0 20.0/ 25.0/ 19.0 31.0/ 36.0/ 37.0 33.0/ 37.0/ 37.0 22.0/ 33.0/ 38.0
Mathematics 29.8/ 32.0/ 33.8 31.4/ 35.8/ 36.4 26.4/ 29.8/ 37.2 30.7/ 33.7/ 33.1 28.4/ 30.6/ 34.2 26.9/ 34.1/ 32.4 25.8/ 32.4/ 35.6 25.2/ 28.8/ 30.6 27.7/ 31.8/ 33.7 28.6/ 31.6/ 30.4 25.6/ 30.4/ 30.8
Media and Communication 34.0/ 35.0/ 33.0 50.0/ 71.0/ 64.0 41.0/ 39.0/ 37.0 32.0/ 43.5/ 44.9 41.0/ 36.0/ 40.0 44.0/ 39.0/ 34.0 29.0/ 39.0/ 40.0 33.0/ 31.0/ 38.0 40.0/ 39.0/ 35.0 38.0/ 41.0/ 41.0 39.0/ 39.0/ 44.0
Music and Performing Arts 31.0/ 37.0/ 38.0 46.0/ 63.0/ 59.0 29.0/ 30.0/ 33.0 29.8/ 37.8/ 36.4 36.0/ 35.0/ 36.0 29.0/ 30.0/ 32.0 34.0/ 42.0/ 38.0 29.0/ 34.0/ 38.0 37.0/ 47.0/ 44.0 30.0/ 30.0/ 35.0 37.0/ 39.0/ 39.0
Physics 35.2/ 36.2/ 39.4 50.6/ 61.4/ 60.8 34.5/ 33.7/ 32.0 31.2/ 38.0/ 39.0 35.2/ 35.6/ 37.6 26.8/ 32.1/ 36.3 30.8/ 40.6/ 40.4 30.9/ 28.8/ 32.3 34.0/ 40.0/ 39.0 30.2/ 35.8/ 34.8 33.2/ 35.4/ 36.8
Politics and Governance 33.7/ 40.4/ 44.3 42.7/ 57.3/ 57.1 29.0/ 35.0/ 33.0 36.8/ 44.6/ 43.2 30.4/ 33.6/ 36.0 37.0/ 29.0/ 32.0 31.5/ 36.8/ 37.6 31.0/ 28.0/ 33.0 28.0/ 30.0/ 30.0 41.9/ 42.7/ 40.2 30.9/ 34.0/ 33.7
Psychology 36.0/ 40.0/ 37.0 57.0/ 71.0/ 64.0 33.0/ 36.0/ 36.0 29.5/ 41.1/ 46.5 42.0/ 37.0/ 41.0 31.0/ 33.0/ 39.0 30.0/ 35.0/ 36.0 34.0/ 45.0/ 41.0 35.0/ 33.0/ 34.0 35.0/ 41.0/ 38.0 34.0/ 37.0/ 40.0
Public Administration 33.0/ 30.0/ 31.0 37.4/ 56.6/ 61.6 31.3/ 35.4/ 28.3 32.3/ 36.4/ 30.3 31.3/ 27.3/ 34.3 25.0/ 33.0/ 33.0 37.4/ 31.3/ 33.3 26.0/ 30.0/ 33.0 37.4/ 38.4/ 32.3 40.0/ 37.0/ 33.0 28.0/ 33.0/ 34.0
Religion and Spirituality 30.0/ 39.0/ 42.0 49.0/ 72.0/ 69.0 39.0/ 45.0/ 45.0 37.9/ 41.9/ 46.0 43.0/ 33.0/ 46.0 39.0/ 36.0/ 44.0 40.0/ 38.0/ 45.0 34.0/ 36.0/ 35.0 38.0/ 33.0/ 30.0 37.0/ 37.0/ 40.0 35.0/ 41.0/ 34.0
Social Welfare and Development 35.0/ 38.0/ 38.0 42.9/ 56.2/ 58.9 35.0/ 38.0/ 40.0 34.7/ 38.3/ 42.0 26.0/ 37.0/ 35.0 25.0/ 31.0/ 30.0 44.0/ 40.0/ 34.0 37.0/ 29.0/ 23.0 30.0/ 28.0/ 33.0 34.0/ 35.0/ 35.0 30.0/ 39.0/ 36.0
Sociology 27.0/ 28.7/ 30.4 36.6/ 44.0/ 45.2 31.0/ 29.0/ 34.0 31.8/ 38.4/ 39.0 28.0/ 31.0/ 32.0 28.0/ 29.0/ 29.0 29.6/ 32.1/ 32.7 25.0/ 29.0/ 25.0 27.0/ 34.0/ 27.0 37.8/ 32.4/ 37.8 28.8/ 24.7/ 23.5
Sports and Recreation 27.2/ 31.7/ 30.7 46.6/ 68.0/ 67.0 25.0/ 35.0/ 34.0 29.0/ 39.4/ 43.6 29.8/ 29.8/ 36.5 37.0/ 30.0/ 31.0 28.8/ 42.4/ 41.8 23.0/ 26.0/ 24.0 29.0/ 33.0/ 34.0 35.3/ 36.5/ 30.1 29.5/ 28.6/ 31.9
Technology and Innovation 41.0/ 44.0/ 49.0 49.4/ 64.2/ 63.8 32.0/ 35.0/ 42.0 38.2/ 43.8/ 45.4 43.0/ 45.0/ 41.0 38.0/ 37.0/ 38.0 36.0/ 40.0/ 40.0 30.0/ 39.0/ 43.0 28.0/ 38.0/ 40.0 35.0/ 38.0/ 42.0 36.0/ 29.0/ 31.0
Transportation and Logistics 24.0/ 31.0/ 29.0 36.9/ 53.8/ 56.1 20.0/ 27.0/ 27.0 32.5/ 35.6/ 35.0 29.3/ 24.2/ 22.2 30.0/ 30.0/ 28.0 22.4/ 25.5/ 31.6 24.2/ 28.3/ 24.2 34.0/ 33.0/ 31.0 34.0/ 31.0/ 32.0 29.0/ 30.0/ 27.0
Table 44: Detailed subject-wise evaluation for meta-llama/Llama-3.1-8B on MILU across different languages. The results reported are X / Y / Z where X denote the 0-shot, Y denotes 1-shot and Z denotes 5-shot performance respectively.
Topic bn en gu hi kn ml mr or pa ta te
Agriculture 30.0/ 24.0/ 31.0 39.4/ 39.4/ 40.6 30.3/ 31.3/ 28.3 33.6/ 34.2/ 31.2 23.0/ 31.0/ 30.0 31.3/ 30.3/ 25.2 35.0/ 33.0/ 27.0 36.0/ 29.0/ 31.0 34.0/ 38.0/ 30.0 20.0/ 28.0/ 31.0 36.0/ 38.0/ 25.0
Anthropology 46.0/ 41.0/ 36.0 50.0/ 52.0/ 44.0 38.0/ 37.0/ 37.0 45.5/ 31.3/ 32.3 42.0/ 40.0/ 41.0 33.0/ 33.0/ 36.0 39.0/ 36.0/ 37.0 38.0/ 40.0/ 35.0 46.0/ 47.0/ 37.0 43.0/ 38.0/ 41.0 36.0/ 35.0/ 36.0
Architecture and Design 37.0/ 31.0/ 25.0 46.0/ 45.0/ 48.0 38.0/ 42.0/ 36.0 38.0/ 30.0/ 37.0 37.0/ 30.0/ 35.0 39.0/ 39.0/ 36.0 35.0/ 29.0/ 32.0 26.0/ 28.0/ 27.0 27.0/ 27.0/ 24.0 30.0/ 31.0/ 29.0 35.0/ 33.0/ 31.0
Arts and Culture 37.6/ 34.5/ 33.9 46.0/ 41.8/ 46.6 32.0/ 42.0/ 30.0 35.5/ 32.5/ 37.1 40.5/ 35.6/ 35.6 34.0/ 35.0/ 31.0 27.9/ 29.0/ 30.0 27.0/ 27.0/ 20.0 40.0/ 37.0/ 33.0 40.0/ 36.4/ 37.6 34.7/ 36.7/ 35.2
Astronomy and Astrophysics 47.0/ 44.0/ 41.0 66.0/ 60.0/ 67.0 52.0/ 41.0/ 42.0 49.2/ 39.1/ 42.2 52.0/ 49.0/ 40.0 53.0/ 47.0/ 39.0 48.0/ 40.0/ 44.0 41.0/ 40.0/ 41.0 43.0/ 37.0/ 37.0 49.0/ 42.0/ 41.0 53.0/ 50.0/ 45.0
Biology 43.7/ 33.8/ 33.8 47.5/ 49.3/ 53.3 34.0/ 31.0/ 33.0 41.9/ 36.5/ 38.7 39.8/ 43.4/ 35.2 35.4/ 32.3/ 35.4 36.1/ 33.4/ 30.8 35.0/ 31.0/ 30.0 37.0/ 36.0/ 37.0 37.3/ 34.4/ 31.9 34.9/ 38.4/ 29.2
Business and Management 32.0/ 33.0/ 27.0 45.8/ 39.6/ 42.4 43.0/ 51.0/ 37.0 37.2/ 31.4/ 32.6 37.0/ 33.0/ 39.0 38.0/ 33.0/ 38.0 38.0/ 40.0/ 37.0 35.0/ 29.0/ 29.0 39.0/ 32.0/ 30.0 36.0/ 31.0/ 32.0 30.2/ 26.4/ 23.6
Chemistry 38.7/ 34.9/ 35.7 50.5/ 49.3/ 53.7 29.3/ 37.1/ 42.2 39.9/ 37.5/ 37.7 35.0/ 33.0/ 35.9 41.0/ 33.0/ 41.0 34.6/ 34.3/ 31.7 34.6/ 34.6/ 30.8 35.0/ 22.0/ 29.0 35.6/ 32.0/ 32.3 37.4/ 37.4/ 37.4
Computer Science 26.0/ 22.9/ 26.6 33.8/ 34.4/ 37.6 36.0/ 38.0/ 41.0 30.4/ 28.0/ 31.4 35.5/ 32.5/ 36.7 42.0/ 40.0/ 41.0 26.0/ 28.4/ 29.6 27.0/ 36.0/ 33.0 32.0/ 28.0/ 31.0 33.1/ 39.5/ 38.4 31.4/ 34.1/ 31.8
Defense and Security 32.0/ 31.0/ 37.0 61.0/ 49.0/ 62.0 40.0/ 32.0/ 35.0 37.9/ 35.1/ 37.9 31.0/ 39.0/ 41.0 37.0/ 32.0/ 34.0 43.0/ 38.0/ 37.0 39.0/ 31.0/ 28.0 38.0/ 45.0/ 37.0 27.0/ 32.0/ 35.0 39.0/ 34.0/ 43.0
Earth Sciences 53.0/ 38.0/ 42.0 50.2/ 50.9/ 52.3 38.0/ 39.0/ 36.0 35.9/ 34.1/ 32.3 37.0/ 40.0/ 45.0 39.0/ 38.0/ 41.0 38.0/ 39.0/ 37.0 28.3/ 31.3/ 30.3 38.0/ 38.0/ 35.0 39.0/ 36.0/ 37.0 35.0/ 34.0/ 32.0
Economics 43.2/ 36.0/ 38.9 48.4/ 41.8/ 42.6 34.0/ 40.0/ 32.0 34.5/ 34.1/ 36.3 43.0/ 40.0/ 42.0 37.0/ 26.0/ 29.0 37.3/ 35.9/ 38.0 28.0/ 35.0/ 28.0 39.0/ 35.0/ 35.0 36.5/ 28.6/ 30.2 37.1/ 35.3/ 33.5
Education 38.0/ 41.0/ 37.0 45.9/ 36.3/ 43.6 33.0/ 36.0/ 28.0 34.3/ 32.8/ 34.6 34.0/ 33.0/ 33.0 31.0/ 32.0/ 29.0 39.5/ 42.0/ 42.0 27.0/ 28.0/ 31.0 32.0/ 28.0/ 27.0 43.0/ 34.0/ 27.0 44.0/ 39.0/ 34.0
Energy and Power 41.0/ 34.0/ 37.0 53.4/ 47.3/ 48.6 43.0/ 42.0/ 38.0 32.2/ 35.2/ 35.6 38.0/ 37.0/ 36.0 35.0/ 32.0/ 33.0 36.0/ 35.0/ 37.0 44.0/ 39.0/ 39.0 43.0/ 37.0/ 41.0 32.0/ 27.0/ 34.0 32.0/ 29.0/ 30.0
Engineering 33.2/ 27.0/ 26.2 35.6/ 34.0/ 36.8 30.8/ 30.6/ 29.0 30.4/ 28.6/ 29.4 36.8/ 32.4/ 32.2 33.3/ 33.3/ 34.2 29.0/ 27.0/ 31.0 31.9/ 28.0/ 30.9 23.0/ 26.0/ 30.0 29.8/ 28.4/ 28.8 35.0/ 29.4/ 31.6
Environmental Science 36.0/ 25.0/ 27.0 48.9/ 42.3/ 49.7 46.0/ 42.0/ 33.0 41.5/ 40.3/ 41.9 48.4/ 34.4/ 38.3 36.0/ 39.0/ 36.0 42.1/ 33.7/ 37.4 29.3/ 26.3/ 30.3 38.0/ 38.0/ 36.0 44.0/ 36.0/ 40.0 41.9/ 40.8/ 30.7
Ethics and Human Rights 40.0/ 41.0/ 38.0 53.0/ 47.0/ 47.0 38.0/ 37.0/ 32.0 42.0/ 37.0/ 42.0 40.0/ 38.0/ 31.0 37.0/ 32.0/ 34.0 40.0/ 43.0/ 40.0 27.0/ 26.0/ 22.0 33.0/ 37.0/ 31.0 40.0/ 38.0/ 34.0 45.0/ 42.0/ 37.0
Finance and Investment 30.7/ 28.7/ 34.2 38.2/ 36.2/ 39.6 31.0/ 34.0/ 33.0 31.4/ 29.0/ 31.2 38.3/ 36.4/ 37.0 35.0/ 30.0/ 35.0 30.5/ 28.8/ 28.8 24.0/ 29.0/ 23.0 22.0/ 25.0/ 27.0 33.0/ 38.0/ 35.2 32.9/ 35.0/ 31.4
Food Science 42.0/ 34.0/ 37.0 46.0/ 48.0/ 54.0 37.0/ 32.0/ 33.0 44.4/ 41.4/ 42.4 33.0/ 34.0/ 32.0 42.0/ 46.0/ 44.0 37.0/ 32.0/ 28.0 30.0/ 34.0/ 33.0 41.0/ 32.0/ 33.0 45.0/ 34.0/ 34.0 41.0/ 35.0/ 38.0
Geography 39.1/ 25.7/ 31.8 47.8/ 42.0/ 47.2 45.0/ 45.0/ 46.0 39.0/ 35.5/ 33.1 41.1/ 39.3/ 34.7 39.4/ 37.4/ 35.4 31.9/ 30.2/ 30.8 31.0/ 27.0/ 32.0 36.0/ 30.0/ 36.0 36.6/ 33.7/ 33.7 42.2/ 37.0/ 38.5
Health and Medicine 40.0/ 48.0/ 39.0 60.3/ 54.9/ 58.5 39.0/ 42.0/ 36.0 48.5/ 43.9/ 43.9 46.0/ 50.0/ 47.0 47.0/ 39.0/ 43.0 33.0/ 34.8/ 31.2 44.0/ 37.0/ 35.0 44.0/ 44.0/ 39.0 41.4/ 39.7/ 34.5 43.0/ 37.4/ 42.1
History 43.5/ 43.5/ 41.9 45.6/ 42.6/ 46.2 46.0/ 50.0/ 37.0 44.4/ 42.7/ 43.4 45.1/ 46.9/ 39.9 36.0/ 37.0/ 40.0 36.1/ 33.5/ 32.7 34.0/ 31.0/ 32.0 46.0/ 42.0/ 39.0 36.7/ 36.7/ 38.6 38.4/ 37.9/ 38.9
Information Technology 49.0/ 42.0/ 42.0 54.0/ 51.0/ 50.0 47.0/ 37.0/ 35.0 50.3/ 45.8/ 46.4 47.0/ 37.0/ 45.0 52.0/ 48.0/ 49.0 44.0/ 40.0/ 41.0 42.0/ 42.0/ 38.0 45.0/ 36.0/ 40.0 47.0/ 40.0/ 43.0 56.0/ 48.0/ 50.0
International Relations 36.0/ 39.0/ 37.0 57.5/ 48.3/ 60.0 45.0/ 43.0/ 40.0 40.1/ 40.9/ 38.9 35.0/ 37.0/ 30.0 35.0/ 34.0/ 32.0 49.0/ 45.0/ 44.0 31.0/ 34.0/ 35.0 44.0/ 42.0/ 41.0 43.0/ 38.0/ 40.0 31.0/ 32.0/ 28.0
Language Studies 29.0/ 29.0/ 37.0 51.4/ 43.8/ 52.2 33.0/ 28.0/ 39.0 31.1/ 30.7/ 31.1 27.9/ 28.7/ 31.0 34.0/ 31.0/ 34.0 32.2/ 28.9/ 27.5 38.0/ 30.0/ 34.0 41.0/ 40.0/ 36.0 30.0/ 26.6/ 25.8 32.8/ 26.4/ 25.9
Law and Ethics 53.4/ 38.5/ 48.6 49.2/ 49.4/ 49.8 31.0/ 37.0/ 29.0 45.2/ 42.4/ 43.6 46.5/ 41.4/ 36.9 28.0/ 32.0/ 33.0 37.9/ 39.7/ 36.2 29.0/ 29.0/ 33.0 46.0/ 40.0/ 42.0 40.2/ 34.6/ 34.6 38.8/ 35.6/ 35.6
Literature and Linguistics 36.0/ 30.6/ 36.0 43.7/ 39.9/ 44.1 33.0/ 39.0/ 35.0 35.1/ 34.1/ 34.9 33.3/ 33.3/ 35.0 32.0/ 29.0/ 31.0 31.9/ 26.6/ 31.9 32.0/ 27.0/ 29.0 30.0/ 34.0/ 33.0 37.6/ 29.1/ 31.2 26.1/ 31.0/ 29.6
Logical Reasoning 36.1/ 33.7/ 34.6 27.4/ 24.0/ 26.6 26.0/ 26.0/ 27.9 30.1/ 28.9/ 28.9 30.7/ 30.3/ 31.5 23.0/ 19.0/ 21.0 29.5/ 27.6/ 25.4 21.0/ 22.0/ 22.0 25.0/ 25.0/ 21.0 33.1/ 30.7/ 30.3 27.5/ 28.2/ 30.2
Materials Science 34.0/ 29.0/ 32.0 45.6/ 41.1/ 46.8 23.0/ 19.0/ 28.0 37.5/ 33.0/ 33.2 44.0/ 42.0/ 36.0 37.0/ 36.0/ 44.0 40.0/ 34.0/ 43.0 31.0/ 36.0/ 34.0 31.0/ 34.0/ 26.0 40.0/ 40.0/ 41.0 34.0/ 42.0/ 37.0
Mathematics 27.8/ 29.2/ 29.8 28.4/ 29.6/ 30.4 28.4/ 27.4/ 28.8 26.5/ 28.9/ 26.1 25.2/ 26.2/ 31.4 26.1/ 27.5/ 29.5 26.0/ 28.4/ 24.4 26.8/ 27.2/ 28.4 25.8/ 26.1/ 26.1 24.0/ 29.8/ 31.4 25.8/ 26.8/ 27.0
Media and Communication 46.0/ 35.0/ 38.0 54.0/ 52.0/ 51.0 37.0/ 36.0/ 40.0 38.8/ 33.3/ 42.2 50.0/ 45.0/ 42.0 47.0/ 41.0/ 38.0 42.0/ 30.0/ 38.0 36.0/ 44.0/ 40.0 40.0/ 32.0/ 36.0 40.0/ 44.0/ 39.0 43.0/ 46.0/ 39.0
Music and Performing Arts 36.0/ 30.0/ 26.0 51.0/ 42.0/ 49.0 35.0/ 33.0/ 37.0 38.4/ 31.1/ 27.2 38.0/ 38.0/ 39.0 29.0/ 25.0/ 20.0 38.0/ 32.0/ 29.0 23.0/ 30.0/ 33.0 32.0/ 33.0/ 38.0 37.0/ 31.0/ 30.0 40.0/ 33.0/ 42.0
Physics 34.6/ 32.2/ 32.2 42.6/ 43.0/ 48.8 30.1/ 31.8/ 31.5 35.6/ 33.4/ 35.6 33.0/ 33.2/ 35.0 33.7/ 37.4/ 33.2 32.0/ 29.2/ 31.8 29.2/ 31.2/ 28.5 32.0/ 36.0/ 39.0 36.4/ 35.6/ 35.8 34.6/ 33.4/ 35.2
Politics and Governance 42.4/ 35.3/ 40.8 48.1/ 44.5/ 47.7 34.0/ 33.0/ 31.0 37.0/ 35.9/ 38.0 37.2/ 37.2/ 33.2 27.0/ 28.0/ 32.0 38.8/ 34.8/ 32.0 30.0/ 28.0/ 29.0 33.0/ 25.0/ 29.0 38.2/ 40.2/ 34.8 33.0/ 33.2/ 30.9
Psychology 41.0/ 28.0/ 35.0 52.0/ 43.0/ 45.0 33.0/ 34.0/ 29.0 36.4/ 34.1/ 35.7 42.0/ 44.0/ 47.0 31.0/ 30.0/ 31.0 37.0/ 41.0/ 38.0 30.0/ 41.0/ 36.0 32.0/ 32.0/ 30.0 40.0/ 42.0/ 37.0 41.0/ 46.0/ 38.0
Public Administration 33.0/ 32.0/ 35.0 39.4/ 35.4/ 36.4 25.2/ 28.3/ 32.3 36.4/ 42.4/ 40.4 24.2/ 26.3/ 34.3 26.0/ 30.0/ 30.0 24.2/ 29.3/ 36.4 24.0/ 27.0/ 25.0 33.3/ 31.3/ 29.3 28.0/ 30.0/ 30.0 31.0/ 27.0/ 33.0
Religion and Spirituality 41.0/ 32.0/ 32.0 49.0/ 44.0/ 51.0 40.0/ 39.0/ 40.0 37.9/ 36.3/ 34.3 53.0/ 49.0/ 44.0 47.0/ 34.0/ 36.0 45.0/ 38.0/ 35.0 31.0/ 22.0/ 25.0 39.0/ 38.0/ 35.0 37.0/ 35.0/ 30.0 42.0/ 35.0/ 29.0
Social Welfare and Development 41.0/ 34.0/ 36.0 44.6/ 47.3/ 50.9 37.0/ 44.0/ 40.0 39.9/ 37.3/ 36.3 37.0/ 34.0/ 41.0 37.0/ 31.0/ 29.0 36.0/ 33.0/ 37.0 31.0/ 29.0/ 29.0 40.0/ 34.0/ 28.0 38.0/ 39.0/ 41.0 37.0/ 34.0/ 36.0
Sociology 37.4/ 30.4/ 30.4 32.6/ 34.8/ 35.0 37.0/ 30.0/ 29.0 30.0/ 30.8/ 32.4 35.0/ 29.0/ 34.0 30.0/ 25.0/ 26.0 37.7/ 34.6/ 32.1 20.0/ 21.0/ 19.0 38.0/ 37.0/ 34.0 30.6/ 35.1/ 34.2 31.2/ 32.4/ 30.0
Sports and Recreation 36.1/ 31.7/ 31.7 53.8/ 46.2/ 50.4 30.0/ 31.0/ 30.0 37.0/ 31.4/ 34.2 30.9/ 33.1/ 29.2 34.0/ 37.0/ 32.0 39.6/ 36.2/ 36.7 26.0/ 28.0/ 28.0 30.0/ 26.0/ 27.0 41.0/ 36.5/ 33.3 35.7/ 35.2/ 30.0
Technology and Innovation 38.0/ 40.0/ 40.0 50.4/ 45.6/ 48.8 33.0/ 39.0/ 39.0 34.4/ 33.8/ 37.4 40.0/ 37.0/ 43.0 35.0/ 38.0/ 40.0 38.0/ 32.0/ 39.0 35.0/ 34.0/ 27.0 40.0/ 43.0/ 41.0 36.0/ 39.0/ 36.0 35.0/ 35.0/ 30.0
Transportation and Logistics 35.0/ 31.0/ 30.0 37.7/ 36.1/ 40.0 24.0/ 31.0/ 24.0 33.4/ 30.9/ 31.9 32.3/ 29.3/ 24.2 33.0/ 24.0/ 29.0 32.6/ 25.5/ 27.6 27.3/ 25.2/ 18.2 26.0/ 25.0/ 27.0 32.0/ 35.0/ 31.0 31.0/ 25.0/ 24.0
Table 45: Detailed subject-wise evaluation for meta-llama/Llama-3.1-8B-Instruct on MILU across different languages. The results reported are X / Y / Z where X denote the 0-shot, Y denotes 1-shot and Z denotes 5-shot performance respectively.
Topic bn en gu hi kn ml mr or pa ta te
Agriculture 42.0/ 44.0/ 48.0 63.8/ 67.4/ 65.8 37.4/ 51.5/ 49.5 49.5/ 52.9/ 58.1 47.0/ 43.0/ 47.0 38.4/ 38.4/ 39.4 31.0/ 38.0/ 43.0 24.0/ 27.0/ 32.0 43.0/ 43.0/ 46.0 34.0/ 32.0/ 35.0 40.0/ 47.0/ 45.0
Anthropology 48.0/ 46.0/ 47.0 60.0/ 64.0/ 62.0 39.0/ 45.0/ 47.0 50.5/ 46.5/ 48.5 44.0/ 41.0/ 47.0 42.0/ 44.0/ 49.0 48.0/ 44.0/ 40.0 42.0/ 36.0/ 40.0 44.0/ 47.0/ 46.0 42.0/ 42.0/ 37.0 40.0/ 45.0/ 46.0
Architecture and Design 43.0/ 46.0/ 42.0 68.0/ 66.0/ 68.0 44.0/ 46.0/ 44.0 46.0/ 47.0/ 47.0 38.0/ 41.0/ 47.0 43.0/ 46.0/ 47.0 44.0/ 51.0/ 47.0 43.0/ 50.0/ 43.0 44.0/ 45.0/ 48.0 39.0/ 37.0/ 42.0 43.0/ 47.0/ 44.0
Arts and Culture 45.5/ 43.0/ 43.6 69.8/ 71.2/ 71.4 50.0/ 50.0/ 51.0 50.9/ 53.1/ 52.7 39.3/ 42.3/ 47.2 40.0/ 49.0/ 45.0 45.4/ 49.7/ 50.8 24.0/ 28.0/ 26.0 41.0/ 44.0/ 41.0 35.1/ 35.8/ 41.2 34.2/ 36.2/ 38.7
Astronomy and Astrophysics 61.0/ 72.0/ 73.0 86.0/ 87.0/ 86.0 63.0/ 72.0/ 73.0 59.4/ 68.0/ 64.8 64.0/ 70.0/ 73.0 62.0/ 67.0/ 65.0 63.0/ 65.0/ 64.0 49.0/ 43.0/ 47.0 58.0/ 57.0/ 59.0 60.0/ 72.0/ 74.0 61.0/ 59.0/ 66.0
Biology 51.1/ 59.9/ 64.4 79.6/ 83.0/ 83.6 57.0/ 59.0/ 58.0 69.3/ 73.4/ 74.6 53.7/ 59.0/ 61.5 57.6/ 59.6/ 59.6 47.8/ 55.2/ 56.7 41.0/ 44.0/ 41.0 69.0/ 68.0/ 70.0 53.8/ 62.1/ 59.2 52.5/ 56.9/ 56.0
Business and Management 38.0/ 43.0/ 43.0 57.3/ 64.7/ 67.2 38.0/ 55.0/ 51.0 50.1/ 51.0/ 52.9 39.0/ 47.0/ 52.0 38.0/ 40.0/ 45.0 43.0/ 43.0/ 41.0 42.0/ 41.0/ 41.0 50.0/ 49.0/ 46.0 42.0/ 40.0/ 46.0 34.0/ 38.7/ 34.0
Chemistry 58.0/ 64.6/ 64.8 82.0/ 83.6/ 85.0 56.0/ 55.2/ 56.9 63.7/ 69.7/ 70.7 49.7/ 54.9/ 57.2 55.0/ 59.0/ 56.0 50.1/ 59.1/ 62.3 43.3/ 48.1/ 46.2 61.0/ 69.0/ 69.0 49.9/ 53.1/ 57.9 54.6/ 57.6/ 64.3
Computer Science 38.0/ 41.1/ 40.6 49.6/ 52.0/ 52.4 51.0/ 53.0/ 51.0 42.8/ 47.4/ 51.0 52.7/ 58.0/ 61.5 39.0/ 48.0/ 43.0 43.2/ 49.7/ 43.2 41.0/ 47.0/ 48.0 51.0/ 54.0/ 52.0 42.4/ 47.1/ 45.4 39.9/ 45.7/ 40.8
Defense and Security 40.0/ 48.0/ 52.0 63.0/ 79.0/ 78.0 41.0/ 40.0/ 37.0 45.5/ 48.8/ 52.1 41.0/ 45.0/ 55.0 36.0/ 42.0/ 41.0 39.0/ 45.0/ 49.0 35.0/ 41.0/ 43.0 42.0/ 46.0/ 41.0 38.0/ 44.0/ 44.0 43.0/ 54.0/ 58.0
Earth Sciences 46.0/ 51.0/ 54.0 73.6/ 77.0/ 77.0 42.0/ 47.0/ 50.0 51.1/ 56.3/ 57.7 55.0/ 57.0/ 65.0 45.0/ 58.0/ 56.0 49.0/ 55.0/ 57.0 36.4/ 32.3/ 35.4 48.0/ 48.0/ 49.0 53.0/ 56.0/ 59.0 44.0/ 49.0/ 45.0
Economics 57.6/ 56.1/ 59.0 64.9/ 72.5/ 73.5 47.0/ 48.0/ 53.0 47.0/ 55.6/ 56.9 57.0/ 58.0/ 62.0 43.0/ 50.0/ 49.0 45.8/ 54.2/ 49.3 36.0/ 33.0/ 34.0 48.0/ 49.0/ 52.0 46.8/ 52.4/ 54.8 47.9/ 45.5/ 47.9
Education 41.0/ 44.0/ 45.0 53.8/ 58.3/ 58.9 33.0/ 35.0/ 39.0 42.3/ 45.6/ 46.5 44.0/ 42.0/ 43.0 42.0/ 43.0/ 45.0 40.8/ 49.0/ 48.4 36.0/ 40.0/ 35.0 36.0/ 40.0/ 40.0 40.0/ 42.0/ 43.0 41.0/ 37.0/ 37.0
Energy and Power 59.0/ 60.0/ 59.0 69.6/ 79.0/ 79.0 56.0/ 66.0/ 66.0 52.9/ 58.6/ 60.0 52.0/ 60.0/ 60.0 46.0/ 60.0/ 58.0 47.0/ 58.0/ 56.0 45.0/ 41.0/ 52.0 51.0/ 52.0/ 54.0 43.0/ 46.0/ 51.0 57.0/ 54.0/ 60.0
Engineering 41.8/ 51.2/ 48.6 54.8/ 54.6/ 59.2 43.0/ 48.0/ 49.6 41.8/ 44.4/ 46.8 41.2/ 46.6/ 48.2 43.6/ 47.0/ 48.7 40.2/ 47.2/ 47.6 33.6/ 30.9/ 34.5 41.0/ 45.0/ 43.0 35.8/ 39.6/ 43.8 40.4/ 46.2/ 49.6
Environmental Science 45.0/ 52.0/ 55.0 69.1/ 74.8/ 74.0 46.0/ 56.0/ 59.0 55.5/ 59.9/ 61.3 51.6/ 53.9/ 54.7 46.0/ 53.0/ 57.0 46.8/ 52.1/ 52.1 40.4/ 44.4/ 40.4 46.0/ 51.0/ 53.0 47.0/ 50.0/ 52.0 40.8/ 44.7/ 48.0
Ethics and Human Rights 49.0/ 57.0/ 56.0 58.0/ 67.0/ 66.0 46.0/ 57.0/ 54.0 51.0/ 56.0/ 51.0 43.0/ 52.0/ 55.0 40.0/ 50.0/ 51.0 37.0/ 46.0/ 49.0 29.0/ 36.0/ 37.0 44.0/ 52.0/ 55.0 45.0/ 51.0/ 53.0 41.0/ 46.0/ 45.0
Finance and Investment 51.0/ 49.0/ 52.0 56.6/ 60.8/ 62.2 47.0/ 48.0/ 56.0 45.6/ 48.4/ 51.4 49.4/ 55.6/ 59.9 47.0/ 52.0/ 53.0 42.9/ 46.3/ 50.8 41.0/ 36.0/ 40.0 44.0/ 45.0/ 41.0 45.8/ 48.6/ 53.1 37.1/ 39.3/ 45.4
Food Science 58.0/ 52.0/ 62.0 72.0/ 79.0/ 85.0 53.0/ 60.0/ 62.0 68.7/ 75.8/ 78.8 49.0/ 58.0/ 59.0 52.0/ 58.0/ 61.0 57.0/ 66.0/ 59.0 52.0/ 51.0/ 51.0 53.0/ 64.0/ 65.0 48.0/ 55.0/ 56.0 52.0/ 62.0/ 63.0
Geography 42.5/ 48.0/ 49.7 64.4/ 70.6/ 72.4 49.0/ 52.0/ 50.0 48.0/ 53.0/ 53.4 48.4/ 52.5/ 59.4 40.4/ 43.4/ 48.5 42.0/ 47.9/ 52.1 43.0/ 42.0/ 43.0 39.0/ 45.0/ 42.0 46.9/ 52.0/ 51.4 42.2/ 42.2/ 49.0
Health and Medicine 59.0/ 60.0/ 69.0 79.0/ 83.0/ 84.0 51.0/ 58.0/ 55.0 67.7/ 73.6/ 73.0 63.0/ 66.0/ 71.0 70.0/ 71.0/ 71.0 49.1/ 53.6/ 60.7 54.0/ 50.0/ 53.0 65.0/ 72.0/ 71.0 56.9/ 61.2/ 59.5 58.9/ 59.8/ 56.1
History 46.1/ 50.1/ 53.3 60.2/ 66.9/ 67.5 53.0/ 56.0/ 53.0 51.4/ 57.3/ 58.7 49.1/ 52.0/ 54.9 36.0/ 42.0/ 48.0 40.5/ 44.2/ 43.9 38.0/ 42.0/ 40.0 44.0/ 47.0/ 46.0 43.6/ 46.6/ 49.5 39.1/ 44.2/ 44.5
Information Technology 62.0/ 64.0/ 70.0 74.0/ 78.0/ 80.0 61.0/ 65.0/ 70.0 62.6/ 67.0/ 72.1 61.0/ 72.0/ 70.0 65.0/ 69.0/ 69.0 61.0/ 68.0/ 73.0 62.0/ 55.0/ 58.0 55.0/ 69.0/ 66.0 69.0/ 60.0/ 64.0 55.0/ 71.0/ 72.0
International Relations 57.0/ 55.0/ 62.0 72.5/ 77.5/ 79.2 48.0/ 57.0/ 61.0 52.5/ 54.9/ 58.0 39.0/ 52.0/ 55.0 52.0/ 50.0/ 58.0 57.0/ 58.0/ 58.0 42.0/ 50.0/ 53.0 52.0/ 50.0/ 51.0 54.0/ 59.0/ 53.0 41.0/ 57.0/ 60.0
Language Studies 33.0/ 34.0/ 28.0 75.0/ 81.0/ 81.8 36.0/ 41.0/ 51.0 38.5/ 43.5/ 42.9 34.9/ 39.5/ 41.9 37.0/ 45.0/ 37.0 36.2/ 39.6/ 41.6 38.0/ 38.0/ 44.0 40.0/ 51.0/ 47.0 35.2/ 37.8/ 41.2 39.7/ 43.1/ 40.2
Law and Ethics 64.9/ 67.6/ 71.0 67.8/ 75.2/ 75.2 56.0/ 60.0/ 57.0 58.2/ 65.2/ 66.4 55.4/ 58.6/ 63.1 47.0/ 52.0/ 53.0 44.0/ 49.6/ 49.1 40.0/ 41.0/ 38.0 48.0/ 56.0/ 54.0 51.2/ 59.8/ 56.7 38.1/ 41.9/ 41.9
Literature and Linguistics 36.0/ 40.5/ 40.5 60.7/ 64.1/ 65.9 34.0/ 39.0/ 41.0 36.5/ 39.7/ 45.3 38.3/ 38.3/ 40.0 27.0/ 30.0/ 33.0 33.6/ 34.5/ 35.4 36.0/ 37.0/ 40.0 37.0/ 45.0/ 43.0 35.6/ 35.6/ 34.8 31.7/ 35.9/ 30.3
Logical Reasoning 33.9/ 38.6/ 39.1 40.8/ 40.4/ 39.8 27.3/ 37.7/ 40.3 34.9/ 41.9/ 40.9 30.3/ 31.9/ 38.6 26.0/ 28.0/ 32.0 31.2/ 31.5/ 32.0 21.0/ 22.0/ 17.0 28.0/ 35.0/ 32.0 33.5/ 33.5/ 35.5 33.9/ 34.9/ 36.6
Materials Science 55.0/ 56.0/ 56.0 70.3/ 71.1/ 73.8 46.0/ 46.0/ 48.0 49.8/ 55.0/ 55.5 45.0/ 45.0/ 46.0 45.0/ 54.0/ 54.0 51.0/ 57.0/ 52.0 34.0/ 40.0/ 40.0 50.0/ 50.0/ 55.0 37.0/ 52.0/ 53.0 50.0/ 55.0/ 59.0
Mathematics 44.6/ 45.4/ 45.8 41.0/ 43.6/ 43.8 42.2/ 42.0/ 45.0 38.1/ 39.7/ 41.3 40.8/ 46.0/ 44.4 42.7/ 51.9/ 49.0 39.2/ 44.0/ 43.6 33.6/ 36.6/ 36.4 36.4/ 36.0/ 35.2 38.0/ 40.4/ 42.2 36.0/ 38.2/ 38.0
Media and Communication 54.0/ 56.0/ 57.0 75.0/ 78.0/ 78.0 54.0/ 54.0/ 56.0 51.7/ 55.8/ 61.2 52.0/ 53.0/ 60.0 50.0/ 52.0/ 59.0 44.0/ 53.0/ 55.0 47.0/ 52.0/ 52.0 50.0/ 42.0/ 38.0 46.0/ 50.0/ 54.0 55.0/ 60.0/ 58.0
Music and Performing Arts 32.0/ 31.0/ 35.0 63.0/ 62.0/ 60.0 39.0/ 40.0/ 41.0 33.1/ 42.4/ 37.1 30.0/ 35.0/ 35.0 28.0/ 35.0/ 39.0 42.0/ 35.0/ 38.0 35.0/ 35.0/ 35.0 37.0/ 38.0/ 42.0 32.0/ 23.0/ 28.0 37.0/ 41.0/ 39.0
Physics 56.0/ 58.6/ 61.4 70.4/ 72.2/ 75.0 50.1/ 50.4/ 54.0 55.2/ 56.8/ 63.6 46.8/ 58.0/ 60.4 48.9/ 57.4/ 57.9 55.4/ 58.2/ 59.8 35.1/ 32.6/ 44.1 53.0/ 52.0/ 52.0 50.8/ 56.6/ 60.2 55.0/ 54.2/ 58.4
Politics and Governance 50.6/ 52.9/ 58.8 60.1/ 63.9/ 65.3 41.0/ 52.0/ 47.0 49.0/ 52.6/ 54.0 47.4/ 50.6/ 52.6 49.0/ 44.0/ 51.0 38.5/ 43.8/ 45.5 40.0/ 36.0/ 40.0 49.0/ 54.0/ 52.0 44.4/ 47.7/ 45.2 39.1/ 39.9/ 42.2
Psychology 53.0/ 57.0/ 62.0 75.0/ 82.0/ 82.0 37.0/ 40.0/ 46.0 51.9/ 55.0/ 62.0 44.0/ 57.0/ 53.0 50.0/ 54.0/ 49.0 50.0/ 52.0/ 52.0 44.0/ 48.0/ 46.0 54.0/ 57.0/ 56.0 41.0/ 45.0/ 48.0 55.0/ 62.0/ 62.0
Public Administration 39.0/ 42.0/ 43.0 60.6/ 62.6/ 63.6 47.5/ 48.5/ 46.5 46.5/ 50.5/ 48.5 42.4/ 40.4/ 42.4 29.0/ 37.0/ 34.0 36.4/ 39.4/ 37.4 36.0/ 45.0/ 39.0 44.4/ 51.5/ 45.5 36.0/ 48.0/ 55.0 36.0/ 54.0/ 50.0
Religion and Spirituality 48.0/ 55.0/ 59.0 73.0/ 81.0/ 80.0 47.0/ 61.0/ 59.0 60.5/ 62.9/ 60.5 49.0/ 62.0/ 67.0 46.0/ 60.0/ 63.0 47.0/ 58.0/ 54.0 37.0/ 34.0/ 31.0 51.0/ 55.0/ 50.0 48.0/ 50.0/ 48.0 48.0/ 58.0/ 62.0
Social Welfare and Development 43.0/ 51.0/ 53.0 62.5/ 71.4/ 69.6 44.0/ 51.0/ 54.0 52.8/ 57.5/ 59.1 36.0/ 42.0/ 46.0 37.0/ 40.0/ 39.0 43.0/ 45.0/ 46.0 32.0/ 36.0/ 29.0 44.0/ 50.0/ 51.0 38.0/ 43.0/ 44.0 42.0/ 46.0/ 50.0
Sociology 28.7/ 34.8/ 34.8 45.8/ 50.4/ 50.8 30.0/ 29.0/ 31.0 36.8/ 42.0/ 43.4 31.0/ 42.0/ 37.0 33.0/ 32.0/ 34.0 32.1/ 40.9/ 39.6 29.0/ 35.0/ 37.0 30.0/ 37.0/ 33.0 40.5/ 42.3/ 46.9 40.6/ 40.6/ 40.0
Sports and Recreation 40.1/ 48.0/ 50.5 71.2/ 74.2/ 75.2 40.0/ 37.0/ 42.0 47.0/ 53.2/ 57.6 38.8/ 48.9/ 55.6 41.0/ 43.0/ 42.0 49.7/ 54.8/ 55.9 30.0/ 30.0/ 38.0 46.0/ 55.0/ 51.0 48.1/ 48.1/ 51.3 41.4/ 43.8/ 45.7
Technology and Innovation 51.0/ 56.0/ 59.0 69.2/ 72.2/ 73.0 47.0/ 47.0/ 52.0 56.6/ 56.2/ 59.4 49.0/ 52.0/ 61.0 50.0/ 48.0/ 52.0 49.0/ 57.0/ 58.0 42.0/ 40.0/ 49.0 49.0/ 49.0/ 51.0 44.0/ 45.0/ 49.0 36.0/ 45.0/ 51.0
Transportation and Logistics 29.0/ 33.0/ 41.0 56.9/ 64.6/ 61.5 39.0/ 40.0/ 41.0 38.8/ 42.6/ 44.5 34.3/ 38.4/ 39.4 33.0/ 33.0/ 33.0 25.5/ 29.6/ 28.6 24.2/ 26.3/ 31.3 33.0/ 37.0/ 39.0 36.0/ 37.0/ 43.0 28.0/ 39.0/ 32.0
Table 46: Detailed subject-wise evaluation for google/gemma-2-9b on MILU across different languages. The results reported are X / Y / Z where X denote the 0-shot, Y denotes 1-shot and Z denotes 5-shot performance respectively.
Topic bn en gu hi kn ml mr or pa ta te
Agriculture 25.0/ 35.0/ 37.0 51.3/ 60.0/ 59.8 30.3/ 42.4/ 31.3 37.6/ 43.9/ 42.0 38.0/ 44.0/ 38.0 30.3/ 35.4/ 36.4 35.0/ 36.0/ 38.0 28.0/ 28.0/ 25.0 35.0/ 36.0/ 32.0 35.0/ 34.0/ 29.0 34.0/ 37.0/ 29.0
Anthropology 37.0/ 41.0/ 44.0 51.0/ 60.0/ 60.0 39.0/ 44.0/ 32.0 41.4/ 46.5/ 39.4 36.0/ 33.0/ 31.0 38.0/ 39.0/ 35.0 37.0/ 41.0/ 27.0 30.0/ 31.0/ 29.0 39.0/ 39.0/ 30.0 32.0/ 36.0/ 35.0 34.0/ 36.0/ 30.0
Architecture and Design 38.0/ 32.0/ 32.0 49.0/ 58.0/ 56.0 31.0/ 35.0/ 31.0 35.0/ 35.0/ 41.0 36.0/ 38.0/ 29.0 39.0/ 40.0/ 41.0 31.0/ 37.0/ 25.0 29.0/ 27.0/ 26.0 40.0/ 34.0/ 31.0 33.0/ 37.0/ 27.0 27.0/ 31.0/ 32.0
Arts and Culture 35.1/ 38.8/ 39.4 55.6/ 65.8/ 67.0 33.0/ 35.0/ 40.0 37.3/ 42.1/ 40.9 36.2/ 38.0/ 35.6 36.0/ 35.0/ 36.0 30.6/ 30.6/ 35.5 31.0/ 32.0/ 22.0 34.0/ 43.0/ 43.0 39.4/ 33.9/ 37.0 32.7/ 31.2/ 34.7
Astronomy and Astrophysics 47.0/ 56.0/ 58.0 70.0/ 83.0/ 78.0 59.0/ 67.0/ 60.0 46.1/ 52.3/ 51.6 59.0/ 63.0/ 65.0 47.0/ 57.0/ 51.0 44.0/ 45.0/ 44.0 36.0/ 44.0/ 44.0 40.0/ 48.0/ 55.0 46.0/ 55.0/ 54.0 53.0/ 54.0/ 54.0
Biology 40.1/ 47.9/ 53.5 65.9/ 78.2/ 80.4 32.0/ 42.0/ 43.0 48.1/ 54.7/ 58.1 41.8/ 43.0/ 42.6 48.5/ 49.5/ 53.5 34.9/ 37.9/ 36.7 35.0/ 37.0/ 28.0 48.0/ 49.0/ 51.0 38.9/ 40.8/ 39.2 40.6/ 39.6/ 43.1
Business and Management 31.0/ 33.0/ 36.0 51.4/ 57.0/ 58.8 31.0/ 32.0/ 36.0 41.9/ 40.8/ 42.6 32.0/ 34.0/ 25.0 42.0/ 41.0/ 40.0 41.0/ 41.0/ 43.0 33.0/ 29.0/ 34.0 40.0/ 42.0/ 41.0 44.0/ 40.0/ 36.0 32.1/ 41.5/ 29.2
Chemistry 43.7/ 52.2/ 56.3 63.5/ 74.6/ 79.4 44.0/ 42.2/ 43.1 49.9/ 54.9/ 53.3 41.2/ 42.2/ 49.0 44.0/ 51.0/ 53.0 37.7/ 42.7/ 42.2 39.4/ 34.6/ 43.3 47.0/ 53.0/ 51.0 44.8/ 48.4/ 45.7 42.1/ 46.6/ 47.6
Computer Science 34.4/ 39.1/ 39.1 43.4/ 48.0/ 48.2 44.0/ 50.0/ 48.0 35.2/ 42.2/ 42.2 39.1/ 48.5/ 48.5 37.0/ 43.0/ 41.0 33.7/ 40.8/ 39.6 41.0/ 43.0/ 44.0 49.0/ 47.0/ 49.0 41.9/ 42.4/ 37.8 40.4/ 37.7/ 36.8
Defense and Security 38.0/ 36.0/ 40.0 53.0/ 63.0/ 64.0 39.0/ 39.0/ 32.0 37.0/ 36.5/ 38.4 46.0/ 42.0/ 44.0 28.0/ 34.0/ 36.0 34.0/ 28.0/ 31.0 39.0/ 30.0/ 37.0 36.0/ 44.0/ 38.0 41.0/ 42.0/ 34.0 49.0/ 45.0/ 36.0
Earth Sciences 43.0/ 43.0/ 44.0 53.6/ 71.6/ 73.2 32.0/ 40.0/ 34.0 42.5/ 46.9/ 47.1 48.0/ 38.0/ 52.0 37.0/ 43.0/ 44.0 44.0/ 43.0/ 42.0 31.3/ 30.3/ 27.3 34.0/ 31.0/ 29.0 40.0/ 52.0/ 43.0 44.0/ 44.0/ 39.0
Economics 41.0/ 46.8/ 48.2 48.8/ 61.0/ 63.0 34.0/ 42.0/ 36.0 42.3/ 47.4/ 48.4 49.0/ 52.0/ 58.0 33.0/ 37.0/ 40.0 33.8/ 39.4/ 43.7 37.0/ 33.0/ 40.0 39.0/ 45.0/ 48.0 45.2/ 49.2/ 38.1 35.9/ 39.5/ 37.1
Education 39.0/ 40.0/ 49.0 43.0/ 51.6/ 53.5 31.0/ 41.0/ 37.0 35.5/ 36.7/ 40.8 44.0/ 38.0/ 38.0 38.0/ 38.0/ 40.0 41.4/ 41.4/ 40.1 27.0/ 27.0/ 30.0 28.0/ 37.0/ 33.0 43.0/ 40.0/ 34.0 38.0/ 40.0/ 42.0
Energy and Power 49.0/ 51.0/ 56.0 54.7/ 69.6/ 69.6 39.0/ 43.0/ 42.0 35.9/ 45.1/ 50.5 36.0/ 47.0/ 48.0 41.0/ 44.0/ 40.0 41.0/ 47.0/ 45.0 27.0/ 32.0/ 37.0 37.0/ 45.0/ 43.0 37.0/ 35.0/ 43.0 42.0/ 36.0/ 39.0
Engineering 30.4/ 38.0/ 41.0 43.6/ 50.0/ 52.8 35.8/ 36.4/ 34.0 33.2/ 32.0/ 38.8 35.0/ 35.8/ 39.2 40.6/ 37.6/ 38.9 34.4/ 33.2/ 33.6 28.9/ 27.1/ 31.6 38.0/ 34.0/ 39.0 32.8/ 32.0/ 36.0 34.2/ 37.4/ 37.8
Environmental Science 34.0/ 37.0/ 42.0 55.7/ 65.3/ 68.1 37.0/ 42.0/ 37.0 46.9/ 53.3/ 50.5 39.1/ 36.7/ 43.8 45.0/ 44.0/ 41.0 41.0/ 41.0/ 43.2 34.3/ 43.4/ 41.4 46.0/ 45.0/ 47.0 42.0/ 45.0/ 48.0 44.7/ 46.9/ 43.0
Ethics and Human Rights 38.0/ 41.0/ 44.0 44.0/ 54.0/ 50.0 44.0/ 47.0/ 47.0 51.0/ 48.0/ 47.0 44.0/ 38.0/ 46.0 39.0/ 39.0/ 40.0 39.0/ 41.0/ 41.0 33.0/ 30.0/ 34.0 40.0/ 39.0/ 43.0 39.0/ 40.0/ 41.0 46.0/ 44.0/ 44.0
Finance and Investment 47.0/ 49.5/ 50.0 49.2/ 57.0/ 58.8 42.0/ 46.0/ 47.0 37.2/ 42.0/ 45.8 45.1/ 49.4/ 52.5 42.0/ 45.0/ 48.0 37.3/ 41.8/ 45.2 35.0/ 29.0/ 36.0 29.0/ 29.0/ 37.0 36.9/ 45.2/ 46.4 35.4/ 37.5/ 36.8
Food Science 36.0/ 49.0/ 51.0 65.0/ 78.0/ 82.0 37.0/ 43.0/ 38.0 37.4/ 54.5/ 54.5 44.0/ 45.0/ 42.0 44.0/ 44.0/ 41.0 36.0/ 39.0/ 42.0 35.0/ 34.0/ 36.0 37.0/ 51.0/ 41.0 40.0/ 45.0/ 37.0 43.0/ 49.0/ 45.0
Geography 34.6/ 48.0/ 45.2 50.4/ 62.6/ 63.6 35.0/ 43.0/ 40.0 38.6/ 43.2/ 44.4 40.6/ 45.2/ 41.1 36.4/ 38.4/ 35.4 30.2/ 31.4/ 34.9 29.0/ 29.0/ 30.0 38.0/ 32.0/ 32.0 36.6/ 43.4/ 36.0 37.5/ 40.1/ 37.5
Health and Medicine 43.0/ 56.0/ 57.0 64.9/ 76.3/ 77.0 41.0/ 52.0/ 48.0 50.5/ 59.5/ 59.3 47.0/ 50.0/ 55.0 48.0/ 54.0/ 51.0 33.0/ 35.7/ 30.4 42.0/ 48.0/ 45.0 47.0/ 55.0/ 50.0 48.3/ 39.7/ 39.7 49.5/ 46.7/ 49.5
History 44.0/ 41.9/ 44.0 52.2/ 57.8/ 59.4 43.0/ 37.0/ 40.0 39.1/ 42.9/ 45.2 45.8/ 44.3/ 47.2 41.0/ 37.0/ 32.0 33.8/ 33.8/ 34.6 33.0/ 28.0/ 30.0 37.0/ 42.0/ 40.0 43.6/ 40.7/ 37.1 37.3/ 35.3/ 39.1
Information Technology 44.0/ 52.0/ 57.0 58.0/ 73.0/ 75.0 46.0/ 48.0/ 49.0 53.1/ 59.2/ 62.6 54.0/ 52.0/ 57.0 55.0/ 58.0/ 55.0 48.0/ 56.0/ 52.0 45.0/ 45.0/ 47.0 47.0/ 50.0/ 53.0 49.0/ 51.0/ 42.0 39.0/ 49.0/ 45.0
International Relations 48.0/ 53.0/ 52.0 57.5/ 72.5/ 74.2 46.0/ 48.0/ 47.0 46.3/ 54.5/ 49.0 35.0/ 47.0/ 47.0 39.0/ 48.0/ 36.0 47.0/ 51.0/ 51.0 41.0/ 44.0/ 45.0 44.0/ 46.0/ 39.0 52.0/ 45.0/ 45.0 39.0/ 37.0/ 30.0
Language Studies 33.0/ 39.0/ 46.0 59.8/ 69.6/ 71.8 34.0/ 37.0/ 38.0 32.3/ 37.9/ 37.7 27.9/ 32.6/ 36.4 36.0/ 36.0/ 33.0 32.9/ 26.9/ 27.5 43.0/ 39.0/ 39.0 36.0/ 45.0/ 45.0 28.5/ 28.8/ 29.6 28.7/ 29.9/ 33.9
Law and Ethics 45.3/ 51.3/ 54.7 57.2/ 63.4/ 64.8 46.0/ 48.0/ 42.0 49.8/ 52.4/ 56.0 43.3/ 47.8/ 54.1 46.0/ 44.0/ 45.0 39.2/ 44.8/ 40.1 27.0/ 28.0/ 31.0 48.0/ 55.0/ 43.0 43.3/ 52.0/ 49.6 34.4/ 35.0/ 33.8
Literature and Linguistics 29.7/ 33.3/ 37.8 50.1/ 59.9/ 61.7 27.0/ 37.0/ 36.0 32.5/ 33.3/ 36.1 37.5/ 35.8/ 40.8 37.0/ 31.0/ 39.0 31.9/ 33.6/ 27.4 36.0/ 34.0/ 40.0 30.0/ 40.0/ 37.0 29.5/ 29.5/ 28.3 33.1/ 40.1/ 33.8
Logical Reasoning 36.1/ 42.8/ 39.3 32.0/ 36.8/ 40.2 33.8/ 34.4/ 36.4 37.1/ 39.7/ 38.1 36.5/ 36.5/ 40.2 42.0/ 36.0/ 40.0 32.9/ 37.3/ 33.4 23.0/ 28.0/ 25.0 42.0/ 32.0/ 35.0 35.9/ 41.4/ 40.6 36.4/ 37.4/ 37.9
Materials Science 39.0/ 38.0/ 46.0 52.5/ 62.4/ 65.4 35.0/ 34.0/ 31.0 41.0/ 43.0/ 42.4 45.0/ 41.0/ 37.0 35.0/ 45.0/ 46.0 43.0/ 39.0/ 39.0 29.0/ 25.0/ 32.0 37.0/ 46.0/ 44.0 36.0/ 37.0/ 34.0 42.0/ 34.0/ 39.0
Mathematics 37.6/ 44.0/ 46.6 37.6/ 43.6/ 49.0 31.6/ 34.6/ 41.0 30.9/ 34.9/ 40.1 30.4/ 37.0/ 41.4 34.7/ 39.5/ 45.0 33.2/ 40.6/ 43.6 32.2/ 34.4/ 39.2 34.5/ 38.3/ 38.3 32.8/ 39.0/ 41.4 29.4/ 30.2/ 32.6
Media and Communication 34.0/ 44.0/ 50.0 61.0/ 71.0/ 72.0 42.0/ 43.0/ 46.0 44.9/ 45.6/ 41.5 40.0/ 43.0/ 41.0 44.0/ 46.0/ 44.0 35.0/ 39.0/ 38.0 34.0/ 37.0/ 38.0 40.0/ 44.0/ 51.0 37.0/ 37.0/ 35.0 50.0/ 51.0/ 46.0
Music and Performing Arts 38.0/ 23.0/ 25.0 40.0/ 52.0/ 55.0 36.0/ 38.0/ 29.0 29.8/ 36.4/ 30.5 24.0/ 22.0/ 20.0 33.0/ 26.0/ 32.0 29.0/ 34.0/ 31.0 28.0/ 25.0/ 22.0 39.0/ 43.0/ 34.0 36.0/ 41.0/ 31.0 45.0/ 35.0/ 31.0
Physics 41.6/ 47.6/ 52.6 54.8/ 67.2/ 68.8 40.9/ 42.6/ 43.2 38.2/ 42.0/ 45.4 39.4/ 42.8/ 48.4 41.0/ 43.2/ 48.9 41.4/ 41.4/ 44.8 26.7/ 30.6/ 37.9 43.0/ 38.0/ 44.0 40.8/ 45.0/ 47.0 42.6/ 44.4/ 48.4
Politics and Governance 41.6/ 53.7/ 52.5 49.5/ 60.5/ 60.3 34.0/ 37.0/ 32.0 34.9/ 46.6/ 42.8 33.6/ 33.6/ 40.1 36.0/ 45.0/ 40.0 30.6/ 39.0/ 33.7 28.0/ 37.0/ 34.0 40.0/ 43.0/ 40.0 44.8/ 45.6/ 45.2 39.4/ 41.5/ 38.0
Psychology 34.0/ 44.0/ 47.0 61.0/ 73.0/ 75.0 38.0/ 39.0/ 32.0 36.4/ 36.4/ 41.1 38.0/ 47.0/ 57.0 36.0/ 36.0/ 37.0 33.0/ 31.0/ 37.0 42.0/ 36.0/ 42.0 42.0/ 45.0/ 43.0 33.0/ 38.0/ 34.0 46.0/ 43.0/ 49.0
Public Administration 26.0/ 34.0/ 36.0 49.5/ 51.5/ 53.5 39.4/ 36.4/ 34.3 38.4/ 46.5/ 43.4 35.4/ 35.4/ 39.4 33.0/ 32.0/ 33.0 31.3/ 32.3/ 34.3 25.0/ 21.0/ 28.0 28.3/ 32.3/ 27.3 31.0/ 34.0/ 29.0 40.0/ 41.0/ 40.0
Religion and Spirituality 44.0/ 49.0/ 48.0 62.0/ 71.0/ 73.0 45.0/ 43.0/ 40.0 44.8/ 49.2/ 46.0 47.0/ 48.0/ 55.0 43.0/ 35.0/ 39.0 38.0/ 38.0/ 38.0 27.0/ 31.0/ 28.0 48.0/ 51.0/ 45.0 34.0/ 36.0/ 31.0 46.0/ 53.0/ 41.0
Social Welfare and Development 33.0/ 39.0/ 46.0 48.2/ 58.0/ 61.6 41.0/ 45.0/ 52.0 38.3/ 39.9/ 39.4 43.0/ 37.0/ 44.0 38.0/ 40.0/ 39.0 37.0/ 34.0/ 33.0 37.0/ 39.0/ 38.0 41.0/ 42.0/ 41.0 37.0/ 48.0/ 43.0 35.0/ 40.0/ 37.0
Sociology 22.6/ 25.2/ 32.2 38.0/ 45.6/ 47.2 35.0/ 28.0/ 33.0 31.2/ 30.0/ 29.4 29.0/ 23.0/ 27.0 22.0/ 19.0/ 24.0 28.3/ 30.2/ 30.8 21.0/ 22.0/ 30.0 27.0/ 34.0/ 34.0 35.1/ 34.2/ 26.1 28.8/ 32.4/ 35.3
Sports and Recreation 36.6/ 43.1/ 46.0 53.4/ 68.2/ 68.4 36.0/ 34.0/ 33.0 39.0/ 44.2/ 38.4 32.0/ 40.5/ 39.3 33.0/ 37.0/ 40.0 39.0/ 45.2/ 33.3 29.0/ 29.0/ 30.0 41.0/ 44.0/ 29.0 28.8/ 39.1/ 31.4 32.9/ 37.1/ 35.7
Technology and Innovation 37.0/ 40.0/ 51.0 54.4/ 64.0/ 67.8 37.0/ 42.0/ 36.0 43.4/ 49.2/ 49.2 40.0/ 43.0/ 45.0 46.0/ 46.0/ 47.0 40.0/ 46.0/ 48.0 34.0/ 38.0/ 37.0 50.0/ 45.0/ 43.0 37.0/ 38.0/ 40.0 39.0/ 36.0/ 45.0
Transportation and Logistics 42.0/ 38.0/ 36.0 43.1/ 52.3/ 53.8 31.0/ 25.0/ 34.0 31.6/ 35.6/ 37.2 29.3/ 30.3/ 31.3 41.0/ 38.0/ 36.0 30.6/ 34.7/ 33.7 25.2/ 30.3/ 26.3 27.0/ 27.0/ 27.0 34.0/ 39.0/ 38.0 30.0/ 27.0/ 27.0
Table 47: Detailed subject-wise evaluation for google/gemma-2-9b-it on MILU across different languages. The results reported are X / Y / Z where X denote the 0-shot, Y denotes 1-shot and Z denotes 5-shot performance respectively.
Topic bn en gu hi kn ml mr or pa ta te
Agriculture 47.0/ 55.0/ 53.0 66.4/ 68.8/ 69.4 46.5/ 58.6/ 54.5 58.1/ 61.6/ 63.6 51.0/ 55.0/ 55.0 50.5/ 59.6/ 60.6 48.0/ 50.0/ 48.0 37.0/ 39.0/ 35.0 51.0/ 55.0/ 59.0 44.0/ 49.0/ 48.0 51.0/ 49.0/ 55.0
Anthropology 54.0/ 62.0/ 56.0 67.0/ 72.0/ 73.0 52.0/ 61.0/ 54.0 55.6/ 55.6/ 57.6 48.0/ 53.0/ 48.0 52.0/ 58.0/ 53.0 54.0/ 60.0/ 60.0 45.0/ 41.0/ 43.0 49.0/ 54.0/ 56.0 49.0/ 57.0/ 56.0 46.0/ 48.0/ 50.0
Architecture and Design 45.0/ 57.0/ 55.0 67.0/ 70.0/ 68.0 49.0/ 55.0/ 56.0 53.0/ 61.0/ 60.0 45.0/ 42.0/ 47.0 52.0/ 54.0/ 54.0 53.0/ 58.0/ 60.0 28.0/ 29.0/ 28.0 48.0/ 45.0/ 49.0 48.0/ 54.0/ 56.0 39.0/ 43.0/ 48.0
Arts and Culture 50.3/ 53.9/ 59.4 70.2/ 78.2/ 77.0 52.0/ 62.0/ 55.0 54.9/ 60.5/ 61.3 52.8/ 59.5/ 63.8 54.0/ 48.0/ 49.0 57.9/ 59.6/ 64.5 30.0/ 25.0/ 24.0 40.0/ 53.0/ 51.0 45.5/ 50.3/ 49.1 45.7/ 47.2/ 51.3
Astronomy and Astrophysics 73.0/ 76.0/ 74.0 88.0/ 94.0/ 93.0 78.0/ 79.0/ 78.0 72.7/ 75.0/ 75.8 68.0/ 75.0/ 78.0 71.0/ 71.0/ 77.0 70.0/ 73.0/ 76.0 63.0/ 61.0/ 69.0 67.0/ 76.0/ 77.0 73.0/ 83.0/ 85.0 70.0/ 76.0/ 78.0
Biology 68.3/ 72.9/ 72.9 84.6/ 86.4/ 86.4 60.0/ 72.0/ 71.0 77.0/ 81.6/ 82.4 66.0/ 73.4/ 72.1 70.7/ 74.8/ 72.7 57.3/ 62.4/ 64.5 53.0/ 52.0/ 47.0 70.0/ 71.0/ 73.0 61.8/ 66.9/ 66.9 61.6/ 64.8/ 64.1
Business and Management 55.0/ 55.0/ 56.0 68.7/ 67.8/ 68.7 49.0/ 62.0/ 56.0 57.6/ 58.3/ 59.0 54.0/ 48.0/ 55.0 57.0/ 54.0/ 59.0 46.0/ 59.0/ 65.0 37.0/ 39.0/ 43.0 56.0/ 55.0/ 60.0 45.0/ 53.0/ 55.0 33.0/ 34.9/ 42.4
Chemistry 65.1/ 71.7/ 73.6 81.8/ 86.2/ 85.0 63.8/ 66.4/ 67.2 73.6/ 77.1/ 77.6 62.4/ 69.6/ 70.6 68.0/ 69.0/ 71.0 62.8/ 68.3/ 73.1 44.2/ 51.0/ 55.8 71.0/ 69.0/ 72.0 60.2/ 65.0/ 67.4 67.6/ 69.1/ 72.3
Computer Science 45.3/ 53.6/ 53.6 49.8/ 57.8/ 57.4 55.0/ 65.0/ 64.0 47.4/ 53.4/ 55.6 51.5/ 59.2/ 61.0 45.0/ 59.0/ 57.0 47.9/ 52.7/ 55.6 42.0/ 56.0/ 54.0 66.0/ 66.0/ 61.0 40.7/ 48.8/ 52.9 41.3/ 48.9/ 53.4
Defense and Security 60.0/ 61.0/ 69.0 81.0/ 86.0/ 86.0 50.0/ 61.0/ 60.0 55.5/ 58.3/ 63.5 51.0/ 60.0/ 66.0 47.0/ 54.0/ 61.0 52.0/ 65.0/ 65.0 37.0/ 38.0/ 43.0 49.0/ 67.0/ 66.0 58.0/ 62.0/ 60.0 53.0/ 66.0/ 69.0
Earth Sciences 58.0/ 67.0/ 66.0 77.0/ 83.0/ 83.4 55.0/ 54.0/ 55.0 60.7/ 63.7/ 67.7 69.0/ 71.0/ 76.0 59.0/ 64.0/ 64.0 63.0/ 65.0/ 63.0 40.4/ 37.4/ 41.4 54.0/ 62.0/ 63.0 64.0/ 64.0/ 67.0 62.0/ 66.0/ 64.0
Economics 63.3/ 69.1/ 74.8 74.5/ 79.3/ 79.1 55.0/ 56.0/ 57.0 64.7/ 68.2/ 69.8 64.0/ 72.0/ 74.0 53.0/ 54.0/ 54.0 52.8/ 56.3/ 54.2 41.0/ 38.0/ 42.0 59.0/ 60.0/ 57.0 57.1/ 59.5/ 63.5 55.7/ 55.7/ 62.3
Education 53.0/ 57.0/ 57.0 60.2/ 65.9/ 67.8 39.0/ 48.0/ 40.0 48.5/ 54.1/ 53.5 45.0/ 48.0/ 50.0 49.0/ 54.0/ 53.0 49.0/ 58.6/ 56.7 36.0/ 37.0/ 44.0 46.0/ 51.0/ 47.0 42.0/ 50.0/ 53.0 47.0/ 53.0/ 58.0
Energy and Power 59.0/ 63.0/ 71.0 72.3/ 80.4/ 83.8 65.0/ 67.0/ 70.0 61.0/ 69.2/ 72.5 62.0/ 73.0/ 70.0 63.0/ 75.0/ 76.0 64.0/ 63.0/ 65.0 42.0/ 53.0/ 56.0 58.0/ 58.0/ 61.0 54.0/ 59.0/ 60.0 63.0/ 67.0/ 66.0
Engineering 49.0/ 55.8/ 58.2 59.6/ 62.2/ 64.6 54.8/ 58.2/ 57.8 49.8/ 51.6/ 54.4 52.4/ 56.2/ 59.4 60.7/ 55.1/ 57.3 53.8/ 56.4/ 57.2 37.2/ 40.2/ 44.6 43.0/ 44.0/ 45.0 48.0/ 50.4/ 52.2 57.2/ 60.0/ 60.8
Environmental Science 52.0/ 67.0/ 64.0 75.0/ 78.0/ 79.4 56.0/ 62.0/ 57.0 65.7/ 68.3/ 69.3 58.6/ 59.4/ 66.4 58.0/ 60.0/ 57.0 54.2/ 65.3/ 66.3 41.4/ 42.4/ 48.5 57.0/ 64.0/ 64.0 51.0/ 55.0/ 58.0 53.1/ 53.6/ 58.7
Ethics and Human Rights 49.0/ 58.0/ 58.0 65.0/ 77.0/ 79.0 58.0/ 63.0/ 60.0 54.0/ 57.0/ 59.0 57.0/ 57.0/ 61.0 57.0/ 62.0/ 62.0 51.0/ 59.0/ 57.0 38.0/ 45.0/ 45.0 49.0/ 54.0/ 60.0 52.0/ 56.0/ 60.0 47.0/ 62.0/ 61.0
Finance and Investment 56.9/ 60.9/ 62.4 62.6/ 65.2/ 65.4 54.0/ 53.0/ 56.0 56.0/ 59.6/ 62.8 64.8/ 69.1/ 66.7 58.0/ 57.0/ 55.0 49.7/ 58.2/ 61.0 41.0/ 51.0/ 53.0 53.0/ 55.0/ 55.0 57.0/ 60.3/ 62.0 45.4/ 48.2/ 45.4
Food Science 56.0/ 66.0/ 70.0 82.0/ 82.0/ 81.0 59.0/ 66.0/ 68.0 69.7/ 74.8/ 75.8 67.0/ 64.0/ 64.0 64.0/ 67.0/ 71.0 65.0/ 72.0/ 72.0 47.0/ 52.0/ 54.0 64.0/ 67.0/ 72.0 63.0/ 69.0/ 66.0 63.0/ 67.0/ 76.0
Geography 59.2/ 60.9/ 67.6 68.6/ 74.4/ 75.6 60.0/ 63.0/ 57.0 58.6/ 64.9/ 64.5 58.5/ 61.2/ 66.2 53.5/ 56.6/ 54.5 54.4/ 56.8/ 58.0 45.0/ 40.0/ 49.0 47.0/ 56.0/ 58.0 54.3/ 57.1/ 59.4 51.6/ 56.8/ 59.9
Health and Medicine 71.0/ 72.0/ 73.0 83.0/ 87.8/ 87.6 65.0/ 70.0/ 70.0 75.1/ 76.5/ 80.6 71.0/ 75.0/ 76.0 71.0/ 71.0/ 75.0 58.0/ 70.5/ 69.6 53.0/ 58.0/ 58.0 70.0/ 73.0/ 70.0 62.1/ 64.7/ 62.9 65.4/ 68.2/ 66.4
History 57.9/ 61.6/ 68.8 66.5/ 70.9/ 70.3 53.0/ 62.0/ 61.0 61.9/ 64.9/ 62.1 63.0/ 65.2/ 70.0 51.0/ 58.0/ 57.0 49.4/ 52.8/ 53.5 31.0/ 41.0/ 46.0 52.0/ 60.0/ 61.0 55.6/ 62.2/ 61.8 48.9/ 53.7/ 54.0
Information Technology 79.0/ 75.0/ 80.0 83.0/ 84.0/ 86.0 67.0/ 79.0/ 77.0 75.4/ 75.4/ 82.1 77.0/ 79.0/ 83.0 77.0/ 81.0/ 76.0 69.0/ 69.0/ 68.0 64.0/ 64.0/ 62.0 73.0/ 72.0/ 74.0 67.0/ 69.0/ 73.0 67.0/ 80.0/ 78.0
International Relations 57.0/ 70.0/ 71.0 72.5/ 80.8/ 83.3 60.0/ 71.0/ 66.0 61.5/ 67.7/ 69.3 54.0/ 67.0/ 66.0 64.0/ 63.0/ 64.0 59.0/ 60.0/ 62.0 51.0/ 52.0/ 50.0 59.0/ 62.0/ 65.0 57.0/ 64.0/ 60.0 59.0/ 66.0/ 63.0
Language Studies 41.0/ 42.0/ 41.0 78.8/ 85.2/ 86.2 43.0/ 43.0/ 47.0 42.3/ 44.9/ 45.1 35.7/ 49.6/ 54.3 41.0/ 44.0/ 48.0 43.0/ 49.7/ 49.0 35.0/ 49.0/ 41.0 56.0/ 47.0/ 55.0 37.8/ 44.6/ 48.7 37.9/ 47.1/ 40.2
Law and Ethics 73.7/ 69.6/ 75.7 71.2/ 78.4/ 78.2 54.0/ 57.0/ 57.0 68.4/ 71.4/ 72.2 66.2/ 70.1/ 72.6 54.0/ 60.0/ 63.0 49.1/ 55.6/ 55.2 38.0/ 49.0/ 48.0 66.0/ 68.0/ 68.0 61.4/ 72.4/ 69.3 51.9/ 53.1/ 56.2
Literature and Linguistics 51.3/ 54.0/ 56.8 67.5/ 72.3/ 74.4 48.0/ 49.0/ 44.0 48.7/ 51.7/ 51.9 44.2/ 48.3/ 50.8 44.0/ 47.0/ 49.0 46.0/ 51.3/ 53.1 34.0/ 38.0/ 38.0 43.0/ 51.0/ 51.0 38.1/ 40.1/ 38.5 43.7/ 45.8/ 45.1
Logical Reasoning 44.5/ 48.4/ 49.6 44.0/ 49.8/ 49.8 41.6/ 46.8/ 46.1 43.7/ 45.9/ 47.7 43.1/ 44.4/ 45.6 35.0/ 39.0/ 36.0 37.6/ 38.2/ 40.1 34.0/ 33.0/ 29.0 35.0/ 42.0/ 38.0 37.9/ 42.2/ 42.6 44.8/ 46.0/ 44.3
Materials Science 62.0/ 61.0/ 68.0 71.5/ 76.8/ 76.8 44.0/ 53.0/ 49.0 56.1/ 58.3/ 61.8 55.0/ 54.0/ 58.0 55.0/ 60.0/ 56.0 58.0/ 51.0/ 62.0 37.0/ 30.0/ 40.0 56.0/ 59.0/ 58.0 52.0/ 58.0/ 59.0 53.0/ 59.0/ 59.0
Mathematics 50.6/ 54.2/ 54.8 49.4/ 50.6/ 51.8 48.8/ 52.2/ 55.8 41.5/ 41.7/ 42.7 46.8/ 50.8/ 51.6 50.4/ 56.2/ 56.2 49.6/ 55.0/ 52.4 42.4/ 46.6/ 47.4 40.9/ 44.3/ 47.0 48.4/ 48.0/ 50.8 40.6/ 40.6/ 44.6
Media and Communication 58.0/ 61.0/ 66.0 75.0/ 84.0/ 82.0 59.0/ 67.0/ 71.0 58.5/ 68.7/ 66.7 60.0/ 64.0/ 66.0 59.0/ 62.0/ 65.0 48.0/ 56.0/ 55.0 42.0/ 42.0/ 45.0 50.0/ 55.0/ 61.0 55.0/ 65.0/ 57.0 63.0/ 67.0/ 68.0
Music and Performing Arts 39.0/ 40.0/ 46.0 66.0/ 74.0/ 68.0 42.0/ 53.0/ 55.0 48.3/ 58.3/ 55.0 49.0/ 62.0/ 60.0 47.0/ 51.0/ 53.0 49.0/ 56.0/ 52.0 32.0/ 36.0/ 44.0 49.0/ 58.0/ 57.0 36.0/ 43.0/ 42.0 44.0/ 50.0/ 53.0
Physics 63.4/ 69.4/ 69.2 74.4/ 79.4/ 81.4 62.9/ 64.9/ 64.3 64.4/ 66.2/ 70.0 60.6/ 65.4/ 68.6 63.7/ 71.0/ 71.6 60.8/ 68.6/ 68.4 43.4/ 47.2/ 50.7 63.0/ 60.0/ 62.0 57.6/ 63.0/ 64.0 60.8/ 62.8/ 66.2
Politics and Governance 57.2/ 65.5/ 70.2 63.9/ 71.1/ 70.7 48.0/ 60.0/ 57.0 60.2/ 63.4/ 65.1 59.5/ 69.2/ 72.5 47.0/ 51.0/ 51.0 48.9/ 52.0/ 53.6 37.0/ 44.0/ 43.0 47.0/ 56.0/ 56.0 57.3/ 63.5/ 63.9 45.1/ 50.9/ 52.6
Psychology 58.0/ 61.0/ 59.0 81.0/ 92.0/ 89.0 55.0/ 54.0/ 58.0 57.4/ 65.1/ 65.1 60.0/ 65.0/ 64.0 60.0/ 61.0/ 60.0 54.0/ 56.0/ 60.0 53.0/ 48.0/ 43.0 62.0/ 66.0/ 65.0 53.0/ 59.0/ 60.0 70.0/ 76.0/ 71.0
Public Administration 47.0/ 41.0/ 46.0 64.6/ 70.7/ 66.7 50.5/ 58.6/ 60.6 55.6/ 54.5/ 53.5 41.4/ 42.4/ 43.4 47.0/ 49.0/ 49.0 40.4/ 46.5/ 46.5 39.0/ 43.0/ 42.0 49.5/ 63.6/ 62.6 47.0/ 44.0/ 45.0 47.0/ 53.0/ 51.0
Religion and Spirituality 58.0/ 64.0/ 64.0 73.0/ 82.0/ 88.0 60.0/ 65.0/ 66.0 63.7/ 70.6/ 68.5 66.0/ 70.0/ 69.0 61.0/ 67.0/ 68.0 62.0/ 64.0/ 67.0 36.0/ 39.0/ 37.0 68.0/ 71.0/ 69.0 64.0/ 63.0/ 62.0 64.0/ 61.0/ 69.0
Social Welfare and Development 57.0/ 56.0/ 58.0 61.6/ 73.2/ 67.9 53.0/ 61.0/ 64.0 56.5/ 62.2/ 64.8 45.0/ 49.0/ 54.0 47.0/ 51.0/ 51.0 43.0/ 56.0/ 55.0 40.0/ 35.0/ 40.0 57.0/ 58.0/ 60.0 47.0/ 56.0/ 53.0 56.0/ 52.0/ 49.0
Sociology 42.6/ 46.1/ 40.9 51.2/ 56.4/ 57.8 33.0/ 40.0/ 43.0 44.8/ 46.0/ 48.8 39.0/ 44.0/ 47.0 40.0/ 40.0/ 46.0 51.6/ 47.8/ 52.8 34.0/ 30.0/ 38.0 38.0/ 42.0/ 45.0 41.4/ 46.0/ 42.3 46.5/ 47.6/ 48.8
Sports and Recreation 56.4/ 68.3/ 69.8 74.2/ 82.0/ 83.0 47.0/ 64.0/ 60.0 57.6/ 63.8/ 69.4 54.5/ 62.9/ 68.0 54.0/ 50.0/ 57.0 65.0/ 69.5/ 71.2 38.0/ 36.0/ 35.0 46.0/ 65.0/ 65.0 55.1/ 59.6/ 59.6 52.4/ 59.1/ 61.4
Technology and Innovation 63.0/ 67.0/ 73.0 76.2/ 79.4/ 78.6 61.0/ 69.0/ 63.0 64.0/ 69.2/ 69.0 57.0/ 63.0/ 64.0 58.0/ 68.0/ 68.0 63.0/ 65.0/ 69.0 47.0/ 48.0/ 48.0 53.0/ 60.0/ 61.0 56.0/ 51.0/ 54.0 52.0/ 56.0/ 56.0
Transportation and Logistics 37.0/ 45.0/ 53.0 59.2/ 70.0/ 68.5 45.0/ 59.0/ 57.0 49.5/ 53.0/ 55.5 43.4/ 40.4/ 37.4 46.0/ 54.0/ 55.0 42.9/ 41.8/ 45.9 31.3/ 31.3/ 31.3 45.0/ 48.0/ 48.0 49.0/ 57.0/ 58.0 44.0/ 44.0/ 42.0
Table 48: Detailed subject-wise evaluation for google/gemma-2-27b on MILU across different languages. The results reported are X / Y / Z where X denote the 0-shot, Y denotes 1-shot and Z denotes 5-shot performance respectively.
Topic bn en gu hi kn ml mr or pa ta te
Agriculture 43.0/ 51.0/ 56.0 60.2/ 68.6/ 69.8 41.4/ 49.5/ 50.5 43.1/ 50.7/ 58.6 45.0/ 42.0/ 47.0 44.4/ 43.4/ 45.5 38.0/ 39.0/ 42.0 40.0/ 34.0/ 38.0 42.0/ 44.0/ 51.0 38.0/ 42.0/ 43.0 47.0/ 39.0/ 52.0
Anthropology 40.0/ 52.0/ 52.0 60.0/ 66.0/ 68.0 49.0/ 50.0/ 51.0 48.5/ 62.6/ 53.5 37.0/ 46.0/ 50.0 44.0/ 51.0/ 53.0 50.0/ 48.0/ 60.0 33.0/ 36.0/ 38.0 40.0/ 50.0/ 53.0 37.0/ 46.0/ 47.0 40.0/ 45.0/ 51.0
Architecture and Design 36.0/ 39.0/ 46.0 63.0/ 69.0/ 73.0 43.0/ 50.0/ 49.0 41.0/ 40.0/ 56.0 35.0/ 35.0/ 40.0 44.0/ 40.0/ 50.0 46.0/ 48.0/ 57.0 29.0/ 28.0/ 32.0 43.0/ 41.0/ 42.0 36.0/ 33.0/ 42.0 32.0/ 38.0/ 37.0
Arts and Culture 44.2/ 47.9/ 52.7 67.2/ 73.8/ 76.8 44.0/ 46.0/ 47.0 43.7/ 51.9/ 57.5 41.7/ 41.7/ 54.0 40.0/ 44.0/ 42.0 32.2/ 48.1/ 52.5 26.0/ 27.0/ 29.0 36.0/ 46.0/ 48.0 43.6/ 43.6/ 47.9 37.7/ 36.7/ 45.7
Astronomy and Astrophysics 58.0/ 64.0/ 73.0 75.0/ 93.0/ 91.0 55.0/ 71.0/ 79.0 52.3/ 66.4/ 72.7 60.0/ 66.0/ 73.0 59.0/ 53.0/ 61.0 53.0/ 60.0/ 77.0 57.0/ 55.0/ 68.0 54.0/ 65.0/ 75.0 54.0/ 61.0/ 74.0 64.0/ 62.0/ 74.0
Biology 53.9/ 62.7/ 68.0 72.8/ 86.6/ 86.4 48.0/ 49.0/ 53.0 54.1/ 66.3/ 81.4 41.4/ 48.8/ 62.3 44.4/ 55.6/ 60.6 43.0/ 49.5/ 58.2 34.0/ 38.0/ 52.0 54.0/ 62.0/ 74.0 41.4/ 55.1/ 61.5 46.2/ 52.8/ 64.8
Business and Management 47.0/ 52.0/ 56.0 57.0/ 62.8/ 69.0 40.0/ 46.0/ 47.0 44.0/ 49.9/ 57.9 34.0/ 44.0/ 54.0 40.0/ 45.0/ 52.0 46.0/ 45.0/ 52.0 38.0/ 39.0/ 45.0 40.0/ 48.0/ 55.0 46.0/ 44.0/ 56.0 34.9/ 34.9/ 47.2
Chemistry 51.9/ 62.1/ 70.9 69.9/ 82.6/ 85.4 49.1/ 58.6/ 69.0 57.9/ 66.5/ 77.1 48.7/ 49.4/ 65.0 49.0/ 57.0/ 70.0 45.1/ 54.4/ 67.5 46.2/ 45.2/ 58.7 50.0/ 59.0/ 66.0 46.0/ 58.2/ 65.0 47.9/ 53.6/ 70.1
Computer Science 39.1/ 47.4/ 57.3 46.8/ 55.2/ 58.4 44.0/ 55.0/ 65.0 42.0/ 50.0/ 51.0 42.0/ 51.5/ 61.0 31.0/ 48.0/ 55.0 41.4/ 47.3/ 53.8 44.0/ 49.0/ 55.0 39.0/ 43.0/ 53.0 40.1/ 50.0/ 61.1 36.3/ 44.8/ 49.8
Defense and Security 46.0/ 57.0/ 66.0 74.0/ 79.0/ 85.0 43.0/ 56.0/ 54.0 43.6/ 46.9/ 53.5 47.0/ 50.0/ 60.0 45.0/ 46.0/ 52.0 44.0/ 44.0/ 53.0 36.0/ 36.0/ 40.0 34.0/ 46.0/ 57.0 49.0/ 53.0/ 53.0 50.0/ 46.0/ 60.0
Earth Sciences 51.0/ 53.0/ 65.0 61.4/ 79.3/ 81.1 34.0/ 56.0/ 58.0 47.3/ 52.7/ 62.9 44.0/ 58.0/ 71.0 37.0/ 52.0/ 56.0 52.0/ 57.0/ 59.0 34.3/ 34.3/ 36.4 39.0/ 44.0/ 52.0 50.0/ 52.0/ 59.0 51.0/ 49.0/ 60.0
Economics 43.2/ 54.7/ 69.1 57.0/ 72.1/ 74.9 40.0/ 53.0/ 59.0 49.8/ 57.3/ 67.3 47.0/ 51.0/ 63.0 36.0/ 41.0/ 51.0 43.7/ 44.4/ 52.1 43.0/ 38.0/ 47.0 42.0/ 54.0/ 55.0 46.0/ 54.8/ 57.9 47.3/ 50.9/ 56.3
Education 43.0/ 43.0/ 54.0 50.3/ 60.2/ 63.1 34.0/ 46.0/ 51.0 36.4/ 42.0/ 49.4 40.0/ 36.0/ 44.0 45.0/ 44.0/ 51.0 42.0/ 42.0/ 44.6 30.0/ 33.0/ 40.0 38.0/ 48.0/ 47.0 39.0/ 41.0/ 47.0 51.0/ 40.0/ 50.0
Energy and Power 46.0/ 55.0/ 62.0 62.2/ 78.4/ 81.8 49.0/ 59.0/ 65.0 47.1/ 53.9/ 68.5 43.0/ 45.0/ 54.0 54.0/ 47.0/ 59.0 41.0/ 54.0/ 61.0 33.0/ 33.0/ 51.0 41.0/ 51.0/ 57.0 45.0/ 45.0/ 60.0 42.0/ 51.0/ 59.0
Engineering 39.0/ 46.0/ 53.2 50.6/ 57.8/ 62.8 41.0/ 46.0/ 52.0 41.4/ 41.0/ 52.4 42.0/ 43.8/ 52.4 41.4/ 45.7/ 51.7 40.4/ 43.0/ 55.8 31.6/ 32.7/ 45.8 38.0/ 42.0/ 52.0 40.2/ 42.0/ 45.6 39.4/ 45.8/ 59.6
Environmental Science 47.0/ 48.0/ 62.0 63.1/ 73.6/ 79.0 49.0/ 54.0/ 58.0 52.1/ 57.5/ 67.9 43.0/ 54.7/ 59.4 49.0/ 52.0/ 60.0 45.8/ 47.9/ 54.7 34.3/ 38.4/ 46.5 54.0/ 65.0/ 71.0 46.0/ 48.0/ 52.0 46.4/ 46.9/ 48.6
Ethics and Human Rights 45.0/ 45.0/ 48.0 54.0/ 70.0/ 76.0 43.0/ 48.0/ 60.0 49.0/ 53.0/ 57.0 45.0/ 42.0/ 50.0 47.0/ 49.0/ 52.0 43.0/ 43.0/ 50.0 37.0/ 33.0/ 35.0 48.0/ 41.0/ 49.0 41.0/ 45.0/ 49.0 51.0/ 47.0/ 51.0
Finance and Investment 42.1/ 52.0/ 59.4 56.6/ 65.2/ 66.8 51.0/ 54.0/ 61.0 44.6/ 48.0/ 58.8 54.3/ 56.2/ 67.9 46.0/ 50.0/ 57.0 45.8/ 48.0/ 56.5 39.0/ 42.0/ 44.0 40.0/ 41.0/ 50.0 45.2/ 48.6/ 57.5 47.5/ 43.9/ 48.9
Food Science 53.0/ 58.0/ 68.0 72.0/ 83.0/ 84.0 48.0/ 57.0/ 62.0 60.6/ 75.8/ 75.8 47.0/ 51.0/ 66.0 51.0/ 55.0/ 66.0 45.0/ 69.0/ 71.0 41.0/ 37.0/ 48.0 45.0/ 49.0/ 62.0 42.0/ 67.0/ 68.0 45.0/ 59.0/ 66.0
Geography 39.1/ 51.4/ 62.6 64.4/ 72.0/ 74.2 44.0/ 66.0/ 62.0 47.8/ 56.4/ 65.1 47.5/ 50.2/ 61.6 39.4/ 39.4/ 49.5 39.1/ 47.3/ 53.8 44.0/ 44.0/ 49.0 39.0/ 49.0/ 56.0 41.1/ 52.6/ 56.0 41.7/ 47.4/ 54.2
Health and Medicine 51.0/ 64.0/ 72.0 70.9/ 83.4/ 85.6 43.0/ 60.0/ 62.0 56.1/ 66.5/ 76.8 48.0/ 53.0/ 64.0 58.0/ 64.0/ 72.0 39.3/ 49.1/ 55.4 41.0/ 48.0/ 54.0 54.0/ 61.0/ 70.0 38.8/ 47.4/ 58.6 47.7/ 57.0/ 62.6
History 52.5/ 54.4/ 61.1 58.4/ 67.9/ 68.9 48.0/ 55.0/ 54.0 49.4/ 55.6/ 61.5 51.3/ 54.2/ 64.5 45.0/ 42.0/ 49.0 39.8/ 45.4/ 50.9 29.0/ 31.0/ 42.0 37.0/ 46.0/ 52.0 45.5/ 56.7/ 58.2 46.3/ 47.6/ 50.9
Information Technology 60.0/ 71.0/ 80.0 67.0/ 81.0/ 81.0 56.0/ 72.0/ 72.0 63.1/ 65.4/ 77.1 52.0/ 51.0/ 74.0 59.0/ 61.0/ 68.0 52.0/ 61.0/ 69.0 44.0/ 48.0/ 58.0 51.0/ 59.0/ 71.0 54.0/ 60.0/ 69.0 51.0/ 64.0/ 76.0
International Relations 51.0/ 57.0/ 67.0 66.7/ 78.3/ 80.8 50.0/ 63.0/ 60.0 48.6/ 56.4/ 63.8 42.0/ 52.0/ 60.0 44.0/ 46.0/ 52.0 46.0/ 57.0/ 60.0 47.0/ 45.0/ 56.0 49.0/ 49.0/ 54.0 54.0/ 54.0/ 64.0 43.0/ 45.0/ 49.0
Language Studies 24.0/ 39.0/ 51.0 59.0/ 77.8/ 83.6 32.0/ 41.0/ 49.0 39.7/ 40.5/ 48.1 27.9/ 26.4/ 43.4 29.0/ 26.0/ 42.0 30.2/ 28.9/ 41.6 41.0/ 40.0/ 47.0 50.0/ 53.0/ 54.0 34.5/ 40.5/ 50.9 32.8/ 37.9/ 39.7
Law and Ethics 58.1/ 63.5/ 73.0 61.2/ 73.6/ 74.8 50.0/ 56.0/ 54.0 52.4/ 57.2/ 63.0 51.0/ 51.0/ 65.6 43.0/ 51.0/ 52.0 41.8/ 43.5/ 46.6 29.0/ 36.0/ 36.0 63.0/ 63.0/ 65.0 55.9/ 54.3/ 59.8 38.8/ 39.4/ 48.8
Literature and Linguistics 40.5/ 46.0/ 50.4 58.5/ 69.3/ 70.3 37.0/ 48.0/ 48.0 39.5/ 44.5/ 51.3 36.7/ 39.2/ 50.0 33.0/ 28.0/ 39.0 38.9/ 44.2/ 45.1 38.0/ 42.0/ 39.0 39.0/ 44.0/ 42.0 38.9/ 37.2/ 38.5 30.3/ 33.8/ 40.8
Logical Reasoning 36.6/ 40.3/ 48.2 41.2/ 45.8/ 56.0 31.8/ 37.0/ 42.9 37.7/ 39.9/ 44.7 42.7/ 41.1/ 48.5 45.0/ 42.0/ 44.0 41.2/ 40.1/ 46.5 38.0/ 39.0/ 41.0 39.0/ 36.0/ 44.0 44.2/ 38.6/ 47.8 38.4/ 38.4/ 44.8
Materials Science 45.0/ 53.0/ 63.0 59.3/ 74.1/ 74.5 35.0/ 40.0/ 51.0 48.2/ 48.5/ 55.0 37.0/ 47.0/ 60.0 49.0/ 49.0/ 61.0 44.0/ 48.0/ 58.0 35.0/ 33.0/ 43.0 38.0/ 47.0/ 56.0 34.0/ 50.0/ 53.0 36.0/ 46.0/ 57.0
Mathematics 39.6/ 48.2/ 54.8 38.4/ 46.2/ 51.4 38.4/ 43.4/ 53.8 35.1/ 38.7/ 44.9 36.2/ 42.6/ 49.2 37.8/ 41.3/ 53.3 35.6/ 39.2/ 49.6 37.0/ 40.0/ 51.6 37.5/ 37.5/ 43.9 35.8/ 40.8/ 48.0 29.2/ 34.8/ 45.8
Media and Communication 50.0/ 54.0/ 59.0 68.0/ 81.0/ 79.0 46.0/ 55.0/ 59.0 47.6/ 59.2/ 68.0 32.0/ 45.0/ 64.0 47.0/ 45.0/ 54.0 34.0/ 44.0/ 52.0 30.0/ 34.0/ 49.0 50.0/ 47.0/ 55.0 41.0/ 44.0/ 52.0 52.0/ 55.0/ 56.0
Music and Performing Arts 40.0/ 42.0/ 48.0 55.0/ 60.0/ 62.0 44.0/ 41.0/ 44.0 37.8/ 39.1/ 47.7 37.0/ 33.0/ 44.0 32.0/ 29.0/ 38.0 40.0/ 41.0/ 49.0 28.0/ 30.0/ 29.0 50.0/ 49.0/ 47.0 30.0/ 40.0/ 38.0 36.0/ 36.0/ 52.0
Physics 46.0/ 55.8/ 65.6 62.0/ 76.0/ 79.6 44.0/ 48.8/ 61.0 51.0/ 58.0/ 68.4 44.0/ 49.0/ 66.4 46.3/ 48.9/ 63.2 43.6/ 49.0/ 67.6 37.1/ 37.9/ 49.6 44.0/ 46.0/ 57.0 44.0/ 49.8/ 59.8 45.2/ 47.8/ 63.0
Politics and Governance 49.4/ 60.4/ 69.0 56.7/ 66.7/ 67.9 38.0/ 42.0/ 45.0 43.8/ 53.4/ 61.2 43.7/ 47.8/ 61.5 42.0/ 46.0/ 43.0 39.3/ 42.4/ 44.7 38.0/ 35.0/ 42.0 43.0/ 41.0/ 53.0 43.1/ 50.2/ 60.6 43.9/ 44.6/ 47.4
Psychology 45.0/ 49.0/ 60.0 76.0/ 89.0/ 89.0 43.0/ 49.0/ 55.0 38.0/ 58.1/ 60.5 55.0/ 53.0/ 71.0 41.0/ 53.0/ 66.0 44.0/ 47.0/ 52.0 44.0/ 43.0/ 56.0 40.0/ 58.0/ 69.0 38.0/ 52.0/ 51.0 44.0/ 64.0/ 71.0
Public Administration 38.0/ 41.0/ 48.0 51.5/ 63.6/ 65.7 40.4/ 44.4/ 40.4 50.5/ 52.5/ 50.5 39.4/ 34.3/ 42.4 39.0/ 37.0/ 43.0 29.3/ 33.3/ 35.4 32.0/ 33.0/ 35.0 38.4/ 50.5/ 47.5 42.0/ 37.0/ 44.0 45.0/ 39.0/ 41.0
Religion and Spirituality 47.0/ 61.0/ 69.0 71.0/ 82.0/ 84.0 48.0/ 62.0/ 66.0 52.0/ 56.9/ 66.9 53.0/ 54.0/ 69.0 51.0/ 46.0/ 62.0 50.0/ 54.0/ 68.0 41.0/ 41.0/ 47.0 45.0/ 59.0/ 62.0 49.0/ 61.0/ 68.0 45.0/ 54.0/ 60.0
Social Welfare and Development 41.0/ 50.0/ 47.0 50.9/ 56.2/ 67.0 47.0/ 52.0/ 54.0 49.2/ 52.8/ 62.2 46.0/ 44.0/ 50.0 42.0/ 45.0/ 50.0 41.0/ 39.0/ 48.0 38.0/ 39.0/ 46.0 37.0/ 51.0/ 59.0 42.0/ 46.0/ 56.0 39.0/ 48.0/ 55.0
Sociology 34.8/ 32.2/ 40.0 44.4/ 56.8/ 56.8 27.0/ 27.0/ 32.0 37.8/ 38.0/ 46.0 24.0/ 35.0/ 41.0 29.0/ 39.0/ 38.0 37.7/ 39.6/ 43.4 34.0/ 35.0/ 39.0 28.0/ 38.0/ 42.0 32.4/ 36.9/ 48.6 31.2/ 41.2/ 41.2
Sports and Recreation 40.1/ 43.6/ 56.9 67.8/ 78.6/ 81.2 37.0/ 57.0/ 59.0 46.4/ 56.6/ 64.6 36.5/ 41.6/ 56.7 43.0/ 42.0/ 49.0 45.2/ 50.3/ 58.2 32.0/ 25.0/ 35.0 34.0/ 40.0/ 49.0 44.2/ 49.4/ 58.3 39.1/ 47.1/ 49.5
Technology and Innovation 55.0/ 52.0/ 65.0 63.6/ 74.0/ 79.4 49.0/ 50.0/ 54.0 47.0/ 53.8/ 66.6 47.0/ 47.0/ 56.0 45.0/ 52.0/ 60.0 55.0/ 54.0/ 63.0 46.0/ 45.0/ 54.0 43.0/ 57.0/ 55.0 34.0/ 36.0/ 55.0 44.0/ 42.0/ 58.0
Transportation and Logistics 45.0/ 35.0/ 38.0 56.1/ 62.3/ 68.5 40.0/ 36.0/ 49.0 43.5/ 45.7/ 54.9 32.3/ 29.3/ 41.4 45.0/ 42.0/ 42.0 37.8/ 44.9/ 45.9 31.3/ 33.3/ 37.4 46.0/ 48.0/ 49.0 40.0/ 50.0/ 55.0 41.0/ 33.0/ 39.0
Table 49: Detailed subject-wise evaluation for google/gemma-2-27b-it on MILU across different languages. The results reported are X / Y / Z where X denote the 0-shot, Y denotes 1-shot and Z denotes 5-shot performance respectively.
Topic bn en gu hi kn ml mr or pa ta te
Agriculture 29.0/ 29.0/ 31.0 49.1/ 48.1/ 48.9 32.3/ 26.3/ 30.3 39.4/ 41.4/ 42.2 37.0/ 29.0/ 30.0 29.3/ 38.4/ 34.3 30.0/ 33.0/ 34.0 30.0/ 29.0/ 33.0 31.0/ 25.0/ 27.0 26.0/ 26.0/ 34.0 24.0/ 27.0/ 25.0
Anthropology 35.0/ 37.0/ 36.0 50.0/ 54.0/ 49.0 35.0/ 37.0/ 34.0 44.4/ 40.4/ 44.4 36.0/ 36.0/ 42.0 32.0/ 36.0/ 36.0 36.0/ 43.0/ 41.0 38.0/ 37.0/ 39.0 35.0/ 32.0/ 32.0 43.0/ 41.0/ 43.0 37.0/ 36.0/ 35.0
Architecture and Design 26.0/ 31.0/ 31.0 48.0/ 51.0/ 50.0 24.0/ 22.0/ 24.0 41.0/ 40.0/ 40.0 35.0/ 37.0/ 32.0 31.0/ 34.0/ 31.0 35.0/ 32.0/ 39.0 24.0/ 26.0/ 28.0 22.0/ 23.0/ 22.0 29.0/ 36.0/ 36.0 34.0/ 24.0/ 29.0
Arts and Culture 23.6/ 24.9/ 23.6 53.6/ 55.6/ 54.0 35.0/ 35.0/ 38.0 43.9/ 39.5/ 40.5 35.6/ 31.3/ 33.1 34.0/ 29.0/ 29.0 37.2/ 35.5/ 38.2 26.0/ 27.0/ 22.0 28.0/ 28.0/ 30.0 33.3/ 29.1/ 27.9 27.1/ 21.1/ 21.6
Astronomy and Astrophysics 40.0/ 35.0/ 40.0 66.0/ 72.0/ 65.0 33.0/ 34.0/ 32.0 50.8/ 49.2/ 50.0 35.0/ 34.0/ 32.0 43.0/ 36.0/ 39.0 48.0/ 48.0/ 51.0 37.0/ 33.0/ 37.0 27.0/ 26.0/ 29.0 38.0/ 39.0/ 38.0 25.0/ 30.0/ 22.0
Biology 27.5/ 26.4/ 23.9 61.3/ 60.3/ 60.3 32.0/ 30.0/ 25.0 45.7/ 48.1/ 49.3 31.1/ 32.0/ 32.0 15.2/ 18.2/ 24.2 33.7/ 34.3/ 35.8 30.0/ 24.0/ 28.0 24.0/ 29.0/ 24.0 33.1/ 30.6/ 31.5 26.4/ 23.9/ 28.3
Business and Management 34.0/ 29.0/ 29.0 48.0/ 53.2/ 58.2 33.0/ 35.0/ 40.0 41.9/ 41.2/ 41.7 33.0/ 26.0/ 27.0 37.0/ 39.0/ 44.0 34.0/ 33.0/ 43.0 32.0/ 30.0/ 28.0 32.0/ 30.0/ 35.0 40.0/ 46.0/ 49.0 30.2/ 26.4/ 27.4
Chemistry 28.8/ 31.3/ 29.7 58.1/ 61.3/ 60.3 28.4/ 29.3/ 30.2 39.5/ 42.7/ 47.7 29.1/ 29.7/ 30.7 28.0/ 28.0/ 28.0 34.6/ 37.7/ 40.9 29.8/ 25.0/ 29.8 25.0/ 24.0/ 28.0 29.4/ 30.3/ 31.4 29.7/ 27.4/ 30.4
Computer Science 35.4/ 32.3/ 33.3 42.8/ 40.8/ 40.0 32.0/ 29.0/ 28.0 34.4/ 35.8/ 38.4 29.6/ 29.0/ 29.0 34.0/ 30.0/ 28.0 34.3/ 30.2/ 32.5 37.0/ 40.0/ 34.0 28.0/ 27.0/ 34.0 31.4/ 32.6/ 34.9 31.4/ 30.9/ 30.9
Defense and Security 32.0/ 40.0/ 36.0 58.0/ 64.0/ 63.0 30.0/ 32.0/ 25.0 39.3/ 40.8/ 42.6 28.0/ 31.0/ 35.0 28.0/ 33.0/ 28.0 31.0/ 38.0/ 36.0 30.0/ 30.0/ 32.0 33.0/ 37.0/ 30.0 26.0/ 28.0/ 25.0 27.0/ 30.0/ 27.0
Earth Sciences 43.0/ 43.0/ 43.0 56.1/ 58.0/ 57.3 22.0/ 28.0/ 30.0 38.9/ 40.3/ 42.3 32.0/ 31.0/ 32.0 36.0/ 42.0/ 35.0 33.0/ 30.0/ 31.0 30.3/ 27.3/ 25.2 26.0/ 21.0/ 18.0 32.0/ 39.0/ 41.0 22.0/ 29.0/ 28.0
Economics 36.0/ 35.2/ 35.2 50.4/ 52.2/ 52.4 39.0/ 35.0/ 29.0 43.8/ 42.5/ 46.6 37.0/ 35.0/ 37.0 31.0/ 28.0/ 32.0 33.1/ 33.8/ 38.0 30.0/ 28.0/ 36.0 27.0/ 19.0/ 24.0 32.5/ 39.7/ 30.9 30.5/ 27.0/ 32.9
Education 37.0/ 42.0/ 43.0 43.6/ 48.4/ 45.9 28.0/ 32.0/ 33.0 33.1/ 37.3/ 38.2 31.0/ 30.0/ 30.0 26.0/ 31.0/ 31.0 40.1/ 37.6/ 39.5 29.0/ 33.0/ 33.0 28.0/ 31.0/ 28.0 40.0/ 42.0/ 43.0 36.0/ 30.0/ 37.0
Energy and Power 27.0/ 33.0/ 31.0 56.1/ 58.1/ 57.4 34.0/ 33.0/ 30.0 39.0/ 42.0/ 43.7 25.0/ 20.0/ 24.0 37.0/ 38.0/ 37.0 38.0/ 41.0/ 46.0 29.0/ 34.0/ 32.0 24.0/ 27.0/ 27.0 24.0/ 27.0/ 25.0 28.0/ 32.0/ 35.0
Engineering 30.2/ 29.8/ 31.8 41.0/ 41.4/ 42.4 25.0/ 27.4/ 27.6 32.2/ 35.4/ 35.8 25.4/ 25.8/ 26.8 24.8/ 26.9/ 25.2 31.2/ 32.6/ 30.6 22.3/ 27.4/ 28.6 21.0/ 21.0/ 18.0 30.2/ 33.4/ 32.2 29.0/ 27.4/ 28.6
Environmental Science 31.0/ 31.0/ 30.0 54.7/ 59.5/ 59.3 25.0/ 34.0/ 33.0 44.3/ 43.3/ 46.5 33.6/ 28.1/ 28.1 39.0/ 33.0/ 35.0 32.1/ 35.8/ 34.7 30.3/ 26.3/ 27.3 34.0/ 31.0/ 35.0 42.0/ 37.0/ 44.0 31.8/ 34.1/ 34.1
Ethics and Human Rights 45.0/ 43.0/ 46.0 48.0/ 52.0/ 56.0 37.0/ 35.0/ 35.0 42.0/ 50.0/ 48.0 29.0/ 33.0/ 30.0 42.0/ 39.0/ 41.0 41.0/ 45.0/ 49.0 44.0/ 38.0/ 39.0 26.0/ 17.0/ 19.0 35.0/ 37.0/ 36.0 32.0/ 32.0/ 36.0
Finance and Investment 34.6/ 40.6/ 38.1 38.0/ 41.6/ 43.0 27.0/ 31.0/ 31.0 33.6/ 36.6/ 38.2 31.5/ 36.4/ 38.3 34.0/ 30.0/ 29.0 26.6/ 30.5/ 29.9 35.0/ 31.0/ 33.0 30.0/ 30.0/ 31.0 40.2/ 41.3/ 39.1 30.7/ 33.9/ 37.1
Food Science 30.0/ 31.0/ 36.0 56.0/ 63.0/ 60.0 27.0/ 27.0/ 31.0 60.6/ 52.5/ 60.6 30.0/ 30.0/ 29.0 39.0/ 37.0/ 33.0 37.0/ 47.0/ 43.0 23.0/ 32.0/ 34.0 18.0/ 29.0/ 25.0 33.0/ 39.0/ 36.0 24.0/ 33.0/ 25.0
Geography 31.8/ 34.6/ 33.0 53.8/ 55.4/ 56.2 27.0/ 29.0/ 30.0 38.6/ 42.2/ 38.4 31.5/ 29.7/ 32.4 25.2/ 27.3/ 25.2 34.9/ 38.5/ 37.3 35.0/ 37.0/ 32.0 25.0/ 30.0/ 29.0 33.1/ 33.7/ 32.6 26.6/ 24.0/ 27.1
Health and Medicine 36.0/ 31.0/ 35.0 62.1/ 63.7/ 63.1 33.0/ 31.0/ 32.0 49.5/ 51.3/ 54.1 29.0/ 33.0/ 29.0 32.0/ 36.0/ 32.0 30.4/ 37.5/ 38.4 33.0/ 25.0/ 34.0 23.0/ 27.0/ 22.0 46.6/ 40.5/ 45.7 32.7/ 26.2/ 33.6
History 29.1/ 28.0/ 31.7 45.6/ 49.4/ 50.0 31.0/ 28.0/ 30.0 44.0/ 45.4/ 44.8 28.6/ 30.4/ 34.1 31.0/ 30.0/ 40.0 35.7/ 34.9/ 36.4 33.0/ 27.0/ 29.0 26.0/ 17.0/ 22.0 36.0/ 38.6/ 36.4 28.1/ 28.4/ 29.2
Information Technology 39.0/ 40.0/ 44.0 57.0/ 56.0/ 62.0 32.0/ 36.0/ 40.0 52.5/ 55.3/ 58.7 33.0/ 31.0/ 28.0 37.0/ 41.0/ 43.0 54.0/ 51.0/ 58.0 40.0/ 41.0/ 39.0 34.0/ 36.0/ 36.0 42.0/ 40.0/ 43.0 39.0/ 40.0/ 39.0
International Relations 37.0/ 35.0/ 40.0 64.2/ 67.5/ 61.7 39.0/ 41.0/ 39.0 49.8/ 49.8/ 52.9 28.0/ 31.0/ 32.0 30.0/ 33.0/ 30.0 34.0/ 40.0/ 37.0 34.0/ 32.0/ 31.0 27.0/ 26.0/ 27.0 39.0/ 38.0/ 36.0 25.0/ 22.0/ 21.0
Language Studies 28.0/ 23.0/ 26.0 60.8/ 64.0/ 62.6 28.0/ 20.0/ 20.0 35.5/ 38.9/ 35.1 30.2/ 29.5/ 29.5 22.0/ 22.0/ 24.0 29.5/ 25.5/ 24.2 34.0/ 36.0/ 36.0 34.0/ 39.0/ 36.0 28.1/ 28.1/ 28.1 29.9/ 24.1/ 25.9
Law and Ethics 46.0/ 42.6/ 43.2 54.2/ 58.8/ 57.6 37.0/ 30.0/ 30.0 49.6/ 51.8/ 53.4 31.9/ 25.5/ 26.8 37.0/ 25.0/ 25.0 37.9/ 39.7/ 40.5 28.0/ 31.0/ 29.0 36.0/ 33.0/ 34.0 42.5/ 40.2/ 38.6 28.1/ 28.7/ 33.1
Literature and Linguistics 27.9/ 33.3/ 28.8 49.7/ 49.5/ 50.3 29.0/ 29.0/ 24.0 38.7/ 38.5/ 39.1 27.5/ 32.5/ 30.8 35.0/ 24.0/ 28.0 41.6/ 40.7/ 47.8 30.0/ 23.0/ 29.0 23.0/ 19.0/ 23.0 27.1/ 25.9/ 28.3 28.9/ 23.9/ 26.1
Logical Reasoning 32.2/ 32.4/ 32.2 33.4/ 32.6/ 35.4 22.7/ 26.0/ 22.7 31.5/ 29.9/ 30.7 31.1/ 29.0/ 28.2 34.0/ 28.0/ 31.0 26.5/ 27.0/ 28.1 25.0/ 24.0/ 25.0 28.0/ 27.0/ 26.0 35.1/ 33.5/ 33.1 26.2/ 27.5/ 27.5
Materials Science 35.0/ 28.0/ 24.0 54.4/ 52.5/ 54.0 21.0/ 22.0/ 27.0 41.9/ 40.4/ 46.7 24.0/ 30.0/ 29.0 31.0/ 29.0/ 34.0 35.0/ 41.0/ 42.0 24.0/ 30.0/ 31.0 34.0/ 26.0/ 24.0 29.0/ 32.0/ 36.0 22.0/ 23.0/ 23.0
Mathematics 29.4/ 34.0/ 32.0 31.0/ 31.8/ 30.2 28.2/ 29.0/ 27.2 32.5/ 28.1/ 29.3 26.4/ 27.2/ 31.2 31.2/ 33.0/ 31.5 27.2/ 28.2/ 29.8 27.6/ 29.2/ 30.0 26.5/ 33.7/ 31.1 29.4/ 32.2/ 32.8 26.6/ 25.8/ 26.8
Media and Communication 36.0/ 33.0/ 39.0 58.0/ 68.0/ 63.0 34.0/ 40.0/ 40.0 41.5/ 38.1/ 47.6 37.0/ 40.0/ 43.0 51.0/ 46.0/ 48.0 36.0/ 41.0/ 45.0 31.0/ 28.0/ 33.0 22.0/ 35.0/ 33.0 38.0/ 38.0/ 38.0 37.0/ 36.0/ 34.0
Music and Performing Arts 29.0/ 27.0/ 27.0 46.0/ 52.0/ 51.0 39.0/ 30.0/ 33.0 38.4/ 34.4/ 32.5 28.0/ 30.0/ 32.0 35.0/ 36.0/ 37.0 41.0/ 40.0/ 41.0 35.0/ 22.0/ 32.0 25.0/ 31.0/ 28.0 27.0/ 27.0/ 27.0 29.0/ 25.0/ 28.0
Physics 32.8/ 31.8/ 33.2 50.4/ 53.6/ 53.0 29.5/ 31.5/ 32.0 38.4/ 40.8/ 42.6 29.4/ 28.2/ 29.6 26.3/ 28.4/ 30.0 32.0/ 35.0/ 35.2 28.8/ 28.5/ 28.1 30.0/ 30.0/ 31.0 31.0/ 32.0/ 32.8 27.4/ 27.2/ 28.2
Politics and Governance 35.7/ 36.1/ 34.9 46.5/ 50.3/ 48.9 26.0/ 25.0/ 27.0 43.4/ 42.0/ 44.6 32.0/ 33.6/ 35.2 40.0/ 40.0/ 34.0 39.0/ 38.8/ 38.8 39.0/ 38.0/ 40.0 24.0/ 27.0/ 29.0 36.5/ 38.6/ 39.8 27.8/ 29.9/ 29.9
Psychology 39.0/ 37.0/ 39.0 63.0/ 65.0/ 67.0 35.0/ 36.0/ 34.0 49.6/ 45.0/ 48.8 37.0/ 35.0/ 41.0 41.0/ 36.0/ 37.0 27.0/ 39.0/ 41.0 35.0/ 34.0/ 34.0 22.0/ 25.0/ 23.0 33.0/ 33.0/ 38.0 32.0/ 29.0/ 23.0
Public Administration 29.0/ 33.0/ 35.0 42.4/ 49.5/ 46.5 34.3/ 38.4/ 37.4 41.4/ 46.5/ 47.5 33.3/ 36.4/ 30.3 26.0/ 35.0/ 32.0 30.3/ 24.2/ 30.3 37.0/ 29.0/ 30.0 29.3/ 29.3/ 27.3 33.0/ 39.0/ 37.0 33.0/ 29.0/ 33.0
Religion and Spirituality 21.0/ 24.0/ 29.0 61.0/ 68.0/ 63.0 26.0/ 27.0/ 25.0 41.9/ 45.2/ 46.8 27.0/ 28.0/ 33.0 23.0/ 28.0/ 28.0 49.0/ 41.0/ 46.0 26.0/ 23.0/ 24.0 26.0/ 22.0/ 23.0 32.0/ 27.0/ 32.0 16.0/ 15.0/ 17.0
Social Welfare and Development 27.0/ 23.0/ 29.0 50.9/ 51.8/ 55.4 34.0/ 38.0/ 30.0 40.9/ 40.9/ 44.0 25.0/ 33.0/ 37.0 28.0/ 28.0/ 34.0 33.0/ 32.0/ 34.0 32.0/ 29.0/ 32.0 17.0/ 15.0/ 21.0 34.0/ 43.0/ 45.0 34.0/ 34.0/ 33.0
Sociology 26.1/ 28.7/ 24.3 40.2/ 42.0/ 41.6 25.0/ 26.0/ 26.0 34.8/ 38.4/ 39.4 29.0/ 25.0/ 25.0 33.0/ 25.0/ 27.0 28.9/ 34.0/ 30.8 23.0/ 29.0/ 31.0 29.0/ 25.0/ 26.0 27.0/ 35.1/ 35.1 28.2/ 27.7/ 27.1
Sports and Recreation 31.2/ 32.7/ 34.6 58.2/ 61.0/ 61.4 25.0/ 33.0/ 31.0 43.8/ 45.8/ 48.0 35.4/ 37.6/ 37.1 35.0/ 35.0/ 36.0 36.2/ 40.1/ 39.0 32.0/ 29.0/ 38.0 28.0/ 28.0/ 30.0 35.3/ 39.7/ 37.8 29.0/ 28.6/ 30.5
Technology and Innovation 34.0/ 41.0/ 36.0 54.2/ 57.2/ 57.6 33.0/ 37.0/ 34.0 41.4/ 43.2/ 46.2 24.0/ 34.0/ 31.0 36.0/ 42.0/ 40.0 45.0/ 47.0/ 48.0 35.0/ 36.0/ 33.0 25.0/ 25.0/ 26.0 37.0/ 42.0/ 40.0 28.0/ 31.0/ 35.0
Transportation and Logistics 30.0/ 36.0/ 27.0 40.8/ 48.5/ 45.4 31.0/ 27.0/ 25.0 30.0/ 34.1/ 36.9 25.2/ 25.2/ 30.3 25.0/ 28.0/ 26.0 26.5/ 30.6/ 33.7 33.3/ 25.2/ 25.2 24.0/ 22.0/ 26.0 28.0/ 30.0/ 32.0 30.0/ 26.0/ 30.0
Table 50: Detailed subject-wise evaluation for CohereForAI/aya-23-35B on MILU across different languages. The results reported are X / Y / Z where X denote the 0-shot, Y denotes 1-shot and Z denotes 5-shot performance respectively.
Topic bn en gu hi kn ml mr or pa ta te
Agriculture 51.0/ 60.0/ 59.0 50.5/ 71.8/ 72.2 43.4/ 52.5/ 54.5 39.6/ 63.8/ 64.2 40.0/ 61.0/ 64.0 54.5/ 61.6/ 62.6 44.0/ 54.0/ 54.0 51.0/ 60.0/ 58.0 46.0/ 59.0/ 58.0 48.0/ 54.0/ 59.0 42.0/ 58.0/ 60.0
Anthropology 53.0/ 67.0/ 69.0 55.0/ 72.0/ 71.0 47.0/ 63.0/ 62.0 49.5/ 62.6/ 67.7 57.0/ 70.0/ 75.0 47.0/ 65.0/ 62.0 58.0/ 70.0/ 71.0 54.0/ 59.0/ 61.0 45.0/ 65.0/ 60.0 53.0/ 61.0/ 60.0 49.0/ 56.0/ 55.0
Architecture and Design 51.0/ 63.0/ 71.0 52.0/ 72.0/ 71.0 47.0/ 64.0/ 58.0 41.0/ 59.0/ 59.0 52.0/ 63.0/ 63.0 49.0/ 65.0/ 62.0 54.0/ 68.0/ 69.0 44.0/ 55.0/ 54.0 45.0/ 59.0/ 65.0 52.0/ 63.0/ 57.0 45.0/ 57.0/ 61.0
Arts and Culture 48.5/ 69.1/ 67.3 59.4/ 81.2/ 79.6 48.0/ 73.0/ 67.0 45.5/ 67.3/ 68.7 54.6/ 73.6/ 74.2 50.0/ 67.0/ 66.0 44.3/ 61.2/ 61.2 50.0/ 54.0/ 56.0 48.0/ 68.0/ 69.0 47.3/ 63.0/ 63.6 46.2/ 64.3/ 66.8
Astronomy and Astrophysics 60.0/ 75.0/ 78.0 60.0/ 96.0/ 88.0 67.0/ 87.0/ 82.0 57.0/ 77.3/ 75.8 63.0/ 80.0/ 82.0 66.0/ 83.0/ 81.0 60.0/ 73.0/ 74.0 67.0/ 77.0/ 76.0 58.0/ 77.0/ 71.0 68.0/ 80.0/ 82.0 57.0/ 74.0/ 73.0
Biology 48.9/ 68.7/ 66.9 61.5/ 88.6/ 86.2 38.0/ 72.0/ 69.0 50.7/ 76.5/ 78.0 51.6/ 75.8/ 78.3 50.5/ 64.6/ 65.7 43.0/ 61.2/ 61.8 47.0/ 51.0/ 52.0 43.0/ 73.0/ 76.0 45.9/ 62.1/ 63.4 45.0/ 60.4/ 63.5
Business and Management 45.0/ 56.0/ 57.0 51.4/ 66.6/ 66.6 36.0/ 65.0/ 64.0 46.1/ 61.6/ 61.8 43.0/ 62.0/ 67.0 49.0/ 57.0/ 59.0 39.0/ 53.0/ 57.0 41.0/ 49.0/ 50.0 48.0/ 68.0/ 63.0 44.0/ 57.0/ 59.0 46.2/ 51.9/ 50.0
Chemistry 58.8/ 72.0/ 71.7 62.1/ 85.4/ 84.8 59.5/ 72.4/ 69.0 52.5/ 75.1/ 76.5 53.3/ 69.6/ 71.2 60.0/ 71.0/ 71.0 53.3/ 66.2/ 68.3 44.2/ 60.6/ 55.8 49.0/ 69.0/ 73.0 51.9/ 61.1/ 63.8 53.9/ 68.3/ 69.8
Computer Science 39.1/ 44.8/ 49.5 45.4/ 52.8/ 53.4 51.0/ 46.0/ 52.0 46.6/ 48.8/ 52.4 40.2/ 51.5/ 54.4 39.0/ 45.0/ 50.0 44.4/ 43.2/ 50.9 38.0/ 38.0/ 47.0 53.0/ 53.0/ 47.0 41.3/ 44.2/ 50.0 36.3/ 40.8/ 47.5
Defense and Security 60.0/ 76.0/ 84.0 70.0/ 87.0/ 82.0 53.0/ 77.0/ 73.0 51.7/ 72.0/ 73.9 54.0/ 65.0/ 75.0 53.0/ 76.0/ 76.0 51.0/ 69.0/ 75.0 51.0/ 61.0/ 67.0 52.0/ 71.0/ 72.0 53.0/ 66.0/ 68.0 60.0/ 76.0/ 76.0
Earth Sciences 46.0/ 61.0/ 68.0 60.9/ 81.8/ 80.0 42.0/ 58.0/ 55.0 42.3/ 67.3/ 68.9 52.0/ 75.0/ 74.0 51.0/ 64.0/ 63.0 47.0/ 63.0/ 70.0 50.5/ 46.5/ 57.6 48.0/ 70.0/ 69.0 55.0/ 71.0/ 69.0 47.0/ 70.0/ 71.0
Economics 61.2/ 72.7/ 75.5 55.6/ 76.7/ 75.5 47.0/ 70.0/ 67.0 49.0/ 69.0/ 69.3 58.0/ 75.0/ 77.0 57.0/ 66.0/ 67.0 50.7/ 59.2/ 58.5 58.0/ 58.0/ 57.0 51.0/ 66.0/ 69.0 54.0/ 69.0/ 69.0 47.3/ 59.9/ 61.7
Education 48.0/ 60.0/ 57.0 49.7/ 67.2/ 65.0 38.0/ 55.0/ 57.0 47.6/ 57.7/ 59.8 54.0/ 62.0/ 61.0 37.0/ 51.0/ 52.0 46.5/ 60.5/ 61.8 43.0/ 48.0/ 53.0 32.0/ 59.0/ 50.0 41.0/ 51.0/ 55.0 43.0/ 61.0/ 57.0
Energy and Power 58.0/ 67.0/ 72.0 56.8/ 83.8/ 79.7 62.0/ 73.0/ 75.0 48.5/ 67.1/ 70.2 51.0/ 78.0/ 77.0 47.0/ 76.0/ 72.0 57.0/ 66.0/ 69.0 46.0/ 62.0/ 65.0 50.0/ 71.0/ 69.0 47.0/ 66.0/ 67.0 50.0/ 69.0/ 67.0
Engineering 39.6/ 54.4/ 58.4 47.6/ 66.0/ 66.4 40.4/ 51.4/ 52.6 38.4/ 52.0/ 53.4 37.8/ 52.8/ 57.4 37.2/ 47.9/ 45.3 36.8/ 52.8/ 57.4 35.7/ 42.9/ 44.4 36.0/ 41.0/ 46.0 39.4/ 44.6/ 45.6 39.6/ 50.4/ 52.6
Environmental Science 47.0/ 65.0/ 74.0 57.1/ 79.4/ 80.0 50.0/ 65.0/ 63.0 49.3/ 69.7/ 70.1 56.2/ 68.8/ 70.3 46.0/ 66.0/ 63.0 45.8/ 54.2/ 62.1 47.5/ 53.5/ 54.5 55.0/ 67.0/ 68.0 54.0/ 62.0/ 58.0 46.4/ 55.9/ 59.2
Ethics and Human Rights 52.0/ 60.0/ 63.0 57.0/ 75.0/ 75.0 49.0/ 61.0/ 61.0 46.0/ 60.0/ 61.0 51.0/ 62.0/ 65.0 51.0/ 69.0/ 71.0 52.0/ 62.0/ 64.0 47.0/ 63.0/ 62.0 44.0/ 58.0/ 60.0 45.0/ 59.0/ 66.0 50.0/ 63.0/ 65.0
Finance and Investment 56.4/ 63.9/ 65.8 56.8/ 69.4/ 68.0 55.0/ 63.0/ 67.0 51.0/ 63.6/ 65.8 54.3/ 66.7/ 66.0 44.0/ 52.0/ 55.0 50.8/ 63.3/ 65.0 42.0/ 54.0/ 52.0 47.0/ 57.0/ 53.0 52.0/ 61.5/ 65.9 45.0/ 50.0/ 53.9
Food Science 41.0/ 70.0/ 77.0 56.0/ 88.0/ 87.0 51.0/ 73.0/ 74.0 50.5/ 74.8/ 70.7 48.0/ 65.0/ 68.0 55.0/ 77.0/ 75.0 50.0/ 68.0/ 72.0 49.0/ 60.0/ 61.0 46.0/ 65.0/ 70.0 47.0/ 69.0/ 68.0 54.0/ 71.0/ 65.0
Geography 53.1/ 69.3/ 71.5 56.0/ 78.2/ 75.8 48.0/ 64.0/ 65.0 45.8/ 67.9/ 68.3 55.2/ 77.6/ 79.0 52.5/ 70.7/ 67.7 53.8/ 61.5/ 65.1 52.0/ 63.0/ 63.0 51.0/ 67.0/ 71.0 52.6/ 59.4/ 67.4 51.6/ 66.1/ 65.1
Health and Medicine 54.0/ 68.0/ 69.0 61.1/ 83.4/ 82.6 52.0/ 71.0/ 66.0 50.9/ 79.0/ 79.8 56.0/ 79.0/ 77.0 59.0/ 74.0/ 76.0 45.5/ 63.4/ 62.5 56.0/ 67.0/ 69.0 68.0/ 84.0/ 83.0 59.5/ 67.2/ 68.1 52.3/ 64.5/ 66.4
History 57.1/ 73.1/ 78.7 55.0/ 75.5/ 70.3 63.0/ 76.0/ 78.0 54.2/ 72.4/ 74.8 59.3/ 77.3/ 80.2 51.0/ 68.0/ 65.0 50.2/ 61.0/ 65.8 55.0/ 58.0/ 59.0 51.0/ 77.0/ 78.0 54.5/ 65.8/ 68.7 51.7/ 61.4/ 61.6
Information Technology 60.0/ 81.0/ 83.0 69.0/ 82.0/ 78.0 60.0/ 75.0/ 73.0 59.8/ 78.8/ 81.0 56.0/ 77.0/ 77.0 57.0/ 83.0/ 83.0 55.0/ 70.0/ 78.0 51.0/ 68.0/ 71.0 64.0/ 75.0/ 78.0 60.0/ 73.0/ 69.0 63.0/ 76.0/ 74.0
International Relations 59.0/ 76.0/ 78.0 65.0/ 82.5/ 82.5 54.0/ 70.0/ 69.0 62.6/ 72.4/ 72.0 56.0/ 63.0/ 66.0 62.0/ 75.0/ 74.0 66.0/ 74.0/ 74.0 55.0/ 64.0/ 64.0 57.0/ 69.0/ 70.0 57.0/ 75.0/ 73.0 51.0/ 68.0/ 71.0
Language Studies 40.0/ 38.0/ 40.0 59.4/ 82.4/ 82.0 33.0/ 41.0/ 47.0 38.5/ 52.5/ 48.9 39.5/ 51.9/ 51.9 30.0/ 41.0/ 41.0 34.9/ 51.0/ 50.3 42.0/ 47.0/ 46.0 44.0/ 50.0/ 53.0 35.2/ 39.3/ 44.2 31.6/ 48.9/ 48.9
Law and Ethics 74.3/ 84.5/ 85.8 65.8/ 80.2/ 80.4 62.0/ 70.0/ 63.0 61.2/ 75.6/ 76.8 58.6/ 78.3/ 74.5 53.0/ 63.0/ 64.0 45.3/ 62.5/ 60.3 56.0/ 66.0/ 68.0 66.0/ 82.0/ 78.0 63.0/ 68.5/ 68.5 60.0/ 63.7/ 68.1
Literature and Linguistics 51.3/ 68.5/ 76.6 55.7/ 74.6/ 75.9 45.0/ 67.0/ 62.0 42.7/ 60.3/ 63.9 48.3/ 66.7/ 65.8 45.0/ 57.0/ 58.0 45.1/ 58.4/ 62.8 45.0/ 53.0/ 56.0 47.0/ 68.0/ 69.0 41.7/ 44.1/ 43.7 38.7/ 52.1/ 58.5
Logical Reasoning 37.4/ 45.2/ 48.9 38.2/ 49.2/ 50.2 33.8/ 50.0/ 48.7 36.1/ 40.1/ 43.7 41.1/ 47.7/ 49.0 38.0/ 44.0/ 46.0 36.2/ 37.6/ 42.3 29.0/ 41.0/ 48.0 33.0/ 41.0/ 36.0 39.8/ 40.2/ 43.0 35.6/ 40.8/ 41.6
Materials Science 54.0/ 67.0/ 67.0 50.6/ 75.3/ 77.2 38.0/ 51.0/ 51.0 46.5/ 61.1/ 61.8 41.0/ 64.0/ 65.0 45.0/ 60.0/ 63.0 44.0/ 66.0/ 67.0 40.0/ 53.0/ 52.0 35.0/ 61.0/ 63.0 45.0/ 51.0/ 56.0 40.0/ 70.0/ 70.0
Mathematics 45.8/ 50.8/ 54.6 42.0/ 50.4/ 51.0 42.8/ 51.0/ 55.0 36.5/ 43.9/ 49.5 41.4/ 49.0/ 52.4 44.1/ 53.9/ 57.0 43.8/ 50.4/ 52.4 41.6/ 48.0/ 53.0 39.0/ 49.6/ 50.8 41.2/ 47.2/ 50.2 39.4/ 46.4/ 47.8
Media and Communication 52.0/ 65.0/ 69.0 63.0/ 77.0/ 80.0 50.0/ 64.0/ 68.0 49.7/ 65.3/ 63.9 47.0/ 70.0/ 71.0 50.0/ 63.0/ 67.0 50.0/ 66.0/ 65.0 55.0/ 61.0/ 63.0 49.0/ 62.0/ 61.0 53.0/ 62.0/ 64.0 45.0/ 66.0/ 65.0
Music and Performing Arts 49.0/ 60.0/ 67.0 58.0/ 82.0/ 78.0 42.0/ 62.0/ 58.0 46.4/ 64.2/ 63.6 48.0/ 59.0/ 68.0 41.0/ 56.0/ 56.0 48.0/ 63.0/ 66.0 49.0/ 58.0/ 62.0 52.0/ 67.0/ 64.0 37.0/ 45.0/ 45.0 46.0/ 63.0/ 66.0
Physics 59.8/ 68.6/ 69.8 60.2/ 81.8/ 78.2 54.0/ 62.7/ 67.1 45.6/ 68.0/ 67.0 56.2/ 70.8/ 71.6 57.9/ 60.5/ 62.6 53.6/ 66.2/ 69.6 53.5/ 57.6/ 63.2 58.0/ 61.0/ 61.0 52.8/ 62.6/ 63.8 55.0/ 64.0/ 67.6
Politics and Governance 54.5/ 77.2/ 79.6 54.1/ 76.1/ 71.5 52.0/ 71.0/ 69.0 52.6/ 71.1/ 74.5 45.8/ 70.9/ 76.1 46.0/ 71.0/ 72.0 43.5/ 55.6/ 57.3 64.0/ 56.0/ 57.0 51.0/ 64.0/ 68.0 45.2/ 63.1/ 61.8 44.6/ 52.8/ 57.1
Psychology 50.0/ 66.0/ 64.0 59.0/ 85.0/ 83.0 40.0/ 65.0/ 66.0 37.2/ 57.4/ 60.5 43.0/ 64.0/ 59.0 46.0/ 62.0/ 60.0 45.0/ 55.0/ 58.0 52.0/ 59.0/ 62.0 51.0/ 68.0/ 69.0 43.0/ 58.0/ 61.0 49.0/ 72.0/ 65.0
Public Administration 41.0/ 52.0/ 62.0 44.4/ 67.7/ 64.6 33.3/ 58.6/ 52.5 44.4/ 68.7/ 68.7 38.4/ 51.5/ 57.6 45.0/ 59.0/ 60.0 43.4/ 47.5/ 50.5 51.0/ 52.0/ 59.0 34.3/ 51.5/ 55.6 42.0/ 53.0/ 56.0 41.0/ 48.0/ 50.0
Religion and Spirituality 59.0/ 74.0/ 77.0 53.0/ 88.0/ 84.0 63.0/ 73.0/ 72.0 53.6/ 74.6/ 74.6 58.0/ 76.0/ 75.0 59.0/ 74.0/ 76.0 57.0/ 73.0/ 75.0 61.0/ 66.0/ 65.0 58.0/ 73.0/ 75.0 54.0/ 71.0/ 73.0 60.0/ 71.0/ 71.0
Social Welfare and Development 44.0/ 61.0/ 60.0 47.3/ 75.0/ 77.7 47.0/ 58.0/ 57.0 48.7/ 62.2/ 63.7 42.0/ 55.0/ 57.0 34.0/ 47.0/ 48.0 47.0/ 61.0/ 60.0 50.0/ 46.0/ 51.0 51.0/ 63.0/ 65.0 49.0/ 59.0/ 59.0 55.0/ 67.0/ 64.0
Sociology 33.0/ 52.2/ 49.6 46.0/ 60.8/ 60.2 37.0/ 47.0/ 45.0 37.0/ 49.2/ 51.0 41.0/ 47.0/ 46.0 26.0/ 47.0/ 46.0 35.2/ 48.4/ 50.3 36.0/ 37.0/ 38.0 36.0/ 46.0/ 50.0 39.6/ 46.0/ 46.0 36.5/ 47.1/ 48.8
Sports and Recreation 49.0/ 69.8/ 77.7 65.0/ 87.2/ 87.0 47.0/ 78.0/ 74.0 54.4/ 73.4/ 77.0 43.8/ 74.2/ 80.3 53.0/ 78.0/ 75.0 52.0/ 74.6/ 74.6 52.0/ 64.0/ 68.0 36.0/ 72.0/ 73.0 53.8/ 70.5/ 73.1 46.2/ 68.1/ 70.0
Technology and Innovation 55.0/ 64.0/ 62.0 58.6/ 81.6/ 80.0 50.0/ 64.0/ 66.0 53.8/ 68.8/ 71.6 48.0/ 71.0/ 69.0 50.0/ 65.0/ 61.0 56.0/ 69.0/ 65.0 48.0/ 56.0/ 60.0 44.0/ 62.0/ 62.0 47.0/ 60.0/ 65.0 52.0/ 57.0/ 58.0
Transportation and Logistics 41.0/ 54.0/ 58.0 56.9/ 65.4/ 66.9 38.0/ 55.0/ 53.0 53.6/ 65.6/ 67.2 47.5/ 53.5/ 59.6 41.0/ 56.0/ 54.0 45.9/ 45.9/ 52.0 45.5/ 40.4/ 49.5 40.0/ 57.0/ 58.0 50.0/ 55.0/ 57.0 43.0/ 47.0/ 54.0
Table 51: Detailed subject-wise evaluation for meta-llama/Llama-3.1-70B on MILU across different languages. The results reported are X / Y / Z where X denote the 0-shot, Y denotes 1-shot and Z denotes 5-shot performance respectively.
Topic bn en gu hi kn ml mr or pa ta te
Agriculture 57.0 52.1 52.5 55.9 58.0 62.6 47.0 54.0 58.0 53.0 59.0
Anthropology 68.0 51.0 57.0 53.5 65.0 60.0 64.0 58.0 61.0 63.0 64.0
Architecture and Design 65.0 48.0 63.0 58.0 59.0 63.0 59.0 55.0 53.0 59.0 63.0
Arts and Culture 61.2 52.2 68.0 61.9 71.2 69.0 53.5 54.0 74.0 58.2 62.3
Astronomy and Astrophysics 80.0 70.0 84.0 71.9 79.0 76.0 67.0 76.0 74.0 74.0 73.0
Biology 64.4 55.7 68.0 70.7 69.3 64.6 57.3 56.0 67.0 65.3 61.3
Business and Management 58.0 53.6 60.0 55.5 52.0 57.0 53.0 43.0 60.0 58.0 56.6
Chemistry 64.6 61.1 69.0 64.3 64.0 56.0 57.3 46.2 64.0 59.4 64.1
Computer Science 41.1 38.0 53.0 41.0 47.3 41.0 42.6 43.0 49.0 45.9 39.5
Defense and Security 70.0 69.0 72.0 70.6 65.0 66.0 70.0 62.0 68.0 65.0 72.0
Earth Sciences 70.0 59.3 57.0 61.7 75.0 67.0 61.0 49.5 69.0 73.0 72.0
Economics 69.1 53.2 65.0 59.3 70.0 58.0 51.4 53.0 61.0 64.3 56.9
Education 57.0 55.4 45.0 56.2 60.0 49.0 50.3 55.0 44.0 51.0 48.0
Energy and Power 74.0 60.1 75.0 62.0 70.0 70.0 59.0 60.0 70.0 62.0 67.0
Engineering 52.0 47.6 52.4 45.6 51.8 45.3 47.0 42.9 43.0 44.2 50.0
Environmental Science 67.0 57.3 64.0 63.9 71.9 67.0 54.2 58.6 63.0 63.0 63.7
Ethics and Human Rights 65.0 55.0 58.0 52.0 56.0 62.0 53.0 59.0 59.0 57.0 60.0
Finance and Investment 57.9 51.6 57.0 55.0 68.5 60.0 55.4 53.0 48.0 55.9 52.1
Food Science 66.0 48.0 57.0 61.6 61.0 61.0 55.0 59.0 68.0 63.0 65.0
Geography 71.5 56.0 65.0 61.7 73.5 70.7 53.8 48.0 68.0 65.1 66.1
Health and Medicine 74.0 63.9 72.0 68.7 78.0 68.0 52.7 69.0 75.0 70.7 62.6
History 74.9 53.0 71.0 68.3 76.6 72.0 59.5 55.0 68.0 64.7 60.4
Information Technology 85.0 52.0 65.0 72.1 76.0 68.0 68.0 66.0 61.0 70.0 69.0
International Relations 69.0 66.7 69.0 65.4 63.0 65.0 66.0 61.0 66.0 68.0 63.0
Language Studies 38.0 57.2 37.0 44.3 40.3 39.0 37.6 43.0 50.0 36.7 39.1
Law and Ethics 81.1 63.4 70.0 72.0 74.5 60.0 57.8 61.0 78.0 76.4 66.2
Literature and Linguistics 64.0 50.1 61.0 52.9 65.8 51.0 62.0 54.0 64.0 47.0 58.5
Logical Reasoning 47.7 31.8 47.4 40.1 46.5 43.0 36.8 30.0 46.0 44.2 41.3
Materials Science 64.0 56.3 51.0 57.0 53.0 55.0 61.0 53.0 59.0 56.0 56.0
Mathematics 46.0 37.4 43.2 35.1 42.6 44.1 44.0 44.6 42.0 41.4 37.8
Media and Communication 68.0 57.0 67.0 59.2 65.0 64.0 53.0 63.0 61.0 68.0 69.0
Music and Performing Arts 62.0 54.0 53.0 57.0 63.0 59.0 58.0 52.0 61.0 53.0 63.0
Physics 60.4 55.8 57.7 57.6 64.4 59.0 56.6 49.6 54.0 58.2 59.0
Politics and Governance 75.7 54.3 66.0 63.0 68.8 62.0 46.4 60.0 63.0 64.7 55.9
Psychology 66.0 59.0 66.0 60.5 66.0 60.0 54.0 51.0 56.0 61.0 60.0
Public Administration 55.0 44.4 54.5 56.6 43.4 53.0 41.4 53.0 51.5 53.0 51.0
Religion and Spirituality 79.0 53.0 76.0 64.9 71.0 73.0 66.0 54.0 68.0 67.0 69.0
Social Welfare and Development 55.0 58.9 60.0 60.6 50.0 48.0 50.0 49.0 64.0 57.0 54.0
Sociology 47.0 42.2 46.0 49.0 50.0 36.0 48.4 39.0 44.0 46.0 46.5
Sports and Recreation 63.4 64.2 68.0 63.8 68.0 61.0 63.8 45.0 66.0 68.6 64.8
Technology and Innovation 65.0 59.2 67.0 59.6 64.0 63.0 63.0 56.0 61.0 58.0 57.0
Transportation and Logistics 48.0 50.8 48.0 53.0 51.5 51.0 44.9 46.5 50.0 43.0 42.0
Table 52: 0-shot subject-wise evaluation for meta-llama/Llama-3.1-70B-Instruct on MILU across different languages.
Topic bn en gu hi kn ml mr or pa ta te
Agriculture 58.0 49.9 61.6 56.7 60.0 55.6 46.0 56.0 62.0 52.0 55.0
Anthropology 68.0 52.0 65.0 55.6 61.0 60.0 69.0 58.0 64.0 62.0 63.0
Architecture and Design 69.0 51.0 61.0 60.0 57.0 62.0 64.0 58.0 60.0 62.0 60.0
Arts and Culture 66.7 50.2 68.0 60.7 73.0 64.0 57.9 53.0 68.0 58.2 63.3
Astronomy and Astrophysics 81.0 73.0 86.0 70.3 76.0 75.0 70.0 74.0 73.0 77.0 71.0
Biology 65.1 54.9 63.0 67.1 68.0 62.6 57.6 56.0 65.0 61.8 62.6
Business and Management 58.0 56.4 57.0 56.9 58.0 55.0 55.0 44.0 61.0 49.0 57.6
Chemistry 63.5 60.9 67.2 61.3 55.9 63.0 64.6 50.0 61.0 60.2 61.9
Computer Science 39.6 41.2 47.0 43.8 47.9 37.0 42.6 42.0 44.0 41.9 31.8
Defense and Security 70.0 70.0 76.0 68.7 64.0 66.0 73.0 61.0 71.0 68.0 73.0
Earth Sciences 64.0 58.2 54.0 59.9 70.0 68.0 59.0 48.5 66.0 72.0 68.0
Economics 66.2 51.0 65.0 61.7 72.0 54.0 54.9 56.0 60.0 64.3 54.5
Education 53.0 51.9 53.0 57.1 53.0 47.0 49.0 54.0 42.0 47.0 53.0
Energy and Power 67.0 62.2 71.0 62.7 66.0 66.0 68.0 60.0 66.0 60.0 66.0
Engineering 51.4 46.2 52.0 46.8 49.0 44.4 51.2 42.9 42.0 44.0 49.6
Environmental Science 63.0 56.3 65.0 62.9 71.1 62.0 55.3 54.5 66.0 69.0 60.3
Ethics and Human Rights 61.0 61.0 59.0 58.0 55.0 56.0 54.0 59.0 58.0 58.0 65.0
Finance and Investment 60.4 56.2 63.0 58.4 63.0 53.0 59.9 51.0 48.0 59.2 51.1
Food Science 66.0 55.0 61.0 60.6 61.0 59.0 56.0 57.0 61.0 62.0 59.0
Geography 72.6 57.0 64.0 61.7 73.1 71.7 58.0 49.0 71.0 64.0 65.6
Health and Medicine 64.0 63.5 74.0 69.3 75.0 71.0 54.5 70.0 76.0 68.1 65.4
History 72.8 51.8 74.0 71.0 75.1 65.0 59.9 60.0 70.0 66.5 59.9
Information Technology 77.0 55.0 66.0 67.0 74.0 68.0 70.0 63.0 66.0 59.0 70.0
International Relations 67.0 65.8 69.0 63.4 66.0 69.0 63.0 61.0 63.0 66.0 61.0
Language Studies 45.0 58.0 38.0 44.7 34.1 40.0 34.9 46.0 53.0 35.6 36.2
Law and Ethics 81.1 62.2 69.0 71.4 66.9 65.0 59.1 65.0 75.0 75.6 64.4
Literature and Linguistics 64.0 47.9 61.0 54.5 62.5 52.0 60.2 53.0 61.0 48.2 60.6
Logical Reasoning 42.5 31.8 47.4 37.5 40.2 42.0 36.8 32.0 36.0 43.4 39.6
Materials Science 65.0 58.2 52.0 57.2 59.0 53.0 62.0 44.0 54.0 52.0 56.0
Mathematics 48.0 45.0 47.6 40.7 40.2 43.5 44.2 46.2 40.5 44.0 37.4
Media and Communication 61.0 59.0 71.0 61.9 60.0 64.0 54.0 60.0 57.0 58.0 70.0
Music and Performing Arts 61.0 55.0 51.0 59.6 65.0 56.0 62.0 50.0 62.0 47.0 65.0
Physics 61.8 60.2 57.1 57.0 58.2 53.2 58.4 55.2 52.0 56.2 58.6
Politics and Governance 74.1 54.9 70.0 66.7 68.4 61.0 48.3 61.0 65.0 63.9 54.5
Psychology 64.0 56.0 65.0 56.6 58.0 61.0 49.0 48.0 57.0 66.0 57.0
Public Administration 53.0 44.4 57.6 54.5 38.4 56.0 42.4 56.0 49.5 50.0 53.0
Religion and Spirituality 77.0 57.0 76.0 67.7 69.0 73.0 67.0 56.0 67.0 70.0 71.0
Social Welfare and Development 54.0 58.0 59.0 60.6 55.0 50.0 52.0 47.0 64.0 55.0 50.0
Sociology 45.2 41.4 48.0 50.0 43.0 37.0 47.2 41.0 47.0 50.4 44.1
Sports and Recreation 68.8 62.0 72.0 65.6 73.6 68.0 67.2 53.0 66.0 66.0 67.6
Technology and Innovation 67.0 60.6 64.0 59.2 58.0 61.0 63.0 55.0 56.0 57.0 53.0
Transportation and Logistics 50.0 50.8 52.0 53.6 52.5 54.0 51.0 50.5 50.0 46.0 41.0
Table 53: 1-shot subject-wise evaluation for meta-llama/Llama-3.1-70B-Instruct on MILU across different languages.
Topic bn en gu hi kn ml mr or pa ta te
Agriculture 60 57.1 54.5 47.1 60 58.6 46 52 56 53 55
Anthropology 69 63 59 53.5 65 62 57 59 61 56 63
Architecture and Design 68 63 64 41 59 65 48 47 62 58 60
Arts and Culture 68.5 66.6 66 46.7 73 77 50.3 62 65 62.4 73.4
Astronomy and Astrophysics 74 70 77 59.4 75 74 60 70 67 75 65
Biology 69 69.9 58 51.5 65.6 70.7 45.1 51 71 60.8 59.8
Business and Management 65 59.4 61 51 63 56 47 53 62 62 48.1
Chemistry 66.2 65.1 57.8 56.1 60.5 69 51.7 39.4 63 60.2 61.9
Computer Science 44.8 44.8 48 46 49.1 49 44.4 38 52 40.1 43.5
Defense and Security 77 68 67 60.7 73 76 52 71 69 68 71
Earth Sciences 65 60 55 46.1 72 69 58 52.5 59 64 67
Economics 69.1 62.6 63 57.3 71 65 57.8 57 69 65.1 61.1
Education 59 54.5 48 48.5 52 49 54.1 47 52 47 53
Energy and Power 73 58.8 71 50.5 72 74 52 58 64 61 67
Engineering 57 59.2 51.2 43 52.2 51.7 42.4 41.4 45 48.4 54.6
Environmental Science 66 64.1 67 49.9 64.8 67 44.2 59.6 71 67 65.9
Ethics and Human Rights 64 62 73 58 66 68 53 61 67 67 68
Finance and Investment 68.8 61.8 75 53.6 67.9 57 62.7 59 51 60.3 55
Food Science 61 66 56 56.6 69 70 49 67 60 58 67
Geography 65.9 60.6 75 46.8 74.4 67.7 53.2 73 69 65.1 63
Health and Medicine 76 68.3 67 56.7 73 78 50 74 71 66.4 68.2
History 74.7 60.2 80 53.6 77.3 74 52.4 66 68 72 63.2
Information Technology 80 70 71 61.5 77 75 71 69 67 64 70
International Relations 75 75 66 64.2 70 72 61 64 69 70 65
Language Studies 43 68 41 38.5 42.6 41 39.6 49 56 41.9 36.2
Law and Ethics 79 66.6 61 61 72.6 67 50.9 64 77 71.7 70.6
Literature and Linguistics 72.1 63.1 67 50.5 65 64 59.3 54 71 46.6 60.6
Logical Reasoning 52.1 45.2 52.6 42.9 53.5 46 37.9 37 42 47.8 46
Materials Science 62 59.3 46 48.5 55 65 45 47 47 46 49
Mathematics 53.2 45.8 46.8 38.3 42.8 44.4 42.6 41.4 41.3 47 41.2
Media and Communication 57 66 71 54.4 72 66 52 56 59 60 68
Music and Performing Arts 56 65 61 51.7 64 63 57 60 66 55 69
Physics 68.2 63.8 60.2 48.2 60.8 55.3 54.8 52.8 56 58.4 59
Politics and Governance 74.9 59.7 67 54.8 69.6 68 48.9 51 56 65.1 61.6
Psychology 67 61 60 41.9 56 63 41 58 68 53 67
Public Administration 61 54.5 51.5 50.5 50.5 55 47.5 48 55.6 56 51
Religion and Spirituality 77 58 76 49.2 81 81 53 67 74 72 75
Social Welfare and Development 62 59.8 68 54.9 55 57 56 51 65 53 64
Sociology 50.4 48 45 39.4 52 45 42.8 45 55 50.4 50
Sports and Recreation 72.8 73.6 70 58.2 71.4 75 61.6 65 65 67.3 67.6
Technology and Innovation 62 63.4 65 54.2 65 69 54 60 55 53 61
Transportation and Logistics 46 53.1 60 51.1 50.5 58 38.8 52.5 58 50 51
Table 54: Detailed subject-wise evaluation for meta-llama/Llama-3.1-405B on MILU across different languages. The results reported are for 0-shot experiments.